A Complete Bibliography of Publications in the C Users Journal (1988–1994) and the C/C++ Users Journal:

1994–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

01 May 2019 Version 2.08

Title word cross-reference

#1 [Ano91b]. #381 [Vol93b]. #define [Jae88c, Pug88e, PH90, PB91]. #defines [Pug90j]. #if [Pug88y].

$19.95 [Phi92a]. 2 [Dee95, Mor98a, Mor98b, Smi92g]. 3 [Bra89, Moo91, O’D90, Smi92g, VG89b, Vil96, Vol95e]. $89 [Gra88]. μ [Gin93a, Lab92]. → [Sch96f, Sch96g, Sch96e, Sch96h].

* [HPS90, PB94, PA91].

-D [Bra89, Dee95, Moo91, Mor98a, Mor98b, O’D90, Smi92g, VG89b, Vil96, Vol95e].

-Like [Ske91].

.EXE [PY90].

12 [PH96]. 128 [Jon88b, Ock89]. 16 [Wri90b]. 1993 [Ano94a]. 1K [War90a]. 1KHz [Cor91].

2 [Ben92, Cor88, Gra93b, Jen96, Kel94, PM89, Wit90c, Wit90d]. 2.0 [Chi89, Spa90]. 2.51 [Fis90a]. 2000 [Efk97]. 2K [Pla99l]. 2nd [Bla90, Phi93b, Wei93b].

3.5in [War92c]. 32-bit [Jen96, Owe95]. 3D [VG89a].

4.x [Whe97]. 488 [And89].

5 [Jae91l]. 51 [Her94].

6.00A [Obe91]. 640K [Bri90, Pug92e].

7F [PF93a].

80x86 [Bec93]. 80x86-Based [Bec93].

90 [Ano99c]. 95 [Dan97, Tam97].

= [Pug92a]. =- [Pug92a].

setvideomode [Pug88b].

A. [Rid91]. Ability [Mat88]. abort [Jae90d]. Abstract [Sak96a, Sak96b, Sak96c]. Abstraction [All95b, Gle90b, Sch97b, Sch97c]. Academic [Pla91e]. Accelerate [Cla88]. Access [Bec98k, Bra92a, Gag97, Hei94, Jae89g, OH94, Pug90a, Pug90h, Pug90m, Pug92q, Sch88b]. Accessible [Pat88b]. Accessing [Lad89a, Pug89a, Pug92g, Spe89a]. Accidental [Pug90b]. According [Pla88d]. Accurate [For91]. Acquisition [Bar88, CM98, Fee93, FM91, San89, VC89]. Across [Lad91, Tom91b, Xia97, SX98, Tur92]. Act [Pla88f]. Action [Bar90b, Gra92b, Phi90b]. Active [Ten98, Ano97-43]. ActiveX [Par98]. ACTLIB [Ste93a]. Adaline [Phi92d]. Adam [Phi92c]. Adaptable [Cam90]. Adapters [Joh90a, Pla97l]. Adapting [Ale99a]. Adaptive [Bra90a, Sti92a]. Adding [Ric99]. Addition [Wil96c]. Addresses [Pla89g]. Addressing [Olh90]. ADU [CD90]. Advanced [AA88, Dos95, Hoh90, KS98, Phi92e, Sha92, BN94]. advantage [Ret96]. Advantages [Pug91i]. Advent [War90f]. Adventure [Cro90, LL90]. Adventurers [Gar88]. Adventures [Sch97c]. Advertiser [Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i]. Advice [War90c]. Affecting [Rib89]. Affiliate [Ano88a]. After [Sak97c, War90f]. Again [Wei93h]. Against [War92a]. Agents [Hof92]. Agley [Pla90p, Pla90q]. Agreements [Rol92]. AI [Phi99c]. AISEARCH [Bou94, Vol94a]. Al [PS92a, Smi89, Phi90c]. Alan [All93f, Bur92a, Rab89]. Alger [Kil96]. Algorithm [And94, Bel89a, Cin95, Erd89, Gag94b, Ign98, Pro92, Rév93, Tre94, dM94, Pla96j]. Algorithmics [Har87, HR92, Rom91a]. Algorithms [Bra88, Cal90, Gra95, Gre89, Ott94, Pla96g, Pla96k, Pla96l, PBS93, Syc91, Tal93, Vol96a, dM94, Lad92, Lad94, Hal93b]. alia [Plu88b]. Alignment [Jae89i, PH91a, Pug94o]. Allen [Eib94, Nel89a]. Allocating [Ast94b, Shi91]. Allocation [Dee95, Pla92n, Pug88q, Pug88t, Sch90d, Wel90]. Allocator [Bat98, Hog90]. Allocators [Pla96h]. Allowed [Deg97]. Allows [Bur89]. Alpha [Hav89b, Pen93]. Alpha-Beta [Hav89b, Pen93]. Alphanumeric [Zwa91]. Alternative [Jag95, Wei94g]. Alternatives [Pug88-30]. Ambiguity [KB99a]. American [Qua88, Sch90b, St.93]. Amiga [Wio88b]. Ammeraal [Kam89, Sch88a]. Amzi [Ano96a]. Analysis [All94a, Coc98a, Coc98b, LaC94, Sil88, Ste92a, Dor95]. Analyzer [Beh96, Bra90a]. Anderson [Hoh90]. Andrew [Rod95a]. Angle [War88h]. Animation [Hea96]. Annealing [Per92a]. Annotated [Pla94e, Sch90b]. Announced [Ano96c]. Announcement [Pug90o]. Announces [Ano88a, Ano96a, Ano97s, Ano97-33, Ano97-50, War89b]. Annual [Vol97e]. Anonymous [Wit90c]. ANSI [Pla94e, Col91, Pla91a, Pla92m, Plu88a, Plu88b, Poo93, Pug88a, Pug88y, Pug92a, Pug93g, Sch90b, TK89]. Answer [TG89]. Answered [Pla99f, Pla99g]. Answers [Bec96f, Bec96g, Bec96h, Bec96j, Bec96i, Bec96k, Bec97a, Bec97f, Bec97e, Bec97b, Bec97c, Bec97g, Bec97h, Bec97d, Bec98b, Bec98i, Bec98c, Bec98e, Bec98d, Bec98j, Bec98f, Bec98g, Bec98h, Bec98k, Bec98l, PF94]. Anthony [How93, Pla95a]. AntiPatterns [Bri98a]. Any [Ben97]. API [Vol90i, Bav99c, Bey98, Mir93]. APIs [BK94a, Bro94]. APL [Cav91]. Apple [And88a, Bel89b, Del88f, Pec89, Pug89c, Pug89s, Str88a]. applets [vSS96]. Application [Fru92, Joh90c, Kal98, Lan96, Pug88s, Ver89, Wei88b]. Applications [Cli92, Fre93, Fre88a, HD95, Irv91, JPM90, Kle97, Lar97, Plu91, Pug88h, Row95, Str88a, Tam97, Vol90f, Vol93a, War90e, Wis97, Boo91, Mar95, Lan91a]. Applied [Zol90b, Zol90a, Zol90c]. Apply [Bat97b]. Applying [Pla93b, War98]. Approach [Col89c, Col90b]. Approved [Bec98c]. Approves [Plu88d]. Approximate [Kie96, Phi94c]. Approximation [Gin93c]. Apps [Kuo97, Mit99]. April [Vol99]. Archie [Wei91a]. Archimedes [Her94]. Architecture [Man91]. Architectures [Bri98a, Pri90a]. Archive [LaB94]. Aren’t [Sut99a]. Argument [All94k, Jae89g, KB90, Kod92, PJ92]. Arguments [Ale98b, Dee96, Pug92m, Pug93s, vH92]. Arithmetic [Bak90, Bak91a, Got94, Heg92, Lab98, Par95, SS89, Sco88]. arounds [Sak97i]. Array [Cru88b, Hog90, Pug89d, Pug90c, Pug92l, Rah97, Sch98j, Sch98d]. Arrays [Ale99a, All93j, Bai91, BG92, Cru88a, Dee95, Hei94, Jae89a, Jae90h, Jae91i, Jae91a, Jae91b, LaB93a, Mar98, McC95, PA93a, Pug89g, Pug90d, Pug90u, Pug90w, PF92a, Pug93q, Pug94m, Sak92a, Shi91]. Arrow [Sch97f]. Art [Gin93b, LM90, Phi90c, Rom91b, Gan92, Pre88, Pre92, Che88]. ART-CEE [Che88]. Artificial [Bou94]. AS8 [Wil90]. asked [Wei92h]. ASN.1 [Blo91]. Assembler [Bal91a, Bal94, Pan91, PD88, PBBD91, RM92, Vol95a, Wil90].

Assemblers

[Ano90b, Ano91c, Bal89, Col88b, Col88c, Col89b, Pec89, Zen91, Col88a]. Assembly [Pat92a, Cla88, Pug89a, Pug89r]. Assembly-Language [Pat92a]. assert [Ray92]. Assertions [Bat92, Bav99a, Mue94a, Pes97b]. Assertive [Fon97, Jae90a]. Assigning [Pug90e]. Assignment [Sak92b]. Assistant [Wei88a]. Assistants [Wei94a]. Associative [McC95, Pla97k, Pla97m]. Assorted [Sak95k, Sak95j, Sak95l]. Assumptions [Bec98b]. ASxxx [Bal94]. ASxxxx [Bal89, Bal91a, Col89b]. Asynchronous [Hei99]. Atari [Hal89, Mat88]. ATL [Fan99]. ATOC [JR94]. atof [Pug89b]. Augenstein [Rid91]. author [Jae86]. Auto [Pug93p]. Auto-Increment [Pug93p]. auto ptr [Sut99b]. AutoFlow [Spe89b]. AutoFlow-C [Spe89b]. Automata [MC88, Web93]. Automated [Car97, Car99, McL91, Mea93, Sch89b]. Automatic [Pan99, Pug92d, Wei93a]. Automating [Del88a]. Automation [Ale99a]. Automatons [Cli92]. AutoQueue [Gut95]. Availability [Bec96a, Har88a, McD88, Pie88, Pug88c, Wei90a, Wio88a]. Available [Cog88b, Fie89e, Gib88]. AVL [Kar94]. Avoid [Bec98b]. Avoiding [New89]. Aware [Kuo97]. awk [Hin88a, Har88a, Bar90d].

B [Fre88a, Swa88]. B-TREE

[Swa88, Fre88a]. Babb [Pla90e]. Back [BD90b, Pug90f, Wei90f, Wei92d, Wei93e]. Backlog [Wei92a]. Backpropagation [Phi96b]. Bad [All99c, Ano91a, KB92, War90b]. Baker [Smi92c]. Balanced [Jar91b, Rie98]. Balancing [And99a, Sch96a]. Ball [Sch98c]. Ballot [Sak97a]. Barclay [Nel89b]. Barton [Swa95a]. Base [For90, Heg92, VE93]. Based [Ben98, Col89a, Col95a, Kru92, Nel96, Nel98, Phi93e, Rua97, Van96, Zol90a, dC90b, Bec93]. Basic [Jae89a, Joh90d, Phi91f, Phi95c, Pug90i, Sch90c]. Basics [All93i]. Basing [Sak98b]. Baysoft [Fer88]. BBS [Wei88c, Wei94e]. BCC [Eng93b]. BCD [Par95]. Be [BC89a]. beauty [Pic90, Hob91b]. Become [Pla94c]. BEGIN [PB91, Pla90h]. Being [Fon97]. Belief [Phi90a]. Ben [Gra89]. Benchmark [Plu88c]. Benchmarked [Cal90]. Benefits [Bru89a]. Best [Cin95, Fie88g, Pla95a, Vol98a, Por93]. Beta [Hav89b, Pen93]. Better [Ale99b, All95a, Bav99a, Bur90a, Fie88c, Hey96, Pla99a, Pla99c, Pla99b, Pla99d, Hel95]. BetterState [Ano97-43]. Between [Bri97, Kru95, Pug89g, Pug90-27, Pug92j, Rib89]. BGI [JPM90]. Bibliography [Lam89, O’D91, Ogg90a]. Bidirectional [Smi92h]. Bigtex [Fie89a]. Bimodal [Nel92a]. Binary [Isa98, Jar91b, Kel92, Pes97a, PF93a, Pug90-27, Rum98, Sch93b, Ter91, Zol90c]. Binding [Bri98g, Haz91, Swa97]. BIOS [BK91, BK94b, Bro94, Kre96]. BISON [HSC89, Coâ94a, Vol89a]. Bit [All93a, All93b, All94c, BG92, Bru91, Hei94, Jae89e, Jen96, Owe95]. Bit-Fields [Jae89e]. Bitmap [Dea93]. Bitmaps [Kie95]. bitset [Pla95b]. Bjarne [Cof88]. Black [Pla99c]. Blaise [McC89, Vol90a]. Blank [Ale91]. Blanker [Pug89n, Tal89]. Blaster [BB94, Wel94a]. blkio [Fro89b]. Bloat [Got95]. Block [Cla88, PG91]. Blow [Tom91a]. Blum [Phi92c]. Bob [Nel90b]. Booch [Mar95, Row95, Lan91a]. Book [All99a, All99b, Bri98a, Bri98b, Bri99a, Bry98, Cla97, Ham98, Lar97, Mil97, Phi99b, TG89, War90g]. Books [Ano97-44, Bar90a, Cla97, Pug90-36, War89j, You96]. Bookstore [Ano97-45, Ano97-46, Ano97-47, Ano97-48, Ano97-49]. Boolean [Bur90a, Phi94b, Sch96d]. Borland [Kel94, Rol92, Wil91, Wil92a]. BOSS [Ano88d]. Both [Pug89m]. Boundaries [SS89]. Box [Wee92, MA88]. BP [CM90]. Brackets [War88h]. Breed [Vol89a]. Brenda [Rom91b]. Bresenham [Kie95]. Brews [Wei90b]. Brief [O’D91]. Brown [Swa95b, Web96]. Browser [NW99]. browsers [Ret96]. Bryan [Hob91a]. bsed [Isa98]. BSPLINE [Vol95f]. Bubble [PN93]. buddy [Gra88]. Budget [Pip95]. Buffer [Gru90, Sti90, War88a]. Buffered [Fro89a]. Buffering [Kel90]. Buffers [BD90a, Lib90c, Lib90d]. Bug [Obe91, Pav90, Pug88b, Mag93]. bug-free [Mag93]. Bugs [Eag88, New89, Pla92a, War88b, Wee92]. Build [Cli92, Pla91b]. Builder [SR99a]. Building [BD90a, BD90b, Bur90a, Col89a, Cox92, Del88e, Sch96b, Sch99a, Vil91, War88a, War89a, Wei96, Blu92]. Builds [Phi90b]. Built [Sch99k]. Built-in [Sch99k]. Bus [Paj88]. Business [Bec95f]. buy [Gra88]. Byte [Erd95, Pla89d, Pug88-29].

C

[All93f, All94j, All95c, All98, All99b, And88b, Pat92a, PS92a, Ano94b, Ano97u, Ban89, Bin88, Bla90, Bri96, Bri99a, Bry98, Bur90b, Cas93, Chi89, Cla90c, Clu89b, Cof88, Cog89, Dos95, Eib94, Fis90a, FD92, Hal93b, Hal93c, Hat94, Haw88, Hob91a, Hoh90, Kel94, Kil96, LK97, Lop93, McL94, Mil97, Mul93, Nel89a, Nel89b, Nel90b, Ogg90b, Par88, Pat96, Pav89, Per89, Phi90c, Phi92c, Phi93c, Phi93b, Phi94a, Pin93, Pla90c, Pla91d, Pla91e, Pla92d, Pla93f, Pla93k, Pla94e, Pla94b, Pla94c, Pla94t, Pla95a, Pla96h, Pla96g, Pla96i, Pla97s, Pla97q, Pla97o, Pla97k, Pla97m, Pla97l, Pla97n, Pla97r, Pla97p, Pla98k, Pla98h, Pla98f, Pla98g, Pla98e, Pla98j, Pla98i, Pla98d, Pla98c, Pla98a, Pla98b, Pla98l, Pla99i, Pla99e, Pla99j, Pla99k, Pla99h, Pla99a]. C [Pla99c, Pla99b, Pla99l, Pla99d, Pla99f, Pla99g, Pug88y, Pug90-31, Pug91j, Rab89, Rod95b, Rod96, Row95, Sak99d, Sch90d, Sch96f, Sch96g, Sch96e, Sch96h, Sic94, Sil88, Smi89, Smi92c, Spa90, Str99, Swa95a, Tho92, TG89, Vol89b, WB88, War90d, Web96, Wil92a, Wio89, vdP93, Abo91, Ake93, Alg95, All88, All89, All93a, All93b, All93c, All93d, All93e, All93m, All94c, All94d, All94l, All94m, All95a, All95d, All95e, All95f, All96a, And88a, AA88, Ano88a, Ano88d, Ano91a, Ano91b, Ano97-33, Ano98b, Ano98a, Ano99a, Ano99b, Ano99c, Art90, Ash90, Bai91, Bak90, Bak91a, BG92, Bak92, Bal89, Bal91b, BS94, Ban94, BD89a, BDB89, Bar88, Bar95a, Bar89, BG94, Bar91b, BN94, Bat97a, Bec95e, Bec97f, Bec97c, Bec98c, Bec99b, Bed91, BT91, BSG92, Bis90]. C [Blu92, Bra92a, Bro89, Bro90a, Bro90b, Bru90, Bru92a, BC89b, BC89c, BC89d, Bur90a, Cam87, Cam93, Cam98, Car92, CE95, CE96a, CE96c, Cas89, Cav91, Che96, Chu91, Cla99, Cla90b, Cog88b, Coh92, Col89b, Col91, Col93a, Col94, Col95b, Cro99, Cru88a, Dac95, Dan98, DM88, Der88, Dic94, Dod88a, Dod88b, Dod91, Doe97, DN93, Dor95, Dug95, Ebn91, Eck89, Erd88, Erd93, Erd95, Eve93, Fea95a, Feu86, Fie88a, Fil88, Fis90b, Fon97, Fos93, Fra88, Fre93, Fre88a, Fre88b, Fro92, Fuk89, GR89, Goo89, Got94, GH99, Gre94, Gui97, Gut94, HJ90, HW92b, Han94, Hav89c, Hav89a, Hav90, Haz91, HD95, Hel95, Hem96, Hen88, Her94, Hin90a, Hol87, Hol92, Hun85, Hun89, Irv91, Jae86, Jae90c, Jae93, Jan93a, Jar90, Jar91b, Jon88b, Kal98, Kam94, KK89]. C [KB92, Kel91, Kel92, Kem87, KR78, KR88, Kis96, KO94, Koc83, Koc88, Koc94, Kod92, Kra89, Kru92, Kwo89b, LaB93b, Lab92, Lad91, Lad92, Lad94, Lam97, Lan91b, Lan91c, LAT96, Lan99a, Lan96, Lan99b, Let91, Lib89b, Lib90a, Lib92, LM94b, LM94a, Low95, Mag93, Mah99, Man91, Man96b, Man89, Man90, Mar92, Mar88b, Mar95, Mas95, Mas97, McC89, McD88, McK98, Mea92, Mer96, Mes96, Mey92, Mey96, Mey98, Mic97, Mil94, MQ87, Mir93, Mue94b, Mue94a, Mul89, Mus90, Nar89, Nol94, Obe91, Ogg90a, Olh90, Oua92, Pas89, PBBD91, Pav90, Pea90, Pen91, Per92b, Per92c, Pet98, Phi91a, Phi92a, Phi93a, Phi94b, Phi95a, Phi95b, Phi95c, Phi96a, Phi96d, Phi98b, Pie88, Pin91, Pip94, Pla88a, Pla88b, Pla88c, Pla88e, Pla88d, Pla88f, Pla88g, Pla89c, Pla89d, Pla89g]. C [Pla89h, Pla90m, Pla91a, PG91, Pla92b, Pla92j, Pla92m, Pla93d, Pla93c, Pla93i, Pla93h, Pla93a, Pla93n, Pla94a, PF94, Pla94n, Pla94q, Pla96b, Pla96a, Pla96j, Pla96k, Pla96l, Pla97j, Pla97i, Pla99i, Plu84, Plu88c, Plu89a, Plu91, Poo93, Por93, Pot88, Pre88, Pre92, Pri89a, Pri89b, Pro89, Pru91, Pug88d, Pug88c, Pug88h, Pug88f, Pug88q, Pug88u, Pug88z, Pug88w, Pug88i, Pug88-31, Pug89c, Pug89i, Pug89j, Pug89r, Pug89s, Pug90d, Pug90i, Pug90v, Pug90-28, Pug90-30, Pug90-35, Pug90-37, PS91a, Pug91a, PN91, Pug91g, PdW92, Pug92s, Pug93c, Pug93a, Pug93g, Pug93n, Pug94a, Pug94g, Pug94h, Pug94n, Pug94p, Pur89b, Pur92, Rag95, RN93, RP91, RY91, Ret96, Ros92b, Sak93f, Sak95d, Sak95g, Sak95c, Sak96a, Sak96b, Sak96c, Sak96e, Sak96d, Sak96f, Sak96g, Sak96j, Sak97d, Sak97a, Sak97b, Sak97f, Sak97e, Sak97g, Sak97h, Sak97c, Sak97i, Sak98a]. C [Sak98b, Sak98d, Sak98e, Sak98f, Sak98c, Sak99c, Sak99h, Sak99f, Sak99e, Sak99g, Sak99a, Sak99b, Sak99d, Sch91a, Sch85, Sch90b, Sch90c, Sch96b, Sch96f, Sch96g, Sch96e, Sch96h, Sch96i, Sch97b, Sch97c, Sch97a, Sch97d, Sch97e, Sch97g, Sch97i, Sch97f, Sch97h, Sch98a, Sch98e, Sch98g, Sch98j, Sch98i, Sch98d, Sch98b, Sch98k, Sch98c, Sch98f, Sch98h, Sch90e, Sch89b, Sch93a, Sch88c, Sch92c, Sco93, Sco92, Ser90, Sha92, SR99a, Sil88, SD89, Siy89, Ske91, Ske93a, Ske93b, Sma95, Smi92d, Smi92e, Smi93c, Smi94a, Spe89b, Spu92, Ste87, Ste90, Ste92b, Str88a, Str86, Str87, Str91b, Str95, Swa95d, TC94, Tal93, Tay92, Tee90, TLA90, TC95, TG89, Tow88, Tri93, TK89, Tuc98, Tur92, VC89, Vig89, Vil91, Vol90a, Vol90b, Vol90i, Vol90l, Vol91a, Vol91b]. C [Vol93a, Vol94a, Vol95g, Vol96e, Vol96d, Vol96c, Vol96b, Vol96f, Vol97a, Vol97d, Vol97c, Vol97b, Vol98a, Vol98b, Vol99, WK92, WPM84, WPM87, Wal89, War86, War89j, War89i, War90b, War90f, War92a, War91c, Wat93, Web96, Wei93j, WF90, WF92, Wel96, WS90, Whi91, Whi98, Whi90b, Wil93c, WR93, Wio88c, Wit95, Yok88, Zei91, Zol90d, dC90a, vdB91, Cad91, Far88, Gin93a, Kou88, Ogg90c, Pri88, Rid91, Tho93, Wal88, Wal89, Wal92]. C-51 [Her94]. C-ACROSS [Tur92]. C-Clearly [Smi93c]. C-INDEX [Fre88a]. C-scape [Ash90]. C-Window [Ebn91]. C/C [FD92, Pla94c, Pla94t, Pla95a, Pla96h, Pla96g, Pla96i, Pla97s, Pla97q, Pla97o, Pla97k, Pla97m, Pla97l, Pla97n, Pla97r, Pla97p, Pla98k, Pla98h, Pla98f, Pla98g, Pla98e, Pla98j, Pla98i, Pla98d, Pla98c, Pla98a, Pla98b, Pla98l, Pla99i, Pla99e, Pla99j, Pla99k, Pla99h, Pla99a, Pla99c, Pla99b, Pla99l, Pla99d, Pla99f, Pla99g, Rod95b, Ano98b, Ano99a, Ano99b, Fon97, Hel95, Mic97, Mue94b, Pla96b, Pla96a, Pla96j, Pla96k, Pla96l, Por93, Ret96, Sch96f, Sch96g, Sch96e, Sch96h, Sch96i, Sch97b, Sch97c, Sch97a, Sch97d, Sch97e, Sch97g, Sch97i, Sch97f, Sch97h, Sch98a, Sch98e, Sch98g, Sch98j, Sch98i, Sch98d, Sch98b, Sch98k, Sch98c, Sch98f, Sch98h, Sch93a, Swa95d, Tal93, Vol96e, Vol96d, Vol96c, Vol97a, Vol97d, Vol97c, Vol97b, Vol98a, Vol98b, Vol99]. C/OS [Lab92, Gin93a]. C9X [Jae94, Sch98b]. C wndw [Str88b]. Caching [Nel96]. CAD [Bra92b]. Calculating [Mah99]. Calculations [Ber92b]. Calculators [Fie89b]. Calendar [dM94]. calendars [O’N75, O’N78, Pla93g]. Call [Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, BT91]. Call-Tree [BT91]. Callbacks [Dug97]. Calling [Mar88b, Pug92b, Pug94r, Vol90b]. Calls [Bec98g, Bed95, Hei99, Pug91h, BK91, BK94b, Bro94]. Calvert [Ast94a]. Campbell [Wio89]. Can [Daw95, Hin88a]. Capabilities [Sma88]. Capable [Fie88d]. Capture [Gra91]. Card [Sch90a]. carefully [Wil96b]. Cargill [Hal93c]. Carl [Ogg90c]. Carroll [Bri96]. Carry [Cog88a]. Carta [Vol91a]. Case [War92a]. Casting [Bat97a]. Castles [Sch99a]. Casts [All94f, Pug90g]. Catching [Bec98d]. Cause [KB99a]. Caution [Bec98l]. CBEMA [Pla93e]. CCITT [Fel90]. CD [Vol99, Bro94, Sak95d, Sak95g, Sak95h, Sak97a, Wei94c]. CD-ROM [Vol99, Bro94]. Ceases [Pla92d]. CEE [Che88]. Cellular [MC88, Web93]. Central [Pla88g]. cerr [Sin93a]. Certain [Jae88c]. Certification [Hed91]. CES [Jar91a]. CForms [Ber94]. CGI [Che96, Kam97, Lam97]. chain [Wil96b]. Challenge [Pla89g]. Challenging [Del88c]. Change [War89b]. Changed [Pla88c]. Changes [Pla90n, Pla90o, Pla93i, Sak95e, Sak95k, Sak95j, Sak95l]. Changing [Mul93, Pug89d]. Chao [Spa90]. Chaos [Hob91b, Pic90]. char [Pug90-42, PF93a, Shi91]. Character [Col89a, Pla89d, Pla90a, Pla91c, Pla93m, Pla93l, Pla93q, Pug88-29, Pug89e, Pug93b]. Character-Based [Col89a]. Characters [Coo92a, MP90, Ock89, Pla90b, Pug90l, PS91b, PS92b, Sea91, Smi94b]. Charles [Ast94a]. Chart [Pat88a]. Charter [Jae94]. Charts [Cel90]. Chebyshef [Pri92a]. Check [Fel90, Pug92f, PBS93]. Checked [Mas95]. Checking [BS94, Joh99b, Pug92r, Pug94n, Whi89b]. Checksums [Mah99]. Chemical [Kar90]. Chest [Nel89a, Hol87]. Chien [Spa90]. Child [Wei96]. Chronograph [Mes96]. Cilwa [Swa95c]. Circular [Jae91g, Jae91h, War88a]. Class [All93c, All93d, Ano96a, Bai99, Bar95a, Bel98, Bou94, Bre96, BS99, Cha97, CM98, Cra97, Cro98b, EU98, Faz98, Fos93, Guf98, Gut95, Har97, Mar92, Mar96, Mas97, Mes96, Mor93b, Per99, Pug94b, Pug94c, Rah97, Roe99, Sak91e, Sak91f, Sak95i, Sak97b, Sch96l, Tay95, Tik98, Tuc98, Van92, VE93, Vol94a, Wal99, Wei95e, Wen97, Wil92a, WL94]. Class-Specific [Sak97b]. Classes [Ada94, Ano97-33, Bec95f, Blu91, CE96a, Edm98, GH99, Jan93a, Kam93, Mor92, Pla88e, Pug93d, Pug93s, Sak91d, Sak93e, Sak94c, Sak94d, Sak94e, Sak94f, Sak94g, Sak94h, Sak98c, Sak99c, Sch91a, Swa97, Van95, Wel96]. Classification [Pla90a]. Cleaning [Sch98i]. Cleanup [Bec99c, Bec99b, Bec99d]. Clearing [Pat92b]. Clearly [Smi93c]. Clement [Dos95]. Client [Col95a, Tik98, Vol96g]. Client/Server [Col95a, Vol96g]. Client/Server-Based [Col95a]. Clients [Bri98g]. Clifford [Hob91b]. Clipper [Sch93a]. Clipping [Rév93, Ste89]. Clones [Cor91]. CMENU [Zol92e]. Code [Ano94a, Ano94b, Ano97-43, Bec96i, Cam88, Com90, CW89, For90, Fri90, Ges91, Got95, Lib89b, McC93, Nol94, Pan99, Pat99, PG91, Pug88c, Pug88v, Pug90o, PS91a, PH91b, Pug92s, Pug94k, Sil88, Sta89, Ush94, Vol96c, Vos90, Wal99, War89g, War89k, Wei89, Wri90b, Yok89, Zol91c, Zol91d, Mag93, Pla96a, Ret96, Hal93a]. CodeRunneR [Vol90h]. Codes [Gib90, PR91]. Coding [BS94, Bri96, Eng93a, Phi92b, Pri89a, Pri90a, Pug91a, Smi94b, Ver89, CE95]. COFF [Pes97b]. Coffeehouse [Vol96g]. collate [Pla98d]. Collection [Col95b]. Colley [Pec89]. Color [Ash95, Edm98, Jos93, Kie96, Pro92, Sch91d, Vil96]. Color-Space [Edm98]. COM [Bri98f, Bri98g, Hei99, Lan99a]. Combination [FB96]. Combining [Mey96]. Coming [Ano99z]. Comma [Jae88a]. Command [Bai99, Bat97b, Bur89, Col90b, Dan93, Kou97, Lib90e, Lib91c, Mal90, McM91, Mea91, Pug92m, PA93b, Smi92f, Wei96]. Command-Line [Kou97, McM91]. Commands [Pas89, Rat91]. Comment [Tsa90]. Comments [Hed91]. Commercial [Ano96b, Van93, War89g]. Committee [Plu88a]. CommNet [Fie88f]. Commodore [Fil88, Jon88b, Ock89, Pug88d]. Commodore-64 [Fil88]. Commodores [Wio88a]. Common [Jae89h]. Communication [Gil90, KS98, Wit90d, Woo97]. Communications [Chu91, Fre90, Gle90c, Law95, PH96, Sak89, Smi92h, Vol91a, Whi89c, Wio89, Cam87, Cam93, Hel94]. comp.sources [Fie88g]. comp.sources.reviewed [Wei91c]. comp.sources.unix [Wei90f]. Compact [All99b]. Company [Fra88]. Compared [Pug89k, Pug89b]. Comparing [Ano99c, Bra91, Cro99, Pla92c]. Comparisons [Zwa91]. Compatible [Her94, Pat88a]. Compendium [PA93b]. Compilation [Sak94a, Sak94b]. Compile [Pec89]. Compiler [And88a, Bro90a, Hed91, McD88, PBBD91, Phi92a, Pie88, Pug88c, Rol92, Tsa94, Vol89b, War89h, Zol92a, Zol92b, Zol92c, Zol92d, Zol92e, Hen88]. Compilers [AT97, Lad91, Pug88-29, Pug89c, PH90, RP91, Vol90l, War89j, Wio88a]. Compiling [Fil88, Pug88z, Pug93c]. Complete [Del88b, Hob91a, War91a, Hel91, Hel94, McC93, WF90, WF92, Ush94]. Completing [Sch95a]. Completion [Man98]. Complex [Bak90, Bak91a, Bec98i, Del88c, Lad89b, Pri93, Wei88a, Wri90a, Pla95c, Pla95i]. Complexity [Dav92]. Compliance [Efk97]. Component [SR99a]. Components [Hal93b, Lad92, Lad94]. Compound [Ale98a, CH88]. Comprehensive [Fre88b, Spu92, Tho93, Wal89, Lad92]. Compression [Dis88, Gag94b, Gag97, Ign98, Iof99, Phi92b, Ros92a, Ros94b]. Computations [Per99]. Computer [Amm87a, Kam89, Rom91b, Sch88a, Amm86, Amm87b, Amm92, LM90]. Computers [Hob91b, Kru95, Pic90, Pug88d]. Computing [McC89, Har87, HR92, Pre88, Pre92, Phi90c, Rom91a]. COMX [CD90]. Concepts [BD89a]. Concerning [Bec98b]. Concerns [Bec98h]. Concise [HPR99]. Concurrent [Pla90c, Wei94b, GR89, McD88, Pie88]. Conditional [CH88, Pla89a, Pug92u]. Conflicts [Pug91f]. Conforming [Erd88]. Confusion [Bec95a]. CONNECT [Hav89b]. Connection [Hem96]. connectionist [Blu92]. Connections [Kis96]. Conquer [Vol91b]. Conquerrent [Vol91b]. Considerations [Ver89]. Console [Mit99, Tam97, Wil92a, Wil92b, WP88]. const [Ano98a, Haw97, Jae89b, Pug88e, Pug91d, Sak96d]. Constants [Cla90a, Pes97a, Pug88-29, Pug90k, Pug93h, Pug94k]. Construct [Pug90n]. Constructing [Mor99]. Construction [Vol90f, McC93, Won93]. Constructor [Dee96, Pug94f]. Constructors [Bec96g, Pug93m, Pug94a]. Consuming [Yue94]. Contact [Sch98h]. Container [Nel98, Phi99a, Pla97l, Sak94c, Sak94d, Sak94e, Sak94f, Sak94g, Sak94h]. Containers [Ale98b, Ben98, Pla97k, Pla97m]. Containing [Bec99a]. Content [Ban89, Clu89b]. Contest [Ano91a, KB92, Lib89b, Nol94, Pug90o, War90b]. Contiguous [Pug90h]. continued [Jae92f]. Contrast [DE98]. Contributing [Ano98b, Ano99a, Ano99b]. Control [All88, All94e, All99c, And89, Bai99, Del88c, Fee93, Har97, Hin88b, Jae90b, Joh88b, Kat99, Man90, Phi93f, Pug90l, Pug93v, Wei88a, Wei88c, Wei93a, Whi88, Pug88w]. Control-C [All88]. Control-Characters [Pug90l]. Control-Z [Joh88b].

Control/PrtSc [Pug88w]. Controller [Wat93]. Controllers [Kra89, OH94]. Controlling [Gab99, Gru90, Pla88c, San93, Sch96c, Whi89a]. Controls [Bug94, Dow96, CD94, Swa95c]. Controversy [Wei90b]. Convention [Hed90]. Conventions [BS94, Pug94e]. Conversation [Sue94]. Conversion [Bec96h, Hea97, Mey95, Pla91d, Pla93o, Pri91a, Pug88f]. Conversions [All94f, Bur93, Edm98, Hic89, Pla92o, Sch96c, Wei90h]. Convert [Vol93a]. Converter [Ret96]. Converting [Pug88h, Pug88g, Pug90i, Pug94d]. Convolution [Fal90a, Fal90b]. Coprocessor [Par95]. Copy [Pri92b, Whi91]. Copy-On-Write [Whi91]. Copyrights [PBW92]. CORBA [And99a, RL98]. CORDIC [Ber92b, HMF94]. Core [Web96]. Coroutines [Ake93, Eng93b]. Corporate [Bru89a]. Correctness [Plu89b]. Corrigendum [Pla94r]. Corrupted [Jan93b]. Corruption [Joh88b, Wil98]. Cortesi [Pro90]. cos [Ber92b]. Cost [Rob95]. Costs [Abo91]. Count [Daw95]. Counting [Ano98m, Mey98, Nga97, PS91b, Rec98, Sch91b]. Couple [Blu91]. Courses [Ano97t]. cout [Sin93a]. CP [Fuk89, Hic89, McD88, Ock89, Pug90-28]. CP/M [Fuk89, Hic89, McD88, Ock89, Pug90-28]. cpio [WH88]. CPP [Rea90b]. Craft [Cam94]. CRC [Mah99, War90a, Wri90b]. CRC-16 [Wri90b]. Create [Bec95a, Joh90d, Kuo97]. Creating [Bec96b, Bec98e, Bug94, Coh92, Cur98, Flo98, Pug92s, Sch96d, Ske91, Ten98, Tha98, vSS96]. Creation [Rin99]. Creative [War92b, Sti92b]. CRecordset [Swa97]. Crisis [Bri98a]. Critical [Pug92c, Wei92i, Wei95e]. Critical-Sections [Wei95e]. Criticism [WC90]. Criticisms [Mar91]. Criticized [WS90]. CROBOTS [Poi89b, Poi89a]. Cross [Ano90b, Ano91c, Bal89, Bal91a, Bal94, Col88b, Col88c, Col89b, Del88e, Pan91, PD88, Pec89, RM92, Smi92g, Vas96, Vol95a, Vol97a, Wil90, Zek95, Zen91]. Cross-Assembler [Vol95a]. Cross-Platform [Vas96, Vol97a, Zek95]. Cross-Reference [Del88e]. Cross-Sections [Smi92g]. Crude [Jae89c]. Crunching [Smi92e]. CTask [Wag90]. Ctrl [Let91]. Ctrl-C [Let91]. ctype.h [Pla90i]. CU [Ber88]. CUG [Vol99, Ano90b, Ano91c, Cam88, Dod88a, Gib88, Hin90b, Pec89, Rup88, Vol90l, Vol93b, Vol93c, Vol97e, Vol98b, War88b, War88c, War89c, War89g, Yok89]. CUG155 [Swa88]. CUG159 [Cro90]. CUG163 [Lan91b]. CUG172 [Ano88b]. CUG173 [Ano88b]. CUG197 [Man88]. CUG198 [Ano88c]. CUG199 [Tea88a, Tea88b]. CUG209 [Ano90a]. CUG220 [Ano88d]. CUG223 [Fuk89]. CUG226 [Che88]. CUG227 [Ger88]. CUG236 [Smi88a]. CUG242 [Col88a]. CUG243 [Art90]. CUG247 [Sco88]. CUG248 [Art89]. CUG252 [Dod88a]. CUG254 [Yok88]. CUG255 [Ced88]. CUG256 [Cog88a]. CUG257 [Dod88b]. CUG258 [Dod88b]. CUG259 [WP88, Wit88]. CUG260 [Ber88]. CUG261 [PD88]. CUG262 [Ram88]. CUG263 [Str88b]. CUG264 [VM88]. CUG265 [WH88]. CUG266 [Pat88a]. CUG267 [Col88b]. CUG268 [Mil88]. CUG270 [BBAH88]. CUG271 [Mar88a]. CUG272 [Mar88a]. CUG273 [Der88]. CUG274 [Cru88a]. CUG275 [MC88]. CUG276 [Col88c]. CUG277 [Flo89]. CUG278 [Sme89]. CUG279 [Poi89b]. CUG280 [Ano89]. CUG281 [Mil89]. CUG282 [SB89]. CUG283 [Osh89]. CUG284 [Sek89]. CUG285 [HSC89]. CUG286 [Kwo89a]. CUG287 [Kwo89b]. CUG288 [LFK89]. CUG289 [CW89]. CUG290 [HP89]. CUG291 [Bro89].

CUG292 [Bal89, Col89b]. CUG293 [VG89a]. CUG294 [VG89a]. CUG295 [Fro89b]. CUG296 [Hav89c]. CUG297 [dF90a]. CUG298 [Dea90]. CUG299 [CM90]. CUG300 [Hug90]. CUG301 [JPM90]. CUG302 [O’D90]. CUG303 [Col90a]. CUG304 [Ber90]. CUG305 [Sch90a]. CUG306 [Col90c]. CUG307 [CD90]. CUG308 [HKB90]. CUG309 [Bro90a]. CUG310 [dF90b]. CUG311 [Har90]. CUG312 [Yeh90]. CUG313 [And90]. CUG314 [Pea90]. CUG315 [Clu90]. CUG316 [Wil90]. CUG317 [Mar90a]. CUG318 [Rea90a]. CUG319 [Rea90b]. CUG320 [Fal90a]. CUG321 [Cur90]. CUG322 [RSSS90]. CUG323 [LL90]. CUG324 [Gie90]. CUG325 [Zia90]. CUG326 [Sch90f]. CUG327 [Bro90b]. CUG328 [Blu90, Blu94]. CUG329 [dFAAW90]. CUG330 [Wag90]. CUG331 [Osb90]. CUG332 [Lar90]. CUG333 [Wit90a]. CUG334 [Kot91]. CUG335 [Zen91]. CUG336 [YEA91]. CUG337 [Pin91]. CUG338 [PBBD91]. CUG339 [Let91]. CUG340 [Ebn91]. CUG341 [Lon91a]. CUG342 [Mil91]. CUG343 [Phi91a]. CUG344 [Ano91b]. CUG345 [Sch91e]. CUG346 [Bal91a]. CUG347 [Hug91]. CUG348 [Pan91]. CUG349 [Pol91]. CUG350 [Ash91a]. CUG351 [Hac91]. CUG352 [Blu91]. CUG353 [Dod91]. CUG354 [Jar91a]. CUG355 [Rea91a]. CUG356 [Rea91b]. CUG358 [Fro92]. CUG359 [FD92]. CUG360 [McC92]. CUG361 [Ekw92]. CUG362 [Coo92b]. CUG363 [RM92]. CUG364 [Tur92]. CUG365 [Kir92]. CUG366 [Law92a]. CUG367 [MO92]. CUG368 [Ohl92]. CUG369 [Whi92]. CUG370 [Lie92]. CUG371 [GC92]. CUG372 [Mor92]. CUG373 [LP92]. CUG374 [Law92b]. CUG375 [Gra93a]. CUG376 [Gra93b]. CUG377 [Gra93c]. CUG378 [Gra93d]. CUG379 [Gra93e]. CUG380 [Joh93]. CUG381 [Lan93]. CUG382 [Gai93]. CUG383 [Mor93a]. CUG384 [Deu93]. CUG385 [Eng93b]. CUG386 [Dav93]. CUG387 [Tal93]. CUG388 [How93]. CUG389 [Gag93]. CUG390 [Ste93a]. CUG391 [Mue94b, Vol95g]. CUG392 [Arc94]. CUG393 [HMF94]. CUG394 [LM94b]. CUG395A [The94]. CUG395B [Kar94]. CUG395C [Saw94]. CUG396 [Ste94]. CUG397 [Nol94]. CUG398 [Bal94]. CUG399 [Wol94]. CUG400 [Swa94]. CUG401 [TC94]. CUG402 [Ber94]. CUG403 [KO94]. CUG404 [Coâ94a]. CUG405 [Coâ94b]. CUG406 [JR94]. CUG407 [Dug94]. CUG408 [Fan94]. CUG409 [Ben94]. CUG410A [All94a]. CUG410B [Hat94]. CUG411 [Moo94]. CUG412 [Bou94]. CUG413A [Wel94a]. CUG413B [BB94]. CUG414 [Hes94]. CUG415 [Ban94]. CUG417 [Vol94b]. CUG418 [Vol94b]. CUG421 [Vol94c]. CUG427 [Vol95b]. CUG428 [Vol95b]. CUG429 [Vol95b]. CUG430 [Vol95c]. CUG431 [Vol95d]. CUG434 [Vol95e]. CUJ [War89b, WC90]. Cure [KB99a]. Current [Lop94]. Curses [Dea90, PM91a, Wei93i]. Curve [Ano90a, Gra92c, Pri92a, Sch96e, Smi93a]. Curves [Bra88, Bra90b]. Custer [Hal93d]. Custom [Swa95c, CD94]. CWEB [Wit95]. CXL [Sme89]. Cyclic [Fel90, Lou92].

D [Bri96, Pla90c, Bra89, Dee95, Moo91, Mor98a, Mor98b, O’D90, Smi92g, VG89b, Vil96, Vol90k, Vol95e]. Daconta [All95c]. Daemon [Kat99]. Dan [Phi93c]. Danger [Bec98j]. Darnell [Wal88]. Data [All95b, Bar88, Bra90a, Bra90b, Cam98, CM98, Cla88, Fee93, Fie88a, FM91, Gag94b, Gag97, Gib99, Gle90b, Haw97, Ign98, Jae91i, Jae91a, Jae91b, Jae91c, Jae91d, Jae91e, Jae91f, Jae91g, Jae91h, Jae92h, Jae92a, Jae92b, Jae92c, Jae92d, Jae92e, Jae92f, Jae92g, Kel91, LAT96, Phi92b, PB92, PF93a, Pug88h, Pug90e, Pug90-29, PY90, PH91a, Pug92h, Rid91, Ros92a, Ros94b, San89, Sch96b, Sch96k, Sto95, TP92, TLA90, Ten98, VC89, War88a, Wel95]. Data-Compression [Ros92a]. Database [Fro92, Rag95, Sch90f, Sch92b, Vol97d, Zol90b, Zol90d]. Databases [All96b]. Date [All93c, All93d, All93m, Bur93, Cla90b, Iof99, Mil94]. d’Autopointer [Sch98e]. David [Per89, Pro90, Rom91a, Sic94, Tho93]. Davis [Dav93]. days [Cal93, Cal95, Ast94a]. DB [Har90]. dBASE [Pug88f, Pug88h]. DBWin32 [Tuc96]. DCOM [Fan99]. DDE [Bav98a]. Deactivating [Pug89f]. Deadlocks [Rog95]. Dealing [Wil94]. Debug [Cha90, Jag95]. Debugger [Lan90, Mat89, Tuc96, Vol88a]. Debugging [Bal91b, Bat92, Cha90, Far89, Kau93, Lan90, PP90, Phi96a, PD93, Rea92, Shi90, Sin93b, Smi92b, Sti92b, Ves95, War91b, War92b, Wel93, Whi90b, Wis99]. Debunker [KB99d, KB99c, KB99a, KB99b]. Decision [Phi98a]. Decision-Making [Phi98a]. Decisions [Sak99a, Sak99b]. Declarations [Pug93g, Sak95c, Sak96f, Sak96g, vdP91]. Declarator [Mer96]. Declarators [Sak96a, Sak96b, Sak96c, Sak96e, Sak96j]. Declaring [Pla88a, Pug92d, Wit90b]. Decoding [Ano93, Blo91, Mar90b]. Decomposition [Rum99]. DECUS [Art90]. Deeper [Sak99e, Sak99g]. Deeply [Sak99f]. Default [Dee96, Sch99k]. Defect [Cla99]. Defined [Pug90-29, Mey95]. Defining [Bec96g, Pen91, Pla89b, Pug90k, PB91]. Definitions [Jae89h]. delete [Sak97d, Sak97b, Sak97e, Sak97g, Sak97h]. delete-expressions [Sak97g, Sak97h]. Demo [Pug90-35]. Dempster [Phi90a]. Dempster-Shafer [Phi90a]. Dependencies [Flo90, PF92b]. Depicting [Pla90b]. Depth [Phi99d]. deque [Pla97o, Pla99a]. Derivatives [HPR99]. Derived [Pug93s]. Description [Pat88a]. Design [BD90d, Bar95b, BC89a, Bur90b, Coc98a, Coc98b, Del88a, Gam95, Jae90e, Lan91a, Nar89, Pat95, Pug94b, Rom91b, Sak98b, Sak99a, Sak99b, Ver89, War89h, Boo91, Dor95, Hen88, LM90, Mul89]. Designer [Ano97s]. Designing [BD90c, Bri96, CE95, CE96a, Mar95, Mir93, Pin91, Row95, Sak94c, Sak94d, Sak94e, Sak94f, Sak94g, Sak94h, Tsa94, Zek95, Lad92]. Designs [Rat90, Mey92]. DeSmet [Cog88a]. Desperately [Vol97b]. DESQview [Pug93f, Vol90i]. Destruction [Gab99]. Destructors [Pug93m]. Detecting [Joh90a, Web93, Wil98]. Detection [Han90, Phi91f, Phi92e, Pug92f]. Determining [Zer89]. Developing [Bec93, Mus90, Pla93f, Mag93]. Development [Dic94, Pri90c, St.93, Wei94b]. Device [All89, Ben92, CD90, Erd88, Joh90b, Law92c, Man90, Nel92b, Nor92, Paj88, Rot91, Wei93b, ET88, ET92]. Devices [And89]. Diagrams [Bar90b, Phi90b]. Dialing [Fie88a]. Dialog [Bar95b, Wel94b, Wel95, Wel99]. DIB [Jos93]. DIBs [Knu98]. Did [Wei94f]. Difference [Gag94a]. Differences [Pug89g, Rib89]. Different [Mat89, Pug90e]. Digit [PBS93]. Digital [Bak91b, Mor99, RP92]. Digits [Pug92f]. Dimensional [Ber92a, Jae91a, Jae91b, LaB93a, Ols93, Pug93q]. Direct [Bra92a, Whi88]. Directing [Pug90-32]. Direction [WB88]. Directly [Spe89b]. Directory [Bel89b, Cra97, Zol90e]. Dirty [Ros97]. Dis [Sch98d]. Dis-Array [Sch98d]. Disadvantages [Pug91i]. Disappear [Pug90t]. Disassembler [Col90a, Pan91]. Disassembler/Z80 [Pan91]. Discipline [Phi96c, Hum95]. Discrete [HJ90, LM94b]. Disk [All94a, Ano94a, Cam90, Mat88, Nel96, Nel98, Pug88u, PG92, Tom91b, War88b, War88c, Whi88, Yok89, Zol90a, All99b]. Disk-Based [Nel96, Nel98]. Disk-Reading [Mat88]. DISKED [JR94]. Disks [Cam88, Hal89, Hin90b, Pug88j, PK92, War89c, War89g, War92c]. Dispatch [Gou98]. Displayable [Pla90b]. Displaying [Phi91c, Phi91d, Pug90l]. Displays [Pot88]. Distinct [Pla88e]. Distributed [RL98, Rod95a, Tan95]. Distribution [Fan99, War89c, Wei93a]. Dithering [Rie98]. Divided [Gag94a]. Divided-Difference [Gag94a]. DIVVY [Vol90c]. DLL [Car99]. DLLs [Bec97g, Pug91b, Smi93g]. DMA [Bra92a, Owe95, Wat93]. Do [Pla88e]. Do-Everything [Pla88e]. Dobb [Hol87, Nel89a]. Doc [Man88]. Doctor [RSSS90, Whi88, Jae86]. Documentation [Phi90b, Pug93d, Rob94, Sil88, War90c]. Dog [Boc96, Wil98]. Doing [Owe95, Zei91]. Domain [Bra91, Vol89d]. Don [vdP93]. Don’t [Bra92c, Lib91a, Tom91a]. Dorfman [Pla94b, Rod96, Rod96]. DOSThread [VE93]. Dot [Pat88b, Phi96d]. Double [Gou98, Pri91a, Pri91b, Pug89h, Pug92k, Pug93e]. doubles [PC91, Pug94d]. Doubly [Hat94, Jae91e, Jae91f]. Doubly-Linked [Jae91e, Jae91f]. Doug [Web96]. Down [Mon97, Ros97]. dprintf [Asa90]. Dr [Nel89a]. Dr. [Hol87]. Draft [Pla94n, Plu88b, Plu88d, Pug88y]. Dragto [Zmo96]. Drawing [Bel89a, Joh99a, Pug89i, SR99a]. Draws [Spe89b]. Drive [Whi88]. Driver [Ben92, CD90, Law95, Law92c, Nel92b, Paj88, Rot91, ET88, ET92, Wei93b]. Drivers [All89, Erd88, Joh90b, Kri97, Nor92]. Drives [Tom91b]. Driving [Sch97b]. Drops [Plu88b]. DSP [SD89]. DSR [Gra93c]. Dual [Ada94, Bec95a]. Dual-Use [Bec95a]. Duntemann [Swa95c]. Duplicate [Tom91b]. DXF [Gag93]. Dynamic [All94g, All94h, All95c, Bec98g, Col95a, GS95, GH99, Hog90, Joh99b, Ker96, LaB93a, Mar98, PM89, Pug93f, Sak92a, Swa97, Van95, Dac95]. Dynamically [Roe99].

E. [Pla94d]. Early [Bri98g]. Eases [GK90]. Easier [Mit99]. Easy [Deg97, Gra92a, Gut95, Har99, Sch85, Haw88]. Ecosoft [Pug88i]. ecvt [Pug89t]. Edge [Phi91f, Phi92e]. Edges [Phi93d]. EDIPAL [YEA91]. Edit [The94, War90d]. Edited [Rom91b]. edition [Bla90, Phi93b, Wei93b, Ano97s, Vol99]. Editor [BD90a, BD90b, BD90c, Dav93, Isa98, Osb90, Pla90h, Sch89a, Vol94d, WO89, Wil93b, Wol94, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Bri98c, Bri98d, Bri99b, Bri99c, Bri99d, Hes94, Pla96b, Pla96a, Pla97i, Pla97a, Pla97b, Pla97c, Pla97d, Pla97e, Pla97f, Pla97g, Pla97h]. Editorial [Ban89, Clu89b, WB88]. Editors [Ano98b, Ano99a, Ano99b, Gib88, Vol93c]. Effective [All99b, McL94, Mey92]. Effectively [Sut99b]. Effects [Bec95g, Bec96g, Blo93]. Efficient [Erd89, Hel95, Joh99b, Mor98a, Mor98b, Pri89a, Pug90m, Rod95b, Sch93b, Tre94]. Effort [Sch89b]. Efforts [Pla92j]. EGA [Ced88]. Egan [Wei93b]. EGAPAL [YEA91]. EGAPAL/EDIPAL [YEA91]. Elegant [Oua92, Phi94a]. Elements [All93f, Pat95, Phi94a, Pug90a, Pug90m, Gam95, Oua92, RN93]. Ellis [Bri96]. Elm [Fie89d]. Elucidated [Pug90a]. Elvis [Kir92]. Emacs [Con88, Joh88a]. Email [Faz98, Jae88f]. Embedded [All94b, Bec93, Cox92, Gin93b, Hal91, Han94, Lab98, Lan96, Lan99b, Nar89, OH94, Pla97j, Vil91, Wel93, Wis99, Gan92, Pla99e]. Embedding [Joh92, Pas89, Pip95]. EMS [Swa95d]. Emulating [Col94, Pug90n]. Encapsulating [Bar98, Bav98a, Bey98, Dow96, Nel94, Par95]. Encapsulation [Haz91, PW91, War88a]. Encoding [Sea91]. Encryption [Bal97, Par98, War98]. END [PB91, Dan98].

Engine [Vas96, WO89]. Engineering [Cad91, Swa95a, Wal88, Wil94, BN94, DM88, Hum95, TK89, Phi96c]. Engineers [Ric91]. English [Wel91]. Enhance [Ger90, Hin88a, Poo93]. Enhanced [Mea91]. Enhancement [DE98, Ges91]. Enhancements [Sak95f, Sak95g, Sak95h]. Enhancing [Lan91c, TR94]. enum [Wal99, Bec97c, Bri98e]. Environment [Joh90c, Lad89a, LFK89, Lam89, Lib89c, Pla92b, Pug88x, Sin93b, KK89, Nel90b, Pug88p]. Environmental [Jae90b]. Environments [Rib89]. Epson [Pat88a]. Epson-Compatible [Pat88a]. Equal [Bec97a]. Equalization [Phi91e]. Equation [Wil96a]. Equations [Wei93j]. Equipment [Gar88]. Equivalents [Mic97]. Eric [Jon88a]. Erich [Pat95]. Eroica [Sch99b]. errno.h [Pla90f]. Error [Ale99b, Bec99i, Bec99k, Cha95, Jae90c, Nad98, Pom98, Pug88k, Pug88s, Pug92c, Pug92f, War88d, Wil96b]. Errors [Ano94a, Hal92, Pom98, Sak96i]. ese [Kis97]. Especially [Plu88c]. Essential [Pro90, Smi88b, Smi93d, Cor88]. Estimating [Fra96, Gil95a]. Evaluating [HPR99, She90]. Evaluation [Pug92i]. Evaluator [Rog96]. Even [Bec96c, Sak95f, Sak99g]. Event [HJ90, Sco93]. Events [Hoo99, Kel90]. ever [Por93, Pla95a]. Every [Bec98f, Wal89]. Everybody [Wei94f]. Everything [Bec96e, Pla88e, Wei92c]. Everywhere [Vol96a]. Evidence [Phi90a]. Evolution [Bec98i, Gra92b, Kel92, Pla89c]. examine [Wil96b]. Example [Gil95b, Gil90, Hav90, Dor95, Per92b, Per92c, Rod96, Tho92]. Examples [Bur90b, Pug88-27, BN94, Mul89]. Exceeding [PF93a, Pug92e]. Exception [Col91, Col94, Guf98, Pla98a, Vid92, WR93, Pla94g]. Exceptional [All99c]. Exceptions [All94d, Ast94b, Bec98d, Pes98, Sch98l, Sch99a, Sch99b, Sch99c, Sch99d, Sch99e, Sch99g, Sch99h, Sch99f, Sch99i, Sch99j, Ves95]. Excess [Sak99h]. Excited [WW88]. Exec [Man88]. Executable [Kuz90]. Executables [Bri97]. Executing [BS99]. Execution [Bur89, Rog89, Wit88]. Executive [All92a, Man89, dC90a, dC90b, GC93]. Exercises [TG89]. Existing [Ric99]. exit [Jae90d]. Exits [PG91]. Expanding [Pla89b, Sue94]. Expansions [Sue93a]. Expect [Lib91b]. Expectations [Sch99c]. Expeditionary [Gar88]. Expensive [Wei94c]. Experienced [Lop93, Jae93]. Experiences [Mar90c]. Expert [Kea90]. Explained [Pug90y]. Explicit [GS95]. Exploration [Mue94b, Vol95g]. Exploratory [Wil94]. Explorer [Sch90a]. Exploring [Jae88b]. Express [Pat92a]. Expression [EU98, Mar98, Pla88c, Rog96]. Expressions [Pla88e, Pug88y, Sak97g, Sak97h]. Extended [Mil94, PS92b, Spe89a]. Extenders [Nel92a]. Extending [Ano98m, Ano99m, Col93a, Har94, Smi92d]. Extensibility [CE96b]. Extensible [Mir93]. Extension [Ste97]. Extensions [Kre96, Pla93h, Pla93n, Sak93f, Sch90e, TC95, Tri93]. Extensive [Fie88f]. extern [Pug88-27, vdP91]. External [Pug90z, Pug93g, Whi90b]. Extracting [LaB94]. Extractor [Mer96]. Extractors [Pla94f]. Extrapolation [Smi93a].

F [Fla90]. Facet [Pla98h, Pla98f, Pla98g, Pla98e, Pla98j, Pla98i, Pla98d]. Facets [Pla98k, Pla98c]. Facilities [BD89a]. Facility [Flo90, Jae89c, Vol88a]. Factorial [Heg92]. Factorial-Base [Heg92]. FAFNIR [Osh89]. Fail [Wis97]. Fail-Safe [Wis97]. Failures [Pug94f]. Faithfully [Sma88]. Family [Car91]. Far [Jae88c, Lib90a]. Fare [Wei95a]. Farewell [Fie88b]. Fast [Ash89, Clu90, Ott94, Ros95, Swa88, Vol88b, Vol89d, Wel90]. Faster [Ber92b, Cli92, Lib89a, Pug95a, Hel95]. Fat [Sak99h]. fcvt [Pug89t]. Features [Wit90d]. Feed [San93]. Feel [Ash90]. Fees [Pla93e]. Fence [KB99d]. Feuer [Rab89]. FFT [LaC94, Pri93]. Field [Pok92]. Fields [Jae89e]. Fight [Pla91c]. File [All93g, All93h, Bur92a, DP93, Fre88a, Fro92, Har94, Hic89, Joh90d, MO92, Phi91b, PF94, Pug88m, Pug89m, PY90, PC92, Pug92o, Rob94, Tha88, Vol95d, WH88, Zol90b, Zol90a, Zol90c, Zol90d, Wei88a, Jae89d]. File-Handling [PF94]. Files [Cog89, Col89c, Fro89a, Fro90, Ger90, Isa98, Joh88b, Lop97, Phi91f, Pla99k, Pla99h, Pug88h, Pug90b, Pug90i, Pug90z, Pug90s, Pug90-27, Pug90-33, Pug93u, Sav96, Sch96i, Sch96j, Spe89b, Tom91b, Tuc98]. Filter [Pen93, Van95]. Filtering [Phi92f, RP92, Sto95]. Filters [Mor99]. Finale [Sak96e]. Finally [Wei92b]. Financial [Smi93e]. Financing [St.93]. Find [Fie88a]. Finding [Hof92, War88d]. Finds [Wio88c]. Finer [All92b]. Finite [KJV98, War89a, War89d]. Finite-State [KJV98]. FIR [Mor99]. Firewalls [Sak94a, Sak94b]. Fireworks [LL90]. First [Plu89b, Sak91e, Sak91f]. First-Order [Plu89b]. Fitting [Ano90a, Bra90b, Gra92c, Pri92a, Smi93a]. Five [Ano97-33, Gib88, Sch99j]. Five-Liners [Sch99j]. Fix [Bec96e, Wei88c]. Fixed [Car93, Fos93, HD95, Lab98, Bak94]. Fixed-Point [Fos93, HD95, Lab98]. Fixed-Size [Car93]. Fixes [Eag88]. Flag [Cog88a]. Flags [Cog88b]. Flak [Bec98d]. Flames [Ban89, Clu89b]. Flamig [Hob91a]. Flash [Cep95]. FLEX [HP89, Coâ94b, Vol89d]. Flex-ible [Vol89d]. Flexible [Asa90, Eme88, Hog90, Vol88a]. Flip [Wei93c, Wei93d]. Flip/Flop [Wei93c]. Float [Pri89b, Pri94]. float-Precision [Pri94]. float.h [Pla91g, Pri90b]. Floating [Man96a, Pla93h, Pug90i, Pug93h, Sch90c, She90, Smi93f, Smi93g, Smi94a, TC95].

Floating-Point [Pla93h, Smi93f, Smi94a, TC95, Man96a]. Floats [Ban88, PC91]. Flood [Tre94]. Flop [Wei93c, Wei93d]. Flop/Flip [Wei93d]. Floppy [PK92, Whi88]. Flow [All99c, Pla89d]. Flowcharts [Spe89b]. Flux [Jer94]. Fly [Eng93a]. Fond [Fie88b]. Font [BG99]. FontLib [Gag93]. Fonts [Sch92a, Wil91]. Forces [Wei92a]. Forked [All90]. Form [Pla88e]. Formal [BD90d, Pla93i]. Format [Pug90n, PP92, Rob94, Wei90h, Pug88i, Ret96, Phi91b]. Formats [Iof99, Sch90c, War88c]. Formatted [Pla89e, Pla90g]. Formatter [Ber90]. Formatting [Pla91f, Pla93p, Pug88j]. Forms [WO89]. Formula [RP92]. FORTRAN [Pug94h, Sch89b, Ano99c, Irv91, Pug90n, Tee90]. FORTRAN-C [Sch89b]. Forum [Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Bri98c, Bri98d, Bri99b, Bri99c, Bri99d, Pla96b, Pla96a, Pla97i, Pla97a, Pla97b, Pla97c, Pla97d, Pla97e, Pla97f, Pla97g, Pla97h]. FOSSIL [Law95]. Foundation [Phi92d, Sch95a, WL94]. Four [Fro92, Pla88e, Hav89b]. Fourier [Clu90, Swa88]. Foxley [Jon88a]. Fractions [Zei91]. Framework [KF97, Lam97, Wam96, Blu92]. Frankenstein [Zen91]. Free [Ano96b, Fie88c, Pug88t, Zer89, Hel94, Mag93, Pug93w]. Freedom [BB94]. Freeware [LS98]. Frequency [Phi92f]. Frequently [Pla99f, Pla99g]. friend [Bec98f]. friendly [WPM84]. Friends [Jae90l]. Frog [Hav93]. Front [Dan98]. fscanf [Clu89a]. fstream [Pla95d, Pla95j]. FTGRAPH [Clu90]. Fun [War88j]. Function [Asa90, Bat97b, Bec98g, Erd91, Hel90, Jae90k, Jag95, Mur93, Pan93, Pat92b, Pug88k, Pug90-38, Pug90-41, Pug91f, Pug91g, Pug92b, Pug93i, Pug93o, Pug94e, Pug94i, Sak91a, Sak93a, Sch91c, Smi92c, Smi93d, War88e, Whi89b, Wri90a, vH92, Bak92, Pug89o]. Functions [BD89b, Bec98j, Bed91, Cla88, Dug97, Gra93c, Gra93f, Jae88f, Jae90d, Mey95, Pla88a, Pla90a, Pla91n, Pla92i, Pla92n, Pla93l, PA93a, Pla93o, Pla93p, PF94, Pri89b, Pri92b, Pri92c, Pri96a, Pri96b, Pug89d, Pug89h, Pug90r, PH91b, Pug92b, PJ92, Pug92h, Pug92k, Pug93l, Pug93s, Pug93w, Pug95c, Sak90a, Sak93j, Sak94i, Sak94k, Ske94, Smi93e, Vol90b, War91b, Wil95b, Wil96b]. Fundamentals [Gra92c, Pug90f]. Further [Sch97c]. Fuzzy [Jan93a, McC94].

G [Gin93b, Pla90e, Tay93]. Gadgets [Ekw92, Fio89]. Gail [Hoh90]. Games [Sch90a, Wei90c]. Gamma [Pat95]. Gang [Pla90p, Pla90q]. Ganssle [Gin93b]. Garbage [Col95b]. Gateway [SDWB96, Ste97]. GATool [Lie92]. gAWK [Wit90a]. gcvt [Pug89t]. GED [Tea88a, Tea88b]. Gehani [Pla90c]. GEM [Pot88]. General [Man98, Wal91, WO89]. General-Purpose [Man98]. Generate [Del88d]. Generates [Ano97-43]. Generating [Bra88, Pri90b, Pug92f, Sch92a]. Generation [Mar96, NW99, Pug88v, PBS93]. Generator [Bar95a, BT91, Com90, FB96, Fru90, Pat99, Vol88b, Wal99, War89d]. Generators [Dwy95, DW96a, DW96b, Pug92s]. Generic [All99a, Ano96b, GW91, Lor99, McM91, PA93a, Sak94c, Sak94d, Sak94e, Sak94f, Sak94g, Sak94h]. Genetic [Gra95, TC94]. Genitor [Whi92]. Geometric [Mor98a, Mor98b, Per99, Phi95a]. Get [Fie89c, Ste93b, Wei90d, Wei92g, Wei95b]. Getting [Bec95b, Coo92a, Sch97i]. Ghostscript [Deu93, Vol93e]. GIFs [Knu98]. GIFSave [HMF94]. Gilbert [War91a]. Give [Plu88c]. Glass [Wee92].

Glass-Box [Wee92]. Global [Joh92, Sta89, Wit90b]. Globally [Pug92g]. GNU [Fie89c, Arc94, FD92, Vol89a, Vol96f]. GNUlib [Ohl92]. GNUPLOT [Kot91, Ste93b]. Go [Wei94f, Gre89]. Goal [BC89a]. Gone [Wei92f]. Good [All99c, Bar90a, Bur92b, Fie88a, Sak97c, War89k]. Goodwin [Cla90c]. goto [Bec95b, EW95]. Gotos [Pug92t]. GPIB [Tay95]. Grabbag [Ano91b]. GRAD [BC90a, Kwo89a, Kwo89b, Pug88l]. Grady [Lan91a]. Grammar [Del88b, Sak95a]. Grammars [Sak96h]. Graphic [Wil91]. Graphical [Dan98, Joh99a]. Graphics [Ash91a, Ash91b, BC90a, Ced88, Clu90, Deg97, Dis88, Ger88, HD95, O’D91, Pat88b, PG92, Sch89a, Sch88b, Sch88a, SB89, Smi88b, Str88a, Vas96, Vol90d, Vol90e, Vol94b, Vol95e, Wal89, Wei90h, Zia90, Amm86, Amm87a, Amm87b, Amm92, Pic90, Kam89]. GraphX [Vol90d]. Gray [Phi93d]. Great [Pug90o, Sch99c, Wei90g]. Greatest [Fie88g]. Greg [Tho92]. Gregory [Web96]. grep [Bar90d]. Ground [Pla90k, Pla93c]. Group [Mar90a, Mar90b, Pla93n, Vol98b, Vol99, Ano88a, Vol96b]. Groups [Pla88b]. GUI [Gui97, Vol90f, Wam96, Zek95]. Guide [Bri98b, Per89, Spa90, Wio89, Cam87, Cam93, Chi89, Hel94, MQ87, Sch88c, Sch92c, WPM84, Ham98]. Guideline [Siy89]. Guidelines [Fre93, Pug92p, Pug93l, WC90, Plu84, Plu89a, Phi93c, Phi93b]. GUIs [Smi92f]. Gun [Pla90r]. GZIP [Gai93, Vol93d].

Habit [Bec96f]. Hack [Bec97c]. Halftoning [Phi91d]. HALO [Vol90e]. Halt [Pug90p]. Handbook [Smi92c, Bak92, Cor88, McC93, Pro90]. Handler [All88]. Handlers [Pom98, Ste90]. Handles [Cro89, Rec98, War88f]. HandleTracker [Mon97]. Handling [All93a, All93b, All94c, Ast94b, Bec97e, Bec99i, Bec99k, Col91, Col94, Fre88b, Jae90c, Kis96, Nel92a, PF94, PH90, Pug92c, Pug94f, Rem97, Vid92, Wei94h, WR93, Yue94]. Happened [Wei93h]. Happens [Bec97d]. Hard [All94a]. Hardware [Cha88, Cha90, Cog88b, Fru92]. Harel [Rom91a]. Hash [Ben98, Pla98b, Pro89]. Hashing [Bur92b]. Haviland [Gra89]. Having [War88j]. Hayes [Dan93]. Hazards [Col93b]. Header [Cog89, Ger90, Jae90e, Pla91g, Pla91h, Pla91i, Pla91j, Pla91k, Pla91l, Pla92e, Pla92f, Pla92g, Pla93j, Pla94g, Pla94i, Pla94j, Pla94h, Pla94k, Pla94l, Pla95b, Pla95c, Pla95d, Pla95e, Pla95h, Pla95f, Pla95g, Pla96c, Pla96d, Pla96e, Pla96i, Pla97s, Pla97q, Pla97o, Pla97r, Pla97p, Pug90z]. Headers [Pla96m, Sak90a, Sak90b]. Heap [Far89, Pug93j, War91b, Whi89a]. Heaps [Man99]. Held [War91a]. Helen [Hal93d]. Heller [Rod95b]. Helm [Pat95]. Help [BT93, Bug98, Vol96e]. Hendrix [Vol89b]. Herbert [Haw88, Pla94e]. Hercules [PG92, Sch88b]. Here [Wei92b]. Heterogeneous [Bec99a]. Hex [Col89c]. Hey [Gra88]. HGA [Sch90a]. Hiding [And88c, BD90e, BD90f, Bec97f, Flo90, Gle90b, Mur93, Phi98b, Phi98c, PW91]. Hierarchy [CM98, Mar96]. High [Hea96, Sak99d, VC89, Vig89, Yor96]. High-Level [Sak99d]. High-Speed [Yor96]. Highly [Smi88a]. Histogram [Phi91e, Phi93e]. Histogram-Based [Phi93e]. Histograms [Phi91e]. Hits [Fie88g]. HLLAPI [Con91]. HOC6 [Pug88p]. Hodgman [Rév93]. Hole [Jae89i]. Holes [Jae89i]. Holub [Eib94, Nel89a]. Hone [Bec98a]. Hooked [Nic96, vSS96]. Hosted [Her94]. hot [vSS96]. Howard [Pla94d]. HP [Flo89, Smi93b]. HTML [Bug98, Col96, Ret96]. Huffman [Eng93a, Phi92b]. Human [Rom91b, LM90]. Human-Computer [Rom91b, LM90]. Humphrey [Phi96c]. hundreds [Jae86]. Hunt [Bin88, Bla90, Cas93]. HyperCard [Rat91]. hypot [Eve93].

I/O [PS92a, Smi89, Kri97, Man98, New91, Pla89c, Pla89h, Pug88m, Ste87, Wil92b, WP88, Zol90b, Zol90a, Zol90c, Zol90d]. I8255 [Mil91]. ible [Vol89d]. IBM [Bin88, Cas93, Sch88a, Amm87a, Her94, Pug91g, PS92b]. IBM-PC [Her94]. IDE [Ano97-31, Fis90a]. Idea [Sak97c]. Identification [Han90]. Identifiers [PF93b, Pug93t]. Identifies [Gim89]. Identity [Jae89a]. IEEE [And89, Cho88]. IEEE-488 [And89]. II [And88a, All92c, MC88, Pla90o, Pla90e, Pla90q, Pug89c, Str88a]. IIgs [Del88f, Pec89]. III [All92d, Fla90, Jae92g, Vol98a]. Illustrate [Pug88-27]. Illustrated [Wal92]. Illustrates [Whi88]. Image [Col95a, Dea93, Fal90a, Fal90b, Hag91, Mar90a, Phi91a, Phi91b, Phi91c, Phi91d, Phi91e, Phi91f, Phi92e, Phi92f, Phi92g, Phi93a, Phi93e, Phi93d, Phi94b, Phi95a, Phi95b, Phi95c, Phi96d, Wei90h]. Imagemagick [Wei93f]. Images [Mar90b, Phi91c, Phi91d, Phi91f, Phi99d, Smi93b, VG89a, Wei93f]. Imaging [VG89a]. Imminent [Pla91d, Pla94n]. Imperial [Ano97s]. Implementation [BD90e, BD90f, Cro99, Mar97, Pug90q, Rua97]. Implementations [Bry98]. Implemented [Whi89b]. Implementing [BD90c, Bec98g, Bra92a, BT93, Cam98, Fan99, Fel90, Lib90b, Lib91c, Pla90i, Pla91m, Pla92h, Pla94m, Pla95i, Pla95j, Pla95l, Pla95k, Pla97m, Vas96, Lad92]. import [All99c, All99d, All99e, All99f, All99g, All99h]. Important [Bru89a]. Improve [Van93, Mey92]. Improved [Wio88c]. Improvement [And94]. Improves [Flo90]. Improving [Lan89, Man99]. Include [Cam88]. including [TK89]. Inclusion [Sch99j]. Incomplete [Sai92]. Inconvenience [Sak93h, Sak93i]. Increases [Pla93e]. Increment [Pla93b, Pug92i, Pug93p]. Indent [Arc94]. Independent [Gui97, WH88]. Index [Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Fre88a]. Indexes [Kar90]. Indirection [Pug88n]. Inexact [Zwa91]. Informal [Fie88f, Pla94n]. Information [Ban94, Cas89, Dan93, Phi98b, Phi98c, Pug88o]. Informix [War91c]. Inheritability [CE96b, CE96c]. Inheritance [Haz91, Sak93b, Sak93c, Van95]. Inherited [Bec95e]. INIT.C [Pug88p]. Initialization [Bec99c, Bec99b, Bec99d, Jae88d, Pug92l, Sak92b, Sch99k]. Initializing [PA93a, Pug89j, Pug90z, PF92a]. Inline [Ale98b]. Input [Ale91, Byb92, Erd91, Fro89a, Pla90g, PM91a, The94, Wei94h]. Input-Edit [The94]. Input/Output [Fro89a]. Inputs [Wil94]. Ins [Jae88d, Roe99]. Inserter [Sch96d]. Inserters [Pla94o]. Installable [Vol95d]. Installation [WH88, Hel94]. Instead [HW92b]. Instructions [RP91]. Instrumentation [Pan99, War91b]. int [PA91]. Integer [Got94, MP90]. Integers [Bre96, Car93, Pug91c, TP92]. Integrated [NW99]. Integration [Sti92a]. Intelligence [Bou94]. Inter [Boy88, Cro98a, Gle90c]. Inter-Object [Cro98a]. Inter-Process [Gle90c]. Inter-Task [Boy88]. Interactive [Coh92, SR99a]. Intercepting [PR91]. Interception [Pug88s]. Interface [Ale99a, Bru92b, Eme88, GK90, MA88, Man89, Mar88b, Mil91, Nel94, Pet98, Pug89u, Rom91b, Wei93i, Wil91, LM90]. Interfaces [Bry98, Cla90c, Col89a, Hin88b, Lan99a, Pin91, Smi92f, Goo89]. Interfacing [Law95]. International [Lib89b, Nol94, Pla89g, Plu88a, Tac94]. Internationalization [Jae95a, Jae95b, LK97, Pla94d, Pug93l, Pug93k, UHP93]. Internet [Kuo97, Vol97d, Vol97c]. Internet-Aware [Kuo97]. Interpolation [For91, Gag94a]. Interpretation [Pla90l]. Interpreter [Spr90]. Interpreting [Gra92d, Pla90j]. Interprocess [Fre90, Gil90, Wit90d]. Interrupt [All90, Bur89, Nel92a, Tay97]. Interrupts [Cor91, Pug90-30, PN91, BK91, BK94a, BK94b, Bro94, Swa95b]. Introduces [Ano97u]. Introduction [Bri98f, Gle90a, Gra95, Lop93, Pla94p, Pla96j, RL98, Sak98a, TC95, BN94, Jae93, UHP93, Pla94d]. Intuition [Fio89]. Intuitive [Hei94]. Inverse [Kie96]. Inverted [Gou98]. IOCTL [Nel94]. iomanip [Pla94h]. ios [Pla94i]. Iostream [Tay95, Pla95e]. Iostreams [Kis97, Pla94p, Pla99j, Kam93]. Iostreams-ese [Kis97]. IPX [Sch92c]. iRMX [Car88, Car91]. Irregular [HP98]. ISAM [Lop97, Mur93]. ISAPI [Ste97]. Isn’t [Cin95, Wil96b]. ISO [Pla91a]. Isolating [Sak99a, Sak99b]. Issue [Wei91e, Wei93g, Wei94d]. Issues [Pug90-39, Siy89]. istream [Pla94j]. Iteration [Bec99e]. Iterator [Cur98, Lor99, Pla96c, Pla96d]. Iterators [Ale98a, Bec98m, Get99, Kos99, Pla96f, Rum98, Sak94e].

J [Bin88, Bla90, Cas93, Gin93a, Pla94b, Rid91, Sic94, War90d, Wei93b]. Jack [Gin93b, Ogg90b]. Jaeschke [Far88, Lop93]. JAM [Col89a]. James [Fla90, Pri88, Vol89b]. Janet [Wei93b]. Japan [Qua88]. Japanese [Ano88a, Sea91, Wat88]. Java [All99d, Ano97s, Ano97-31, Ano97-50, Bri99a, All99e, Ano96a, Ano96b, Ano96c, Ano97t, Ano97-33, Ano97-43, Bri98e, Cla97, Cro99, Dan98, Hem96, Kal98, McK98, Mic97, Nic96, Pat99, Pla96b, Pla96a, Pla97i, Tuc98, Wel99, vSS96]. java.* [All99c, All99d, All99e, All99f, All99g, All99h]. Jay [All93f]. Jean [Gin93a]. Jeff [Swa95c, Kil96]. JJB [Bro89]. JMODEM [Joh93, Vol93d]. Joe [Wio89]. John [Pat95, Swa95a]. Johnson [Pat95]. Journal [Pla92d, Zol91b, Hol87, Ano94b, Ban89, Clu89b, Cog89, Mul93, Nel89a, Pla91d, Pla91e, Pla93k, Pla94c, Pla94t, WB88, War90d]. Journeyman [Bec98a, Bec99a, Bec99c, Bec99b, Bec99d, Bec99e, Bec99i, Bec99f, Bec99g, Bec99h, Bec99j, Bec99k]. JoyStick [Jar91a]. JPEG [Lan93, Vol93b, Vol93d]. Judy [Nel90b]. Jump [Jae90l]. Jumping [All99d]. Just [KB99b, Sch99d]. JYACC [Col89a].

Kanji [Fie89a]. Kansas [Ben97]. Kay [Nel90b]. KBFake [Vol95b]. Keeping [Pla93k, Pug92p]. Keith [Gra89, Hob91a]. Kempf [Pri88]. Kenneth [Nel89b]. Kermit [bG91]. Kernel [Gin93a, Hal91, Lab92]. Key [Pug89f, PN91]. Keyboard [Coo92a, Gib90, Gru90]. Keypress [Pug90p]. Keys [Pug90-41, Pug91g]. Keyword [Pug88-27]. Keywords [Bec95a, Car93]. Kicking [Bec96f]. Kind [Mat89]. Kit [Ben94, WH88, Won93]. Known [Bec96g]. Kochan [Kou88]. Korn [TR94]. Kruglinksi [Sic94]. Kummerfeld [Nel90b]. Kyocera [War88g].

L [Bur92a]. Label [Ano93]. Labrosse [Gin93a]. Lack [Cog89, Pat92b, Pug93m]. Lacks [Qua88]. Ladd [Hal93b]. Lambert [HP98]. Langsam [Rid91]. Language [Pat92a, BD89a, Cla88, GK90, Irv91, KR88, Kru92, Lib90e, Pat88a, Pat88b, Pla90c, Pla92d, Pla94a, Pri90c, Pug89r, Sak93f, Sak99d, Str91b, Str99, Sue93a, Sue93b, War92a, Wel91, Wio88d, GR89, Hen88, KR78, MQ87, Str86, Str87, Str95, TK89, WPM84, TG89, Cof88, Web96]. Languages [Sch90b]. Large [Bai91, Bre96, HP98, Pla93m, Pla93l, TP92, Wei94g]. Laser [PdW92, Smi93b, War88g]. Last [Pla96k, Pla98c, San93]. Late [Bri98g, Haz91]. Late-Binding [Haz91]. Laurel [Rom91b]. LCDs [Lan99b]. Leak [Doe97]. Leaks [Mon97]. Learning [Ros92b, Sch96f, Sch96g, Sch96e, Sch96h, Sch96i, Sch97b, Sch97c, Sch97a, Sch97d, Sch97e, Sch97g, Sch97i, Sch97f, Sch97h, Sch98a, Sch98e, Sch98g, Sch98j, Sch98i, Sch98d, Sch98b, Sch98k, Sch98c, Sch98f, Sch98h, Str99, Ano97t]. Lee [Swa95a]. Leendert [Kam89, Sch88a]. Legacy [Gib99, Kle97]. Legal [Yok89]. Len [Rod96, Pla94b, Rod96]. Length [All94k, Jae89g, Kod92, PJ92]. Leor [Wal92]. Less [KB99d]. LessTif [LS98]. Let [Fra88, Sch97e, Wei88a]. Letters [PLBL92]. Level [Ric99, Sak99d]. Level-2 [Ric99]. Levels [Pug88n]. lex [Ano88b, Eag88, Pug91k, Vol89c, Vol89d, Lib90f]. Lexer [McK98]. Lexical [Ste92a]. Liana [Kru92]. Lib [Vol95b]. Libes [vdP93]. LIBFTP [Vol94b]. Librarian [Pug91j]. Librarians [Wei89]. Libraries [Bru92a, Fro92, Mur93, Pat96, PM89, Pug94p, Ros92b, SX98, Swa95d, Van93, Vol96d, Wei88d]. Library [All94j, All95d, All95e, All95f, Ano91c, Ano96b, Ano97u, Ash91a, Ban94, Bru92b, BC90a, BC90b, Cam88, Ced88, Col89c, Cur90, EU98, Eve93, Flo89, Fre88b, Fro89a, Fro89b, Ger88, Gib88, Guf98, Hin90b, Jae90f, Jar91a, JR91b, Kam93, LS98, Lan99b, Let91, MA88, Man91, Mar88a, Mic96, Mic97, Mil94, Mil88, Mil89, Mil91, Mor93b, Mus90, New91, Pea90, Pla90k, Pla93d, Pla93c, Pla93f, Pla94a, Pla95m, Pla97n, Pol91, Pri94, Pug88l, Pug88-31, Pug90r, Pug91f, Sch90f, Sch92b, Sco88, She90, Smi88b, Smi93d, Smi93e, Vol90c, Vol90d, Vol90e, Vol90i, Vol93c, Vol94a, Vol98b, War88b, War88c, War89c, WO89, War89g, WL94, Wri90a, Yok89, Zia90, Ber88, Wil96b]. License [Rol92]. Licensing [Par98]. Life [Lib90c, Lib90d]. Lightweight [Col90d, Lan99b]. Like [Sch99g, Ske91]. Likes [Smi88b]. Limit [Pug92e]. Limitations [Pug90d]. limits [Ano99m, Pla97p]. limits.h [Pla91h, Sak90b]. Line [Ale91, Bel89a, Byb92, Col90b, Erd91, Kou97, Mal90, McM91, Mea91, Pug92m, San93, Smi92f, Vol90k]. Line-Drawing [Bel89a]. Linear [For91, MC88, Wei93j]. Liners [CH88, Sch99j]. LINES [Sil88, Pug89i, Wei96]. LINES/C [Sil88]. Lingua [SR99b]. Lingual [PG93]. Link [PM89, Wil96b]. Linkage [Fea95b]. Linked [Erd89, Hat94, Jae91c, Jae91d, Jae91e, Jae91f, Jar90, Pug92h, Pug92q, PC93, Roe99]. Linking [Cav91, GS95, Pug88l, Pug90r]. Lint [Pug93n, Pug94g, Gim89, Wio88b]. Linux [Rob95]. List [Pla99b, Pug88a, Pug92h, Wel94b, Wel95, Wel99, Pla97q, HKB90]. Listbox [Tha98]. Listings [Cog89, Zol91c, Zol91d]. Lists [All94k, Ano98a, Erd89, Gro91, Hat94, Heg91, Jae89g, Jae91c, Jae91d, Jae91e, Jae91f, Jae91g, Jae91h, Jar90, Kod92, Lop97, Mar97, PJ92, Pug92q, PC93, Whi98]. Literal [Jae88e]. Literate [Wit95]. Little [Bec96g, dF90b]. Little-Known [Bec96g]. Living [Bri98e, Sch95b]. LL [HMF94]. Load [And99a]. locale.h [Pla91i, Pla91m]. Locales [Pla91b]. Localization [Pla94d, Pug93k, UHP93]. Locating [New89]. Locator [All94b]. Locking [Pug90s, Rog95]. Locks [Fru92]. Logging [Hoo99]. Logic [Ano96a, Jan93a, Lad89b, McC94]. Login [Zol91a]. Long [Pri91b, Wei95a]. Look [Ash90, Jae89e]. Looking [Sak93d]. Lookup [DE98, Vil96]. Loop [RP91]. Looping [Bec99e]. Lost [Tal93]. Louis [Smi92c]. Low [Rob95]. Lowly [EW95]. LUP [Rum99]. Lvalues [Jae89j, Pug90g].

M [Rid91, Fuk89, Hic89, McD88, Ock89, Pug90-28]. M.I.R.A.C.L. [Sco88]. m68kdis [Vol95c]. Mac [Rat90, Rat91]. Machine [Kno90, Nel92b, Pla92m, War89a, War89d]. Machine-Readable [Pla92m]. Machinery [Del88a]. Machines [Fis90b, KJV98, She92].

Macintosh [War86, Bru89b, Hea97, Pug90-27, Rea91b, War92c, And88b]. Macintosh-to-Windows [Hea97]. Macro [Bec96f, Cru88b, Jae88g, Jag95, Pug90y, Ray92, Smi92b, Vol88a]. Macros [Col93b, Ges91, Pla89b, PD93, WO89]. Madaline [Phi92d]. Made [Har99, Haw88, Sch85]. Magic [Wei93f, Wil95a]. Magna [Vol91a]. Maguire [Hal93a]. Mail [Ano97-51, Ano97-52, Ano97-53, Ano97-54, Ano97-55, Ano97-56, Ano97-57, Ano97-58, Ano97-59, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-34, Ano98-35, Ano98-36, Ano98-37, Ano98-38, Ano98-39, Ano99-27, Ano99-28, Ano99-29, Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Ano99-37, Ano99-38, Kam97, Wei94a]. Mailer [Hat96]. Mailing [Ano93]. main [Pug91h, Pug92n]. Mainframe [Con91]. Maintenance [Phi90a, Pug92p]. Make [Cog88b, Yeh90, Wei88b, Wil96b, Cru89, Gil92, Wei88a, Wei88d]. Make-Maker [Yeh90]. MakeFont [Gag93]. Maker [Yeh90]. Makes [Pat88b, Pla88f, Vol90h]. Making [All96a, Del88d, JR91c, Phi98a, Pla93a]. malloc [KB90, Lib89a, Pug90t, Hal92, Pug93w, Whi89a]. malloc-Related [Hal92]. Man [Beh96, Kna89]. Management [Ada94, All94g, All94h, Bec99f, Bec99g, Bec99h, Fre88a, Fro92, Jae90e, Jae90f, Jen96, Sin94, Wel96, Dac95, DN93, All95c, Pla94b]. Manager [Cha95, Har94, Pro89, Zol91b]. Mandelbrot [Gar88, Sch90a]. Manifest [Pug90k]. Manipulating [Col89c, Pet99, Phi93a]. Manipulation [Fal90b, Hag91]. Manipulator [Bec96b]. Manipulators [Wil92b]. Manual [Phi94a, Oua92]. Manufacture [Del88a]. Map [Low95]. MAPI [Tha98]. Mapping [BD89a, Gra93f, Gre94, Kie96, Van92]. Marc [Pla94b]. Marching [Pug95a].

Margaret [Bri96]. Margolis [Wal88].

Marilyn [Bra92c]. Mark

[Fra88, Bur90b, Cla90c]. Market [Ano97-34, Ano97-35, Ano97-36, Ano97-37, Ano97-38, Ano97-39, Ano97-40, Ano97-41, Ano97-42]. Marketplace [Pla89g]. Martin [Bri96, Pav89, Row95, War92b]. Masked [Bra89]. Master [Lad89a, LFK89]. Mastering [Mey95]. masters [Alg95, Kil96]. MAT LIB [Hug90]. Matching [Cor95, Ott94, Phi94c, Pro92, Sch91d, Vil96]. Math [Ano97u, Eve93, Par95, Pla91n, Pla91o, Pla91p, Pri92b, Pri92c, Pri94, Pri96a, Pri96b, Wil96b]. Math.h [BC90b]. Mathematical [Smi92c, Bak92]. MathWorks [Ano97u]. MATLAB [Ano97u]. Matrices [Bak90, Bak91a, Pet99]. Matrix [KO94, Mas97, Pat88b, Pri89a, Rum99, Sch96l, Wil95a, Wil96c]. Matter [Pla90l]. Matthew [Pla93g]. Maximum [Pug91c]. Maybe [Sak97c]. MC68K [Col90a]. McConnell [Ush94]. MD5 [Par98]. MDI [Wel96]. Me [Sch97e, Sch97f]. Meal [Hav90]. Medical [VG89a]. MEL [CM90, Cre90]. Member [BD89b, BS99, Dug97, Ske94]. Members [Haw97, Sak95i]. Membership [Bec96a, Rup88]. memcpy [Pug89k]. memmove [Pug89k]. Memories [Sav96]. Memory [All94g, All94h, All95c, PS92a, Ast94b, Bat98, Bec97d, Bec98b, Bec99f, Bec99g, Bec99h, Bra92a, Bri90, Cep95, Doe97, Jae90f, Jen96, Nel96, Pla88f, Pla94b, Pug88q, Pug88t, Pug88u, Pug89l, Pug90d, Pug90h, PH91a, PF92b, SS89, Sch90d, Sin94, Smi89, Spe89a, Wel90, Wil98, Dac95, DN93, Ste87, Pla96i]. memory-resident [Ste87]. MenaC [Sch99h]. Menu [Pug88r, Ste92b, War88i, Zol92a, Zol92b, Zol92c, Zol92d, Zol92e]. Menuet [Vol90f]. Menus [Bug91, Mat91]. Merge [Nel97]. Mess [Bra92c]. Message [Cha95]. Messages [Ale99b, Sea98]. Messaging [Boy88, Cro98a, Kle93].

Metalanguage [Cre90]. Metaprogramming [Pes97a]. MetaWare [Vig89]. Method [Ber92b, HPR99, Sch91d, Sch93b, Mar95, Row95]. Methods [Fla90, Hag91, Pri92a, Ric91, Pet88]. METRIC [Ver90]. Meyers [All99b, McL94]. MFC [Ano97-33, Bav99a, Joh99a, Kos99, Swa97]. Mice [Bug91, Paj88]. Michael [All95c]. Micro [Art89, Joh88a, Vol96f]. Micro-C [Vol96f]. Microcontroller [vdB91]. MicroEMACS [Law92a, LP92, Ano88c, Man88]. MicroPlox [WH89, Pat88a]. Microprocessor [RP91]. Micros [Pri90d]. Microsoft [Cru89, Hun89, Mag93, Mar88b, Obe91, Paj88, Sch90c, Sch90d, WL94, Yok88, Pat92a]. MicroSpell [Law92b]. MIDI [Ben92, Law92c]. Migration [Sak91b]. Migrator [Hav89c, Hav89a]. MINED [Vol94d, Wol94]. Mini [Zol90b, Zol90d]. Mini-Database [Zol90b, Zol90d]. Minimalist [Web94]. Minimum [Pug91c]. Minor [Sak95f, Sak95g, Sak95h]. Miscellaneous [Wei92c]. Miscellany [BBAH88]. mismatche [Pug88k]. Mistake [Sak97i]. Mitchell [Pav89]. MIX [Phi92a]. Mixed [LaB93b, Pug89l]. Mixing [Pru91, War91c]. MKMK [Vol90g]. MKS [Vol89c]. MNP [Pea90]. Mode [Olh90, Pas88]. Model [And88c, Pla89c, Pug89l, PF92b]. Modeling [Moo91, Ols93]. Modem [PA93b, War91a, Hel91, Hel94]. Moderator [Wei91b, Wei91f]. Modes [Olh90]. Modifications [War88b]. Modifier [Bak94]. Modifying [Lam89, PS92b]. Modula [Bat91]. Modula-2 [Bat91]. Modular [Phi98b]. Modules [Eme88, Sak91d]. Molecular [Vol94b]. Monetary [Pla91f]. Money [Mar92]. money get [Pla98e]. money put [Pla98g]. moneypunct [Pla98f]. Monitor [Pin93, Rog89]. Monitors [Shi90]. Monthly [War89b].

Morphing [Phi95b]. Morte [Sch98e]. Mortem [KB99d, KB99c, KB99a, KB99b]. Motif [JR91e, LS98]. Mouse [Cur90, Han90, Jar91a, Kel90, Mar88b, Paj88, Mor92]. Moves [Cla88]. Moving [Pug94h, Pug95b]. MS [All88, Ano88d, Bar95b, Bro90a, Bur89, Erd88, Fie88e, Fre88a, Ger88, Gra93b, Hal89, HSC89, Hic89, Joh90b, Kle97, LaB94, Lad89a, Lad91, Lam89, Lan89, Lan91b, MO92, Mat88, Ohl92, PD88, Phi93f, Pin93, PLBL92, Pug88j, Pug88s, Pug88t, Pug88x, Pug89m, Pug90b, Pug90r, Pug90-28, Pug90-33, Pug90-32, PN91, Pug92v, Rea91a, Vol90f, War88i, Wei90a, Wit88, Zer89, Zol90e, Zol91a, dF90b]. MS-DOS [All88, Bro90a, Bur89, Erd88, Fie88e, Fre88a, Ger88, Gra93b, Hal89, HSC89, Hic89, Joh90b, Kle97, LaB94, Lad89a, Lad91, Lam89, Lan89, Lan91b, MO92, Mat88, Ohl92, PD88, Phi93f, Pin93, PLBL92, Pug88j, Pug88s, Pug88t, Pug88x, Pug89m, Pug90b, Pug90r, Pug90-28, Pug90-33, Pug90-32, PN91, Pug92v, Rea91a, Ver89, Vol90f, War88i, Wei90a, Wit88, Zer89, Zol90e, Zol91a, dF90b]. MS-Windows [Bar95b]. MSC [Kwo89a, War88d]. MSI [Vol90h]. MSU [HKB90]. Much [Wio88c]. Mullin [Bur90b]. Multi [Bed91, Car88, Got94, Jae91a, Jae91b, Pla89d, Pri90c, PG93, Pug93q, Sch90e, Sco88, Vol94d, Wei88a]. Multi-Byte [Pla89d]. Multi-Dimensional [Jae91a, Jae91b, Pug93q]. Multi-file [Wei88a]. Multi-Language [Pri90c]. Multi-Lingual [PG93]. Multi-platform [Vol94d]. Multi-Precision [Got94, Sco88]. Multi-Tasking [Car88]. Multi-Threaded [Bed91]. Multi-Threading [Sch90e]. Multibyte [Pla92i, Pla99h]. Multidimensional [Rah97]. Multijoy [Vol95b]. Multilanguage [SR99b]. Multiple [HW92a, Kat99, Kis96, PG91, Pri92b, Pug92h, Pug92t, Sch99j, Shi91]. Multiply [Wil95a]. Multiprocess [SDWB96]. Multitasking [Col90d, Hal91, She92, Vol90c, Vol90i, Vol91b, dC90a]. Multithreaded [Sut99a, Vol93a, VE93]. Multithreading [Dug95, Har99, Kel94, KF97, Man91, Ten98, Lar97]. Multiuser [Fro92]. Must [BC89a, Wil96b]. Mutable [Sak95i]. Mutations [Sch96f, Sch96g, Sch96e, Sch96h]. My [Sch97f]. mysteries [Lib92, vdP93].

Nackman [Swa95a]. Name [Bar91b, Mul93, Pla88h, Pla93k, Pug90o, Sak91a]. Name/Obscure [Pug90o]. Named [Dan97]. Names [Pug90c, Pug91f, Sak93d]. Namespaces [Sak98a, Sak98d, Sak98e, Sak98f, Sak98c]. Naming [Hed90, Pug90v, Pug91f, Sta89]. Namir [Dos95]. Narain [Pla90c]. Nash [All93f]. Nasties [Pla90j]. National [Sch90b]. Natural [Sue93a, Sue93b, Wei88c, Wio88d]. Nature [PS91a]. Navigation [Zol90e]. Nearest [Pro92]. Need [Han94]. Needed [Car93, PA93b, Wei91b]. Needs [And94, Bec98f, Vol90k]. Neighbor [Pro92]. Neophyte [BC89c]. Nested [Bel96, Lea96, Sak93e]. Net [Wei90d]. NetAcquire [Ano97-50]. NetBIOS [Bed95, Sch88c, Sch92c, Per89]. Nets [Wei90b]. Network [BK94a, Col89d, Nad98, Phi96b, Rua97, Sea98, Ste94, Wei91e, Wei93g, Wei94d, Bro94]. Networks [Heg93, Phi92d, Phi92c, Wei93g, Blu92]. Neuberger [Pla94b]. Neural [Blu92, Col89d, Phi92d, Phi92c, Phi96b, Ste94]. new- [Sak97g, Sak97h]. NEWMAT [Gra93d]. News [Fie89b, Wei91e, Wei93g, Wei94d]. NeXT [Tay92, War89e]. Niche [War89e]. NNUTILS [Ste94]. No [Deg97, Sch99d, War89f, Plu88b]. Nodes [Wal91]. Noisy [Sto95]. Non [Pla90b, dC90a, Pug92a]. Non-ANSI [Pug92a]. Non-Displayable [Pla90b]. Non-Preemptive [dC90a]. Nonrandom [Bec95d]. Normal [Wei92e]. Notation [Sak95a]. Note [Pla90m, Pug90g, Pug93p, Tsa90]. Noted [Rob94]. Notes [HPS90, Lop94, Pip94, Pla92j, PBW92, PA93a, PN93, Pug88v, Pug90w, Pug90x, Pug90-37, Pug90-38, Pug93f, Pug93o, Pug93s, Pug93t, Pug93v, Pug94b, Pug94i, Pug94k, War88h]. NRO [VM88]. NT [Bel98, Cus93, Guf98, Hoo99, Hal93d]. Null [Jae88g, War88d, vH92, Jae88g]. num get [Pla98h]. num put [Pla98k]. Number [Dwy95, DW96a, DW96b, Fru90, Gin93c, Mar96, Smi92e, Sue93a]. Numbers [LaB93b, Phi91c, Pug88q, Pug92v]. Numeric [Fos93, Kru95, Pla92o]. Numerical [Irv91, Kem87, Phi90c, Pla93n, Pre88, Pre92, Pri88, Ric91, Sti92a]. numpunct [Pla98k].

O [PS92a, Smi89, Kri97, Man98, New91, Pla89c, Pla89h, Pug88m, Ste87, Wil92b, WP88, Zol90b, Zol90a, Zol90c, Zol90d]. Obfuscated [Lib89b, Lib92, Nol94, vdP93]. Object [Ada94, All96b, All99e, Ano97-32, Ano97-33, Ano98z, BD89a, BD90d, BD91, Bec99f, Bec99g, Bec99h, Boo91, Bru90, Bur90b, Cla90b, Cla90c, Coc98a, Coc98b, Col90b, Col93a, Cro98a, Hav90, Hav93, KJV98, Lan91a, Mul89, Pat95, PH96, Rog96, Row95, Sch89a, Str91a, Wei88d, Wel95, Wel99, Whi90a, dC90b, Blu92, Dor95, Gam95, Goo89, Mar95, Won93]. Object-Based [dC90b]. Object-Oriented [All99e, Ano97-32, Ano97-33, Ano98z, BD89a, BD90d, Bru90, Bur90b, Cla90c, Coc98a, Coc98b, Col90b, Col93a, Hav90, Hav93, KJV98, Lan91a, Pat95, PH96, Rog96, Row95, Sch89a, Str91a, Whi90a, Blu92, Dor95, Gam95, Goo89, Mar95, Won93]. Objective [Tay92]. Objective-C [Tay92]. Objects [All99f, All99g, Bec97e, Bec98l, Col96, Eck89, Eib94, Flo98, Gab99, Jae89d, Lan90, Mey98, Nel96, Nel98, RL98, Ric99, Ske91, Smi92g, Whi91, Wil93a, Hol92]. ObjectSpace [Ano96b, Ano97-33]. Obscure [Pug90o]. Observer [Bar98]. Obstacles [CE96c]. October [Ano94a]. Octree [Ash95, Sch91d]. ODBC [Bav99c, Bey98, Rag95]. Odd [JR91c]. Odd-Shaped [JR91c]. Offers [Ano97t, Bru89a, Fre88a, Vol88a]. Official [Wel91]. offsetof [Pug90y]. Old [Boc96]. Olympic [Sto95]. Omission [Eve93]. On-Line [Vol90k]. On-The-Fly [Eng93a]. One [CH88, Lib90f, Pug90e, Sch99e]. One-Liners [CH88]. O’Neill [Pla93g]. Online [Ign98]. OOP [BD90c, BSG92, HW92b, Jer91, Mur93, Tsa94]. OOPs [Rat90]. Open [Pug93l]. OpenGL [Bav98b, Bav99b]. Opening [Pug89m, Pug90b]. Operating [Bus91, Car88, Car91, Rod95a, Tan95]. Operations [Mor98a, Mor98b, Phi92g, Phi94b, Phi95a, Phi95c]. Operator [Jae88a, Jae88b, Pla93b, Pug92a, Pug92u, Pug93p, Pug94j, Sak92c, Sak92d, Sak92e, Sak92f, Sak93a]. operator= [Bec97a]. Operators [Bec96h, Jae90g, Pug93a, Pug93r, Pug94a]. Optimizations [Sut99a]. Optimized [Knu98]. Optimizing [PN93, Pug89n, Sch93b, Ter91]. Options [Col90b, Mal90]. Oracle [Bru89a]. Orbit [Lon91a, Lon91b]. Order [Gab99, Plu89b, PH91b, Pug92i]. Ordering [Erd95, Pla96l]. Organization [Bur92a, PS91a, Tha88]. Oriented [All99e, Ano97-32, Ano97-33, Ano98z, BD89a, BD90d, Bru90, Bur90b, Cla90c, Coc98a, Coc98b, Col90b, Col93a, Hav90, Hav93, KJV98, Lan91a, Pat95, PH96, Pug88m, Rog96, Ros94b, Row95, Sch89a, Str91a, Whi90a, Blu92, Boo91, Dor95, Gam95, Goo89, Mar95, Mul89, Won93]. Original [Sma88]. Origins [Lop94]. OS/2 [Pro90, Ben92, Cor88, Gra93b, Jen96, Kel94, PM89, Wit90c, Wit90d]. OSE [Pat96]. ostream [Pla94k, Pla94m]. Othello [CW89, WC89]. Other [Bec98h, Nel89a, Pri92a, Pug88j, Sak95k, Sak95j, Sak95l, VM88, vdP93, Hol87, Lib92]. Oualline [Phi94a, Web96]. Our [Sch97a]. Output [Del88d, Del88f, Fro89a, Pin93, Pla89e, Pug89e, Pug90f, Pug90-32, PG93]. Outs [Jae88d]. Overflow [Pug90-30]. Overhead [Dee95]. Overlay [Phi94b]. Overlays [Pug88-30]. Overloaded [Bec96h, Pug94a]. Overloading [Sak91a, Sak92c, Sak92d, Sak92e, Sak92f, Sak94j]. Overriding [Sak94j]. Overview [Bal91b, Pla96m]. OWL [Chi97]. Own [Bec96b, BS99, Nel90a, Pla91b, Poi89a, Sak90b].

P. [War90d]. Package

[Bra92b, Cru88b, Har90, Pug88-28]. Packages [Mea92, Pug88f, Pug88-31]. Packaging [All99f]. Packetized [Sti90]. Painter [Fer88]. Panel [Hin88b]. Panels [Bro90b]. Panning [Cha88]. Paper [Sch99f]. Papers [Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r]. Parallel [Pla90m, Pla90e, Pla90d, Smi92h, Yor96, Bab87, Bab88]. Parameter [Ano98a]. Parameters [Wei88c]. Parentheses [Bec96e, Pug94j]. Parser [And99b, Clu89a, Kou97, Vol88b]. Parsers [Sak96h, Vol88b]. Parsing [Del88a, Sak96f, Sak96g, Sak96i, Wio88d]. Part [All93b, All93c, All93d, All93h, All93i, All93j, All93k, All94c, All94g, All94h, All95d, All95e, All95f, Bal91a, Bal94, BD90a, BD90b, BD90c, BD90e, BD90f, Bec99c, Bec99d, Bec99g, Bec99h, Coc98a, Coc98b, DW96b, Jae90i, Jae90j, Jae91i, Jae91a, Jae91b, Jae91c, Jae91d, Jae91e, Jae91f, Jae91g, Jae91h, Jae91j, Jae91k, Jae91l, Jae92h, Jae92a, Jae92b, Jae92c, Jae92d, Jae92e, Jae92f, Jae92g, Jae95a, Jae95b, JR91a, JR91b, JR91c, JR91d, JR91e, Lib90d, Mor98a, Mor98b, Phi91b, Phi91c, Phi91d, Phi91e, Phi91f, Phi92e, Phi92f, Phi92g, Phi93a, Phi93e, Phi93d, Phi94b, Phi95a, Phi95b, Phi95c, Phi96d, Pla90n, Pla90o, Pla90q, Pla96c, Pla96d, Pri96a, Pri96b, RJ91, Sak92d, Sak92e, Sak92f, Sak93b, Sak93c, Sak93h, Sak93i, Sak94a, Sak94b]. Part [Sak94c, Sak94d, Sak94e, Sak94f, Sak94g, Sak94h, Sak95k, Sak95j, Sak95l, Sak96a, Sak96b, Sak96c, Sak96f, Sak96g, Sak97g, Sak97h, Sak98d, Sak98e, Sak98f, Sak99a, Sak99b, Sch96f, Sch96g, Sch96e, Sch96h, Sch96i, Ske93a, Ske93b, Zol91d, Zol92b, Zol92c, Zol92d, Zol92e, Vol98a]. Partition [All94a]. Partitioning [Sak98d, Sak98e, Sak98f, Sak99c]. Party [Sch99g, BK91, BK94b, Bro94]. Pascal [Cam98]. Passing [Eck89, PB92, Pug90u, Pug92j, vH92]. Passwords [Bec97f]. Patch [Gib99]. Path [Sak91b]. Pattern [Ano98m, Bar98, Bat97b, Gou98, Hob91b, Ott94, Pic90]. Pattern-Matching [Ott94]. Patterns [Pat95, Rog89, Gam95]. Paul [Hoh90, Swa95c]. Paving [Sak91b]. PC [Bin88, Cas93, Amm87a, BK91, BK94b, Con91, Dea90, Efk97, Han90, Her94, Hin90a, Kea90, Lan89, Lop94, New91, Owe95, Pug90-27, Pug90t, Pug91g, PF92b, Rot91, Sma88, Smi92h, Ver90, Vol90j, War92c, Whi88, Yor96, dFAAW90, Sch88a]. PC-METRIC [Ver90]. PC-Tags [Vol90j]. PC/VI [Lop94, Sma88]. PCcurses [Lar90]. PCX [Ash91a, Ash91b, Bar90c, Rob94]. Pennies [Pri91b]. Penumbra [Sch98f]. Perfect [Bur92b]. Perfectly [Ban88]. Perform [Pug88m]. Performance [Man99, Pri93]. Performer [Fie88d]. Perinotti [Pla94d]. Periodicals [Ogg90a]. Perl [Wei90g]. Perry [Tho92]. Persistence [All96b]. Persistent [Flo98, Get99, Lop97]. Perspectives [Sak96h]. Peter [Wal88]. Peters [Fla90]. Phantom [Sch99h]. Phases [Pla89f]. Philip [Wal88]. Phoenix [Sch98l]. Pi [Pen91]. PIC1650 [Col88c]. Pickover [Hob91b]. Pictor [Vol95b]. PID [Har97]. Piecewise [DE98]. Pipeline [Pri90d]. Pipes [Dan97, Hey96, Wit90c]. Pitfall [Bec95e]. Pivot [Blo93]. Place [Bec98a, DP93]. Placement [PS91a, Pug94k, Sak97f, Sak97e]. Plain [Phi98c]. Plane [Sch88b]. Plastic [Sch99f]. Platform [Gui97, Vas96, Vol97a, Zek95, Vol94d]. Platform-Independent [Gui97]. Platforms [SX98, Xia97]. Plauger [War90d]. Play [Nel95, Pla88g, Won93]. Player [WC89, Wei94c]. Players [Poi89a]. Plots [Ste93b]. Plotter [Flo89, War88g]. Plotting [Bra89, Chi97, Pat88b, Smi92a]. Plug [Nel95, Roe99, Won93]. Plug-Ins [Roe99]. Plum [Phi93c, Phi93b]. plus [WPM84, WPM87, Par88, Pav89]. Point [Blo93, Fos93, HD95, Lab98, Mat91, Pla93h, Pug90i, Pug93h, Sch90c, Sch97i, She90, Smi93f, Smi93g, Smi94a, TC95, Man96a, Wil93b]. Point-And-Shoot [Mat91]. Pointer [Met92, New89, Pug90e, Pug90c, Pug91e, Pug92d, Pug94l, RY91, SS89, Sch97h, Ske93a, Ske93b, War88d, Pug89o]. Pointer-Pointers [RY91]. pointer-to-a-function [Pug89o]. Pointers [All93i, All93j, All93k, All95c, Ano99z, BD89b, Col95b, Jae88g, Jae89a, Jae89d, Jae90h, Jae90l, Jae91a, Jae91b, Mas95, Pug89d, Pug89g, Pug89h, Pug89l, Pug90w, Pug90-38, PA91, Pug92k, Pug93b, Pug93e, Pug93o, Pug93q, Pug94m, Pug94r, Pug94q, Pug95a, Pur89a, RY91, Ske94, WK92, vH92, Dac95, Jae86]. Pointers-To-Functions [Pug89d]. Points [All92b, Jae89f]. Polled [Wen97]. Polygons [Rév93]. Polyphase [Nel97]. Pond [Hav93]. Poor [Beh96, Kna89]. POP3 [Hat96]. PopMail [Hat96]. Port [Joh88a, KS98, Smi92h, Yor96]. Portability [Bec98h, BC89a, DP93, Flo90, Ger90, Gim89, Lad91, Poo93, Pug93l, Pur91, Sak89, Siy89, Van93, Ver89, TK89]. Portable [Byb92, Coo90, Dwy95, Erd95, Fri90, Ger88, Jae89g, Kat99, KF97, Kri97, MA88, Mic96, New91, Pla90d, Plu91, Rem97, SDWB96, Sek89, SR99b, Smi88a, Sta89, Wam96, Wei90h, Wei93i, Zol91b, Zol92a, Zol92b, Zol92c, Zol92d, Zol92e, vdB91]. Portably [Whi89b]. Porter [Pla95a]. Porting [Bru92a, Kal98, Pug90-27, Pug90-28, Smi92f, WH89, WL94]. Ports [Man98, Rib89]. Positioning [Wal91]. POSIX [Cho88]. Possibility [Pug90-29]. Post [KB99d, KB99c, KB99a, KB99b, Pug92i]. Post-Increment [Pug92i]. Post-Mortem [KB99d, KB99c, KB99a, KB99b]. Postfix [Pug93r]. Postman [Ram92]. PostScript [Fie89e]. Potential [Gim89, Wil96b]. Power [Jon88b, Phi92a, Ske93a, Ske93b, Swa95c, CD94]. Powerful [Del88e, Mue94a, Ske94, Vol94d]. Practical [Del88e, HW92a, War90e, Web96, Zol90d, McC93]. Practice [Ano98a, Phi99b, Sak96a, Sak96b, Sak96c, Sak96e, Sak96d, Sak97d, Sak97a, Sak97b, Sak97f, Sak97e, Sak97g, Sak97h, Sak97c, Sak97i, Sak98a, Sak98b, Sak98d, Sak98e, Sak98f, Sak98c, Sak99c, Sak99h, Sak99f, Sak99e, Sak99g, Sak99a, Sak99b, Sak99d]. Practices [McC89]. Pragma [Bec97h]. Prata [Pav89]. Precedence [Jae90g, Pug90-39]. Precise [Pla88f]. Precision [Got94, Pri94, Pri96a, Pri96b, Sco88, Smi93f]. Predictor [TR94]. Preemptive [dC90a]. Prefix [Pug93r]. Preprocessing [Pug89p, Pug90x]. Preprocessor [All94i, Art90, Man90, PT93, Pri89b]. Preserving [Pla88d]. Press [Phi90c]. Preventing [Wee92]. Price [Wei90g]. Pricing [Hav90]. Prickly [Bec97h]. Primer [Hob91a, Jae95a, Jae95b, Par88, Pav89, WPM84, WPM87, WF90, WF92]. Primitive [All99h]. Primitives [Pla89h, Pla91o, Pla92k]. Principles [Kam89, Pug94b, Sak98b, Amm86, Amm87b, Amm92]. Print [Ric91, Ste90]. Printable [Smi94b]. Printer [Ock89, Pug88w, Pug90-32, PM91b, War88g]. Printers [Pat88a, Pug90f, PdW92, Smi93b]. printf [PP90, Pug89e, Pug90-40, All92c, PF92b, San93]. Printing [Deg97, Phi91c, Phi91d, Pug90p, Pug90-35, PdW92, Smi93b, Tam97]. PrintScreen [Pug89f]. Priorities [Bec99l]. Private [Man99]. PRO [Ano97-43]. Problem [Fis90a, Haw97, Nel89b, Pug90-31, San93, Wei88c, Bar89, BG94, Kna89]. Problems [Bec95d, Dic94, Far89, Gim89, Pug88z, PM89, Pug90t, Pug91e, Pug92l, Pur89a, Sch90d, Wat88, Wri90b, Yok89]. Procedure [Bed95, Ber92a]. Procedures [Smi92d]. Process [Car99, Fal90a, Gle90c, Hed91, PB92, PR91, PJ92]. Processes [Kat99, Wei92i, Wei96]. Processing [All92b, All92c, All92d, All93g, All93h, All93m, Bak91b, Col95a, Cru88b, Fre93, Kas96, Ker96, Mar90a, Mea91, Nad98, Phi91a, Phi91b, Phi91c, Phi91d, Phi91e, Phi91f, Phi92e, Phi92f, Phi92g, Phi93a, Phi93e, Phi93d, Phi94b, Phi95a, Phi95b, Phi95c, Phi96d, Pug90-30, Zol91c, Zol91d, Tha88, Bur92a]. Processor [Cre90, Sue93b, Sue94]. Processors [Pri90d, Bab87, Bab88, Pla90e, Pla90d]. Produce [WO89]. Produces [Eme88, Vol88b]. Production [Phi98a]. Products [Ano97-31, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano97-29, Ano97-30, Ano98n, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano99n, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y]. Professional [Kea90, Ste93b, Swa95d]. Profiler [Wit88]. Program [All94b, Bur90b, Kam97, Lib90f, Mey96, Pug93f, Ros94a, SD89, Ste92b, Zol90d, Mul89]. Programmer [Ano97-34, Ano97-35, Ano97-36, Ano97-37, Ano97-38, Ano97-39, Ano97-40, Ano97-41, Ano97-42, Bri98b, Chi89, Hin90a, Ogg90a, Ogg90b, Per89, Phi94a, Pro90, Sch89b, Spa90, Wal89, Wio89, Zmo96, BK91, BK94a, BK94b, Bro94, Cam87, Cam93, Cor88, Pur89b, Pur92, Sch88c, Sch92c, Wei88a]. Programmers [Bro89, Fre88a, Gin93c, Gut94, Hin88a, Plu88c, Sco93, Jae93, Oua92, Bri99a, Lop93]. Programming [All93f, All99a, All99e, Amm86, Amm87b, Amm92, And88b, Ano97-32, Ano98z, Ast94a, Bab87, Bab88, Bin88, Bru90, Bru89b, Cam94, Cas93, Cep95, Che96, Cof88, Col96, Col93a, Eib94, Fio89, Fla90, Gin93b, Hal93c, Jae90a, JR91a, JR91b, JR91c, JR91d, JR91e, Kam89, KR88, Ker96, Koc83, Koc88, Koc94, Kou88, Kra89, Kre96, Kru92, LS98, Nel90b, Phi93c, Phi93b, Phi98b, Pla90m, Pla90c, Pla90e, Pot88, Pug88p, Pug88o, Rag95, RJ91, Ros97, Sch90b, Str91a, Str91b, Swa95c, TC94, Tay92, Tom91c, TG89, Wal89, War86, War90f, Wat93, Wel94b, Whi90a, Wit95, Bar89, BG94, Cal93, Cal95, Car92, CD94, Dor95, Gan92, GR89, Goo89, HS87, Hel95, Hol92, Jae86, KK89, KR78, MQ87, Pet88, Plu84, Plu89a, RN93, Ste87, Str86, Str87, Str95]. programming [TK89, WPM84, Won93, PS92a, Cla90c, Gra89, Nel89b, Nel95, Phi99b, Rod95b, Smi89, Web96]. Programs [BT91, Bur89, Cog88b, Dan98, Fie88a, Pla90d, Plu88c, Pug88s, Pug91g, Pug92j, Vol93a, VE93, Hun89, Mag93, Mey92, Oua92, Phi94a]. Prohibited [Pug88y]. Project [BB94, Ham98, Fie89c]. Projections [Low95]. Projects [Bri98a, Wei88a]. Prolog [dF90a, Spr90]. Promise [Sak96d]. Promotes [Pla88d]. Promotion [Bec95g, Bec98h]. Prompting [Pan93]. Propagation [Lon91a, Lon91b, Man96b]. Proper [PH93]. Proposal [St.93]. Proposed [Bak94, Bru91, Tri93]. Pros [Bat91]. Protected [Olh90]. Protocol [Beh96, bG91, Joh90d, Tay93]. Prototypes [Pug94i, War88e]. Prototyping [Ble98, GK90, Hin88b, Mar90c]. Prove [Del88c]. PrtSc [Pug88w]. Pruning [Hav89b]. Pseudo [Clu89a, Cro98b, Fru90]. Pseudo-Parser [Clu89a]. Pseudo-Random [Fru90]. Pseudo-Terminal [Cro98b]. Pssssttt [Gra88]. PSUtils [Dug94]. Public [Plu88d, Vol89d]. Publication [Pla92d, Zol91c, Zol91d]. Publications [Vol90k]. Publisher [War88h]. Publishers [Hof92, War90c]. Pulse [Gil95a]. Pun [KB92, War90b]. Punctuator [Jae88a]. Puns [Ano91a]. Purdum [Ogg90b]. Purpose [Man98, WO89]. Put [Lib91a]. Puzzle [Mar91]. Puzzles [Jae90i, Jae90j, Jae91j, Jae91k, Jae91l].

QNX [Bar91a, Boy88, Bus91]. qsort [Pug90-31, Pug95c, All93l]. Quad [Pri96a, Pri96b]. Quad-Precision [Pri96a, Pri96b]. Quadcode [Van92]. Quadrature [Sti92a]. Qualifier [Jae89b]. Quantization [Ash95]. Queries [Cla99]. Questions [Bec96f, Bec96g, Bec96h, Bec96j, Bec96i, Bec96k, Bec97a, Bec97f, Bec97e, Bec97b, Bec97c, Bec97g, Bec97h, Bec97d, Bec98b, Bec98i, Bec98c, Bec98e, Bec98d, Bec98j, Bec98f, Bec98g, Bec98h, Bec98k, Bec98l, PF94, Pla99f, Pla99g, Pug89q]. Queues [Jae92c, Jae92d]. Quick [Abo91, Ano88d, Bar95b, Bat98, Bro89, Dwy95, Pri90d, Ste93b, Fis90a, Hun89]. Quicksort [Nel90a, Van96, Cel91, Blo93, Jae90k]. Quiero [Sch98k]. Quiet [Pla90n, Pla90o]. Quip [SB89]. Quirk [KB90].

R [Pla94d, Ano97-43, Vol90k]. R&D

[Ano97-45, Ano97-46, Ano97-47, Ano97-48, Ano97-44, Ano97-49]. R-Active [Ano97-43]. Radiosity [Ash92]. Ralf [Swa95b]. Ralph [Pat95]. Ranade [All93f]. rand [Bec95c, Bec95d, Bec96c]. Random [Dwy95, DW96a, DW96b, Fru90, Gag97, Mar96, Phi96d, Sco93]. Ranges [Phi99a]. Rapid [Ble98]. Rasmol [Vol94b]. Rational [Ano96c]. Raw [Pas88]. Ray [Smi92g]. RDCF [Vol95f]. Re [Pug90-32]. Re-Directing [Pug90-32]. Readability [Ges91]. Readable [Pla92m, Ret96]. Reader [Ban89, Clu89b, Vol98a]. Reader/Writer [Ban89]. Readers [Bec98d]. Reading [Gib90, Mat88, Phi91b, PF93a, Pug88u, Pug89m, Pug90-33, Pug91g, Pug92m, PC92, TP92, Tuc98, War92c]. Reads [Hal89]. Ready [Pat92a, Hun89]. Ready-To-Run [Pat92a, Hun89]. Real [All92a, Car88, Fee93, FM91, Gin93c, Gin93a, Hal91, HJ90, Kra89, Man89, Nad98, VC89, War90e, Wio88b, dC90b, Lab92]. Real-Number [Gin93c]. Real-Time [All92a, Car88, Fee93, FM91, Gin93a, Hal91, HJ90, Kra89, Man89, VC89, War90e, dC90b, Nad98, Lab92]. Realism [Ash92]. realloc [Bec96k]. Really [Pug95a]. Reasonably [MA88]. Recipes [Phi90c, Pre88, Pre92]. Recognition [Gle90a]. Recommendations [War90g]. Recommended [McC89]. Reconsidering [Sak93g]. Reconstructions [VG89b]. Record [Pug88m, Ros94b]. Record-Oriented [Pug88m, Ros94b]. Records [Kas96, PC92]. Recovering [Jan93b, Sak96i]. Recursion [Syc91]. Recursive [Bri99e, Pug91h]. RED [Rea90a, Pla99c]. Red-Black [Pla99c]. Reducing [Dee95]. Reduction [Kam94]. Redundancy [Fel90, Lou92]. Reel [Pla90r]. Ref [Del88b]. Refactoring [Bri98a]. Reference [Ano98m, Del88e, Nga97, Ogg90a, Pug94r, Rec98, Sak91c, Vos90, War91a, BK91, BK94a, BK94b, Bro94, Hel91, Hel94, Lad92]. Reference-Counting [Ano98m, Rec98]. References [BDB89, Bar90d, Gre89, War89h]. Refines [Plu88a]. Reflective [Sch98g]. Refresh [PM91a]. Regions [Web93]. Register [MS91]. Registers [Cha90, Cog88b]. Registration [Bal97, Sak95d, Sak95g, Sak95h]. Regular [EU98]. Relate [Wei95c]. Related [Hal92]. Relational [All96b]. Releases [Ano96b, Ano97-31]. Reliability [Phi96a]. Religion [War89f]. Remediation [Iof99]. Remote [Bed95, Vol95d]. Removal [Ale91]. Removing [Syc91, Tom91b]. REMZ [HKB90]. Rendering [Smi92g]. repair [Sti92b, War92b]. Repeat [Pug90n]. Repetitive [Gra93f]. Replacement [Cha95]. Replicates [Sma88]. Replies [Pug90-34]. Report [NW99]. Reports [Cla99, WC90]. Representing [Bre96]. Request [Fan99]. Requests [Vol98a]. requirements [Wil96b]. Resampling [Hag91]. Rescue [Bec96j]. Resident [PS92a, Smi89, Ste90, Ste87]. Resolution [Lan89]. Resolving [Bec97b]. Resource [Hea97, Mon97, Wal89]. Resources [Plu91, Vol97a, Vol97d]. Response [Coh92, Gil95a]. Rest [All93k, Jer91]. Restricted [Ano99z]. Restrictions [Bec98k, War89g]. Results [Ano91a, KB92, Plu88c, War90b]. Resumé [Lib91a]. Retrospective [Hin90b]. Return [Jae88f, Pug89o, PR91, Pug92n, Pug93i, Pug93m, Pug93s, Pug94e, Sak94k, Pug88k]. Returning [Eck89]. Returns [Hel90, Pug92t]. Reusable [Bat97b, Bri96, Cra97, Har97, Pat95, CE95, Gam95]. Reveals [Rog89]. Review [All99a, All99b, Bri98a, Bri98b, Bri99a, Bry98, Cla97, Ham98, Lar97, Mil97, Phi99b, Plu88d, Plu89b]. Reviewer [Wio88c]. Reviews [Wei91d]. Revised [Plu88d]. Revising [Pla99i]. Revision [Del88c, Phi93f, Pla94q]. Revisited [Lea96]. Reviving [Sch91c]. Rewriting [Sak91d, Sak93g]. Rex [Far88, Lop93]. RFVDEMO [Vol94c]. Richard [Pat95]. Rid [Bec95b]. Rights [War89c]. Ring [Sti90, War88a]. Rising [Sch98l]. Risky [Bec95f]. RMAXTask [Coo92b]. Road [Gra92a]. Robert [Hal93b, Pla90e, Row95]. Robust [War88a]. ROFF5 [Ber90]. Role [Pla88g]. ROM [Vol99, Bro94]. ROMable [Pug88v]. ROMLDR [All94b]. Roome [Pla90c]. Rose [Ano96c]. Rounded [Ban88]. Rounding [And94]. Routine [Byb92, GW91, Hoo99, PM91b, Sch90d, Smi92a]. Routines [Pat92a, Cog88a, Gra91, Pug88r, Vos90, Zol90a]. Row [San93]. RPC [Joh90d]. RTI [Ano97-50]. Rules [Pla88c, Pla88d, Pla90k, Pla93c, Sak95e, Wei92a]. Ruminations [Mil97]. Run [Pat92a, Ban94, Pug94n, SR99b, Vol90h, Hun89]. Run-Time [Ban94, Pug94n, SR99b]. Runtime [Joh90a, Pug90r]. Rvalues [Pla93b].

S [Phi96c]. S6 [Col88b]. Safe [Wei94h, Wis97, Rin99]. Safer [Erd93, Joh99a]. Safety [Ano99m, Pla98a, Pla98l, Ric99, Sch96a]. Saks [Phi93c]. Salama [Gra89]. Salesman [Kna89, LFK89]. Same [Mey96]. Sampler [Pla91p]. Sand [Sch99a]. Sargasso [Sch97g]. Satir [Web96]. Save [Pla93a]. Saves [Sch89b]. Saving [PG92]. Say [Sch97e, Sch99d]. sbrk [Sch91c]. Scaffolding [BC89b]. Scaling [Kie95, Web94]. Scan [Gib90]. scanf [All92b]. scape [Ash90]. sccs [Wei88d]. Schedule [War89b]. Schedulers [War90e]. Scheme [Wel90]. Schildt [Haw88, Pla94e]. Schwaderer [Per89]. Science [Cam94]. Scientific [BN94, Chi97, Phi90c, Sch96k, Swa95a, Wel94b, Pre88, Pre92]. Scientists [Ric91]. SCO [Fie88d, Joh88a]. Scope [Sak95e]. Scott [All99b, Hal93b, McL94]. Screen [PS92a, Fre88b, Gra91, Kno90, Mic96, New91, Pat92b, Pin91, Pug89n, Pug90l, PM91a, San93, Sch92a, Smi89, Tal89, WO89, Ste87, Fer88]. Screen-Clearing [Pat92b]. Screens [PG92]. Script [Hin88a]. Scrolling [Cha88, Wel94b]. SCSI [Bri98b].

SD [LFK89]. SDI [PH96]. SDI-12 [PH96]. SDK [Kuo97]. SE [Osb90]. Sea [Sch97g]. Search [Bou94, GW91, Mic96, Rum98, Ste92a, Vol94a, Vol96d]. Searches [Pop92]. Searching [Gro91, Phi99c, Pla92l, Ros95]. Second [Plu88d, Sak97a, Sch98h]. Secondary [Pin93]. Secrets [Alg95, BG99, Kil96]. Section [Tac94]. Sections [Smi92g, Wei92i, Wei95e]. Sectors [Pug88u]. sed [Hin88a, Bar90d]. Seeking [Vol97b]. Segment [SS89]. Segmentation [Phi93e, Phi93d]. Segments [PY90]. Selectable [Dee96]. Selection [Blo93]. Self [LaB94, MQ87]. Self-Extracting [LaB94]. self-teaching [MQ87]. Semantic [Pug89q]. Semaphore [Bec99l]. Semaphores [Fro90]. Semi [Vol97e]. Semi-Annual [Vol97e]. Sending [Faz98, Kam97, Ock89, Pug88w]. Sensible [Sak95a]. SENTINEL [Sin93b]. Separating [Eme88]. September [Sch99i]. Sequence [Jae89f]. Serial [Beh96, Chu91, Eme88, KS98, Pug92v, Whi89c, Wio89, Woo97, Cam87, Cam93]. Series [Gua90]. Serious [Bin88, Cas93, Eve93]. Server [Fan99, Joh90d, Ste97, Vol96g, Ano96a]. Server-Based [Col95a]. Service [Qua88, Wei91a]. Services [Bel98, Sin94, Vol90k]. Serving [Wat88]. Servo [McC94]. Set [Dan93, Phi99a, Pla91c, Pla93m, Pla93l, Vol90f, Gua90]. setjmp [Erd93]. setjmp.h [Pla91j]. Sets [Pla89d]. Setting [Pas88, Pug88x, Pug92o]. Setup [PM91b]. Seven [Cla97]. Several [CH88, Fie89e]. SGPC [TC94]. Shades [Phi93d]. Shading [HP98, VG89b]. Shafer [Phi90a]. Shammas [Dos95]. Shaped [JR91c]. Shapes [Phi93a]. Shared [Sav96, SX98]. Shareware [Swa95d]. Sharing [Bri97]. Shell [Kle97, TR94, Zol91a]. Shells [Fie89b]. Sherlock [Rea92, Mat89, Rea91b, Rea91a]. Shoot [Mat91]. Shop [Bec98a, Bec99a, Bec99c, Bec99b, Bec99d, Bec99e, Bec99i, Bec99f, Bec99g, Bec99h, Bec99j, Bec99k]. Short [Smi94a]. Should [Cam88, Pug93j]. show [Ret96]. Showcase [Ano97-32]. Shutdown [Car99]. Side [Jae88b, Lib90a]. Sight [Phi98c]. Signal [Bai99, Bak91b, Fre93, Ker96, Rem97]. signal.h [Pla92e]. Signals [Bra91, Wei88c]. Signed [Pug90-42]. Silent [Sch96c]. Silicon [Nar89]. SIM [LM94b, LM94a]. Simple [All92a, And88c, And99b, Bat98, Bav99b, Del88e, Del88f, Dis88, Erd91, Flo98, Gut95, Joh90c, Pla99j, Plu88c, Pug90f, Pug90-35, RP92, Ros94a, Ros92a, Sch96l, TC94, War88i, Wil96a]. Simpler [Pug92p]. Simplex [Ano90a]. Simplifying [GS95, Lad89b]. Simulated [Per92a]. Simulating [Pom98, Sma95]. Simulation [Ake93, HJ90, Mus90, Pol91, Sco93]. Simulations [LM94b]. Simultaneous [Wil96a]. sin [Ber92b]. Single [Coo92a]. Singleton [Gab99, Rin99]. Singletons [Sch99j]. Singly [Jae91c, Jae91d]. Singly-Linked [Jae91c, Jae91d]. Sites [Fie89e, vSS96]. Six [Plu88c]. Size [Bec97e, Car93, Bec98e]. sizeof [Pok92, Pug88y, Jae88b, PT93]. Sizes [PA91, Pug92o]. Skeleton [Pat99]. S’kill [Wei95d]. Skills [Bec98a]. Skip [Gro91, Heg91, Mar97, Whi98]. Slap [EW95]. Sloops [BD90a]. Small [Bra92b, FB96, Fuk89, Gil92, Hen88, KO94, Lan91b, Lan91c, Pug88z, Pug89s, Sch91a, Sch98j, Sch99e, Spr90, Vol89b, dF90a]. smaller [Hel95]. Smalltalk [dF90b]. Smart [Bec98m, Col95b, WK92]. Smooth [Cha88]. SMTP [Faz98, Ste97]. SNews [Fan94]. S’nice [Wei95d]. Sociability [Sch96a]. Socket [Swa94]. Sockets [Gil95b, Gle90c, Ros97, Tom91c]. SoftC [Sch90f, Sch92b]. Software [And89, Ano89, Ano98-27, Bal97, Bec99j, Bri98a, Cad91, DM88, Fie88c, Ham98, Lib90b, McL91, Phi92a, Phi96c, PBW92, Pla94d, Pla94s, Pri88, Sak89, San89, St.93, UHP93, Vol91a, Vol93b, Wal88, War90c, War92b, Wee92, Wei90d, Wil94, Wis97, Zim91, Gam95, Hum95, Kem87, McC93, Sti92b, TK89, Pat95]. Solid [Hal93a, Moo91, Mag93]. Solution [Kna89]. Solutions [Far88, Jae86, Phi99c, TG89]. Solver [Wil96a]. Solving [Nel89b, Wei93j, Bar89, BG94]. Some [Bar90a, BC89c, Cal90, Cel91, Gra91, Mar91, Pip94, PBW92, PLBL92, PF94, Pug88v, Pug88-27, Pug89q, Pug90-36, Pug90-37, Pug90-38, Pug90-39, Pug92p, Pug93f, Pug93s, Pug93t, Pug93v, Pug94b, Pug94k, Pur91, Sch91a, Tri93, War89h, War89i, WC90, War90g, Wio88d]. Somewhat [Pla88c]. Song [Sch99i]. Sort [Cal90, PN93, Pla99d, Pri90d, Ram92, Wil95b]. Sorting [All93l, Erd89, Har88b, Heg93, Nel97, PC93]. Sortlist [Kar94]. Sound [BB94, Wel94a]. Soundex [Ros94a]. Source [Ano88c, Ano89, Ano94b, Cam88, CW89, Fie89c, Pas89, Pla92m, Pri90b, Pug88c, Pug90-28, PH91b, Ret96, Rob94, Spe89b, VG89a, Vos90, Wei89, Wei95b, Wil91, Wil93b, Gra88]. Sources [Ano94b, Dan93, Gra91, LaC94, Pug88o, Vol96e, Vol96d, Vol96c, Vol97a, Vol97d, Vol97c, Vol97b, Vol98a, War89i, Wei92f, Wei92g, Wio88d, dM94, Wil96b]. Space [Edm98, Pug88t, Zer89]. Spaces [PS91b]. Sparse [Pet99]. Spatial [Phi92f]. Spawned [PB92]. Speaking [Kis97]. Special [Ock89, Wei91e, Wei93g, Wei94d]. Specific [Sak97b, Mey92]. Specifications [BD90d]. Specified [Pug93l]. Specifier [HPS90, PP92]. Specifying [Pok92]. Spectral [LaC94]. Speech [Gle90a]. Speed [Hea96, VC89, Yor96]. Speeding [Pug92q]. Speeds [Col89a]. Spell [Art89]. Spiffier [Ser90]. Spin [Bug94]. Spinnaker [Jon88b]. Spirit [Rom91a, Har87, HR92]. SPLASH [Mor93b]. Sponsors [Sch97a]. Spring [Sch98i]. Spuler [Tho93]. SPX [Sch92c]. SQL [Pas89, Pru91, Smi92d, Tik98, Vol90b]. sscanf [PP92]. sstream [Pla95f]. SSX [GC93]. ST [Mat88, Hal89]. Stack [GC93, Pes97b, Pes98, Pug90-30, Pug93j, Tom91a, Whi89b]. Stacks [Jae92h, Jae92a, Jae92b]. Standard [Ale99a, All96a, Bra88, Cho88, Col89c, Fea95a, Fea95b, Pat92b, Pla88e, Pla91a, Pla91c, Pla92m, Pla94n, Pla94q, Pla95m, Pla96j, Pla96k, Pla96l, Pla99i, Plu88b, Pot88, Pug91f, Sak90a, Sak90b, Sch90b, All94j, All95d, All95e, All95f, All98, Bec98c, Bug98, Cla99, Eve93, LK97, Pip94, Pla88a, Pla88b, Pla88c, Pla88d, Pla88f, Pla88g, Pla89d, Pla89g, Pla89h, Pla93f, Pla96h, Pla96g, Pla96i, Pla97s, Pla97q, Pla97o, Pla97k, Pla97m, Pla97l, Pla97n, Pla97r, Pla97p, Pla98k, Pla98h, Pla98f, Pla98g, Pla98e, Pla98j, Pla98i, Pla98d, Pla98c, Pla98a, Pla98b, Pla98l, Pla99i, Pla99e, Pla99j, Pla99k, Pla99h, Pla99a, Pla99c, Pla99b, Pla99l, Pla99d, Pla99f, Pla99g, Pug88y, Sak99d, Str99, Tri93, War90f, Pla94e]. Standardization [Pla92j]. Standardizing [Pla90m]. Standards [Pla93e, Pug88a]. Stanford [Hed90]. State [Cli92, Fis90b, KJV98, Nel92b, She92, War89a, War89d]. Statement [CH88, Jae88f]. Statements [Bel96, Lea96, PT93, PF92b, Wei94g]. Static [Lib90c, Lib90d, Pug92d, Pug95c]. stdarg [Pug88-28]. stdarg.h [Jae89g, Pla91k, Pug90q, Sak90b]. stddef.h [Jae89h, Pla91l, Sak90b]. stdio.h [Pla92h, Pla92k]. stdlib.h [Pla92f, Sak90b]. Steganography [Phi98c]. Step [Sch99e]. Stephen [Pav89, Kou88]. Stereograms [Phi96d]. Steve [Hal93a, Phi94a, Rod95b, Ush94, Web96]. Steven [Mar88a]. Stevens [PS92a, Smi89]. STEVIE [And90]. Still [Col88a, Ric91]. Stitt [War92b]. STL [Ale98a, Bec98m, Ben98, Kas96, Kos99, Pla96m, Pla98a, Pla99g, Vol96e, You96, All99a]. STL-Style [Kos99]. Storage [Bis90, Pla92n, Pug94c]. Stored [Smi92d]. Storing [PY90]. Storm [Wei95c]. Story [All93k]. Straddled [KB99d]. Strategies [MS91]. Strategy [Poi89a]. Streak [KB99c]. Stream [Ale91, Isa98, Kau93, Pin93, War98, Wil92a]. streambuf [Pla94l]. Streams [All93e, Pla89h, Pla93q]. String [Phi94c, Pri91a, Pug88g, PS91b, PA93b, Pug94c, RP91, Ros95, Sak90a, Smi93d, Bec98i, Mor92, Pla95g, Pla95k, Blu91]. string.h [Pla92g]. Stringizing [Cla90a, Pug90-34]. Strings [Jae88e, Kuz90, Lor99, Pla92c, Pla92l, Pug88q, PJ92, Pug92g, Pug94d, Wei94h]. Stroked [Wil91]. Strong [Pug92r]. Stroustrup [Cof88]. strstream [Pla95h, Pla95l]. strtod [Pug89b]. struct [Pug89a]. structs [Bec98e]. Structure [Cel90, Gre94, Pok92, Pug90a, Pug90m, Pug94o]. Structured [Fro89a]. Structures [All94e, Gra93f, Jae89i, Jae91i, Jae91a, Jae91b, Jae91c, Jae91d, Jae91e, Jae91f, Jae91g, Jae91h, Jae92h, Jae92a, Jae92b, Jae92c, Jae92d, Jae92e, Jae92f, Jae92g, Joh92, Pug89d, Pug90d, Pug90h, PC92, Pug93u, Rid91, TP92, LAT96, TLA90]. Studying [Ros92b]. Stuffing [MP90]. Style [Kos99, Mar91, Phi94a, Pug90-37, Pug91a, Sak95b, Sak98b, Tsa90, Whi90a, Car92, Oua92, RN93, Byb92, All93f, Hal93c]. Subdirectory [Pop92]. Subexpressions [Pla88b]. Subroutine [Pol91]. Substrings [All92d]. Subsumption [Man91]. Subtle [Jae88b, Jae89j]. Subtractions [Wil96c]. Such [Sak97c]. Suggestions [Poo93]. Suicidal [Bec95f]. Summary [Ano90b]. Summation [Man96a]. Summer [Wei90e]. SUN [Har88b]. Super [Olh90, Fox85, Jon88a]. super-users [Fox85, Jon88a]. Superior [Jae88c]. Superscalar [Pri90a]. SuperVGA [Ben94]. Support [Fie89a, Han94, Jae89g, Lop94, Pla93m, Pla94a, Plu88a, SR99b]. Supports [Cru88b, Hin88b, Pla89d, WH88]. sure [Wil96b]. Surface [SR99a]. Surfaces [And88c, HP98]. Surprising [Bec95g].

Survey [Rib89, Vol90l]. Survival [Ham98]. Suspension [Pug93f]. Sutherland [Rév93]. Sutherland-Hodgman [Rév93]. Swamp [Sch97d]. Swap [GC93, PA93a]. SWIG [Ble98]. Switch [McM91, Wei94g, HW92b]. SwitchBoard [Cro98a, Cro99]. Symbolic [OH94]. Synapsys [Col90c, Col89d]. Synchronization [Rec98, Wei95e]. Synchronizing [Wei92i]. Syntax [Blo91, Sak95b, Sak97g, Sak97h]. System [All94b, Bal97, Bus91, Car88, Cox92, Fie88d, Fro92, Gra89, Gua90, Kle93, Nar89, Pas88, Phi91a, Phi93f, Pug90-33, Rib89, Sek89, Vol95d, WH88, War88i, Zol90b, HS87, JR91a, Pug90-36, Wei90a]. Systems [All90, Bec93, Bel89b, Car91, Coh92, Del88c, Fie88f, HJ90, Han94, Lab98, Phi98a, Vil91, Wel93, Wis99, Blu92, Gan92, Tan95, Gin93b, Rod95a].

T [Pla94d]. Tab [Ale91]. Table [All94a, Ano91c, Jae90g, Jos93, Pro89, Swa97, Vil96]. Tables [Ben98, Ber92a, DE98, Jae90l, Pla98b]. Tag [Phi91b]. Tags [Wil98, Vol90j]. Tanenbaum [Rod95a]. Target [Met92]. Task [Boy88, Eme88, Rog89]. Tasking [Car88]. Tasks [Yue94]. TAVL [Hug91]. TCP [Kis96]. Teach [Ast94a, Cal93, Cal95]. Teaching [War92a, MQ87]. Technical [Bar90a, Ber90, Pla94r]. technician [Hel94]. Technique [Dis88, Rie98, Ros92a]. Techniques [Bec99k, Gre94, HW92b, Kam94, PG93, Ste89, Tam97, TR94, War92b, Wee92, AA88, BN94, DN93, Mag93, Ste87, Sti92b, Hoh90, Zim91]. Teixeira [Wei93b]. Telephony [Pet98]. Telling [Jae89i]. Template [Ale99b, Bec98f, Bri99e, Edm98, Kam93, Mar97, Mas97, Nga97, Pes97a, Pla95m, Rua97, Sea98, Van96]. Template-Based [Rua97, Van96]. Templates [AT97, Bav99b, Bec97b, Bec97g, Got95, Joh99b, Mar98, Pug93c, Rum99, Sak94h, Sma95, Wil93c, Xia97]. Temporary [Ada94, Bec98l, Sak93h, Sak93i]. Tenenbaum [Rid91]. Tense [Sue93a]. Term [Ekw92]. termcap [Flo90]. Terminal [Cro98b, Flo90, Pas88, Wei88c]. Territory [Bec96d]. Terry [And88b]. Test [Ben94, Plu89b, Pug88-29]. Testing [Bec99j, Car97, DW96a, DW96b, Efk97, McL91, Mea93, Phi96a, Pla94s, Pom98, Wee92, Wil93a, Wil95b, Wil96b, Zim91, Hel94]. Text [All92b, All92c, All92d, BD90a, BD90b, BD90c, Ber90, Cha95, Del88f, Dis88, Gut94, MO92, Pla92o, Pug90f, Vol94d, Web93, Zol90a]. Text-Based [Zol90a]. Texture [Phi95c]. Textview [Gra93a]. Tharp [Bur92a]. Their [Dic94, Poi89a, Pug93e]. Them [Hel90]. Theme [War88j]. Theory [Ano98a, Phi90a, Sak96a, Sak96b, Sak96c, Sak96e, Sak96d, Sak97d, Sak97a, Sak97b, Sak97f, Sak97e, Sak97g, Sak97h, Sak97c, Sak97i, Sak98a, Sak98b, Sak98d, Sak98e, Sak98f, Sak98c, Sak99c, Sak99h, Sak99f, Sak99e, Sak99g, Sak99a, Sak99b, Sak99d, Hen88]. There [Fie88e]. These [Cla88]. Thinking [All99g, Sak99f, Sak99e, Sak99g]. third [BK91, BK94b, Bro94]. third-party [BK91, BK94b, Bro94]. Thomas [Wei93b, Phi93c, Phi93b]. Thomson [Dav93]. Thomson-Davis [Dav93]. those [Ret96]. Thoughts [Pla94s, Pur91]. Thread [BS99, Col90c, Man99, Pla98l, Rec98, Ric99, Rin99]. Thread-Private [Man99]. Thread-safe [Rin99]. Threaded [Bed91]. Threading [Sch90e]. Threads [Col90d, HW92a, Man98, Pla93a]. Three [Ols93, Pug91c, Tam97]. Three-Dimensional [Ols93]. Thunking [Tay97]. TIFF [Vos90]. TIGA [Fri90]. tight [Wil96b]. Time [All92a, All93m, Ban94, Bra91, Car88, Dav92, Fee93, Fra96, FM91, Gil95a, Gin93a, Hal91, HJ90, KB99b, Kra89, Man89, O’N75, O’N78, Pla93o, Pla93p, Pla93g, Pug92i, Pug92q, Pug94n, SR99b, Sue94, VC89, War90e, Yue94, dC90b, Lab92, Nad98].

Time-Consuming [Yue94]. Time-To-Transfer [Fra96]. time.h [Pla93j]. time get [Pla98i]. time put [Pla98j]. Timer [Wen97]. Timers [Lib90b]. Timing [Lan89]. Tips [BC89c, Cel91, Hoh90, Pla95a, Pug90-40, AA88, Jae86, Por93]. TLC [Sch91e]. TLC/TLP [Sch91e]. TLIB [Pug91j]. TLP [Sch91e]. Today [AT97, Nie94]. Tom [Hal93c]. Tongues [Wei90c]. Too [Pla91e, Sch98a]. Tool [Ano97-32, BS94, Car97, Gut94, Hav89a, Hin88b, Lad89b, Lib90e, Phi90b]. Toolbox [Bin88, Bla90, Cas93, KO94, Hun85, Hun89, Hin90a]. Toolbox/PC [Hin90a]. Toolkit [Bar90c, Col89a, JR91d, Str88b, Vol91a, Pur89b, Pur92, Ano97-50, Ogg90b]. Toolkits [Gui97]. Tools [Ano89, Ano98-27, Dic94, Gra93b, How93, Jar91a, McC89, Mue94b, Pri88, Ram88, RSSS90, VM88, Vol90a, Vol95g, War92b, Wel94a, Whi90b, WP88, Zol90e, dFAAW90, Kem87, Pet88, Sti92b, Fla90]. Topological [Kar90]. Torque [McC94]. Touches [Gra92d]. Touching [Pla88f]. Touchscreen [Gra92d]. toupper [Pug88g]. Tour [Cla97, Gib88]. Townsend [Ogg90c]. tr [Bar90d]. Trace [Cha97, Jae89c, Pes97b, Pes98]. Tracing [Smi92g]. Tracking [Coo90, Met92, Mon97]. Trailing [Ale91]. Trainer [Rab89, Feu86]. Training [War89i]. Traits [Bec96j]. Transaction [Coo90]. Transfer [Blo91, Fra96, WH88]. Transferring [Kru95]. Transfers [Yor96]. Transform [Ash89, Clu90, Swa88]. Transformations [Dea93]. Transforming [CH88]. Transforms [O’D90]. Transition [Cli92]. Translating [Pug89r, Tee90]. Translation [Del88a, Gut94, Pla89a, Pla89f, Pri90c, Pug88f, Sch89b, Pla92d]. translator [Pla97i]. Transliterating [Pla96a]. Transporting [Pug89s]. Trap [Cur90, Lib91c]. Trapping [Pug90-41]. Traveling [LFK89, Kna89]. Treasures [Nel89a, Hol87]. Treatment [Pug88-29]. Tree [BT91, Hug91, Kel92, Pla99c, Rum98, Sil88, Swa88, Ano97t, Fre88a]. Trees [Jae92e, Jae92f, Jae92g, Jar91b, Sch93b, Ter91, Wal91]. Trial [Cho88]. Trial-Use [Cho88]. Trick [Boc96]. Tries [Ste92a]. Trig [Pri89b]. Trimming [Sak99h]. trouble [Hel94]. trouble-free [Hel94]. TrueType [BG99]. Truncating [DP93]. Truth [Jae89i]. TSQRT [Boc96]. TSR [McC89, Pug88o]. TSRs [Vol90h]. tty [Ber88]. Tuning [Pri92c, Pri93]. Turbo [Pat92a, PS92a, Pav90, Smi89, Spa90, Vol90a, Abo91, All89, Bai91, Bro89, Cas89, Chi89, Chu91, Der88, Dod88b, Hun89, Kel92, Kwo89b, Lan90, McC89, Pug89i, Pug90d, Pug90-31, Pug91j, Ser90, Ste87, Ste90, Ste92b, Vil91]. Tutor [Dod88a, Dod88b, Dod91, Saw94]. Tutorial [Bak91b, Zol90b, Zol90a, Zol90c, Zol90d]. Tutorials [Web96]. Twice [Lib90f, Pug90b]. Two [Ber92a, Cor95, LaB93a, MP90, Ott94, Pug88-29, Sch95b, Sch96i, Shi90, War89j, Web96, You96]. Two-Byte [Pug88-29]. Two-Dimensional [Ber92a, LaB93a]. Type [Ano99m, Ban94, BD91, Bec97b, Jae89b, Joh99b, Kel91, Pug90e, Pug90-42, Pug92r, Pug94n, Sch96b, Sch96a, Smi94a]. Type-Checking [Pug92r]. typedef [Jae88c, Pug91i]. typedefs [Pug90j, Pug92r]. Types [All99h, Bec99a, Cam98, Pla88c, Pla88d, Pla88g, Pug89o, Pug90-29, Pug92h, Pug92n, Pug94l, Sai92, Sak91c, Sak94k, Sch99k, Smi93g, Ten98]. Typical [Wei88b]. Typing [Saw94].

u [PF92b]. U.S. [Wat88, Wel91]. UART [Rot91]. ubiquitous [Ret96]. UI2 [Com90]. UltraWin [Hac91]. Uncaught [Sch98l, Sch99a, Sch99b, Sch99c, Sch99d, Sch99e, Sch99g, Sch99h, Sch99f, Sch99i, Sch99j]. Uncertain [Wil94]. Uncertainty [Man96b]. Undeclared [Bec98j]. Undefined [Bec96d]. Understanding [Bec98k, Fea95a, Fea95b, Ogg90c, Sak95c, Sak96j, Tow88]. Underway [Pla94q]. Undocumented [Bur89]. Unfreed [Bec97d]. Unicode [Pla99k]. Unicorn [Mil88, Mil89]. Uninterrupted [Bro94, Swa95b]. Unions [Pug89j]. Unit [Mea93]. Universal [Bel89b, Ben94, Kel91]. UNIX [Fla90, Gra89, Jon88a, Nel90b, Pug88p, Wei93b, Bai99, Cro98b, ET88, ET92, Fie88f, Flo90, Fox85, Fre90, Gra88, HS87, KK89, Paj88, Pas88, Pet88, PLBL92, Pug88a, Pug88-30, Pug90s, Pug90-33, Pug90t, Pug92o, Rem97, Rib89, Rob95, Sch91c, Sma88, TR94, Ver89, Vol94b, Vol97d, War90g, Wei88c, Wei94e, Whi89c, WL94, Xia97, dFAAW90, Nad98, War88i]. UNIX/MS [Ver89]. UNIX/MS-DOS [Ver89]. unseen [Pic90]. Unsigned [Pug90-42]. Update [Law92a, Pla97n, Pla99e, Vol95g, Vol97e, Vol98b, Wei93g, Wei94d, Vol96b]. Updating [Lib89c]. Upgrade [Cas89]. Uppercase [Pug88g]. Uproar [Pla96b]. Uren [Pla94d]. urve [Sch96f, Sch96g, Sch96h, Sch96i, Sch97b, Sch97c, Sch97a, Sch97d, Sch97e, Sch97g, Sch97i, Sch97f, Sch97h, Sch98a, Sch98e, Sch98g, Sch98j, Sch98i, Sch98d, Sch98b, Sch98k, Sch98c, Sch98f, Sch98h]. Usage [MS91, Hen88]. Use [Ano96b, Bec95a, Bec98l, Cho88, HW92a, Hel90, PH93, Pug88-27, Pug90q, Pug93j, Pug94r, Pug94q, RP91, Ret96, War88h, War89g]. Useful [Del88d, Plu88c]. Usefulness [Rup88]. USENET [Fie88f, Fie89e, Wei93g, Wei94d, Fie88g]. User [Bru89a, Cla90c, Col89a, Goo89, GK90, Mey95, Pug90-29, Smi88b, Smi92f, Vol89c, Vol96b, WPM84, WC90, Wei93i]. User-Defined [Pug90-29, Mey95]. Users [Ano88a, Ano94b, Ban89, Clu89b, Cog89, Fie88f, Mul93, Pla91d, Pla91e, Pla93k, Pla94c, Pla94t, Plu88a, Vol98b, Vol99, WB88, War90d, Wio88b, Fox85, Jon88a]. Uses [Jae88c, Pug93e]. Using [All99h, Bai91, BDB89, BD90c, Bar88, Bec97c, Bec97g, Bed95, BT93, Cor91, Cro98a, Del88e, Fro90, Ger90, Ges91, Got94, HW92b, Hei99, Hin88a, Jae90k, Joh90d, Jos93, Kno90, Kod92, LK97, Lan99a, Lib90f, Lop97, Man90, McC89, McC95, NW99, Pes97a, Pes97b, Phi90a, Phi91d, Phi92b, Phi93d, PF93a, PT93, PB94, Pug88g, Pug89e, Pug89i, PM89, Pug89t, Pug90-40, Pug91j, Pug92q, Pug92s, Pug92u, PP92, Pug92t, Pug93s, Pug94p, Rat90, Rid91, Row95, Rum99, Sch93a, SX98, Sin94, Sin93b, Smi92h, Ste90, Ste92a, Sti92a, Str88a, Sut99b, Tac94, TR94, Van93, Van95, Web93, Wei88d, Wei94e, Wei92i, Wei93i, Wei93j, Wei95e, Whe97, Xia97, Yok89, Zol90c, LAT96, Mar95, TLA90, Wit95]. Uspell [McC92]. Utilities [PS92a, Cor95, Der88, Fie89d, Fie89e, Hin88a, MO92, Sil88, Smi88a, Smi89, Swa95d, Ste87]. Utility [Bel89b, Cam90, Del88b, Del88e, Vol90g, Zmo96, Pla96e]. uucp [Fie88e, Tay93].

V [Fie88d, Wam96]. v1.00B [Vol91a]. v1.04 [Pug90-31]. v1.5 [Kir92]. v1.8 [Arc94]. V2.0 [Lan91c, Vol90a, GC92, Law92b, Pav90]. v2.12 [Mue94b]. v2.15C [Bus91]. v3.52 [Bar90c]. v3.9 [Ano88c, Man88]. v4.0 [Mil88, Sme89]. v5.1 [Mil89]. v5.10 [Sch90d]. v5.3 [Rea90b]. v7.0 [Rea90a]. va arg [Pug90q]. Vacation [Wei90e]. valarray [Mor99, Pla97r]. Valarrays [Ano99c]. Value [Bec96i, Pla88d, Pug90e, Pug94e]. Values [Kru95, Pla91f, Pug92j, Pug93i, Pug93m, Pug93s, Pug88k]. Variable [Ale98b, All94k, Bec97e, Bec98e, Jae89g, Kod92, PS91a, PJ92, Sta89, vdP91]. Variable-Length [All94k, Kod92]. Variable-Size [Bec97e, Bec98e]. Variables [Bri97, Bru91, Joh92, Lib89c, Pug88x, Pug90z, Pug92d, Wit90b]. Variant [Kas96]. Variations [Sch97h, War88j]. Varieties [Jae89j]. Varying [Smi93f]. VAX [Gil90, Rib89]. VAX/VMS [Gil90]. VBE [Ben94]. VC [Whe97]. Vector [Get99, Per99, Ale99a, Pla97s]. Vendor [Qua88]. Vendors [Bar90a, Wat88]. Ventura [War88h]. Ver [Fra88, Sil88]. Versatile [Smi92a, Ste92b, Wil96c]. Version [Bro89, Bru92b, Pla92m, Pug93v]. Versions [Obe91]. Versus [Abo91, PLBL92, PC91, Pug92t]. Very [Sch98j, TP92]. VESA [Ben94, Ben94, Kre96]. VGA [Gag93, Olh90, Zia90]. VGL [Mor93a, Vol93e]. VI [Lop94, Sma88]. via [Bat97b, Phi98b, Ten98, Wei93f]. Vibe [Ano97-31]. Video [Joh90a, Vol95b]. View [Vol89c]. Viewer [Gag93]. VIM [Moo94]. Virtual [Bri90, Pug89u, Sak93j, Sak94i, Sak94k, Tha98]. Visibility [All94l, All94m]. VisiBroker [And99a]. Visit [Bri99e, Tre94, Vol96g]. Visitor [Gou98]. Visix [Ano97-31]. Visual [Sic94]. Visualizing [Phi99d, Sch96k]. Vitamin [Fre88b]. VLAs [Sch98b, Sch98k]. Vlissides [Pat95]. Vlist [Blu91]. VMS [Byb92, Gil90]. VMS-Style [Byb92]. Voice [Coh92]. Void [Jae90l, PB94, PA91]. Volume [Pug92v, Gua90]. Volunteers [Pla93e]. Vs [Jae91a, Jae91b, Pug90c, Pug90j, Bec96a, Irv91, Pug88e, Sak92b, Sak98c].

w [Jar91a]. W. [Per89]. w/ [Jar91a]. Waite [Pav89]. Walk [Cra97]. Walks [Bec96i]. Walsh [Ash89]. wanta [Gra88]. Ward [And88b]. Warping [Phi95b]. Was [Wei90e]. Wasn’t [Sak97c]. Watts [Phi96c]. Waveforms [Jan93b]. Way [Rin99, Wei95c]. ways [Mey92]. weak [Wil96b]. Weaknesses [Pip94]. Web [NW99, Ret96, Ros97, Ste97, Vol96a, Vol96e, Vol96c, Vol97a, vSS96]. Web-Server [Ste97]. Weight [Kam94]. Weiskamp [Hob91a]. Welcome [Pla94t, WW88, Wei91f]. Well [Wei95a]. We’re [Ben97]. WGCONIO [Gie90]. Wha [Pla90p, Pla90q]. Where [Fie88a, Wei92f, Wei92g, Wei94f, Wei95b].

Which [And94, Pug93j]. Whirlwind [Cla97]. Wide [Pla93q]. Wildcard [Cor95, Pop92]. William [Pla90c, Bin88, Bla90, Cas93, Nel95, Pla93g]. Williams [Fra88]. WIMP [MA88]. WIN.INI [Tac94]. Win16 [Car97]. Win32 [Bav98b, Bec99l, Bri97, Dug97, Har99, Hea96, KS98, Lar97, Mit99, Pes98, Woo97]. WinCGI [Whe97]. WindosIO [GC92]. Window [Ano88d, Gua90, JR91a, Pug88-31, Pug90-36, Wei90a, Ebn91, JR91a, JR91b, JR91c, JR91d, JR91e, RJ91]. Windows [Ast94a, Cus93, Gra92a, Hal93d, Smi93g, Swa95c, Vol95g, Bar95b, Bar91a, Bav99c, BT93, Bug94, Cal93, Cal95, CD94, Dan97, Deg97, Dow96, Fee93, Fie89a, Gil95a, Gil95b, Guf98, Har94, Hat96, Hea97, Hey96, Hoo99, JR91c, Jos93, Kau93, Kle93, Kle97, Knu98, Kru92, LP92, Nor92, Ols93, Pin93, Ser90, Sin93a, Sin94, Tam97, Tuc96, Vol96c, Wel95, Wel96, Yue94, Zmo96]. WinJES [Gil95b]. Within [Bur89, PG91, Pug92b]. Without [BSG92, Got95, Pug89e, PK92, Rog95, Bri98e]. WLIST [Hat94]. Wonderful [Wei92e]. Wong [Nel95]. Won’t [Bec96e]. Word [Cur98, Sch91b, Sch97a]. Word97 [Lan99a]. Work [Pug92h, Sak94i, Sak97i]. Work-arounds [Sak97i]. Works [Dis88]. Workshop [Wio88c]. World [Mal90, Sut99a, Pic90]. Worlds [Sch95b]. Wrapper [Bav98b, Bel98, Pug93w, Roe99]. Wrappers [All99h, Sea98, Smi92b, Van93, War91b]. Wrapping [Bav99c, Bec97h, Bec99k]. Write [Erd88, Poi89a, Whi91, PH93]. Writer [Ban89]. Writers [Hin88a]. Writing [All89, Bra92b, Bri98g, ET88, ET92, Fri90, Hal93a, Joh90b, Mag93, Man89, Nel90a, Paj88, Phi91f, Pug89m, PK92, Sak90a, Sak90b, Sak91e, San89, War89k, Wei93b]. WTWG [Blu90, Blu94, Vol94e].

X [Gua90, Ano97s, BBAH88, Col88a, Del88b, Fie89a, JR91a, JR91b, JR91c, JR91d, JR91e, Pug90-36, Pug93l, Rat91, RJ91, Vol96c, Wei90a, Wei92h]. X-Assemblers [Col88a]. X-Designer [Ano97s]. X-Ref [Del88b]. X-Windows [Fie89a]. X.25 [Wri90b]. X/Open [Pug93l]. X3J11 [Plu88d]. Xenix [Fie88d, Fie88f, Law92c, WH89, Con88, Joh88a]. Xerox [War88h]. Xlib [RJ91]. XML [And99b]. Xmodem [Flo89, War90a]. XMODEM/CRC [War90a]. XOR [Pug89i]. xref [Del88d]. XYZ [Vol95e].

Y2K [Gib99, Iof99]. Yacc [Del88e, Yok88, Pug91k, Vol89a, Vol89c, Lib90f]. Yachts [BD90a]. Year [Efk97]. Yo [Sch98k]. Yourself [Ast94a, Cal93, Cal95].

Z [Joh88b, Mor92]. Z80 [Col88c, Pan91]. Zinc [Bru92b]. Zmodem [Ber88]. Zolman [Wal92]. ZOO [Gra93e]. Zortech [BC89d].

References

Anderson:1988:ACT

[AA88] Paul Anderson and Gail Anderson. Advanced C: tips and techniques. Hayden Books C library. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1988. ISBN 0-672-48417-X (paperback). xi + 446 pp. LCCN QA76.73.C15 A48 1988. US$24.95.

Aboshanab:1991:CQC

[Abo91]  Belinda Aboshanab.  Costs:

Quick C versus Turbo C. C Users Journal, 9(9):140–??, September 1991. ISSN 0898-9788.

Adams:1994:TOM

[Ada94] Robert M. Adams. Temporary object management through dual classes. C Users Journal, 12(5): 27–??, May 1994. ISSN 0898-9788.

Akerbaek:1993:CCS

[Ake93] Trond Akerbaek. C++, coroutines, and simulation. C Users Journal, 11(3):74–??, March 1993. ISSN 0898-9788.

Alessi:1991:TBT

[Ale91] Joseph S. Alessi. Trailing blank, trailing tab, and blank line removal from input stream. C Users Journal, 9(6):123–??, June 1991. ISSN 0898-9788.

Alexandrescu:1998:CIS

[Ale98a] Andrei Alexandrescu. Compound iterators for STL. C/C++ Users Journal, 16(10):79–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Alexandrescu:1998:ICV

[Ale98b] Andrei Alexandrescu. Inline containers for variable arguments. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCU-JEX. ISSN 1075-2838.

Alexandrescu:1999:AAA

[Ale99a] Andrei Alexandrescu. Adapting automation arrays to the standard vector interface. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Alexandrescu:1999:BTE

[Ale99b] Andrei Alexandrescu. Better template error messages. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Alger:1995:SCM

[Alg95]  Jeff Alger.  Secrets of the

C++ masters. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-049940-1. xii + 388 pp. LCCN QA76.73.C153 A44 1995 Reserves. See book review [Kil96].

Allen:1988:CHM

[All88] Robert Allen. Control-C handler for MS-DOS. C Users Journal, 6 (2):12–??, February 1988. ISSN 0898-9788.

Allen:1989:WDD

[All89] Robert Allen. Writing device drivers with Turbo C. C Users Journal, 7(2):59–??, February 1989. ISSN 0898-9788.

Allen:1990:FIS

[All90] Marc L. Allen. Forked interrupt systems. C Users Journal, 8(4): 113–??, April 1990. ISSN 0898-9788.

Allison:1992:SRE

[All92a] Charles B. Allison. A simple real-time executive. C Users Journal, 10(11):33–??, November 1992. ISSN 0898-9788.

Allison:1992:TPF

[All92b] Chuck Allison. Text processing I — the finer points of scanf. C Users Journal, 10(10):23–??, October 1992. ISSN 0898-9788.

Allison:1992:TPIa

[All92c] Chuck Allison. Text processing II: printf. C Users Journal, 10(11):117–??, November 1992. ISSN 0898-9788.

Allison:1992:TPIb

[All92d] Chuck Allison. Text processing III: Substrings. C Users Journal, 10(12):73–??, December 1992. ISSN 0898-9788.

Allison:1993:BHCa

[All93a] Chuck Allison. Bit handling in C. C Users Journal, 11(11):89– ??, November 1993. ISSN 0898-9788.

Allison:1993:BHCb

[All93b] Chuck Allison. Bit handling in C++, part 1. C Users Journal, 11 (12):71–??, December 1993. ISSN 0898-9788.

Allison:1993:CDCa

[All93c] Chuck Allison. A C++ date class, part 1. C Users Journal, 11 (2):123–??, February 1993. ISSN 0898-9788.

Allison:1993:CDCb

[All93d] Chuck Allison. A C++ date class, part 2. C Users Journal, 11(3): 93–??, March 1993. ISSN 0898-9788.

Allison:1993:CS

[All93e] Chuck Allison. C++ streams. C Users Journal, 11(7):101–??, July 1993. ISSN 0898-9788.

Allison:1993:TEC

[All93f] Chuck Allison. The Elements of C Programming Style, by Jay Ranade and Alan Nash. C Users Journal, 11(7):129–??, July 1993. ISSN 0898-9788. See [RN93].

Allison:1993:FP

[All93g] Chuck Allison. File processing. C Users Journal, 11(5):97–??, May 1993. ISSN 0898-9788.

Allison:1993:FPP

[All93h] Chuck Allison. File processing, part 2. C Users Journal, 11(6): 125–??, June 1993. ISSN 0898-9788.

Allison:1993:PPB

[All93i] Chuck Allison. Pointers, part 1:

The basics. C Users Journal, 11 (8):107–??, August 1993. ISSN 0898-9788.

Allison:1993:PPP

[All93j] Chuck Allison. Pointers, part 2:

Pointers and arrays. C Users Journal, 11(9):109–??, September 1993. ISSN 0898-9788.

Allison:1993:PPR

[All93k] Chuck Allison. Pointers, part 3:

The rest of the story. C Users Journal, 11(10):83–??, October 1993. ISSN 0898-9788.

Allison:1993:STQ

[All93l] Chuck Allison. Sorting with qsort. C Users Journal, 11(4): 107–??, April 1993. ISSN 0898-9788.

Allison:1993:TDP

[All93m] Chuck Allison. Time and date processing in C. C Users Journal, 11(1):115–??, January 1993. ISSN 0898-9788.

Allen:1994:CPT

[All94a] Gary A. Allen, Jr. CUG410A — partition table and hard disk analysis. C Users Journal, 12(6): 100–??, June 1994. ISSN 0898-9788.

Allison:1994:TRE

[All94b] Charles B. Allison. ROMLDR, an embedded system program locator. C Users Journal, 12(3):35– ??, March 1994. ISSN 0898-9788.

Allison:1994:BHC

[All94c] Chuck Allison. Bit handling in C++, part 2. C Users Journal, 12(1):91–??, January 1994. ISSN 0898-9788.

Allison:1994:CE

[All94d] Chuck Allison. C++ exceptions. C/C++ Users Journal, 12(7):69– ??, July 1994. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1994:CS

[All94e] Chuck Allison. Control structures. C Users Journal, 12(6):81– ??, June 1994. ISSN 0898-9788.

Allison:1994:CC

[All94f] Chuck Allison. Conversions and casts. C/C++ Users Journal, 12 (9):67–??, September 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Allison:1994:DMMa

[All94g] Chuck Allison. Dynamic memory management, part 1. C/C++ Users Journal, 12(10):101–??, October 1994. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1994:DMMb

[All94h] Chuck Allison. Dynamic memory management, part 2. C/C++ Users Journal, 12(11):71–??, November 1994. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1994:P

[All94i] Chuck Allison. The preprocessor. C Users Journal, 12(3):101– ??, March 1994. ISSN 0898-9788.

Allison:1994:SCL

[All94j] Chuck Allison. The Standard C++ library. C/C++ Users Journal, 12(12):69–??, December 1994. CODEN CCUJEX. ISSN 1075-2838.

Allison:1994:VAL

[All94k] Chuck Allison. Variable-length argument lists. C Users Journal, 12(2):103–??, February 1994. ISSN 0898-9788.

Allison:1994:VCa

[All94l] Chuck Allison. Visibility in C. C Users Journal, 12(4):91–??, April 1994. ISSN 0898-9788.

Allison:1994:VCb

[All94m] Chuck Allison. Visibility in C++.

C Users Journal, 12(5):81–??, May 1994. ISSN 0898-9788.

Allison:1995:BC

[All95a] Chuck Allison. A better C.

C/C++ Users Journal, 13(4):67– ??, April 1995. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1995:DA

[All95b] Chuck Allison. Data abstraction.

C/C++ Users Journal, 13(5):75– ??, May 1995. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1995:TCP

[All95c] Chuck Allison. C++ Pointers and Dynamic Memory Management by Michael Daconta. C/C++ Users Journal, 13(11): 79–??, November 1995. CODEN CCUJEX. ISSN 1075-2838. See [Dac95].

Allison:1995:SCLa

[All95d] Chuck Allison. The Standard C library, part 1. C/C++ Users Journal, 13(1):69–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Allison:1995:SCLb

[All95e] Chuck Allison. The Standard C library, part 2. C/C++ Users Journal, 13(2):89–??, February 1995. CODEN CCUJEX. ISSN 1075-2838.

Allison:1995:SCLc

[All95f] Chuck Allison. The Standard C library, part 3. C/C++ Users Journal, 13(3):59–??, March 1995. CODEN CCUJEX. ISSN 1075-2838.

Allison:1996:CMS

[All96a] Chuck Allison. C++: The making of a standard. C/C++ Users Journal, 14(10):39–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Allison:1996:OPR

[All96b] Chuck Allison. Object persistence with relational databases. C/C++ Users Journal, 14(5):37– ??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1998:WNS

[All98] Chuck Allison. What’s new in Standard C++? C/C++ Users Journal, 16(12):??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:BRG

[All99a] Chuck Allison. Book review:

Generic Programming and the STL. C/C++ Users Journal, 17 (11):??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:BRS

[All99b] Chuck Allison. Book review:

Scott Meyers’ Effective C++ Compact Disk. C/C++ Users Journal, 17(5):??, May 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJC

[All99c] Chuck Allison. import java.*:

Control flow — the bad, the good, the exceptional. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJJ

[All99d] Chuck Allison. import java.*:

Jumping into java. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJO

[All99e] Chuck Allison. import java.*:

Object-oriented programming in Java. C/C++ Users Journal, 17 (11):??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJP

[All99f] Chuck Allison. import java.*: Packaging your objects. C/C++ Users Journal, 17(9):42–??, September 1999. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1999:IJT

[All99g] Chuck Allison. import java.*: Thinking in objects. C/C++ Users Journal, 17(7):79–??, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJU

[All99h] Chuck Allison. import java.*: Using primitive types and wrappers. C/C++ Users Journal, 17 (3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Ammeraal:1986:PPC

[Amm86] L. (Leendert) Ammeraal. Programming principles in computer graphics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1986. ISBN 0-471-90989-0 (paperback). viii + 168 pp. LCCN QA76.6.A466 1986. US$17.88. See book review [Kam89].

Ammeraal:1987:CGI

[Amm87a] L. (Leendert) Ammeraal. Computer graphics for the IBM PC. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1987. ISBN 0-471-91501-7 (paperback). viii + 141 pp. LCCN QA76.8.I2594 A48 1987 Bar. US$21.00. See book review [Sch88a].

Ammeraal:1987:PPC

[Amm87b] L. (Leendert) Ammeraal. Programming principles in computer graphics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, reprinted with corrections edition, 1987. ISBN 0-471-90989-0 (paperback). viii + 168 pp. LCCN QA 76.6 A466 1987. See book review [Kam89].

Ammeraal:1992:PPC

[Amm92] L. (Leendert) Ammeraal. Programming principles in computer graphics. Wiley professional computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 1992. ISBN 0-471-93128-4. vii + 233 pp. LCCN QA76.6 .A466 1992.

Andersen:1988:AIC

[And88a] Larry Andersen. Apple II C compiler. C Users Journal, 6(2):13– ??, February 1988. ISSN 0898-9788.

Anderson:1988:TPC

[And88b] Jay Martin Anderson. Programming C on the Macintosh, by Terry A. Ward. C Users Journal, 6(6):46–??, June 1988. ISSN 0898-9788. See [War86].

Anderson:1988:SMH

[And88c] Jay Martin Anderson. A simple model for hiding surfaces. C Users Journal, 6(8):51–??, August 1988. ISSN 0898-9788.

Anderson:1989:SCI

[And89] Jay Martin Anderson. Software control of IEEE-488 devices. C Users Journal, 7(11):27– ??, November 1989. ISSN 0898-9788.

Andrews:1990:CS

[And90] Tony Andrews. CUG313 — STE-VIE. C Users Journal, 8(5):123– ??, May 1990. ISSN 0898-9788.

Andresen:1994:RAW

[And94] Eric Andresen. A rounding algorithm which needs improvement. C Users Journal, 12(4):129–??, April 1994. ISSN 0898-9788.

Anderson:1999:CLB

[And99a] Marc Anderson. CORBA load balancing with VisiBroker. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Andrivet:1999:SXP

[And99b] Sebastien Andrivet. A simple XML parser. C/C++ Users Journal, 17(7):22, 24, 26–28, 30, 32, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1988:CUG

[Ano88a] Anonymous. C Users Group announces Japanese affiliate. C Users Journal, 6(6):81–??, June 1988. ISSN 0898-9788.

Anonymous:1988:CCL

[Ano88b] Anonymous. CUG172 and CUG173 — lex. C Users Journal, 6(2):57–??, February 1988. ISSN 0898-9788.

Anonymous:1988:CMV

[Ano88c] Anonymous. CUG198 — MicroEmacs v3.9 source. C Users Journal, 6(9):92–??, September 1988. ISSN 0898-9788.

Anonymous:1988:CWB

[Ano88d] Anonymous. CUG220 — window BOSS for MS C 5.2 and quick C. C Users Journal, 6(2):57–??, February 1988. ISSN 0898-9788.

Anonymous:1989:CST

[Ano89] Anonymous. CUG280 — software tools source. C Users Journal, 7(2):103–??, February 1989. ISSN 0898-9788.

Anonymous:1990:CSC

[Ano90a] Anonymous. CUG209 — simplex curve fitting. C Users Journal, 8(5):123–??, May 1990. ISSN 0898-9788.

Anonymous:1990:SCC

[Ano90b] Anonymous. Summary of CUG cross assemblers. C Users Journal, 8(6):128–??, June 1990. ISSN 0898-9788.

Anonymous:1991:BCP

[Ano91a] Anonymous. Bad C puns contest results. C Users Journal, 9 (5):26–??, May 1991. ISSN 0898-9788.

Anonymous:1991:CCG

[Ano91b] Anonymous. CUG344 — C grabbag #1. C Users Journal, 9(6): 132–??, June 1991. ISSN 0898-

Anonymous:1991:TCA

[Ano91c] Anonymous. Table of cross assemblers in the CUG library. C Users Journal, 9(7):126–??, July 1991. ISSN 0898-9788.

Anonymous:1993:DYM

[Ano93] Anonymous. Decoding your mailing label. C Users Journal, 11(8): 112–??, August 1993. ISSN 0898-9788.

Anonymous:1994:OCD

[Ano94a] Anonymous. October 1993 code disk errors. C Users Journal, 12 (1):116–??, January 1994. ISSN 0898-9788.

Anonymous:1994:SCU

[Ano94b] Anonymous. Sources of C Users Journal source code. C Users Journal, 12(6):5–??, June 1994. ISSN 0898-9788.

Anonymous:1996:AAJ

[Ano96a] Anonymous. Amzi! announces Java class for Amzi! Logic Server. C/C++ Users Journal, 14(10): 94, October 1996. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1996:ORJ

[Ano96b] Anonymous. ObjectSpace releases Java generic library for free commercial use. C/C++ Users Journal, 14(10):96, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1996:RAR

[Ano96c] Anonymous. Rational announced Rational Rose for Java. C/C++ Users Journal, 14(9):94, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:AIa

[Ano97a] Anonymous. Advertiser index.

C/C++ Users Journal, 15(1):56– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:AIb

[Ano97b] Anonymous. Advertiser index.

C/C++ Users Journal, 15(2):64– ??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:AIc

[Ano97c] Anonymous. Advertiser index.

C/C++ Users Journal, 15(3):96– ??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:AId

[Ano97d] Anonymous. Advertiser index.

C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:AIe

[Ano97e] Anonymous. Advertiser index.

C/C++ Users Journal, 15(5):72– ??, May 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:AIf

[Ano97f] Anonymous. Advertiser index.

C/C++ Users Journal, 15(6):73– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:AIg

[Ano97g] Anonymous. Advertiser index.

C/C++ Users Journal, 15(7):72– ??, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:AIh

[Ano97h] Anonymous. Advertiser index. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:AIj

[Ano97i] Anonymous. Advertiser index. C/C++ Users Journal, 15(9):64– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:CPa

[Ano97j] Anonymous. Call for papers. C/C++ Users Journal, 15(1):77– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:CPb

[Ano97k] Anonymous. Call for papers. C/C++ Users Journal, 15(2):78– ??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:CPc

[Ano97l] Anonymous. Call for papers. C/C++ Users Journal, 15(3):80– ??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:CPd

[Ano97m] Anonymous. Call for papers. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:CPe

[Ano97n] Anonymous. Call for papers. C/C++ Users Journal, 15(5):78– ??, May 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:CPf

[Ano97o] Anonymous. Call for papers.

C/C++ Users Journal, 15(6):86– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:CPg

[Ano97p] Anonymous. Call for papers.

C/C++ Users Journal, 15(7):84– ??, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:CPh

[Ano97q] Anonymous. Call for papers.

C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:CPi

[Ano97r] Anonymous. Call for papers.

C/C++ Users Journal, 15(9):88– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:IAX

[Ano97s] Anonymous. Imperial announces X-Designer: Java Edition. C/C++ Users Journal, 15 (2):88, February 1997. CODEN CCUJEX. ISSN 1075-2838. X-Designer is a GUI builder for Motif with Java, X/Motif, and Windows code generators.

Anonymous:1997:LTO

[Ano97t] Anonymous. Learning Tree offers Java courses. C/C++ Users Journal, 15(1):78, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:MIM

[Ano97u] Anonymous. The MathWorks introduces MATLAB C++ math library. C/C++ Users Journal, 15 (9):97, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:NPa

[Ano97v] Anonymous. New products. C/C++ Users Journal, 15(1):78– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:NPb

[Ano97w] Anonymous. New products. C/C++ Users Journal, 15(2):85– ??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:NPc

[Ano97x] Anonymous. New products. C/C++ Users Journal, 15(3): 102–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:NPd

[Ano97y] Anonymous. New products. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:NPe

[Ano97z] Anonymous. New products. C/C++ Users Journal, 15(5):96– ??, May 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:NPf

[Ano97-27] Anonymous. New products. C/C++ Users Journal, 15(6):94– 97, June 1997. CODEN CCU-JEX. ISSN 1075-2838. DigitalThink Offers Online C++ and Java Courses (http://www.

digitalthink.com/). Software

Research Announces TCAT (coverage analyser) for Java (http:/

/www.soft.com/).

Anonymous:1997:NPg

[Ano97-28] Anonymous. New products.

C/C++ Users Journal, 15(7):94– 97, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:NPh

[Ano97-29] Anonymous. New products.

C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:NPi

[Ano97-30] Anonymous. New products.

C/C++ Users Journal, 15(9):95– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:NPV

[Ano97-31] Anonymous. New products:

Visix releases Vibe Java IDE. C/C++ Users Journal, 15(7):94, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:AIi

[Ano97-32] Anonymous. Object-oriented programming tool showcase. C/C++ Users Journal, 15(9):51– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:OAF

[Ano97-33] Anonymous. ObjectSpace announces five new object-oriented classes for C++, Java, and MFC. C/C++ Users Journal, 15(9):98, September 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:PMa

[Ano97-34] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (1):83–??, January 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMb

[Ano97-35] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (2):90–??, February 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMc

[Ano97-36] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (3):107–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMd

[Ano97-37] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMe

[Ano97-38] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (5):101–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMf

[Ano97-39] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (6):99–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMg

[Ano97-40] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (7):100–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMh

[Ano97-41] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:PMi

[Ano97-42] Anonymous. Programmer’s market. C/C++ Users Journal, 15 (9):100–??, September 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:RAB

[Ano97-43] Anonymous. R-Active’s Better-State PRO generates Java code. C/C++ Users Journal, 15(1):79, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:RBe

[Ano97-44] Anonymous. R&D books.

C/C++ Users Journal, 15(5):62– ??, May 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:RBa

[Ano97-45] Anonymous. R&D bookstore.

C/C++ Users Journal, 15(1): 17a–??, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:RBb

[Ano97-46] Anonymous. R&D bookstore.

C/C++ Users Journal, 15(2):??, February 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:RBc

[Ano97-47] Anonymous. R&D bookstore.

C/C++ Users Journal, 15(3):??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:RBd

[Ano97-48] Anonymous. R&D bookstore.

C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:RBf

[Ano97-49] Anonymous. R&D bookstore. C/C++ Users Journal, 15(7):64– ??, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:RAN

[Ano97-50] Anonymous. RTI announces NetAcquire Java Toolkit. C/C++ Users Journal, 15(9):96, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:WMa

[Ano97-51] Anonymous. We have mail. C/C++ Users Journal, 15(1):82– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:WMb

[Ano97-52] Anonymous. We have mail. C/C++ Users Journal, 15(2):89– ??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:WMc

[Ano97-53] Anonymous. We have mail. C/C++ Users Journal, 15(3): 106–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:WMd

[Ano97-54] Anonymous. We have mail. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:WMe

[Ano97-55] Anonymous. We have mail. C/C++ Users Journal, 15(5): 100–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:WMf

[Ano97-56] Anonymous. We have mail.

C/C++ Users Journal, 15(6):98– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:WMg

[Ano97-57] Anonymous. We have mail.

C/C++ Users Journal, 15(7):98– ??, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:WMh

[Ano97-58] Anonymous. We have mail.

C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:WMi

[Ano97-59] Anonymous. We have mail.

C/C++ Users Journal, 15(9):99– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:CTP

[Ano98a] Anonymous. C++ theory and practice: const in parameter lists. C/C++ Users Journal, 16 (9):??, September 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:CCE

[Ano98b] Anonymous. C/C++ contributing editors. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFa

[Ano98c] Anonymous. Editor’s forum.

C/C++ Users Journal, 16(1):??, January 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:EFb

[Ano98d] Anonymous. Editor’s forum. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:EFc

[Ano98e] Anonymous. Editor’s forum. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFd

[Ano98f] Anonymous. Editor’s forum. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFe

[Ano98g] Anonymous. Editor’s forum. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFf

[Ano98h] Anonymous. Editor’s forum. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFg

[Ano98i] Anonymous. Editor’s forum. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFh

[Ano98j] Anonymous. Editor’s forum. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:EFi

[Ano98k] Anonymous. Editor’s forum.

C/C++ Users Journal, 16(9):??, September 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:EFl

[Ano98l] Anonymous. Editor’s forum.

C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:ERC

[Ano98m] Anonymous. Extending the reference-counting pattern. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPa

[Ano98n] Anonymous. New products.

C/C++ Users Journal, 16(1):??, January 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:NPb

[Ano98o] Anonymous. New products.

C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:NPc

[Ano98p] Anonymous. New products.

C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPd

[Ano98q] Anonymous. New products.

C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPe

[Ano98r] Anonymous. New products. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPf

[Ano98s] Anonymous. New products. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPg

[Ano98t] Anonymous. New products. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPh

[Ano98u] Anonymous. New products. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPi

[Ano98v] Anonymous. New products. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:NPj

[Ano98w] Anonymous. New products. C/C++ Users Journal, 16(10): 94–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPk

[Ano98x] Anonymous. New products. C/C++ Users Journal, 16(11): 95–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:NPl

[Ano98y] Anonymous. New products.

C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:OOP

[Ano98z] Anonymous. Object-oriented programming. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:ST

[Ano98-27] Anonymous. Software tools.

C/C++ Users Journal, 16(1):??, January 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:WMa

[Ano98-28] Anonymous. We have mail.

C/C++ Users Journal, 16(1):??, January 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:WMb

[Ano98-29] Anonymous. We have mail.

C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:WMc

[Ano98-30] Anonymous. We have mail.

C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMd

[Ano98-31] Anonymous. We have mail.

C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMe

[Ano98-32] Anonymous. We have mail. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMf

[Ano98-33] Anonymous. We have mail. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMg

[Ano98-34] Anonymous. We have mail. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMh

[Ano98-35] Anonymous. We have mail. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMi

[Ano98-36] Anonymous. We have mail. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1998:WMj

[Ano98-37] Anonymous. We have mail. C/C++ Users Journal, 16(10): 98–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMk

[Ano98-38] Anonymous. We have mail. C/C++ Users Journal, 16(11): 99–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1998:WMl

[Ano98-39] Anonymous. We have mail.

C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:CCCa

[Ano99a] Anonymous. C/C++ contributing editors. C/C++ Users Journal, 17(2):??, February 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:CCCb

[Ano99b] Anonymous. C/C++ contributing editors. C/C++ Users Journal, 17(3):??, March 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:CFC

[Ano99c] Anonymous. Comparing Fortran 90 and C++ valarrays. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:EFa

[Ano99d] Anonymous. Editor’s forum.

C/C++ Users Journal, 17(1):??, January 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:EFb

[Ano99e] Anonymous. Editor’s forum.

C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:EFc

[Ano99f] Anonymous. Editor’s forum.

C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:EFd

[Ano99g] Anonymous. Editor’s forum. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:EFe

[Ano99h] Anonymous. Editor’s forum. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:EFf

[Ano99i] Anonymous. Editor’s forum. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:EFj

[Ano99j] Anonymous. Editor’s forum. C/C++ Users Journal, 17(10): ??, October 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:EFk

[Ano99k] Anonymous. Editor’s forum. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:EFl

[Ano99l] Anonymous. Editor’s forum. C/C++ Users Journal, 17(12): ??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:ELT

[Ano99m] Anonymous. Extending <limits> for type safety. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPa

[Ano99n] Anonymous. New products.

C/C++ Users Journal, 17(1):??, January 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:NPb

[Ano99o] Anonymous. New products.

C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:NPc

[Ano99p] Anonymous. New products.

C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPd

[Ano99q] Anonymous. New products.

C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPe

[Ano99r] Anonymous. New products.

C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPf

[Ano99s] Anonymous. New products.

C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPg

[Ano99t] Anonymous. New products.

C/C++ Users Journal, 17(7):95– ??, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:NPh

[Ano99u] Anonymous. New products. C/C++ Users Journal, 17(8):94– ??, August 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:NPi

[Ano99v] Anonymous. New products. C/C++ Users Journal, 17(9):95– ??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPj

[Ano99w] Anonymous. New products. C/C++ Users Journal, 17(10): ??, October 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:NPk

[Ano99x] Anonymous. New products. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:NPl

[Ano99y] Anonymous. New products. C/C++ Users Journal, 17(12): ??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:RPC

[Ano99z] Anonymous. Restricted pointers are coming. C/C++ Users Journal, 17(7):55–??, July 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMa

[Ano99-27] Anonymous. We have mail. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:WMb

[Ano99-28] Anonymous. We have mail.

C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:WMc

[Ano99-29] Anonymous. We have mail.

C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMd

[Ano99-30] Anonymous. We have mail.

C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMe

[Ano99-31] Anonymous. We have mail.

C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMf

[Ano99-32] Anonymous. We have mail.

C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMg

[Ano99-33] Anonymous. We have mail.

C/C++ Users Journal, 17(7):99– ??, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:WMh

[Ano99-34] Anonymous. We have mail.

C/C++ Users Journal, 17(8):98– ??, August 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:WMi

[Ano99-35] Anonymous. We have mail. C/C++ Users Journal, 17(9):99– ??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMj

[Ano99-36] Anonymous. We have mail. C/C++ Users Journal, 17(10): ??, October 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1999:WMk

[Ano99-37] Anonymous. We have mail. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1999:WMl

[Ano99-38] Anonymous. We have mail. C/C++ Users Journal, 17(12): ??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Arceneaux:1994:CGI

[Arc94] Joseph Arceneaux. CUG392 — GNU indent v1.8. C Users Journal, 12(1):88–??, January 1994. ISSN 0898-9788.

Artigas:1989:CMS

[Art89] Robert Artigas, Jr. CUG248 — micro spell. C Users Journal, 7 (2):104–??, February 1989. ISSN 0898-9788.

Artigas:1990:CDC

[Art90] Robert Artigas, Jr. CUG243 — DECUS C preprocessor. C Users Journal, 8(6):127–??, June 1990. ISSN 0898-9788.

Asaf:1990:FDF

[Asa90] Arkin Asaf. A flexible dprintf function. C Users Journal, 8(9): 37–??, September 1990. ISSN 0898-9788.

Ashdown:1989:FWT

[Ash89] Ian Ashdown. The fast Walsh transform. C Users Journal, 7(7): 101–??, July 1989. ISSN 0898-9788.

Ashdown:1990:CLF

[Ash90] Ian Ashdown. C-scape and look and feel. C Users Journal, 8(12): 106–??, December 1990. ISSN 0898-9788.

Ashdown:1991:CPG

[Ash91a] Ian Ashdown. CUG350 — PCX graphics library. C Users Journal, 9(8):120–??, August 1991. ISSN 0898-9788.

Ashdown:1991:PG

[Ash91b] Ian Ashdown. PCX graphics. C Users Journal, 9(8):89–??, August 1991. ISSN 0898-9788.

Ashdown:1992:RR

[Ash92] Ian Ashdown. Radiosity and realism. C Users Journal, 10(8):33– ??, August 1992. ISSN 0898-9788.

Ashdown:1995:OCQ

[Ash95] Ian Ashdown. Octree color quantization. C/C++ Users Journal, 13(3):31–??, March 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Astle:1994:TTY

[Ast94a] Randel Dale Astle. Teach Yourself Windows Programming in 21 Days, by Charles Calvert. C/C++ Users Journal, 12(9):47– ??, September 1994. CODEN CCUJEX. ISSN 1075-2838. See [Cal93].

Astle:1994:HEA

[Ast94b] Randel Dale Astle. Handling exceptions to allocating memory. C Users Journal, 12(1):118–??, January 1994. ISSN 0898-9788.

Admal:1997:TTC

[AT97] Anil Admal and Chris Tarr. Templates and today’s compilers. C/C++ Users Journal, 15(1):27– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Babb:1987:PPP

[Bab87] Robert G. Babb, II. Programming parallel processors. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-11721-5. xiii + 380 pp. LCCN QA76.6.P75168 1988. See book review [Pla90e].

Babb:1988:PPP

[Bab88]  Robert G. Babb, II.  Pro-

gramming parallel processors. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-11721-5. xiii + 380 pp. LCCN QA76.6 .P75168 1988. US$34.95. See book review [Pla90e].

Baird:1991:ULA

[Bai91] Stuart T. Baird. Using large arrays in Turbo C. C Users Journal, 9(1):31–??, January 1991. ISSN 0898-9788.

Bailey:1999:SCC

[Bai99] Patrick Bailey. A signal command and control class for unix. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Baker:1990:CAM

[Bak90] Louis Baker. Complex arithmetic and matrices in C. C Users Journal, 8(5):59–??, May 1990. ISSN 0898-9788.

Baker:1991:CAM

[Bak91a] Louis Baker. Complex arithmetic and matrices in C++. C Users Journal, 9(1):123–??, January 1991. ISSN 0898-9788.

Baker:1991:DSP

[Bak91b] Louis Baker. Digital signal processing: A tutorial. C Users Journal, 9(5):47–??, May 1991. ISSN 0898-9788.

Baker:1992:CMF

[Bak92] Louis Baker. C mathematical function handbook. McGraw-Hill programming tools for scientists and engineers. McGraw-Hill, Inc., New York, NY, USA, 1992. ISBN 0-07-911158-0. xviii + 757 pp. LCCN QA 351 B17 1992. See book review [Smi92c]. System requirements for computer disk: PC; C or C++ compiler.

Baker:1994:TFM

[Bak94] Woody Baker. fixed modifier proposed. C Users Journal, 12 (1):124–??, January 1994. ISSN 0898-9788.

Baldwin:1989:CAC

[Bal89] Alan R. Baldwin. CUG292 — ASxxxx C cross assemblers. C Users Journal, 7(10):122–??, October 1989. ISSN 0898-9788.

Baldwin:1991:CAC

[Bal91a] Alan R. Baldwin. CUG346 ASxxxx cross assembler — part 2. C Users Journal, 9(7):126–??, July 1991. ISSN 0898-9788.

Baldwin:1991:DCO

[Bal91b] Wahhab Baldwin. Debugging in C — an overview. C Users Journal, 9(10):50–??, October 1991. ISSN 0898-9788.

Baldwin:1994:CAC

[Bal94] Alan R. Baldwin. CUG398 — ASxxx cross assembler — part 3. C Users Journal, 12(3):118– ??, March 1994. ISSN 0898-9788.

Ball:1997:ESS

[Bal97] Steve Ball. An encryption system for software registration. C/C++ Users Journal, 15(2):55– ??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Banning:1988:PRF

[Ban88]  Craig Banning.  Perfectly

rounded floats. C Users Journal, 6(2):14–??, February 1988. ISSN 0898-9788.

Banning:1989:RWF

[Ban89] Craig Banning. Reader/writer flames on C Users Journal editorial content. C Users Journal, 7 (1):111–??, January 1989. ISSN 0898-9788.

Banerji:1994:CRT

[Ban94] Arindam Banerji. CUG415 — run-time type information library for C++. C/C++ Users Journal, 12(8):88–??, August 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Barabash:1988:DAU

[Bar88] W. Barabash. Data acquisition using C. C Users Journal, 6 (2):13–??, February 1988. ISSN 0898-9788.

Barclay:1989:CPS

[Bar89] Kenneth A. Barclay. C, problem solving and programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1989. ISBN 0-13-115510-5. xii + 465 pp. LCCN QA76.73.C15 B365 1989 Bar. US$28.50. See book review [Nel89b].

Baran:1990:SGV

[Bar90a] Fuat C. Baran. Some good vendors of technical books. C Users Journal, 8(11):134–??, November 1990. ISSN 0898-9788.

Barnett:1990:AD

[Bar90b] Louis Barnett. Action diagrams. C Users Journal, 8(6):35– ??, June 1990. ISSN 0898-9788.

Barrett:1990:PTV

[Bar90c] Robert Barrett. PCX toolkit, v3.52. C Users Journal, 8(5):103– ??, May 1990. ISSN 0898-9788.

Barrett:1990:RGA

[Bar90d] Robert Barrett. References

on grep, awk, sed, and tr.

C Users Journal, 8(3):138–??, March 1990. ISSN 0898-9788.

Barnes:1991:QW

[Bar91a] Lan Barnes. QNX Windows. C Users Journal, 9(4):33–??, April 1991. ISSN 0898-9788.

Bart:1991:CN

[Bar91b] Randall Bart. C++: The name. C Users Journal, 9(12):130–??, December 1991. ISSN 0898-9788.

Barbu:1995:CCG

[Bar95a] Adrian Barbu. A C++ class generator. C/C++ Users Journal, 13(7):45–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Barbu:1995:QMD

[Bar95b] Adrian Barbu. Quick MS-Windows dialog design. C/C++ Users Journal, 13(9):51–??, September 1995. CODEN CCU-JEX. ISSN 1075-2838.

Bartosik:1998:EOP

[Bar98] Mark Bartosik. Encapsulating the observer pattern. C/C++ Users Journal, 16(10):72–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Bates:1991:MP

[Bat91] Rodney M. Bates. Modula-2: The pros. C Users Journal, 9 (1):137–??, January 1991. ISSN 0898-9788.

Bates:1992:DA

[Bat92] Rodney M. Bates. Debugging with assertions. C Users Journal, 10(10):40–??, October 1992. ISSN 0898-9788.

Batten:1997:CC

[Bat97a] Sean Batten. Casting in C++.

C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Batten:1997:RAF

[Bat97b] Sean Batten. A reusable apply function via the command pattern. C/C++ Users Journal, 15 (6):57–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Batov:1998:QSM

[Bat98] Vladimir Batov. A quick and simple memory allocator. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Bavestrelli:1998:ED

[Bav98a] Giovanni Bavestrelli. Encapsulating DDE. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Bavestrelli:1998:OWW

[Bav98b] Giovanni Bavestrelli. An OpenGL wrapper for Win32. C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Bavestrelli:1999:BAM

[Bav99a] Giovanni Bavestrelli. Better assertions for MFC. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Bavestrelli:1999:STO

[Bav99b] Giovanni Bavestrelli. Simple templates for OpenGL. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Bavestrelli:1999:WWO

[Bav99c] Giovanni Bavestrelli. Wrapping the Windows ODBC API. C/C++ Users Journal, 17(7):39– ??, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Bird:1994:CSB

[BB94] Jeffrey Bird and Christopher M. Box. CUG413B — sound blaster freedom project. C/C++ Users Journal, 12(8):88–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Bosser:1988:CMX

[BBAH88] Marc Bosser, Ralf E. Brendler, Les Aldridge, and Scott Holland. CUG270 — miscellany X. C Users Journal, 6(9):91– ??, September 1988. ISSN 0898-9788.

Burk:1989:PMB

[BC89a] Ron L. Burk and Helen Custer. Portability must be a design goal. C Users Journal, 7(1):89– ??, January 1989. ISSN 0898-9788.

Burk:1989:SCC

[BC89b] Ron L. Burk and Helen Custer. Scaffolding for C and C++. C Users Journal, 7(2):27–??, February 1989. ISSN 0898-9788.

Burk:1989:STC

[BC89c] Ron L. Burk and Helen Custer. Some tips for the C++ neophyte. C Users Journal, 7(4):51– ??, April 1989. ISSN 0898-9788.

Burk:1989:ZC

[BC89d] Ron L. Burk and Helen Custer.

Zortech C++. C Users Journal, 7 (5):51–??, May 1989. ISSN 0898-9788.

Burk:1990:GGL

[BC90a] Ron L. Burk and Helen Custer.

GRAD graphics library. C Users Journal, 8(1):127–??, January 1990. ISSN 0898-9788.

Burk:1990:ML

[BC90b] Ron L. Burk and Helen Custer.

Math.h++ library. C Users Journal, 8(10):107–??, October 1990. ISSN 0898-9788.

Bar-David:1989:MOC

[BD89a] Tsvi Bar-David. Mapping objectoriented concepts into C++ language facilities. C Users Journal, 7(7):9–??, July 1989. ISSN 0898-9788.

Bar-David:1989:PMF

[BD89b] Tsvi Bar-David. Pointers to member functions. C Users Journal, 7(4):75–??, April 1989. ISSN 0898-9788.

Bar-David:1990:BTEa

[BD90a] Tsvi Bar-David. Building a text editor: Part 2 — buffers, sloops, and yachts. C Users Journal, 8(3):43–??, March 1990. ISSN 0898-9788.

Bar-David:1990:BTEb

[BD90b] Tsvi Bar-David. Building a text editor: Part 3 — back to the editor. C Users Journal, 8(5):43–??, May 1990. ISSN 0898-9788.

Bar-David:1990:DIT

[BD90c] Tsvi Bar-David. Designing and implementing a text editor using OOP — part 1. C Users Journal, 8(1):105–??, January 1990. ISSN 0898-9788.

Bar-David:1990:FSO

[BD90d] Tsvi Bar-David. Formal specifications and object-oriented design. C Users Journal, 8(7):69– ??, July 1990. ISSN 0898-9788.

Bar-David:1990:HIPa

[BD90e] Tsvi Bar-David. Hiding the implementation — part 1. C Users Journal, 8(9):117–??, September 1990. ISSN 0898-9788.

Bar-David:1990:HIPb

[BD90f] Tsvi Bar-David. Hiding the implementation — part 2. C Users Journal, 8(11):83–??, November 1990. ISSN 0898-9788.

Bar-David:1991:TO

[BD91] Tsvi Bar-David. Type as object. C Users Journal, 9(7):55–??, July 1991. ISSN 0898-9788.

Bar-David:1989:URC

[BDB89] Tsvi Bar-David and David Bern. Using references in C++. C Users Journal, 7(10):21–??, October 1989. ISSN 0898-9788.

Beck:1993:DES

[Bec93] Andrew P. Beck. Developing 80x86-Based embedded systems. C Users Journal, 11(3):45–??, March 1993. ISSN 0898-9788.

Becker:1995:DKC

[Bec95a] Pete Becker. Dual-use keywords create confusion. C/C++ Users Journal, 13(10):93–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Becker:1995:GRG

[Bec95b] Pete Becker. Getting rid of goto.

C/C++ Users Journal, 13(12): 83–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Becker:1995:MR

[Bec95c] Pete Becker. More on rand().

C/C++ Users Journal, 13(11): 95–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Becker:1995:NRP

[Bec95d] Pete Becker. Nonrandom rand()

problems. C/C++ Users Journal, 13(9):91–??, September 1995. CODEN CCUJEX. ISSN 1075-2838.

Becker:1995:PIC

[Bec95e] Pete Becker. A pitfall inherited from C. C/C++ Users Journal, 13(7):91–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Becker:1995:SCR

[Bec95f] Pete Becker. Suicidal classes are risky business. C/C++ Users Journal, 13(8):91–??, August 1995. CODEN CCUJEX. ISSN 1075-2838.

Becker:1995:SPE

[Bec95g] Pete Becker. Surprising promotion effects. C/C++ Users Journal, 13(6):89–??, June 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1996:AVM

[Bec96a] Pete Becker. Availability vs. membership. C/C++ Users Journal, 14(3):55–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Becker:1996:CYO

[Bec96b] Pete Becker. Creating your own manipulator. C/C++ Users Journal, 14(1):83–??, January 1996. CODEN CCUJEX. ISSN 1075-2838.

Becker:1996:EMR

[Bec96c] Pete Becker. Even more on rand( ). C/C++ Users Journal, 14(5): 79–??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1996:UT

[Bec96d] Pete Becker. In undefined territory. C/C++ Users Journal, 14 (2):89–??, February 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1996:PWF

[Bec96e] Pete Becker. Parentheses won’t fix everything. C/C++ Users Journal, 14(4):87–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Becker:1996:QAK

[Bec96f] Pete Becker. Questions and answers: Kicking the macro habit. C/C++ Users Journal, 14(10): 89–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Becker:1996:QAL

[Bec96g] Pete Becker. Questions and answers: Little-known effects of defining constructors. C/C++ Users Journal, 14(9):77–??, September 1996. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1996:QAO

[Bec96h] Pete Becker. Questions and answers: Overloaded conversion operators. C/C++ Users Journal, 14(6):83–??, June 1996. CODEN CCUJEX. ISSN 1075-2838.

Becker:1996:QAV

[Bec96i] Pete Becker. Questions and answers: The value of code walks. C/C++ Users Journal, 14(7):87– ??, July 1996. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1996:QAT

[Bec96j] Pete Becker. Questions and answers: Traits to the rescue. C/C++ Users Journal, 14(11): 87–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Becker:1996:QAW

[Bec96k] Pete Becker. Questions and answers: What about realloc()? C/C++ Users Journal, 14(8):83– ??, August 1996. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1997:QAA

[Bec97a] Pete Becker. Questions & answers: Not all operator=’s are equal. C/C++ Users Journal, 15 (5):91–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Becker:1997:QAR

[Bec97b] Pete Becker. Questions & answers: Resolving type inside templates. C/C++ Users Journal, 15(2):79–??, February 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1997:QAUa

[Bec97c] Pete Becker. Questions & answers: Using the enum hack in C. C/C++ Users Journal, 15(1):75– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1997:QAWb

[Bec97d] Pete Becker. Questions & answers: What happens to unfreed memory? C/C++ Users Journal, 15(4):??, April 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1997:QAHb

[Bec97e] Pete Becker. Questions and answers: Handling variable-size objects. C/C++ Users Journal, 15 (8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Becker:1997:QAHa

[Bec97f] Pete Becker. Questions and answers: Hiding passwords in C. C/C++ Users Journal, 15(7):85– ??, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1997:QAUb

[Bec97g] Pete Becker. Questions and answers: Using templates in DLLs. C/C++ Users Journal, 15(9):89– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Becker:1997:QAWa

[Bec97h] Peter Becker. Questions & answers: Wrapping the prickly pragma. C/C++ Users Journal, 15(3):93–??, March 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1998:JSP

[Bec98a] Pete Becker. The journeyman’s shop: A place to hone your skills. C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAA

[Bec98b] Pete Becker. Questions & answers: Assumptions to avoid concerning memory. C/C++ Users Journal, 16(7):??, July 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1998:QACb

[Bec98c] Pete Becker. Questions & answers: C++ Standard approved. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1998:QACd

[Bec98d] Pete Becker. Questions & answers: Catching exceptions, and flak, from readers. C/C++ Users Journal, 16(11):91–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QACc

[Bec98e] Pete Becker. Questions & answers: Creating variable-size structs. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAE

[Bec98f] Pete Becker. Questions & answers: Every template needs a friend (now and then). C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAI

[Bec98g] Pete Becker. Questions & answers: Implementing dynamic function calls. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAP

[Bec98h] Pete Becker. Questions & answers: Portability, promotion, and other concerns. C/C++ Users Journal, 16(10):85–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QACa

[Bec98i] Pete Becker. Questions & answers: The complex evolution of <string>. C/C++ Users Journal, 16(1):??, January 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAD

[Bec98j] Pete Becker. Questions & answers: The danger of undeclared functions. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAUa

[Bec98k] Pete Becker. Questions & answers: Understanding access restrictions. C/C++ Users Journal, 16(3):??, March 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1998:QAUb

[Bec98l] Pete Becker. Questions & answers: Use caution with temporary objects. C/C++ Users Journal, 16(5):??, May 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1998:SIS

[Bec98m] Thomas Becker. Smart iterators and STL. C/C++ Users Journal, 16(9):??, September 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSC

[Bec99a] Pete Becker. The journeyman’s shop: Containing heterogeneous types. C/C++ Users Journal, 17 (12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSIb

[Bec99b] Pete Becker. The journeyman’s shop: Initialization and cleanup in C++. C/C++ Users Journal, 17(4):??, April 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSIa

[Bec99c] Pete Becker. The journeyman’s shop: Initialization and cleanup, part 1. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSIc

[Bec99d] Pete Becker. The journeyman’s shop: Initialization and cleanup, part 3. C/C++ Users Journal, 17 (5):??, May 1999. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1999:JSL

[Bec99e] Pete Becker. The journeyman’s shop: Looping and iteration. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSMb

[Bec99f] Pete Becker. The journeyman’s shop: Memory and object management. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSMc

[Bec99g] Pete Becker. The journeyman’s shop: Memory and object management, part 2. C/C++ Users Journal, 17(8):71–??, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSMd

[Bec99h] Pete Becker. The journeyman’s shop: Memory and object management, part 3. C/C++ Users Journal, 17(10):??, October 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSMa

[Bec99i] Pete Becker. The journeyman’s shop: More on error handling. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCU-JEX. ISSN 1075-2838.

Becker:1999:JSS

[Bec99j]  Pete Becker.  The journey-

man’s shop: Software testing 101. C/C++ Users Journal, 17(9):83– ??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:JSW

[Bec99k] Pete Becker. The journeyman’s shop: Wrapping up error handling techniques. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCUJEX. ISSN 1075-2838.

Becker:1999:SPW

[Bec99l] Thomas Becker. A semaphore with priorities for Win32. C/C++ Users Journal, 17(8):18–??, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Beddow:1991:MCF

[Bed91] A. J. M. Beddow. Multi-threaded C functions. C Users Journal, 9 (1):57–??, January 1991. ISSN 0898-9788.

Beddow:1995:RPC

[Bed95] Mark Beddow. Remote procedure calls using NetBIOS. C/C++ Users Journal, 13(8):67– ??, August 1995. CODEN CCU-JEX. ISSN 1075-2838.

Behrendsen:1996:PMS

[Beh96] Tim Behrendsen. The poor man’s serial protocol analyzer. C/C++ Users Journal, 14(2):29– ??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Bell:1989:LA

[Bel89a] John T. Bell. A line-drawing algorithm. C Users Journal, 7(8): 81–??, August 1989. ISSN 0898-9788.

Bell:1989:UDU

[Bel89b] John T. Bell. A universal directory utility for Apple systems. C Users Journal, 7(5):37–??, May 1989. ISSN 0898-9788.

Bell:1996:NS

[Bel96] James M. Bell. Nested for statements. C/C++ Users Journal, 14(7):47–??, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Bell:1998:WCN

[Bel98] John T. Bell. A wrapper class for NT services. C/C++ Users Journal, 16(8):??, August 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Benda:1992:OMD

[Ben92] Carl M. Benda. An OS/2 MIDI device driver. C Users Journal, 10(1):64–??, January 1992. ISSN 0898-9788.

Bennett:1994:CSV

[Ben94] Kendall Bennett. CUG409 — SuperVGA VESA VBE test kit and universal VESA VBE. C Users Journal, 12(6):98–??, June 1994. ISSN 0898-9788.

Benzinger:1997:WKM

[Ben97] Mike Benzinger. We’re not in Kansas any more. C/C++ Users Journal, 15(5):35–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Benzinger:1998:SCB

[Ben98] Mike Benzinger. STL containers based on hash tables. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCUJEX. ISSN 1075-2838.

Bergeon:1988:CZC

[Ber88] Francis Bergeon. CUG260 — zmodem, CU and tty library. C Users Journal, 6(3):65–??, March 1988. ISSN 0898-9788.

Bergmann:1990:CRT

[Ber90] Ernest E. Bergmann. CUG304 — ROFF5: Technical text formatter. C Users Journal, 8(3):130– ??, March 1990. ISSN 0898-9788.

Berens:1992:TPT

[Ber92a] Tim Berens. Two-dimensional procedure tables. C Users Journal, 10(7):47–??, July 1992. ISSN 0898-9788.

Bertrand:1992:CMF

[Ber92b] Michael Bertrand. The CORDIC method for faster sin and cos calculations. C Users Journal, 10 (11):57–??, November 1992. ISSN 0898-9788.

Berntzon:1994:CC

[Ber94] Lars Berntzon. CUG402 — CForms. C Users Journal, 12(4): 122–??, April 1994. ISSN 0898-9788.

Beyer:1998:EOA

[Bey98] Malcolm Beyer. Encapsulating the ODBC API. C/C++ Users Journal, 16(7):??, July 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Gangwere:1991:KP

[bG91] Don bandit Gangwere. Kermit protocol. C Users Journal, 9(9): 25–??, September 1991. ISSN 0898-9788.

Baker:1992:BAC

[BG92] Louis Baker and Anthony Giancola. Bit arrays with C++. C Users Journal, 10(7):21–??, July 1992. ISSN 0898-9788.

Barclay:1994:CPS

[BG94] Kenneth A. Barclay and Brian J.

Gordon. C++ problem solving and programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-126673-X. xiii + 552 pp. LCCN QA76.73.C153 B347 1994.

Bertrand:1999:TFS

[BG99] Michael Bertrand and Dave Grundgeiger. TrueType font secrets. C/C++ Users Journal, 17(8):32, 34–36, 38–40, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Binstock:1988:TCT

[Bin88] Andrew Binstock. C Toolbox — Serious C Programming for the IBM PC, The, by William J. Hunt. C Users Journal, 6(1):41– ??, January 1988. ISSN 0898-9788. See [Hun85].

Bishop:1990:SC

[Bis90] Matt Bishop. Storage in C. C Users Journal, 8(5):73–??, May 1990. ISSN 0898-9788.

Brown:1991:PIP

[BK91] Ralf Brown and Jim Kyle. PC interrupts: a programmer’s reference to BIOS, DOS, and thirdparty calls. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-57797-6. various pp. LCCN QA76.6 .B7734 1991. US$32.95, CDN$42.95. See [Bro94].

Brown:1994:NIP

[BK94a] Ralf Brown and Jim Kyle. Network interrupts: a programmer’s reference to network APIs. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-62644-6. vi + 730 pp. LCCN TK5105.5 .B77 1994. US$32.95, CDN$42.95. See [Bro94].

Brown:1994:PIP

[BK94b] Ralf Brown and Jim Kyle. PC interrupts: a programmer’s reference to BIOS, DOS, and thirdparty calls. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-57797-6. vi + 1210 pp. LCCN QA76.6 .B7734 1991. US$39.95, CDN$51.95. See [Bro94].

Blain:1990:TCT

[Bla90] Derrel R. Blain. C Toolbox, The, 2nd edition, by William J. Hunt. C Users Journal, 8(11):62– ??, November 1990. ISSN 0898-9788.

Blezek:1998:RPS

[Ble98] Daniel Blezek. Rapid prototyping with SWIG. C/C++ Users Journal, 16(11):61–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Blossom:1991:DAT

[Blo91] Eric Blossom. Decoding ASN.1 transfer syntax. C Users Journal, 9(9):57–??, September 1991. ISSN 0898-9788.

Bloodworth:1993:PPS

[Blo93] Carey Bloodworth. Pivot point selection effects in Quicksort. C Users Journal, 11(2):136–??, February 1993. ISSN 0898-9788.

Blum:1990:CW

[Blu90]  David Blum.  CUG328 —

WTWG. C Users Journal, 8(11): 126–??, November 1990. ISSN 0898-9788.

Blum:1991:CSV

[Blu91] David Blum. CUG352 — String and Vlist: A couple of classes. C Users Journal, 9(9):126–??, September 1991. ISSN 0898-9788.

Blum:1992:NNC

[Blu92] Adam Blum. Neural networks in C++: an object-oriented framework for building connectionist systems. Wiley professional computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1992. ISBN 0-471-55201-1 (book/disk set), 0-471-53847-7. x + 213 pp. LCCN QA76.87 .B58 1992 Accompanying disk shelved in Reserve. US$74.90 (book + disk), US$34.95 (paperback). See book review [Phi92c]. Includes one computer disk inserted in pocket.

Blum:1994:CW

[Blu94]  David Blum.  CUG328 —

WTWG 2.0. C/C++ Users Journal, 12(7):83–??, July 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Barton:1994:SEC

[BN94] John J. Barton and Lee R. Nackman. Scientific and engineering C++: an introduction with advanced techniques and examples. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-53393-6. xv + 671 pp. LCCN QA76.73.C153 B38 1994. See book review [Swa95a].

Bockenfeld:1996:TNT

[Boc96] Don Bockenfeld. TSQRT: A new trick for an old dog. C/C++ Users Journal, 14(3):39– 41, March 1996. CODEN CCU-JEX. ISSN 1075-2838. Implements a small table-driven square root function in C, using exclusively integer operations.

Booch:1991:OOD

[Boo91]  Grady Booch.  Object ori-

ented design with applications. Benjamin/Cummings series in Ada and software engineering. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1991. ISBN 0-8053-0091-0. xix + 580 pp. LCCN QA76.64 .B66 1991. See book review [Lan91a].

Bouthoorn:1994:CAA

[Bou94] Peter M. Bouthoorn. CUG412 — AISEARCH: Artificial intelligence search class. C/C++ Users Journal, 12(7):84–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Boyle:1988:IMU

[Boy88] William Boyle. Inter-task messaging under QNX. C Users Journal, 6(11):65–??, November 1988. ISSN 0898-9788.

Brannigan:1988:AGS

[Bra88] Michael Brannigan. Algorithms for generating standard curves. C Users Journal, 6(8):87–??, August 1988. ISSN 0898-9788.

Brannigan:1989:MP

[Bra89] Michael Brannigan. Masked 3-D plotting. C Users Journal, 7(8): 39–??, August 1989. ISSN 0898-9788.

Brannigan:1990:ADA

[Bra90a] Michael Brannigan. An adaptive data analyzer. C Users Journal, 8(5):113–??, May 1990. ISSN 0898-9788.

Brannigan:1990:FCD

[Bra90b] Michael Brannigan. Fitting curves to data. C Users Journal, 8(1):53–??, January 1990. ISSN 0898-9788.

Brandt:1991:CST

[Bra91] Michael E. Brandt. Comparing signals in the time domain. C Users Journal, 9(5):58–??, May 1991. ISSN 0898-9788.

Bradley:1992:IDM

[Bra92a] Don C. Bradley. Implementing direct memory access (DMA) in C. C Users Journal, 10(5):53–??, May 1992. ISSN 0898-9788.

Brailsford:1992:WSC

[Bra92b] Steve Brailsford. Writing a small CAD package. C Users Journal, 10(8):53–??, August 1992. ISSN 0898-9788.

Brandt:1992:DMM

[Bra92c] Michael E. Brandt. Don’t mess with marilyn! C Users Journal, 10(7):62–??, July 1992. ISSN 0898-9788.

Breitzman:1996:CRL

[Bre96] Anthony Breitzman. A class for representing large integers. C/C++ Users Journal, 14(11): 61–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Bright:1990:VM

[Bri90] Walter Bright. Virtual memory for 640K DOS. C Users Journal, 8(7):37–??, July 1990. ISSN 0898-9788.

Briand:1996:DCR

[Bri96] Marc Briand. Designing and Coding Reusable C++, by Martin D. Carroll and Margaret A. Ellis. C/C++ Users Journal, 14 (6):63–??, June 1996. CODEN CCUJEX. ISSN 1075-2838. See [CE95].

Brill:1997:SVB

[Bri97] Gregory Brill. Sharing variables between Win32 executables. C/C++ Users Journal, 15 (3):57–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Briand:1998:BRA

[Bri98a] Marc Briand. Book review: AntiPatterns — Refactoring Software, Architectures, and Projects in Crisis. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Briand:1998:BRP

[Bri98b] Marc Briand. Book review: The Programmer’s Guide to SCSI. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Briand:1998:EFj

[Bri98c] Marc Briand. Editor’s forum. C/C++ Users Journal, 16(10): 6, October 1998. CODEN CCU-JEX. ISSN 1075-2838.

Briand:1998:EFk

[Bri98d] Marc Briand. Editor’s forum. C/C++ Users Journal, 16(11):6, November 1998. CODEN CCU-JEX. ISSN 1075-2838.

Bridges:1998:LEJ

[Bri98e] Michael Bridges. Living without enum in Java. C/C++ Users Journal, 16(11):82–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Brill:1998:IC

[Bri98f] Gregory Brill. An introduction to COM. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Brill:1998:WCC

[Bri98g] Gregory Brill. Writing COM clients with late and early binding. C/C++ Users Journal, 16 (10):37–??, October 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Briand:1999:BRC

[Bri99a] Marc Briand. Book review: C++ for Java Programmers. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Briand:1999:EFg

[Bri99b] Marc Briand. Editor’s forum. C/C++ Users Journal, 17(7):6, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Briand:1999:EFh

[Bri99c] Marc Briand. Editor’s forum.

C/C++ Users Journal, 17(8):6, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Briand:1999:EFi

[Bri99d] Marc Briand. Editor’s forum.

C/C++ Users Journal, 17(9):6, September 1999. CODEN CCU-JEX. ISSN 1075-2838.

Briand:1999:RVT

[Bri99e] Marc Briand. A recursive visit template. C/C++ Users Journal, 17(10):??, October 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Broderick:1989:CJV

[Bro89] John James Broderick. CUG291 — JJB — version 1.2 for quick C and Turbo C programmers. C Users Journal, 7(8):115–??, August 1989. ISSN 0898-9788.

Brown:1990:CCC

[Bro90a] Brian Brown. CUG309 — 6809 C compiler for MS-DOS. C Users Journal, 8(4):123–??, April 1990. ISSN 0898-9788.

Brown:1990:CPC

[Bro90b] J. Brown. CUG327 — panels for C. C Users Journal, 8(11):125– ??, November 1990. ISSN 0898-9788.

Brown:1994:UIP

[Bro94] Ralf Brown. Uninterrupted interrupts: a programmer’s CD-ROM reference to network APIs, and to BIOS, DOS, and third-party calls. Addison-Wesley, Reading, MA, USA, 1994. ISBN ???? LCCN ???? See book review [Swa95b]. One CD-ROM. Based on [BK91, BK94b, BK94a].

Brumbaugh:1989:OOB

[Bru89a] David Brumbaugh. Oracle offers benefits important to corporate user. C Users Journal, 7(11):21– ??, November 1989. ISSN 0898-9788.

Bruton:1989:PM

[Bru89b] Allan Bruton. Programming the Macintosh. C Users Journal, 7 (10):97–??, October 1989. ISSN 0898-9788.

Brumbaugh:1990:OPC

[Bru90]  David Brumbaugh.  Object-

oriented programming in C. C Users Journal, 8(7):113–??, July 1990. ISSN 0898-9788.

Brungs:1991:BVP

[Bru91] Michael Brungs. Bit variables proposed. C Users Journal, 9(8): 126–??, August 1991. ISSN 0898-9788.

Brumbaugh:1992:PCL

[Bru92a] David Brumbaugh. Porting C libraries to C++. C Users Journal, 10(1):21–??, January 1992. ISSN 0898-9788.

Brumbaugh:1992:ZIL

[Bru92b] David Brumbaugh. Zinc interface library, version 2.0. C Users Journal, 10(4):115–??, April 1992. ISSN 0898-9788.

Bryson:1998:BRC

[Bry98] Donald Bryson. Book review: C Interfaces and Implementations. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCU-JEX. ISSN 1075-2838.

Ballay:1994:TCC

[BS94] Helene Ballay and Rainer Storn.

A tool for checking C coding conventions. C/C++ Users Journal, 12(7):41–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Broadman:1999:ECM

[BS99] Allen Broadman and Eric Shaw.

Executing a class member in its own thread. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Bingham:1992:OC

[BSG92] William Bingham, Tom Schlintz, and Greg Goslen. OOP without C++. C Users Journal, 10(3):31– ??, March 1992. ISSN 0898-9788.

Bergman-Terrell:1991:CGC

[BT91] Eric Bergman-Terrell. A calltree generator for C programs. C Users Journal, 9(10):33–??, October 1991. ISSN 0898-9788.

Bugg:1993:IUW

[BT93] Keith E. Bugg and Jack Tackett, Jr. Implementing and using Windows help. C Users Journal, 11 (9):63–??, September 1993. ISSN 0898-9788.

Bugg:1991:MM

[Bug91] Keith E. Bugg. Of mice and menus. C Users Journal, 9(4):46– ??, April 1991. ISSN 0898-9788.

Bugg:1994:CSC

[Bug94] Keith E. Bugg. Creating spin controls for Windows. C Users Journal, 12(4):39–??, April 1994. ISSN 0898-9788.

Bugg:1998:NHH

[Bug98] Keith Bugg. The new HTML Help Standard. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Burki:1989:MUI

[Bur89] David Burki. MS-DOS undocumented interrupt allows command execution from within programs. C Users Journal, 7(5):23– ??, May 1989. ISSN 0898-9788.

Burk:1990:BBB

[Bur90a] Ron L. Burk. Building a better Boolean with C++. C Users Journal, 8(4):57–??, April 1990. ISSN 0898-9788.

Burk:1990:TOP

[Bur90b] Ron L. Burk. Object-Oriented Program Design With Examples In C++, by Mark Mullin. C Users Journal, 8(6):105–??, June 1990. ISSN 0898-9788. See [Mul89].

Burk:1992:TFO

[Bur92a] Ron L. Burk. File Organization and Processing, by Alan L. Tharp. C Users Journal, 10 (2):52–??, February 1992. ISSN 0898-9788. See [Tha88].

Burk:1992:HGP

[Bur92b] Ron L. Burk. Hashing: From good to perfect. C Users Journal, 10(2):41–??, February 1992. ISSN 0898-9788.

Burki:1993:DC

[Bur93] David Burki. Date conversions.

C Users Journal, 11(2):29–??, February 1993. ISSN 0898-9788.

Busigin:1991:QOS

[Bus91] Nick Busigin. The QNX operating system v2.15C. C Users Journal, 9(9):106–??, September 1991. ISSN 0898-9788.

Bybee:1992:PVI

[Byb92] Robert D. Bybee. A portable VMS-Style input line routine. C Users Journal, 10(1):36–??, January 1992. ISSN 0898-9788.

Cady:1991:TSE

[Cad91] Robert E. Cady. Software Engineering in C. C Users Journal, 9 (1):133–??, January 1991. ISSN 0898-9788.

Callaway:1990:SSA

[Cal90] David X. Callaway. Some sort algorithms benchmarked. C Users Journal, 8(12):128–??, December 1990. ISSN 0898-9788.

Calvert:1993:TYW

[Cal93] Charlie Calvert. Teach yourself Windows programming in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 1993.

ISBN 0-672-30344-2. xvii + 904 pp. LCCN QA76.76.W56 C35 1993. US$29.95, CDN$37.95. See book review [Ast94a].

Calvert:1995:TYW

[Cal95] Charlie Calvert. Teach yourself Windows 95 programming in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1995. ISBN 0-672-30531-3. xxiv + 1259 pp. LCCN QA76.76.O63 C352 1995.

Campbell:1987:CPG

[Cam87] Joe Campbell. C programmer’s guide to serial communications. Howard W. Sams, Indianapolis, IN 46268, USA, 1987. ISBN 0-672-22584-0 (paperback). xii + 655 pp. LCCN QA76.73.C15 C36 1987 Bar. US$22.95. See [Wio89].

Cameron:1988:CLD

[Cam88] Alex Cameron. CUG library disks should include source code. C Users Journal, 6(11):93–??, November 1988. ISSN 0898-9788.

Cameron:1990:ADU

[Cam90] Alex Cameron. An adaptable disk utility. C Users Journal, 8(5):135–??, May 1990. ISSN 0898-9788.

Campbell:1993:CPG

[Cam93] Joe Campbell. C programmer’s guide to serial communications. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1993. ISBN 0-672-30286-1. xxxiii + 939 pp. LCCN QA76.73.C15C36 1993.

US$39.95, CDN$49.95. Folded ASCII wall chart laid in.

Campbell:1994:PSC

[Cam94] Frank Campbell. Programming:

Science or craft. C Users Journal, 12(1):118–??, January 1994. ISSN 0898-9788.

Campbell:1998:IPD

[Cam98] Brian Campbell. Implementing Pascal data types in C++. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Carver:1988:IRM

[Car88] Richard Carver. iRMX: A realtime multi-tasking operating system. C Users Journal, 6(11):35– ??, November 1988. ISSN 0898-9788.

Carver:1991:IFO

[Car91] Richard Carver. The iRMX family of operating systems. C Users Journal, 9(1):109–??, January 1991. ISSN 0898-9788.

Cargill:1992:CPS

[Car92] Tom Cargill. C++ programming style. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56365-7. xi + 233 pp. LCCN QA76.73.C153 C37 1992. See book review [Hal93c].

Cargill:1993:KNF

[Car93] Ian Cargill. Keywords needed for fixed-size integers. C Users Journal, 11(6):138–??, June 1993. ISSN 0898-9788.

Carlson:1997:ATT

[Car97] Paul Carlson. An automated testing tool for Win16. C/C++ Users Journal, 15(7):25–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Carlson:1999:APS

[Car99] Paul Carlson. An automated process shutdown DLL. C/C++ Users Journal, 17(8):59–??, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Casey:1989:UIT

[Cas89] Tamara Casey. Upgrade information on Turbo C. C Users Journal, 7(2):111–??, February 1989. ISSN 0898-9788.

Casey:1993:TCT

[Cas93] William Casey. C Toolbox — Serious C Programming for the IBM PC, The, by William J. Hunt. C Users Journal, 11(1): 136–??, January 1993. ISSN 0898-9788.

Cavallito:1991:LCA

[Cav91] Frank Cavallito. Linking C and APL. C Users Journal, 9(6):39– ??, June 1991. ISSN 0898-9788.

Cameron:1990:CAC

[CD90] Alex Cameron and Hugh Daschbach. CUG307 — ADU and COMX (device driver). C Users Journal, 8(4):123–??, April 1990. ISSN 0898-9788.

Cilwa:1994:WPP

[CD94] Paul Cilwa and Jeff Duntemann. Windows programming power with custom controls. Coriolis Group Books, Scottsdale, AZ, USA, 1994. ISBN 1-883577-00-4. xviii + 492 pp. LCCN QA76.76.W56 C49 1994. See book review [Swa95c].

Carroll:1995:DCR

[CE95] Martin D. Carroll and Margaret A. Ellis. Designing and coding reusable C++. Addison-Wes-ley, Reading, MA, USA, 1995. ISBN 0-201-51284-X. xvi + 317 pp. LCCN QA 76.73 C153 C39 1995. See book review [Bri96].

Carroll:1996:DCC

[CE96a] Martin D. Carroll and Margaret A. Ellis. Designing C++ classes. C/C++ Users Journal, 14(6):55–??, June 1996. CODEN CCUJEX. ISSN 1075-2838.

Carroll:1996:IE

[CE96b] Martin D. Carroll and Margaret A. Ellis. Inheritability and extensibility. C/C++ Users Journal, 14(7):37–??, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Carroll:1996:OIC

[CE96c] Martin D. Carroll and Margaret A. Ellis. Obstacles to inheritability in C++. C/C++ Users Journal, 14(8):67–??, August 1996. CODEN CCUJEX. ISSN 1075-2838.

Cedola:1988:CEG

[Ced88] Kent Cedola. CUG255 — EGA graphics library. C Users Journal, 6(1):53–??, January 1988. ISSN 0898-9788.

Celko:1990:SC

[Cel90] Joe Celko. Structure charts. C Users Journal, 8(6):113–??, June 1990. ISSN 0898-9788.

Celko:1991:STQ

[Cel91] Joe Celko. Some tips for Quick-Sort. C Users Journal, 9(1):63– ??, January 1991. ISSN 0898-9788.

Cepek:1995:PFM

[Cep95] Mike Cepek. Programming flash memory. C/C++ Users Journal, 13(2):35–??, February 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Cohen:1988:TCC

[CH88] Stanley Cohen and Mark Harris. Transforming a compound conditional statement into several oneliners. C Users Journal, 6(6):91– ??, June 1988. ISSN 0898-9788.

Chalk:1988:SHS

[Cha88] Andrew J. Chalk. Smooth hardware scrolling and panning. C Users Journal, 6(8):43–??, August 1988. ISSN 0898-9788.

Chappell:1990:DHD

[Cha90] Jon Chappell. Debugging with the 80386 hardware debug registers. C Users Journal, 8(10): 69–??, October 1990. ISSN 0898-9788.

Chapman:1995:EMM

[Cha95] David Chapman. An error manager with message text replacement. C/C++ Users Journal, 13 (4):23–??, April 1995. CODEN CCUJEX. ISSN 1075-2838.

Chaudhry:1997:NTC

[Cha97] Puneesh Chaudhry. A new trace class. C/C++ Users Journal, 15 (6):51–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Cheatum:1988:CA

[Che88] Daniel Cheatum. CUG226 — ART-CEE. C Users Journal, 6 (1):53–??, January 1988. ISSN 0898-9788.

Cheng:1996:CPC

[Che96] Harry H. Cheng. CGI programming in C. C/C++ Users Journal, 14(11):17–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Chien:1989:PGT

[Chi89] Chao C. Chien. Programmer’s guide to Turbo C 2.0. Bantam Doubleday Dell Publishing Group Inc., 666 Fifth Avenue, New York, NY 10130, USA, 1989. ISBN 0-553-34757-8. xii + 299 pp. LCCN ???? See book review [Spa90].

Chiaia:1997:SPO

[Chi97] Gualtiero Chiaia. Scientific plotting with OWL. C/C++ Users Journal, 15(7):17–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Choma:1988:IPT

[Cho88] Peter Choma. The IEEE POSIX trial-use standard. C Users Journal, 6(6):89–??, June 1988. ISSN 0898-9788.

Chursenoff:1991:SCT

[Chu91] Greg Chursenoff. Serial communications with Turbo C. C Users Journal, 9(4):63–??, April 1991. ISSN 0898-9788.

Cindric:1995:WBA

[Cin95] Blaise B. Cindric. When the “best” algorithm isn’t. C/C++ Users Journal, 13(3):27–??, March 1995. CODEN CCUJEX. ISSN 1075-2838.

Clark:1988:ABD

[Cla88] David D. Clark. Accelerate block data moves with these assembly language functions. C Users Journal, 6(3):37–??, March 1988. ISSN 0898-9788.

Clamage:1990:SC

[Cla90a] Steve Clamage. Stringizing constants. C Users Journal, 8(6):75– ??, June 1990. ISSN 0898-9788.

Clark:1990:DOC

[Cla90b] David D. Clark. A date object in C++. C Users Journal, 8(6):57– ??, June 1990. ISSN 0898-9788.

Clark:1990:TUI

[Cla90c]  David D. Clark.  User In-

terfaces In C++ and Object-Oriented Programming, by Mark Goodwin. C Users Journal, 8(7): 131–??, July 1990. ISSN 0898-9788. See [Goo89].

Clarke:1997:BRS

[Cla97] G. Miller Clarke. Book review: Seven Java books — A whirlwind tour. C/C++ Users Journal, 15 (5):37–39, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Clamage:1999:QDR

[Cla99] Stephen D. Clamage. Queries, defect reports, and the C++ Standard. C/C++ Users Journal, 17(9):71–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Cline:1992:BAF

[Cli92] Alan Cline. Build applications faster with state transition automatons. C Users Journal, 10 (12):93–??, December 1992. ISSN 0898-9788.

Clune:1989:TFP

[Clu89a] Thomas R. Clune. fscanf() as a pseudo-parser. C Users Journal, 7(4):83–??, April 1989. ISSN 0898-9788.

Clune:1989:RFT

[Clu89b] Thomas R. Clune. Reader flames on C Users Journal editorial content. C Users Journal, 7(1):111– ??, January 1989. ISSN 0898-9788.

Clune:1990:CFF

[Clu90] Thomas R. Clune. CUG315 — FTGRAPH (fast Fourier transform graphics). C Users Journal, 8(6):127–??, June 1990. ISSN 0898-9788.

Crews:1990:CMB

[CM90] George Crews and Ronald Michaels. CUG299 — MEL and BP. C Users Journal, 8(2):125– ??, February 1990. ISSN 0898-

Chiaia:1998:CHD

[CM98] Gualtiero Chiaia and Marco Marcon. A class hierarchy for data acquisition. C/C++ Users Journal, 16(11):39–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Coatmeur:1994:CB

[Coâ94a] Alain Coâtmeur. CUG404 — Bison++. C Users Journal, 12(5): 97–??, May 1994. ISSN 0898-9788.

Coatmeur:1994:CF

[Coâ94b] Alain Coâtmeur. CUG405 — Flex++. C Users Journal, 12 (5):97–??, May 1994. ISSN 0898-9788.

Cockburn:1998:OOAa

[Coc98a] Alistair Cockburn. Objectoriented analysis and design, part 1. C/C++ Users Journal, 16(5): ??, May 1998. CODEN CCU-JEX. ISSN 1075-2838.

Cockburn:1998:OOAb

[Coc98b] Alistair Cockburn. Objectoriented analysis and design, part 2. C/C++ Users Journal, 16(6): ??, June 1998. CODEN CCU-JEX. ISSN 1075-2838.

Coffin:1988:TCP

[Cof88]  Bernard E. Coffin.  The

C++ Programming Language, by Bjarne Stroustrup. C Users Journal, 6(6):63–??, June 1988. ISSN 0898-9788. See [Str86].

Cogar:1988:CDC

[Cog88a] Phil E. Cogar. CUG256 — DeSmet carry flag routines. C Users Journal, 6(2):57–??, February 1988. ISSN 0898-9788.

Cogar:1988:HMH

[Cog88b] Phil E. Cogar. How to make hardware registers and flags available to C programs. C Users Journal, 6(2):65–??, February 1988. ISSN 0898-9788.

Cogar:1989:LHF

[Cog89] Phil E. Cogar. Lack of header files in C Users Journal listings. C Users Journal, 7(10):138–??, October 1989. ISSN 0898-9788.

Cohn:1992:CIV

[Coh92] Michael W. Cohn. Creating interactive voice response systems in C++. C Users Journal, 10(6):41– ??, June 1992. ISSN 0898-9788.

Colley:1988:CSM

[Col88a] William C. Colley III. CUG242 — still more X-Assemblers. C Users Journal, 6(1):53–??, January 1988. ISSN 0898-9788.

Colley:1988:CSC

[Col88b] William C. Colley III. CUG267 — 8085, 2650, and S6 cross assemblers. C Users Journal, 6(8): 63–??, August 1988. ISSN 0898-9788.

Colley:1988:CZP

[Col88c] William C. Colley III. CUG276 — Z80, PIC1650, and 6804 cross assemblers. C Users Journal, 6 (9):92–??, September 1988. ISSN 0898-9788.

Colby:1989:JJT

[Col89a] Lance Colby. JYACC’s JAM toolkit speeds the building of character-based user interfaces. C Users Journal, 7(10):114–??, October 1989. ISSN 0898-9788.

Colley:1989:CAC

[Col89b] William C. Colley III. CUG292 — ASxxxx C cross assemblers. C Users Journal, 7(11):97–??, November 1989. ISSN 0898-9788.

Colley:1989:MHF

[Col89c] William C. Colley III. Manipulating hex files — A ‘standard library’ approach. C Users Journal, 7(5):27–??, May 1989. ISSN 0898-9788.

Colvin:1989:SNN

[Col89d] Gregory Colvin. SYNAPSYS: A neural network. C Users Journal, 7(4):59–??, April 1989. ISSN 0898-9788.

Collins:1990:CMD

[Col90a] John M. Collins. CUG303 — MC68K disassembler. C Users Journal, 8(3):130–??, March 1990. ISSN 0898-9788.

Colner:1990:OAC

[Col90b] Don Colner. An object-oriented approach to command line options. C Users Journal, 8(7):43– ??, July 1990. ISSN 0898-9788.

Colvin:1990:CTS

[Col90c] Gregory Colvin. CUG306 thread and synapsys. C Users Journal, 8(3):131–??, March 1990. ISSN 0898-9788.

Colvin:1990:MLT

[Col90d] Gregory Colvin. Multitasking with lightweight threads. C Users Journal, 8(3):55–??, March 1990. ISSN 0898-9788.

Colvin:1991:EHA

[Col91] Gregory Colvin. Exception handling in ANSI C. C Users Journal, 9(8):77–??, August 1991. ISSN 0898-9788.

Colvin:1993:ECO

[Col93a] Gregory Colvin. Extending C for object-oriented programming. C Users Journal, 11(7):51–??, July 1993. ISSN 0898-9788.

Colvin:1993:HM

[Col93b] Gregory Colvin. The hazards of macros. C Users Journal, 11(7): 70–??, July 1993. ISSN 0898-9788.

Colvin:1994:ECE

[Col94] Gregory Colvin. Emulating C++ exception handling. C/C++ Users Journal, 12(12):51–??, December 1994. CODEN CCUJEX. ISSN 1075-2838.

Colton:1995:DCS

[Col95a] Paul Colton. Dynamic client/ server-based image processing. C/C++ Users Journal, 13(5):33– ??, May 1995. CODEN CCU-JEX. ISSN 1075-2838.

Colvin:1995:SPC

[Col95b] Gregory Colvin. Smart pointers for C++ garbage collection. C/C++ Users Journal, 13(12): 37–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Colner:1996:HPO

[Col96] Don Colner. HTML programming with objects. C/C++ Users Journal, 14(7):17–18, 21–24, 26– 28, 29–30, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Combellick:1990:UCG

[Com90] Paul Combellick. The UI2 code generator. C Users Journal, 8 (2):107–??, February 1990. ISSN 0898-9788.

Connor:1988:EX

[Con88] Marty Connor. Emacs for Xenix. C Users Journal, 6(6):84–??, June 1988. ISSN 0898-9788.

Connolly:1991:PMH

[Con91] Douglas Connolly. From PC to mainframe with HLLAPI. C Users Journal, 9(9):65–??, September 1991. ISSN 0898-9788.

Cook:1990:PTT

[Coo90] Russell Cook. Portable transaction tracking. C Users Journal, 8 (10):27–??, October 1990. ISSN 0898-9788.

Cook:1992:GSC

[Coo92a] James Cook. Getting single characters from the keyboard. C Users Journal, 10(3):126–??, March 1992. ISSN 0898-9788.

Cooper:1992:CR

[Coo92b] Russ Cooper. CUG362 — RMAXTask. C Users Journal, 10(5):121–??, May 1992. ISSN 0898-9788.

Cortesi:1988:PEO

[Cor88] David E. Cortesi. The programmer’s essential OS/2 handbook. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1988. ISBN 0-934375-82-8, 0-934375-89-5 (book + disk), 0-558-51000-1 (??invalid ISBN??) (disk). 709 pp. LCCN QA76.76.O63 C68 1988 Reserves. US$24.95 (book), US$39.95 (book + disk), US$25.00 (disk). See book review [Pro90].

Cornelius:1991:UIA

[Cor91] Paul A. Cornelius. Using 1KHz interrupts on AT clones. C Users Journal, 9(12):53–??, December 1991. ISSN 0898-9788.

Cornelison:1995:TWM

[Cor95] Mike Cornelison. Two wildcard matching utilities. C/C++ Users Journal, 13(4):55–??, April 1995. CODEN CCUJEX. ISSN 1075-2838.

Cox:1992:BES

[Cox92] Keith W. Cox. Building an embedded system. C Users Journal, 10(3):23–??, March 1992. ISSN 0898-9788.

Crabtree:1997:RDW

[Cra97] Chris Crabtree. A reusable directory walk class. C/C++ Users Journal, 15(5):57–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Crews:1990:MMP

[Cre90] George Crews. MEL: A metalanguage processor. C Users Journal, 8(2):33–??, February 1990. ISSN 0898-9788.

Crosser:1989:H

[Cro89]  Kim C. Crosser.  Handles.

C Users Journal, 7(2):113–??, February 1989. ISSN 0898-9788.

Crowther:1990:CA

[Cro90] Willie Crowther. CUG159 — adventure. C Users Journal, 8(11): 125–??, November 1990. ISSN 0898-9788.

Crowe:1998:IOM

[Cro98a] William Crowe. Inter-object messaging using a SwitchBoard. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Crowe:1998:PTC

[Cro98b] William L. Crowe. A pseudoterminal class for unix. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Crowe:1999:CCJ

[Cro99] William L. Crowe. Comparing C++ and Java: A Java Switch-Board implementation. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Cruse:1988:CAC

[Cru88a] James A. Cruse. CUG274 — arrays for C. C Users Journal, 6 (9):91–??, September 1988. ISSN 0898-9788.

Cruse:1988:NMP

[Cru88b] James A. Cruse. New macro package supports array processing. C Users Journal, 6(9):97– ??, September 1988. ISSN 0898-9788.

Cruse:1989:MTM

[Cru89] James A. Cruse. Microsoft’s make. C Users Journal, 7(5):128– ??, May 1989. ISSN 0898-9788.

Curran:1990:CMT

[Cur90] James M. Curran. CUG321 mouse trap library. C Users Journal, 8(8):111–??, August 1990. ISSN 0898-9788.

Curran:1998:CWI

[Cur98] James M. Curran. Creating a word iterator. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Custer:1993:IWN

[Cus93] Helen Custer. Inside Windows NT. Microsoft Press, Bellevue, WA, USA, 1993. ISBN 1-55615-481-X. xxiv + 385 pp. LCCN QA76.76.O63 C89 1993. US$24.95, CDN$32.95. See book review [Hal93d].

Culp:1989:COS

[CW89] Gary Culp and Jonathan Ward.

CUG289 — Othello source code. C Users Journal, 7(7):119–??, July 1989. ISSN 0898-9788.

Daconta:1995:CPD

[Dac95]  Michael C. Daconta.  C++

pointers and dynamic memory management. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-04998-0 (paper/disk). xxvii + 446 pp. LCCN QA76.73.C153 D33 1995 Accompanying disk shelved in Reserves. See book review [All95c].

Daniels:1993:HAC

[Dan93] Scott Daniels. Hayes AT command set information sources. C Users Journal, 11(12):130–??, December 1993. ISSN 0898-9788.

Danilko:1997:NPU

[Dan97] Brian Danilko. Named pipes under Windows 95. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Danilko:1998:GJF

[Dan98] Brian Danilko. A graphical Java front end to C++ programs. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCU-JEX. ISSN 1075-2838.

Davis:1992:TC

[Dav92] Wilbon Davis. Time complexity. C Users Journal, 10(9):29– ??, September 1992. ISSN 0898-9788.

Davis:1993:CTE

[Dav93]  Frank Davis.  CUG386 —

Thomson-Davis editor. C Users Journal, 11(11):120–??, November 1993. ISSN 0898-9788.

Dawes:1995:YCC

[Daw95] Beman Dawes. You can count on it. C/C++ Users Journal, 13(7): 23–??, July 1995. CODEN CCU-JEX. ISSN 1075-2838.

deChamplain:1990:NPM

[dC90a] Michel de Champlain. A nonpreemptive multitasking executive in C++. C Users Journal, 8 (11):63–??, November 1990. ISSN 0898-9788.

deChamplain:1990:OBR

[dC90b] Michel de Champlain. An objectbased real-time executive. C Users Journal, 8(12):73–??, December 1990. ISSN 0898-9788.

Dare-Edwards:1998:CEP

[DE98] Conrad Dare-Edwards. Contrast enhancement with piecewise lookup tables. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Dean:1990:CPC

[Dea90] Jeffrey S. Dean. CUG298 — PC curses. C Users Journal, 8 (1):123–??, January 1990. ISSN 0898-9788.

Dean:1993:BIT

[Dea93] Christopher Dean. Bitmap image transformations. C Users Journal, 11(12):49–??, December 1993. ISSN 0898-9788.

Deen:1995:RAO

[Dee95] Glen W. Deen. Reducing allocation overhead in 2-D arrays. C/C++ Users Journal, 13(12): 25–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Deen:1996:SDC

[Dee96]  Glen Deen.  Selectable de-

fault constructor arguments.

C/C++ Users Journal, 14(2): 57–??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Degani:1997:EPU

[Deg97] Adi Degani. Easy printing under Windows — no graphics allowed. C/C++ Users Journal, 15(3):68– ??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Deloria:1988:ADM

[Del88a] Dennis Deloria. Automating the design and manufacture of parsing and translation machinery. C Users Journal, 6(3):25–??, March 1988. ISSN 0898-9788.

Deloria:1988:CGX

[Del88b] Dennis Deloria. A complete grammar for the X-Ref utility. C Users Journal, 6(6):17–??, June 1988. ISSN 0898-9788.

Deloria:1988:CRC

[Del88c] Dennis Deloria. Complex revision control systems prove challenging. C Users Journal, 6(2):31–??, February 1988. ISSN 0898-9788.

Deloria:1988:MXG

[Del88d] Dennis Deloria. Making ‘xref’ generate useful output. C Users Journal, 6(9):63–??, September 1988. ISSN 0898-9788.

Deloria:1988:PYB

[Del88e] Dennis Deloria. Practical yacc — building A simple cross-reference utility using this powerful. C Users Journal, 6(1):44–??, January 1988. ISSN 0898-9788.

Deluca:1988:STO

[Del88f] Thomas A. Deluca. Simple text output on an Apple IIgs. C Users Journal, 6(11):83–??, November 1988. ISSN 0898-9788.

Derr:1988:CTC

[Der88] James Derr. CUG273 — Turbo C utilities. C Users Journal, 6 (9):91–??, September 1988. ISSN 0898-9788.

Deutsch:1993:CG

[Deu93] L. Peter Deutsch. CUG384 — ghostscript. C Users Journal, 11 (10):112–??, October 1993. ISSN 0898-9788.

deFeraudy:1990:CSP

[dF90a] Henri de Feraudy. CUG297 — small prolog. C Users Journal, 8(1):123–??, January 1990. ISSN 0898-9788.

deFeraudy:1990:CLS

[dF90b] Henri de Feraudy. CUG310 — little Smalltalk for MS-DOS. C Users Journal, 8(5):123–??, May 1990. ISSN 0898-9788.

deFeraudy:1990:CUT

[dFAAW90] Henri de Feraudy, Robert Artigas, Jr., Arkin Asaf, and Martin D. Winnick. CUG329 UNIX tools for PC. C Users Journal, 8(11):126–??, November 1990. ISSN 0898-9788.

Dickey:1994:PCC

[Dic94] Bruce Dickey. Problems in C and C++ and their development tools. C Users Journal, 12(3): 137–??, March 1994. ISSN 0898-

Disque:1988:SCT

[Dis88] Tom Disque. Simple compression technique works with text and graphics. C Users Journal, 6(8): 23–??, August 1988. ISSN 0898-9788.

Darnell:1988:SEC

[DM88] Peter A. Darnell and Philip E. Margolis. Software engineering in C. Springer books on professional computing. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96574-2 (paperback). xx + 612 pp. LCCN QA76.76.D47 D371 1988. See book review [Wal88].

deMagalhaes:1994:CAA

[dM94] Joao C. de Magalhaes. Calendar algorithms and algorithm sources. C Users Journal, 12 (1):116–??, January 1994. ISSN 0898-9788.

Dorfman:1993:CMM

[DN93] Len Dorfman and Marc J. Neuberger. C memory management techniques. The Len Dorfman practical programming series. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1993. ISBN 0-8306-4058-4 (paperback). xii + 304 pp. LCCN QA 76.73 C15 D673 1993. US$32.95. See book review [Pla94b]. System requirements for computer disk: IBM PC; DOS; C programming language (programs tested with Borland’s Turbo C++ and Microsoft’s C 6.0 and 7.0 compilers); hard disk recommended.

Dodrill:1988:CCCa

[Dod88a] Gordon Dodrill. CUG252 and CUG 253 — C tutor. C Users Journal, 6(1):53–??, January 1988. ISSN 0898-9788.

Dodrill:1988:CCCb

[Dod88b] Gordon Dodrill. CUG257 and CUG258 — C tutor for Turbo C. C Users Journal, 6(2):57–??, February 1988. ISSN 0898-9788.

Dodrill:1991:CCT

[Dod91] Gordon Dodrill. CUG353 — C++ tutor 2.0. C Users Journal, 9(10):131–??, October 1991. ISSN 0898-9788.

Doe:1997:HLM

[Doe97] Rodney B. Doe, P.E. How to leak memory in C++. C/C++ Users Journal, 15(3):69–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Dorfman:1995:CEO

[Dor95] Len Dorfman. C++ by example: object-oriented analysis, design and programming. McGraw-Hill, Inc., New York, NY, USA, 1995. ISBN 0-07-911954-9. xii + 300 pp. LCCN QA76.64 .D67 1995. US$39.95.

Doshi:1995:TAC

[Dos95] Nimish R. Doshi. Advanced C++ by Namir Clement Shammas. C/C++ Users Journal, 13 (1):59–??, January 1995. CO-DEN CCUJEX. ISSN 1075-2838. See [Sha92].

Downs:1996:EWC

[Dow96] Chris Downs. Encapsulating Windows controls. C/C++ Users Journal, 14(11):31–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Dragovich:1993:TFP

[DP93] Jeff Dragovich and Kenneth Pugh. Truncating a file in place and portability. C Users Journal, 11(1):106–??, January 1993. ISSN 0898-9788.

Duggan:1994:CP

[Dug94] Angus Duggan. CUG407 — PSUtils. C Users Journal, 12(6):97– ??, June 1994. ISSN 0898-9788.

Dugger:1995:MC

[Dug95] Jim Dugger. Multithreading in C++. C/C++ Users Journal, 13 (11):23–??, November 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Duggan:1997:MFC

[Dug97] Tim Duggan. Member functions as callbacks in Win32. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Dwyer:1996:TRNa

[DW96a] Jerry Dwyer and K. B. Williams. Testing random number generators. C/C++ Users Journal, 14 (6):39–??, June 1996. CODEN CCUJEX. ISSN 1075-2838.

Dwyer:1996:TRNb

[DW96b] Jerry Dwyer and K. B. Williams. Testing random number generators, part 2. C/C++ Users Journal, 14(8):57–??, August 1996.

CODEN CCUJEX. ISSN 1075-2838.

Dwyer:1995:QPR

[Dwy95]  Jerry Dwyer.  Quick and

portable random number generators. C/C++ Users Journal, 13 (6):33–??, June 1995. CODEN CCUJEX. ISSN 1075-2838.

Eager:1988:LBF

[Eag88] Michael J. Eager. LEX bugs and fixes. C Users Journal, 6(1):65– ??, January 1988. ISSN 0898-9788.

Ebnet:1991:CC

[Ebn91] Josef Ebnet. CUG340 — C-Window. C Users Journal, 9(5): 135–??, May 1991. ISSN 0898-9788.

Eckel:1989:PRO

[Eck89] Bruce Eckel. Passing and returning objects in C++. C Users Journal, 7(8):33–??, August 1989. ISSN 0898-9788.

Edmunds:1998:TCC

[Edm98] Cyril Edmunds. Template classes for color-space conversions. C/C++ Users Journal, 16 (7):??, July 1998. CODEN CCU-JEX. ISSN 1075-2838.

Efken:1997:TYC

[Efk97] Charles Efken. Testing Year 2000 compliance on the PC. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Eiben:1994:TCP

[Eib94] Rob Eiben. C + C++: Programming With Objects in C and C++ by Allen I. Holub. C/C++ Users Journal, 12(12):57–??, December 1994. CODEN CCUJEX. ISSN 1075-2838. See [Hol92].

Ekwall:1992:CGT

[Ekw92] Jack E. Ekwall. CUG361 — gadgets and term. C Users Journal, 10(4):114–??, April 1992. ISSN 0898-9788.

Emerson:1988:SIT

[Eme88] Phillip L. Emerson. Separating interface from task modules produces more flexible serial interface. C Users Journal, 6(5):68– ??, May 1988. ISSN 0898-9788.

Engbert:1993:OHC

[Eng93a] Ludger Engbert. On-the-fly Huffman coding. C Users Journal, 11 (12):29–??, December 1993. ISSN 0898-9788.

English:1993:CBC

[Eng93b] John English. CUG385 — BCC+ coroutines. C Users Journal, 11(11):119–??, November 1993. ISSN 0898-9788.

Erdelsky:1988:HWC

[Erd88] Philip J. Erdelsky. How to write conforming MS-DOS device drivers in C. C Users Journal, 6 (9):69–??, September 1988. ISSN 0898-9788.

Erdelsky:1989:EAS

[Erd89] Philip J. Erdelsky. An efficient algorithm for sorting linked lists. C Users Journal, 7(5):89–??, May 1989. ISSN 0898-9788.

Erdelsky:1991:SLI

[Erd91] Philip J. Erdelsky. A simple line input function. C Users Journal, 9(7):36–??, July 1991. ISSN 0898-9788.

Erdelsky:1993:STS

[Erd93] Philip J. Erdelsky. A safer setjmp in C++. C Users Journal, 11(1):41–??, January 1993. ISSN 0898-9788.

Erdelsky:1995:PBO

[Erd95] Philip J. Erdelsky. Portable byte ordering in C++. C/C++ Users Journal, 13(1):25–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Egan:1988:WUD

[ET88] Janet I. Egan and Thomas J.

Teixeira. Writing a UNIX device driver. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1988. ISBN 0-471-62859-X (paperback), 0-471-62811-5. viii + 357 pp. LCCN QA76.76.O63 E35 1988.

Egan:1992:WUD

[ET92] Janet I. Egan and Thomas J.

Teixeira. Writing a UNIX device driver. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 1992. ISBN 0-471-53575-3 (cloth), 0-471-53574-5 (paper). ix + 358 pp. LCCN QA76.76.D49 E42 1992. See book review [Wei93b].

Ellis:1998:REC

[EU98] Duncan Ellis and Sameer Udeshi. A regular expression class library. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Evenden:1993:THS

[Eve93] Gerald I. Evenden. hypot as serious omission from Standard C math library. C Users Journal, 11(12):127–??, December 1993. ISSN 0898-9788.

Enright:1995:ASA

[EW95] Aaron Garth Enright and Linda M. Wilkens. Another slap at lowly goto. C/C++ Users Journal, 13 (12):49–??, December 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Faler:1990:CCI

[Fal90a] Wesley Faler. CUG320 — convolution image process. C Users Journal, 8(8):111–??, August 1990. ISSN 0898-9788.

Faler:1990:IMC

[Fal90b] Wesley Faler. Image manipulation by convolution. C Users Journal, 8(8):95–??, August 1990. ISSN 0898-9788.

Fandrich:1994:CS

[Fan94] Daniel Fandrich. CUG408 — SNews. C Users Journal, 12(6): 97–??, June 1994. ISSN 0898-9788.

Fan:1999:IRD

[Fan99] James J. Fan. Implementing a request distribution server with DCOM and ATL. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Farrell:1988:TSC

[Far88] Paul F. Farrell. Solutions in C, by Rex Jaeschke. C Users Journal, 6(2):53–??, February 1988. ISSN 0898-9788. See [Jae86].

Farah:1989:DHP

[Far89] Avraham Farah. Debugging heap problems. C Users Journal, 7 (10):105–??, October 1989. ISSN 0898-9788.

Fazio:1998:SES

[Faz98] Bob Fazio. Sending email with an SMTP class. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Flacy:1996:SCG

[FB96] Mark Flacy and Larry Brunelle.

A small combination generator. C/C++ Users Journal, 14(3):43– ??, March 1996. CODEN CCU-JEX. ISSN 1075-2838.

Foundation:1992:CGC

[FD92] Free Software Foundation and D. J. Delorie. CUG359 — GNU C/C++ for 386. C Users Journal, 10(3):117–??, March 1992. ISSN 0898-9788.

Feather:1995:UCS

[Fea95a] Clive D. W. Feather. Understanding the C standard. C/C++ Users Journal, 13(5):69–??, May 1995. CODEN CCUJEX. ISSN 1075-2838.

Feather:1995:USL

[Fea95b] Clive D. W. Feather. Understanding the standard — linkage. C/C++ Users Journal, 13(7):71– ??, July 1995. CODEN CCU-JEX. ISSN 1075-2838.

Feeney:1993:RDA

[Fee93] James A. Feeney. Real-time data acquisition and control under Windows 3.1. C Users Journal, 11(11):19–??, November 1993. ISSN 0898-9788.

Felice:1990:ICC

[Fel90] Robert Felice. Implementing the CCITT cyclic redundancy check. C Users Journal, 8(9):61– ??, September 1990. ISSN 0898-9788.

Ferreira:1988:SPB

[Fer88] Pete Ferreira. Screen Painter, from Baysoft. C Users Journal, 6(1):38–??, January 1988. ISSN 0898-9788.

Feuer:1986:CT

[Feu86] Alan R. Feuer. The C trainer. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1986. ISBN 0-13-109752-0 (paperback), 0-13-109745-8 (paperback). viii + 268 pp. LCCN QA76.73.C15 F484 1986. See book review [Rab89].

Fiedler:1988:DDW

[Fie88a] David Fiedler. Dialing for data, or where to find good C programs. C Users Journal, 6(9):93– ??, September 1988. ISSN 0898-9788.

Fiedler:1988:FF

[Fie88b] David Fiedler. A fond farewell and …. C Users Journal, 6(1): 6–??, January 1988. ISSN 0898-9788.

Fiedler:1988:MBF

[Fie88c] David Fiedler. More and better free software. C Users Journal, 6 (11):91–??, November 1988. ISSN 0898-9788.

Fiedler:1988:NSX

[Fie88d] David Fiedler. New SCO Xenix System V is a capable performer. C Users Journal, 6(3): 41–??, March 1988. ISSN 0898-9788.

Fiedler:1988:NTU

[Fie88e] David Fiedler. Now there’s ‘uucp’ for MS-DOS! C Users Journal, 6 (6):83–??, June 1988. ISSN 0898-9788.

Fiedler:1988:UIE

[Fie88f] David Fiedler. USENET: An informal but extensive CommNet for UNIX and XENIX systems and users. C Users Journal, 6(1): 54–??, January 1988. ISSN 0898-9788.

Fiedler:1988:UGH

[Fie88g] David Fiedler. Usenet’s greatest hits — the best of ‘comp. sources’. C Users Journal, 6 (2):58–??, February 1988. ISSN 0898-9788.

Fiedler:1989:BKS

[Fie89a] David Fiedler. ‘bigtex,’ Kanji support, and X-Windows. C Users Journal, 7(2):107–??, February 1989. ISSN 0898-9788.

Fiedler:1989:CSN

[Fie89b] David Fiedler. Calculators, shells and news. C Users Journal, 7(11):122–??, November 1989. ISSN 0898-9788.

Fiedler:1989:HGS

[Fie89c] David Fiedler. How to get source from The GNU Project. C Users Journal, 7(1):109–??, January 1989. ISSN 0898-9788.

Fiedler:1989:NEM

[Fie89d] David Fiedler. New elm, more utilities. C Users Journal, 7(8): 123–??, August 1989. ISSN 0898-9788.

Fiedler:1989:SPU

[Fie89e] David Fiedler. Several PostScript utilities now available from USENET sites. C Users Journal, 7(5):121–??, May 1989. ISSN 0898-9788.

Fillinger:1988:CCC

[Fil88] Fred Fillinger. Commodore-64 C compiling. C Users Journal, 6 (2):13–??, February 1988. ISSN 0898-9788.

Fiore:1989:PIG

[Fio89] James Fiore. Programming with intuition gadgets. C Users Journal, 7(1):79–??, January 1989. ISSN 0898-9788.

Fischer:1990:PTI

[Fis90a] David W. Fischer. Problem with the IDE in Quick C 2.51. C Users Journal, 8(12):128–??, December 1990. ISSN 0898-9788.

Fischer:1990:SMC

[Fis90b] Paul Fischer. State machines in C. C Users Journal, 8(12):119– ??, December 1990. ISSN 0898-9788.

Flatters:1990:TUP

[Fla90] Chris Flatters. UNIX Programming: Methods and Tools, by James F. Peters III. C Users Journal, 8(11):119–??, November 1990. ISSN 0898-9788. See [Pet88].

Florence:1989:CHP

[Flo89] Ronald Florence. CUG277 — HP plotter library and xmodem. C Users Journal, 7(1):101–??, January 1989. ISSN 0898-9788.

Florence:1990:UTF

[Flo90] Ronald Florence. UNIX ‘termcap’ facility improves portability by hiding terminal dependencies. C Users Journal, 8(1):93–??, January 1990. ISSN 0898-9788.

Florentin:1998:CSP

[Flo98] Alberto Florentin. Creating simple persistent objects. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Fugelso:1991:RDA

[FM91] David Fugelso and Michael Michnovicz. Real-time data acquisition. C Users Journal, 9(3):48– ??, March 1991. ISSN 0898-9788.

Fong:1997:BAC

[Fon97] Earl Fong. Being assertive in C/C++. C/C++ Users Journal, 15(6):29–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Forcier:1990:CB

[For90] Darren Forcier. Code base 4. C Users Journal, 8(4):109–??, April 1990. ISSN 0898-9788.

Forkosh:1991:MAL

[For91] John Forkosh. More accurate linear interpolation. C Users Journal, 9(5):77–??, May 1991. ISSN 0898-9788.

Fosner:1993:FNC

[Fos93] Ron Fosner. A fixed-point numeric class for C++. C Users Journal, 11(9):73–??, September 1993. ISSN 0898-9788.

Foxley:1985:US

[Fox85] Eric Foxley. UNIX for superusers. International computer science series. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-14228-7. xiv + 213 pp. LCCN QA76.76.O63 F69 1985 Bar. UK£11.95. See book review [Jon88a].

Frazier:1988:LCV

[Fra88] R. A. Frazier. Let’s C (ver. 4.0.6), from mark williams company. C Users Journal, 6(1):67–??, January 1988. ISSN 0898-9788.

Frazier:1996:ET

[Fra96] George F. Frazier. Estimating time-to-transfer. C/C++ Users Journal, 14(3):25–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Freeman:1988:COB

[Fre88a] Neil Freeman. C-INDEX+ offers B-Tree file management to MS-DOS applications programmers. C Users Journal, 6(2):17– ??, February 1988. ISSN 0898-9788.

Freeman:1988:VCC

[Fre88b] Neil Freeman. Vitamin C: A comprehensive screen handling library. C Users Journal, 6(5):57– ??, May 1988. ISSN 0898-9788.

Freda:1990:UIC

[Fre90] William J. Freda. UNIX interprocess communications. C Users Journal, 8(11):49–??, November 1990. ISSN 0898-9788.

Freed:1993:GSP

[Fre93] Adrian Freed. Guidelines for signal processing applications in C. C Users Journal, 11(9):85– ??, September 1993. ISSN 0898-9788.

Friend:1990:WPT

[Fri90] Tom Friend. Writing portable TIGA code. C Users Journal, 8(8):55–??, August 1990. ISSN 0898-9788.

Frost:1989:BIO

[Fro89a] Lyle Frost. A buffered input/ output library for structured files. C Users Journal, 7(10):83–??, October 1989. ISSN 0898-9788.

Frost:1989:CTB

[Fro89b] Lyle Frost. CUG295 — blkio library. C Users Journal, 7(11):

117–??, November 1989. ISSN 0898-9788.

Frost:1990:UFS

[Fro90]  Lyle Frost.  Using files as

semaphores. C Users Journal, 8 (4):83–??, April 1990. ISSN 0898-9788.

Frost:1992:CFL

[Fro92] Lyle Frost. CUG358 — four libraries for a multiuser C database file management system. C Users Journal, 10(1):137–??, January 1992. ISSN 0898-9788.

Fruit:1990:PNG

[Fru90] Robert Fruit. A pseudo-random number generator. C Users Journal, 8(5):83–??, May 1990. ISSN 0898-9788.

Fruit:1992:AHL

[Fru92] Robert Fruit. Application hardware locks. C Users Journal, 10 (6):23–??, June 1992. ISSN 0898-9788.

Fukuda:1989:CSC

[Fuk89] Shigeo Fukuda. CUG223 — small C for CP/M. C Users Journal, 7(4):108–??, April 1989. ISSN 0898-9788.

Gabrilovich:1999:CDO

[Gab99] Evgeniy Gabrilovich. Controlling the destruction order of singleton objects. C/C++ Users Journal, 17(10):??, October 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Gagliano:1993:CVF

[Gag93] Joseph V. Gagliano. CUG389 — VGA FontLib, MakeFont and DXF viewer. C Users Journal, 11(12):111–??, December 1993. ISSN 0898-9788.

Gage:1994:DI

[Gag94a] Philip Gage. Divided-difference interpolation. C Users Journal, 12(6):23–??, June 1994. ISSN 0898-9788.

Gage:1994:NAD

[Gag94b] Philip Gage. A new algorithm for data compression. C Users Journal, 12(2):23–??, February 1994. ISSN 0898-9788.

Gage:1997:RAD

[Gag97] Philip Gage. Random access data compression. C/C++ Users Journal, 15(9):23–??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Gailly:1993:CG

[Gai93] Jean-Loup Gailly. CUG382 — GZIP. C Users Journal, 11(9): 120–??, September 1993. ISSN 0898-9788.

Gamma:1995:DPE

[Gam95] Erich Gamma. Design patterns:

elements of reusable objectoriented software. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-63361-2. xv + 395 pp. LCCN QA76.64 .D47 1995. See book review [Pat95].

Ganssle:1992:APE

[Gan92] Jack G. Ganssle. The art of programming embedded systems. Academic Press Inc., New York, USA, 1992. ISBN 0-12-274880-8. xiii + 279 pp. LCCN QA76.6 .G334 1992. See book review [Gin93b].

Garrison:1988:EEM

[Gar88] Peter Garrison. Expeditionary equipment for Mandelbrot adventurers. C Users Journal, 6(6):75– ??, June 1988. ISSN 0898-9788.

Graubert-Cervone:1992:CWV

[GC92] Jeff Graubert-Cervone. CUG371 — WindosIO v2.0. C Users Journal, 10(12):121–??, December 1992. ISSN 0898-9788.

Green:1993:SSS

[GC93] Tom Green and Dennis Cronin. SSX — stack swap eXecutive. C Users Journal, 11(3):24–??, March 1993. ISSN 0898-9788.

Gerhards:1988:CPG

[Ger88] Rainer Gerhards. CUG227 — portable graphics library for MS-DOS. C Users Journal, 6(2):57– ??, February 1988. ISSN 0898-9788.

Gerhards:1990:UHF

[Ger90] Rainer Gerhards. Using header files to enhance portability. C Users Journal, 8(1):33–??, January 1990. ISSN 0898-9788.

Gessling:1991:CRE

[Ges91] Donald Gessling. Code readability enhancement using macros. C Users Journal, 9(1):135–??, January 1991. ISSN 0898-9788.

Getov:1999:PVI

[Get99] Radoslav Getov. Persistent vector iterators. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Gray:1999:DCC

[GH99] Robert S. Gray and Gísli Hjálmtysson. Dynamic C++ classes. C/C++ Users Journal, 17(10):??, October 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Gibson:1988:TFE

[Gib88] David Gibson. A tour of five editors available from the CUG library. C Users Journal, 6(1):59– ??, January 1988. ISSN 0898-9788.

Gibson:1990:RKS

[Gib90] Joseph W. Gibson. Reading the keyboard scan codes. C Users Journal, 8(8):86–??, August 1990. ISSN 0898-9788.

Gibbs:1999:YPL

[Gib99] Tanton Gibbs. A Y2K patch for legacy data. C/C++ Users Journal, 17(1):??, January 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Giel:1990:CW

[Gie90] William Giel. CUG324 — WG-CONIO. C Users Journal, 8(9): 130–??, September 1990. ISSN 0898-9788.

Gilson:1990:ICV

[Gil90] Michael J. Gilson. Interprocess communication: A VAX/VMS example. C Users Journal, 8 (11):70–??, November 1990. ISSN 0898-9788.

Gilson:1992:STM

[Gil92] Michael J. Gilson. A small make. C Users Journal, 10(6):73–??, June 1992. ISSN 0898-9788.

Gilhooly:1995:PEW

[Gil95a] Kevin Gilhooly. Pulse — estimating Windows response time. C/C++ Users Journal, 13(11): 69–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Gilhooly:1995:WWS

[Gil95b] Kevin Gilhooly. WinJES — A Windows sockets example. C/C++ Users Journal, 13(5):23– ??, May 1995. CODEN CCU-JEX. ISSN 1075-2838.

Gimpel:1989:LIP

[Gim89] James F. Gimpel. ‘lint’ identifies potential portability problems. C Users Journal, 7(1):51–??, January 1989. ISSN 0898-9788.

Gingrich:1993:TCR

[Gin93a] Mark Gingrich. μC/OS, the realtime kernel, by Jean J. Labrosse. C Users Journal, 11(11):87–??, November 1993. ISSN 0898-9788. See [Lab92].

Gingrich:1993:TAP

[Gin93b] Mark Gingrich. The Art of Programming Embedded Systems, by Jack G. Ganssle. C Users Journal, 11(3):113–??, March 1993. ISSN 0898-9788. See [Gan92].

Gingrich:1993:RAR

[Gin93c] Mark Gingrich. Real-number approximation for real programmers. C Users Journal, 11(2):35– ??, February 1993. ISSN 0898-9788.

Guarna:1990:UIL

[GK90] Vincent Guarna and James Krause. User interface language eases prototyping. C Users Journal, 8(2):17–??, February 1990. ISSN 0898-9788.

Gleason:1990:ISR

[Gle90a] B. J. Gleason. An introduction to speech recognition. C Users Journal, 8(3):83–??, March 1990. ISSN 0898-9788.

Glennon:1990:DHA

[Gle90b] Bryan Glennon. Data hiding and abstraction. C Users Journal, 8(6):101–??, June 1990. ISSN 0898-9788.

Glennon:1990:ICS

[Gle90c] Bryan Glennon. Inter-process communications with sockets. C Users Journal, 8(12):111–??, December 1990. ISSN 0898-9788.

Goodwin:1989:UIC

[Goo89] Mark D. (Mark Dennis) Goodwin. User interfaces in C++ and object-oriented programming. Advanced computer books. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel:

(503) 282-5215, 1989. ISBN 1-55828-023-5. xiv + 394 pp. LCCN QA76.73.C153 G66 1989. US$26.95 (book), US$53.95 (book + disk). See book review [Cla90c].

Gotwals:1994:MIA

[Got94] John K. Gotwals. Multi-precision integer arithmetic using C++. C Users Journal, 12(6):35–??, June 1994. ISSN 0898-9788.

Gottner:1995:TCB

[Got95] Dave Gottner. Templates without code bloat. C/C++ Users Journal, 13(8):57–??, August 1995. CODEN CCUJEX. ISSN 1075-2838.

Gould:1998:DDI

[Gou98] David Gould. Double dispatch with an inverted visitor pattern. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Gehani:1989:CCP

[GR89] Narain Gehani and William D. Roome. The Concurrent C programming language. Silicon Press, Summit, NJ, USA, 1989. ISBN 0-929306-00-7. xv + 303 pp. LCCN QA76.73.C15 G46 1989 Bar. See book review [Pla90c].

Graham:1988:PHB

[Gra88] Ken Graham. Pssssttt! hey buddy, you wanta buy UNIX source for $89? C Users Journal, 6(6):43–??, June 1988. ISSN 0898-9788.

Graham:1989:TUS

[Gra89] Ken Graham. UNIX System Programming, by Keith Haviland and Ben Salama. C Users Journal, 7(10):36–??, October 1989. ISSN 0898-9788.

Graham:1991:SCR

[Gra91] Ken Graham. Screen capture routines: Some sources. C Users Journal, 9(6):121–??, June 1991. ISSN 0898-9788.

Graham:1992:ERW

[Gra92a] Steven K. Graham. An easy road to windows? C Users Journal, 10(12):123–??, December 1992. ISSN 0898-9788.

Graham:1992:EA

[Gra92b] Steven K. Graham. Evolution in action. C Users Journal, 10 (11):75–??, November 1992. ISSN 0898-9788.

Graham:1992:FCF

[Gra92c] Steven K. Graham. Fundamentals of curve fitting. C Users Journal, 10(2):104–??, February 1992. ISSN 0898-9788.

Gray:1992:ITT

[Gra92d] W. Harvey Gray. Interpreting touchscreen touches. C Users Journal, 10(8):88–??, August 1992. ISSN 0898-9788.

Graham:1993:CT

[Gra93a] Steven K. Graham. CUG375 — textview. C Users Journal, 11 (1):121–??, January 1993. ISSN 0898-9788.

Graham:1993:COM

[Gra93b] Steven K. Graham. CUG376 — OS/2 and MS-DOS tools. C Users Journal, 11(1):121–??, January 1993. ISSN 0898-9788.

Graham:1993:CDF

[Gra93c] Steven K. Graham. CUG377 — DSR functions. C Users Journal, 11(1):122–??, January 1993. ISSN 0898-9788.

Graham:1993:CN

[Gra93d] Steven K. Graham. CUG378 — NEWMAT. C Users Journal, 11 (1):122–??, January 1993. ISSN 0898-9788.

Graham:1993:CZ

[Gra93e] Steven K. Graham. CUG379 — ZOO. C Users Journal, 11 (1):122–??, January 1993. ISSN 0898-9788.

Graham:1993:MFR

[Gra93f] Steven K. Graham. Mapping functions for repetitive structures. C Users Journal, 11(4):38– ??, April 1993. ISSN 0898-9788.

Grant:1995:IGA

[Gra95]  Keith Grant.  An introduc-

tion to genetic algorithms. C/C++ Users Journal, 13(3):45– ??, March 1995. CODEN CCU-JEX. ISSN 1075-2838.

Green:1989:RGA

[Gre89] T. M. Green. References for GO algorithms. C Users Journal, 7 (4):111–??, April 1989. ISSN 0898-9788.

Greissman:1994:SMT

[Gre94] Adam Greissman. Structure mapping techniques in C++. C/C++ Users Journal, 12(7):51– ??, July 1994. CODEN CCU-JEX. ISSN 1075-2838.

Grogan:1991:SSL

[Gro91] Kenneth L. Grogan, Jr. Searching with skip lists. C Users Journal, 9(2):71–??, February 1991. ISSN 0898-9788.

Gruel:1990:CKB

[Gru90] Steven Gruel. Controlling the keyboard buffer. C Users Journal, 8(7):85–??, July 1990. ISSN 0898-9788.

Giglio:1995:SED

[GS95] Paul Giglio and David Schwartz.

Simplifying explicit dynamic linking. C/C++ Users Journal, 13(9):43–??, September 1995. CODEN CCUJEX. ISSN 1075-2838.

Guarna:1990:XWS

[Gua90] Vincent Guarna. X Window System Series (8 volume set). C Users Journal, 8(12):109–??, December 1990. ISSN 0898-9788.

Gufford:1998:WNE

[Guf98] Eric Gufford. A Windows NT exception class library. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Guinan:1997:PIC

[Gui97]  Jamie Guinan.  Platform-

independent C++ GUI toolkits. C/C++ Users Journal, 15(1):19– ??, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Guthrie:1994:TTT

[Gut94] R. Scott Guthrie. A text translation tool for C programmers.

C/C++ Users Journal, 12(12): 33–??, December 1994. CODEN CCUJEX. ISSN 1075-2838.

Guthrie:1995:SEA

[Gut95] R. Scott Guthrie. A simple, easy AutoQueue class. C/C++ Users Journal, 13(7):55–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Gilson:1991:GSR

[GW91] Michael J. Gilson and Matt Weisfeld. A generic search routine. C Users Journal, 9(10):87–??, October 1991. ISSN 0898-9788.

Hack:1991:CU

[Hac91] Kevin Hack. CUG351 — Ultra-Win. C Users Journal, 9(9):126– ??, September 1991. ISSN 0898-9788.

Hagan:1991:RMI

[Hag91] Girish T. Hagan. Resampling methods for image manipulation. C Users Journal, 9(8):53–??, August 1991. ISSN 0898-9788.

Haldenrain:1989:ASR

[Hal89] Arthur M. H. Haldenrain. Atari ST reads MS-DOS disks. C Users Journal, 7(7):120–??, July 1989. ISSN 0898-9788.

Halbert:1991:ERM

[Hal91] Joel Halbert. Embedded realtime multitasking kernel. C Users Journal, 9(3):33–??, March 1991. ISSN 0898-9788.

Halladay:1992:TME

[Hal92] Steve Halladay. malloc-related errors. C Users Journal, 10(5):

72–??, May 1992. ISSN 0898-9788.

Hale:1993:TWS

[Hal93a] A. Donnie Hale, Jr. Writing Solid Code, by Steve Maguire. C Users Journal, 11(12):121–??, December 1993. ISSN 0898-9788. See [Mag93].

Halladay:1993:TCC

[Hal93b] Steve Halladay. C++ Components and Algorithms, by Scott Robert Ladd. C Users Journal, 11(2):121–??, February 1993. ISSN 0898-9788. See [Lad92].

Halladay:1993:TCP

[Hal93c] Steve Halladay. C++ Programming Style, by Tom Cargill. C Users Journal, 11(4):117–??, April 1993. ISSN 0898-9788. See [Car92].

Halladay:1993:TIW

[Hal93d] Steve Halladay. Inside Windows NT, by Helen Custer. C Users Journal, 11(9):129–??, September 1993. ISSN 0898-9788. See [Cus93].

Hamilton:1998:BRS

[Ham98] Dave Hamilton. Book review:

The Software Project Survival Guide. C/C++ Users Journal, 16 (6):??, June 1998. CODEN CCU-JEX. ISSN 1075-2838.

Haney:1990:MDI

[Han90] Daniel R. Haney. Mouse detection and identification on a PC. C Users Journal, 8(7):95–??, July 1990. ISSN 0898-9788.

Hansberry:1994:NES

[Han94] Russell Hansberry. The need for embedded systems support in C++. C Users Journal, 12(2): 121–??, February 1994. ISSN 0898-9788.

Harel:1987:ASC

[Har87] David Harel. Algorithmics: the spirit of computing. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-19240-3 (paperback). xiv + 425 pp. LCCN QA76 .H28331 1987. UK£14.95.

Harris:1988:AA

[Har88a] Sandy Harris. AWK availability. C Users Journal, 6(3):12–??, March 1988. ISSN 0898-9788.

Harris:1988:SS

[Har88b] Sandy Harris. Sorting on a SUN. C Users Journal, 6(3):13– ??, March 1988. ISSN 0898-9788.

Harris:1990:CDP

[Har90] Ken Harris. CUG311 — DB package. C Users Journal, 8(5): 123–??, May 1990. ISSN 0898-9788.

Harmon:1994:EWF

[Har94]  Trevor Harmon.  Extending

the Windows file manager. C/C++ Users Journal, 12(9):37– ??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Hartmann:1997:RPC

[Har97] Steve Hartmann. A reusable PID control class. C/C++ Users Journal, 15(2):19–??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Harrington:1999:WMM

[Har99] John Harrington. Win32 multithreading made easy. C/C++ Users Journal, 17(8):48, 50–52, 54–56, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Hatch:1994:CWC

[Hat94] William E. Hatch. CUG410B — WLIST C++ doubly linked lists. C Users Journal, 12(6):100–??, June 1994. ISSN 0898-9788.

Hathcock:1996:PPM

[Hat96]  Arvel Hathcock.  PopMail!–

A POP3 mailer for Windows. C/C++ Users Journal, 14(5):25– ??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Havener:1989:CCM

[Hav89a] Charles D. Havener. A C to C++ migrator tool. C Users Journal, 7(10):65–??, October 1989. ISSN 0898-9788.

Havener:1989:CFA

[Hav89b] Charles D. Havener. ‘CONNECT FOUR’ with alpha-beta pruning. C Users Journal, 7(4):33–??, April 1989. ISSN 0898-9788.

Havener:1989:CCC

[Hav89c] Charles D. Havener. CUG296 — C to C++ migrator. C Users Journal, 7(11):117–??, November 1989. ISSN 0898-9788.

Havener:1990:PMO

[Hav90] Charles D. Havener. Pricing a meal: An object-oriented example in C++. C Users Journal, 8(8):47–??, August 1990. ISSN 0898-9788.

Havener:1993:OFP

[Hav93] Charles D. Havener. An objectoriented frog pond. C Users Journal, 11(5):52–??, May 1993. ISSN 0898-9788.

Hawkins:1988:TCM

[Haw88] James Hawkins. C Made Easy, by Herbert Schildt. C Users Journal, 6(1):33–??, January 1988. ISSN 0898-9788. See [Sch85].

Hawes:1997:PCD

[Haw97]  Jack Hawes.  The problem

with const data members. C/C++ Users Journal, 15(3):67– ??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Hazzah:1991:EIL

[Haz91] Ali Hazzah. Encapsulation, inheritance and late-binding in C++. C Users Journal, 9(7):63– ??, July 1991. ISSN 0898-9788.

Heinrich:1995:FCG

[HD95] Peter Heinrich and Nathan Dwyer. Fixed-point C for graphics applications. C/C++ Users Journal, 13(8):37–??, August 1995. CODEN CCUJEX. ISSN 1075-2838.

Heaton:1996:HSW

[Hea96] Jeff Heaton. High speed Win32 animation. C/C++ Users Journal, 14(8):47–??, August 1996. CODEN CCUJEX. ISSN 1075-2838.

Heaton:1997:MWR

[Hea97]  Jeff Heaton.  Macintosh-to-

Windows resource conversion. C/C++ Users Journal, 15(6):61– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Hedrickson:1990:SNC

[Hed90] William T. Hedrickson. The Stanford naming convention. C Users Journal, 8(5):140–??, May 1990. ISSN 0898-9788.

Hedquist:1991:CCC

[Hed91] Barry E. Hedquist. Compiler certification: Comments on the process. C Users Journal, 9(3):138– ??, March 1991. ISSN 0898-9788.

Hegeman:1991:SL

[Heg91] Frederick W. Hegeman. Skip lists. C Users Journal, 9(2):33– ??, February 1991. ISSN 0898-9788.

Hegeman:1992:AF

[Heg92] Frederick W. Hegeman. Arithmetic in factorial-base. C Users Journal, 10(2):73–??, February 1992. ISSN 0898-9788.

Hegeman:1993:SN

[Heg93] Frederick W. Hegeman. Sorting networks. C Users Journal, 11 (2):89–??, February 1993. ISSN 0898-9788.

Heintze:1994:IAB

[Hei94] Siegfried Heintze. Intuitive access to bit arrays. C Users Journal, 12 (2):67–??, February 1994. ISSN 0898-9788.

Heidel:1999:UAC

[Hei99] Greg Heidel. Using asynchronous calls in COM. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Held:1990:FRH

[Hel90] Arthur Held. Function returns: How to use them. C Users Journal, 8(10):93–??, October 1990. ISSN 0898-9788.

Held:1991:CMR

[Hel91] Gilbert Held. The complete modem reference. Wiley professional computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-52911-7. xii + 273 pp. LCCN TK5103 .H44 1991. US$26.95. See book review [War91a].

Held:1994:CMR

[Hel94] Gilbert Held. The complete modem reference: the technician’s guide to installation, testing, and trouble-free communications. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 1994. ISBN 0-471-00852-4 (paperback). xiv + 412 pp. LCCN TK5103.H44 1994. US$34.95.

Heller:1995:ECC

[Hel95]  Steve Heller.  Efficient C/

C++ programming: smaller, faster, better. AP Professional, Boston, MA, USA, second edition, 1995. ISBN 0-12-339095-8, 0-12-339096-6 (disk). xxi + 415 pp. LCCN QA76.73.C15 H45 1995. See book review [Rod95b].

Hemrajani:1996:JCC

[Hem96] Anil Hemrajani. The Java and C connection. C/C++ Users Journal, 14(9):43–46, 48–50, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Hendrix:1988:SCC

[Hen88] James E. Hendrix. A Small C compiler: language, usage, theory, and design. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1988. ISBN 0-934375-88-7 (book), 1-55851-007-9 (disk), 0-934375-97-6 (book + disk). 586 pp. LCCN QA76.73.S58H45 1988. US$23.95 (book), US$25.00 (disk), US$38.95 (book and disk). See book review [Vol89b].

Hershey:1994:ACH

[Her94] Charles Hershey. Archimedes C-51 hosted on an IBM-PC compatible. C Users Journal, 12(5):49– ??, May 1994. ISSN 0898-9788.

Hessling:1994:CHE

[Hes94] Mark Hessling. CUG414 — THE — The Hessling Editor. C/C++ Users Journal, 12(8):88–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Heyman:1996:BPW

[Hey96] Bill Heyman. Better pipes for Windows. C/C++ Users Journal, 14(7):49–??, July 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Hicks:1989:CMF

[Hic89] Donald Hicks. CP/M to MS-DOS file conversions. C Users Journal, 7(7):121–??, July 1989. ISSN 0898-9788.

Hino:1988:PCE

[Hin88a] Kenji Hino. Programmers can ‘enhance’ utilities by using ‘awk’ and ‘sed’ as script writers. C Users Journal, 6(6):65–??, June 1988. ISSN 0898-9788.

Hino:1988:TSP

[Hin88b] Kenji Hino. Tool supports prototyping and ‘control panel’ interfaces. C Users Journal, 6(8):29– ??, August 1988. ISSN 0898-9788.

Hino:1990:CPT

[Hin90a] Kenji Hino. C programmer’s Toolbox/PC. C Users Journal, 8 (2):97–??, February 1990. ISSN 0898-9788.

Hino:1990:RCL

[Hin90b] Kenji Hino. Retrospective on the CUG library disks. C Users Journal, 8(2):127–??, February 1990. ISSN 0898-9788.

Halladay:1990:DES

[HJ90] Steve Halladay and Steve Johnson. Discrete event simulation in C for real-time systems. C Users Journal, 8(4):33–??, April 1990. ISSN 0898-9788.

Ho:1990:CMR

[HKB90] Dinghuei Ho, Michael Kelly, and Robert Briggs. CUG308 — MSU, REMZ, and LIST. C Users Journal, 8(4):123–??, April 1990. ISSN 0898-9788.

Huseby:1994:CLG

[HMF94] Sverre H. Huseby, George Matas, and Timothy M. Farnum. CUG393 — LL, GIFSave, and Cordic++. C Users Journal, 12 (1):88–??, January 1994. ISSN 0898-9788.

Hobbs:1991:TCC

[Hob91a] Sam Hobbs. Complete C++ Primer, The, by Keith Weiskamp and Bryan Flamig. C Users Journal, 9(4):96–??, April 1991. ISSN 0898-9788. See [WF90].

Hobbs:1991:TCP

[Hob91b] Sam Hobbs. Computers, Pattern, Chaos and Beauty, by Clifford A. Pickover. C Users Journal, 9(9): 121–??, September 1991. ISSN 0898-9788. See [Pic90].

Hoffman:1992:FAP

[Hof92] Sheldon Hoffman. Finding agents or publishers. C Users Journal, 10(1):138–??, January 1992. ISSN 0898-9788.

Hogaboom:1990:FDA

[Hog90] Richard Hogaboom. A flexible dynamic array allocator. C Users Journal, 8(11):114–??, November 1990. ISSN 0898-9788.

Hohl:1990:TAC

[Hoh90] Randy Hohl. Advanced C: Tips and Techniques, by Paul and Gail Anderson. C Users Journal, 8(4): 46–??, April 1990. ISSN 0898-9788.

Holub:1987:CCO

[Hol87] Allen I. Holub. C chest and other C treasures from Dr. Dobb’s Journal. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1987. ISBN 0-934375-40-2 (paperback). 536 pp. LCCN QA76.73.C15 C15 1987. US$24.95.

Holub:1992:CCP

[Hol92] Allen I. Holub. C + C++: programming with objects in C and C++. McGraw-Hill, Inc., New York, NY, USA, 1992. ISBN 0-07-029662-6. xiv + 427 pp. LCCN QA76.64.H65 1992. US$24.95. See book review [Eib94].

Hooker:1999:LRW

[Hoo99] David Hooker. A logging routine for Windows NT events. C/C++ Users Journal, 17(7):65–??, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Howe:1993:CAT

[How93] Anthony C. Howe. CUG388 — Anthony’s tools. C Users Journal, 11(11):121–??, November 1993. ISSN 0898-9788.

Haney:1989:CF

[HP89] Daniel R. Haney and Vern Paxson. CUG290 — FLEX. C Users Journal, 7(8):115–??, August 1989. ISSN 0898-9788.

Huss:1998:LSL

[HP98] R. E. Huss and M. A. Pumar. Lambert shading of large irregular surfaces. C/C++ Users Journal, 16(1):??, January 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Huss:1999:CME

[HPR99] Ronald E. Huss, Mark A. Pumar, and Robert L. Rudin. A concise method for evaluating derivatives. C/C++ Users Journal, 17 (11):??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Hansen:1990:NS

[HPS90] David Hansen, Kenneth Pugh, and David A. Schmitt. Notes on the * specifier. C Users Journal, 8 (7):95–??, July 1990. ISSN 0898-9788.

Harel:1992:ASC

[HR92] David Harel and Roni Rosner.

Algorithmics: the spirit of computing. Addison-Wesley, Reading, MA, USA, second edition, 1992. ISBN 0-201-50401-4. xv + 476 pp. LCCN QA76.9.A43 H37 1992. See book review [Rom91a].

Haviland:1987:USP

[HS87] Keith Haviland and Ben Salama.

UNIX system programming. International computer science series. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-12919-1 (paperback). xv + 339 pp. LCCN QA76.76.O63 H38 1987 Bar. UK£15.95.

Haney:1989:CBM

[HSC89] Daniel R. Haney, Richard Stallman, and Robert Corbett. CUG285 — BISON for MS-DOS. C Users Journal, 7(7):119–??, July 1989. ISSN 0898-9788.

Hughes:1990:CM

[Hug90] John J. Hughes III. CUG300 — MAT LIB. C Users Journal, 8 (2):125–??, February 1990. ISSN 0898-9788.

Hughes:1991:CTT

[Hug91] Bert C. Hughes. CUG347 — TAVL tree. C Users Journal, 9(7):126–??, July 1991. ISSN 0898-9788.

Humphrey:1995:DSE

[Hum95] Watts S. Humphrey. A discipline for software engineering. The SEI series in software engineering. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-54610-8. xxvi + 789 pp. LCCN QA76.758 .H857 1995. See book review [Phi96c].

Hunt:1985:CT

[Hun85] William James Hunt. The C toolbox. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-11111-X. xviii + 413 pp. LCCN QA76.73.C15 H85 1985. US$19.95. See book review [Bin88].

Hunt:1989:CTR

[Hun89] William James Hunt. The C toolbox: ready-to-run programs in Turbo C, Microsoft C, and Quick C. Addison-Wesley, Reading, MA, USA, second edition, 1989. ISBN 0-201-51815-5. xxi + 503 pp. LCCN QA 76.73 C15 H85 1989. US$22.95.

Halladay:1992:PUM

[HW92a] Steve Halladay and Michael Wiebel. A practical use for multiple threads. C Users Journal, 10(1):73–??, January 1992. ISSN 0898-9788.

Halladay:1992:UOT

[HW92b] Steve Halladay and Michael Wiebel. Using OOP techniques instead of switch in C++. C Users Journal, 10(10):105–??, October 1992. ISSN 0898-9788.

Ignatchenko:1998:AOD

[Ign98] Sergey Ignatchenko. An algorithm for online data compression. C/C++ Users Journal, 16 (10):63–??, October 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Iofin:1999:DCF

[Iof99] Leon Iofin. Date compression formats for Y2K remediation. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Irvine:1991:FVC

[Irv91] Ron Irvine. Fortran vs. C as language for numerical applications. C Users Journal, 9(11):132–??, November 1991. ISSN 0898-9788.

Isaacson:1998:BSE

[Isa98] Steven G. Isaacson. bsed: A stream editor for binary files. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCU-JEX. ISSN 1075-2838.

Jaeschke:1986:SCH

[Jae86] Rex Jaeschke. Solutions in C:

hundreds of programming tips by the author of “Doctor C’s pointers”. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-15042-5 (paperback). xii + 254 pp. LCCN QA76.73.C15 J34 1986 Bar. US$17.95. See book review [Far88].

Jaeschke:1988:COP

[Jae88a] Rex Jaeschke. The comma as operator and punctuator. C Users Journal, 6(8):75–??, August 1988. ISSN 0898-9788.

Jaeschke:1988:ESS

[Jae88b] Rex Jaeschke. Exploring the subtle side of the ‘sizeof’ operator. C Users Journal, 6(2):46–??, February 1988. ISSN 0898-9788.

Jaeschke:1988:CUT

[Jae88c]  Rex Jaeschke.  For certain

uses typedef is far superior to #define. C Users Journal, 6(3): 34–??, March 1988. ISSN 0898-9788.

Jaeschke:1988:IOI

[Jae88d] Rex Jaeschke. The ins and outs of initialization. C Users Journal, 6 (11):29–??, November 1988. ISSN 0898-9788.

Jaeschke:1988:LS

[Jae88e] Rex Jaeschke. Literal strings. C Users Journal, 6(6):48–??, June 1988. ISSN 0898-9788.

Jaeschke:1988:RSE

[Jae88f] Rex Jaeschke. The return statement: Email for functions. C Users Journal, 6(5):21–??, May 1988. ISSN 0898-9788.

Jaeschke:1988:NMN

[Jae88g] Rex Jaeschke. The ‘NULL’ macro and null pointers. C Users Journal, 6(9):29–??, September 1988. ISSN 0898-9788.

Jaeschke:1989:APB

[Jae89a] Rex Jaeschke. Arrays and pointers: A basic identity. C Users Journal, 7(1):69–??, January 1989. ISSN 0898-9788.

Jaeschke:1989:TCT

[Jae89b] Rex Jaeschke. The const type qualifier. C Users Journal, 7(4): 45–??, April 1989. ISSN 0898-9788.

Jaeschke:1989:CTF

[Jae89c] Rex Jaeschke. A crude trace facility. C Users Journal, 7(11):9–??, November 1989. ISSN 0898-9788.

Jaeschke:1989:FOP

[Jae89d] Rex Jaeschke. ‘FILE’ objects and pointers. C Users Journal, 7 (2):21–??, February 1989. ISSN 0898-9788.

Jaeschke:1989:LAB

[Jae89e] Rex Jaeschke. A look at bitfields. C Users Journal, 7(5):45– ??, May 1989. ISSN 0898-9788.

Jaeschke:1989:SP

[Jae89f] Rex Jaeschke. Sequence points.

C Users Journal, 7(7):51–??, July 1989. ISSN 0898-9788.

Jaeschke:1989:SPS

[Jae89g] Rex Jaeschke. stdarg.h: Portable support for access to variable length argument lists. C Users Journal, 7(1):33–??, January 1989. ISSN 0898-9788.

Jaeschke:1989:SCD

[Jae89h] Rex Jaeschke. stddef.h — common definitions. C Users Journal, 7(5):115–??, May 1989. ISSN 0898-9788.

Jaeschke:1989:SAH

[Jae89i] Rex Jaeschke. Structures, alignment and holes (or telling the hole truth). C Users Journal, 7(8):27–??, August 1989. ISSN 0898-9788.

Jaeschke:1989:SVL

[Jae89j] Rex Jaeschke. Subtle varieties of lvalues. C Users Journal, 7(10): 49–??, October 1989. ISSN 0898-9788.

Jaeschke:1990:AP

[Jae90a] Rex Jaeschke. Assertive programming. C Users Journal, 8 (7):23–??, July 1990. ISSN 0898-9788.

Jaeschke:1990:EC

[Jae90b] Rex Jaeschke. Environmental control. C Users Journal, 8(11): 25–??, November 1990. ISSN 0898-9788.

Jaeschke:1990:EHC

[Jae90c] Rex Jaeschke. Error handling in C. C Users Journal, 8(4):23–??, April 1990. ISSN 0898-9788.

Jaeschke:1990:TET

[Jae90d] Rex Jaeschke. The exit and abort functions. C Users Journal, 8(6):23–??, June 1990. ISSN 0898-9788.

Jaeschke:1990:HDM

[Jae90e] Rex Jaeschke. Header design and management. C Users Journal, 8 (2):59–??, February 1990. ISSN 0898-9788.

Jaeschke:1990:MML

[Jae90f] Rex Jaeschke. The memory management library. C Users Journal, 8(1):21–??, January 1990. ISSN 0898-9788.

Jaeschke:1990:OPT

[Jae90g] Rex Jaeschke. Operators and the precedence table. C Users Journal, 8(8):23–??, August 1990. ISSN 0898-9788.

Jaeschke:1990:PA

[Jae90h] Rex Jaeschke. Pointers to arrays. C Users Journal, 8(5):25– ??, May 1990. ISSN 0898-9788.

Jaeschke:1990:PPa

[Jae90i] Rex Jaeschke. Puzzles — part 1. C Users Journal, 8(10):23–??, October 1990. ISSN 0898-9788.

Jaeschke:1990:PPb

[Jae90j] Rex Jaeschke. Puzzles — part 2. C Users Journal, 8(12):18–??, December 1990. ISSN 0898-9788.

Jaeschke:1990:UQF

[Jae90k] Rex Jaeschke. Using the Quicksort function. C Users Journal, 8 (9):23–??, September 1990. ISSN 0898-9788.

Jaeschke:1990:VPJ

[Jae90l] Rex Jaeschke. Void pointers, jump tables, and friends. C Users Journal, 8(3):17–??, March 1990. ISSN 0898-9788.

Jaeschke:1991:DSPb

[Jae91a] Rex Jaeschke. Data structures — part 2: Arrays of pointers vs. multi-dimensional arrays. C Users Journal, 9(5):20–??, May 1991. ISSN 0898-9788.

Jaeschke:1991:DSPc

[Jae91b] Rex Jaeschke. Data structures — part 3: Arrays of pointers vs. multi-dimensional arrays. C Users Journal, 9(6):17–??, June 1991. ISSN 0898-9788.

Jaeschke:1991:DSPd

[Jae91c] Rex Jaeschke. Data structures — part 4: Singly-linked lists. C Users Journal, 9(7):17–??, July 1991. ISSN 0898-9788.

Jaeschke:1991:DSPe

[Jae91d] Rex Jaeschke. Data structures — part 5: Singly-linked lists. C Users Journal, 9(8):21–??, August 1991. ISSN 0898-9788.

Jaeschke:1991:DSPf

[Jae91e]  Rex Jaeschke.  Data struc-

tures — part 6: Doubly-linked lists. C Users Journal, 9(9):17– ??, September 1991. ISSN 0898-9788.

Jaeschke:1991:DSPg

[Jae91f] Rex Jaeschke. Data structures — part 7: Doubly-linked lists. C Users Journal, 9(10):18–??, October 1991. ISSN 0898-9788.

Jaeschke:1991:DSPh

[Jae91g] Rex Jaeschke. Data structures — part 8: Circular lists. C Users Journal, 9(11):21–??, November 1991. ISSN 0898-9788.

Jaeschke:1991:DSPi

[Jae91h] Rex Jaeschke. Data structures — part 8: Circular lists. C Users Journal, 9(12):18–??, December 1991. ISSN 0898-9788.

Jaeschke:1991:DSPa

[Jae91i] Rex Jaeschke. Data structures — part I: Arrays. C Users Journal, 9 (4):21–??, April 1991. ISSN 0898-9788.

Jaeschke:1991:PPa

[Jae91j] Rex Jaeschke. Puzzles — part 3. C Users Journal, 9(1):23–??, January 1991. ISSN 0898-9788.

Jaeschke:1991:PPb

[Jae91k] Rex Jaeschke. Puzzles — part 4. C Users Journal, 9(2):17–??, February 1991. ISSN 0898-9788.

Jaeschke:1991:PPc

[Jae91l] Rex Jaeschke. Puzzles — part 5. C Users Journal, 9(3):23–??, March 1991. ISSN 0898-9788.

Jaeschke:1992:DSPb

[Jae92a] Rex Jaeschke. Data structures — part 10 — more on stacks. C Users Journal, 10(3):18–??, March 1992. ISSN 0898-9788.

Jaeschke:1992:DSPc

[Jae92b] Rex Jaeschke. Data structures — part 11 — yet more on stacks. C Users Journal, 10(4):89–??, April 1992. ISSN 0898-9788.

Jaeschke:1992:DSPd

[Jae92c] Rex Jaeschke. Data structures — part 12: Queues. C Users Journal, 10(5):20–??, May 1992. ISSN 0898-9788.

Jaeschke:1992:DSPe

[Jae92d] Rex Jaeschke. Data structures — part 13: Queues. C Users Journal, 10(6):93–??, June 1992. ISSN 0898-9788.

Jaeschke:1992:DSPf

[Jae92e] Rex Jaeschke. Data structures — part 14: Trees. C Users Journal, 10(7):67–??, July 1992. ISSN 0898-9788.

Jaeschke:1992:DSPg

[Jae92f] Rex Jaeschke. Data structures — part 15: Trees (continued). C Users Journal, 10(8):107–??, August 1992. ISSN 0898-9788.

Jaeschke:1992:DSPh

[Jae92g] Rex Jaeschke. Data structures — part 16: Trees III. C Users Journal, 10(9):103–??, September 1992. ISSN 0898-9788.

Jaeschke:1992:DSPa

[Jae92h] Rex Jaeschke. Data structures — part 9: Stacks. C Users Journal, 10(2):20–??, February 1992. ISSN 0898-9788.

Jaeschke:1993:CIE

[Jae93] Rex Jaeschke. C++: an introduction for experienced C programmers. CBM Books, A Division of Cardinal Business Media Inc., 101 Witmer Road, Horsham, PA 19044, USA, 1993. ISBN 1-878956-27-2. xii + 236 pp. LCCN QA76.73.C15J3354 1993. See book review [Lop93].

Jaeschke:1994:CC

[Jae94] Rex Jaeschke. The C9X charter. C/C++ Users Journal, 12 (11):47–??, November 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Jaeschke:1995:IPPa

[Jae95a] Rex Jaeschke. Internationalization: A primer, part 1. C/C++ Users Journal, 13(9):67– ??, September 1995. CODEN CCUJEX. ISSN 1075-2838.

Jaeschke:1995:IPPb

[Jae95b] Rex Jaeschke. Internationalization, A primer, part 2. C/C++ Users Journal, 13(10):69–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Jagger:1995:ADF

[Jag95] Jon Jagger. An alternative debug function macro. C/C++ Users Journal, 13(10):29–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Janzen:1993:CCF

[Jan93a] Thomas E. Janzen. C++ classes for fuzzy logic. C Users Journal, 11(11):55–??, November 1993. ISSN 0898-9788.

Janzen:1993:RCW

[Jan93b] Thomas E. Janzen. Recovering corrupted waveforms. C Users Journal, 11(6):39–??, June 1993. ISSN 0898-9788.

Jarvis:1990:LLC

[Jar90] Robert Jarvis. Linked lists in C++. C Users Journal, 8(3):107– ??, March 1990. ISSN 0898-9788.

Jarrett:1991:CCM

[Jar91a] John F. Jarrett. CUG354 — CES mouse tools library w/ JoyStick.

C Users Journal, 9(11):128–??, November 1991. ISSN 0898-9788.

Jarvis:1991:BBT

[Jar91b] Robert Jarvis. Balanced binary trees in C++. C Users Journal, 9(1):65–??, January 1991. ISSN 0898-9788.

Jensen:1996:MMO

[Jen96] Jens A. Jensen. 32-bit memory management in OS/2. C/C++ Users Journal, 14(5):55–??, May 1996. CODEN CCUJEX. ISSN 1075-2838.

Jervis:1991:ORU

[Jer91] Robert Jervis. OOP for the rest of us. C Users Journal, 9(7):41– ??, July 1991. ISSN 0898-9788.

Jervis:1994:AF

[Jer94] Bob Jervis. All is flux. C/C++ Users Journal, 12(10):39–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Johnson:1988:MES

[Joh88a] Marcus W. Johnson. Micro Emacs to SCO Xenix port. C Users Journal, 6(6):87–??, June 1988. ISSN 0898-9788.

Johnson:1988:CCF

[Joh88b] Susan M. Johnson. Control-Z corruption of files. C Users Journal, 6(1):67–??, January 1988. ISSN 0898-9788.

Johnson:1990:DVA

[Joh90a] Marcus W. Johnson. Detecting video adapters at runtime. C Users Journal, 8(8):113–??, August 1990. ISSN 0898-9788.

Johnson:1990:WMD

[Joh90b] Marcus W. Johnson. Writing MS-DOS device drivers. C Users Journal, 8(12):41–??, December 1990. ISSN 0898-9788.

Johnson:1990:SAE

[Joh90c]  Mark A. Johnson.  A sim-

ple application environment. C Users Journal, 8(1):67–??, January 1990. ISSN 0898-9788.

Johnston:1990:URP

[Joh90d] Richard Johnston. Using an RPC protocol to create a basic file server. C Users Journal, 8(9):65– ??, September 1990. ISSN 0898-9788.

Johns:1992:EGV

[Joh92] David Johns. Embedding global variables in structures. C Users Journal, 10(7):100–??, July 1992. ISSN 0898-9788.

Johnson:1993:CJ

[Joh93] Richard B. Johnson. CUG380 — JMODEM. C Users Journal, 11(9):119–??, September 1993. ISSN 0898-9788.

Johnson:1999:SGD

[Joh99a] Ivan J. Johnson. Safer graphical drawing with MFC. C/C++ Users Journal, 17(10):??, October 1999. CODEN CCUJEX. ISSN 1075-2838.

Johnson:1999:TED

[Joh99b] Ivan J. Johnson. Templates for efficient dynamic type checking. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Jones:1988:TUS

[Jon88a] Darrell Jones. UNIX for Super-Users, by Eric Foxley. C Users Journal, 6(1):42–??, January 1988. ISSN 0898-9788. See [Fox85].

Jones:1988:SPC

[Jon88b] Thomas Jones. Spinnaker’s Power C on the Commodore 128. C Users Journal, 6(9):13– ??, September 1988. ISSN 0898-9788.

Joslin:1993:UWD

[Jos93] Philip Joslin. Using the Windows DIB color table. C Users Journal, 11(9):27–??, September 1993. ISSN 0898-9788.

Johnson:1990:CBA

[JPM90] Mark A. Johnson, Henry M. Pollock, and John Muczynski. CUG301 — BGI applications. C Users Journal, 8(2):125–??, February 1990. ISSN 0898-9788.

Johnson:1991:XWPa

[JR91a] Eric F. Johnson and Kevin Reichard. X Window programming — part 1: The X Window System. C Users Journal, 9(3):87– ??, March 1991. ISSN 0898-9788.

Johnson:1991:XWPb

[JR91b] Eric F. Johnson and Kevin Reichard. X Window programming — part 2: The X library. C Users Journal, 9(5):30–??, May 1991. ISSN 0898-9788.

Johnson:1991:XWPc

[JR91c] Eric F. Johnson and Kevin Reichard. X Window programming — part 4: Making odd-shaped Windows. C Users Journal, 9 (10):103–??, October 1991. ISSN 0898-9788.

Johnson:1991:XWPd

[JR91d] Eric F. Johnson and Kevin Reichard. X Window programming — part 5: X toolkit programming. C Users Journal, 9(11):59– ??, November 1991. ISSN 0898-9788.

Johnson:1991:XWPe

[JR91e] Eric F. Johnson and Kevin Reichard. X Window programming — part 6: Motif programming. C Users Journal, 9(12):91–??, December 1991. ISSN 0898-9788.

Jennings:1994:CAD

[JR94] Gregg Jennings and Mike Rejsa. CUG406 — ATOC and DISKED. C Users Journal, 12 (5):98–??, May 1994. ISSN 0898-9788.

Kalev:1998:PCA

[Kal98] Danny Kalev. Porting a C++ application to Java. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCUJEX. ISSN 1075-2838.

Kamp:1989:TPP

[Kam89] William Kamp. Programming Principles in Computer Graphics, by Leendert Ammeraal. C Users Journal, 7(2):66–??, February 1989. ISSN 0898-9788. See [Amm86, Amm87b].

Kamradt:1993:TCT

[Kam93] Randal Kamradt. Template classes for the iostreams library. C Users Journal, 11(1):45–??, January 1993. ISSN 0898-9788.

Kamradt:1994:WRT

[Kam94] Randal Kamradt. Weight reduction techniques in C++. C Users Journal, 12(3):70–??, March 1994. ISSN 0898-9788.

Kamins:1997:SMC

[Kam97] Bob Kamins. Sending mail from a CGI program. C/C++ Users Journal, 15(9):45–??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Karn:1990:TCI

[Kar90] Phil Karn. Topological chemical indexes. C Users Journal, 8(3): 142–??, March 1990. ISSN 0898-9788.

Karas:1994:CSA

[Kar94] Walter Karas. CUG395B — sortlist AVL. C Users Journal, 12 (2):112–??, February 1994. ISSN 0898-9788.

Kasparek:1996:PVR

[Kas96]  Linda Kasparek.  Process-

ing variant records with STL. C/C++ Users Journal, 14(9):19– ??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Kathan:1999:PCM

[Kat99] Joseph Kathan. Portable control of multiple daemon processes. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Kauffmann:1993:WDS

[Kau93] Kit Kauffmann. A Windows debugging stream. C Users Journal, 11(12):45–??, December 1993. ISSN 0898-9788.

Kelly-Bootle:1990:QTM

[KB90] Stan Kelly-Bootle. Quirk of malloc() with argument of 0. C Users Journal, 8(9):144–??, September 1990. ISSN 0898-9788.

Kelly-Bootle:1992:BCP

[KB92] Stan Kelly-Bootle. Bad C pun contest results. C Users Journal, 10(5):83–??, May 1992. ISSN 0898-9788.

Kelly-Bootle:1999:PMDc

[KB99a] Stan Kelly-Bootle. Post-mortem debunker: Ambiguity: Its cause and cure. C/C++ Users Journal, 17(10):??, October 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:1999:PMDd

[KB99b] Stan Kelly-Bootle. Post-mortem debunker: Just in time? C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:1999:PMDb

[KB99c] Stan Kelly-Bootle. Post-mortem debunker: Streak on! C/C++ Users Journal, 17(8):104, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:1999:PMDa

[KB99d] Stan Kelly-Bootle. Post-mortem debunker: The fence less straddled. C/C++ Users Journal, 17 (6):??, June 1999. CODEN CCU-JEX. ISSN 1075-2838.

Keathly:1990:PEP

[Kea90] David Keathly. PC expert and PC expert professional. C Users Journal, 8(9):125–??, September 1990. ISSN 0898-9788.

Kelly:1990:BME

[Kel90] Michael Kelly. Buffering mouse events. C Users Journal, 8(12): 59–??, December 1990. ISSN 0898-9788.

Kelly:1991:UCD

[Kel91] Michael Kelly. A universal C++ data type. C Users Journal, 9 (6):51–??, June 1991. ISSN 0898-9788.

Kelly:1992:EBT

[Kel92] Michael Kelly. The evolution of a binary tree in Turbo C. C Users Journal, 10(10):121–??, October 1992. ISSN 0898-9788.

Kelly:1994:MOB

[Kel94] Michael Kelly. Multithreading with OS/2 and Borland C++. C/C++ Users Journal, 12(8):67– ??, August 1994. CODEN CCU-JEX. ISSN 1075-2838.

Kempf:1987:NST

[Kem87] James Kempf. Numerical software tools in C. Prentice-Hall software series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1987. ISBN 0-13-627274-6 (paperback). x + 261 pp. LCCN QA297.K43 1987.

Kerzner:1996:SPD

[Ker96] Mark Kerzner. Signal processing with dynamic programming. C/C++ Users Journal, 14(6):17– ??, June 1996. CODEN CCU-JEX. ISSN 1075-2838.

Kougiouris:1997:PMF

[KF97] Panos Kougiouris and Marco Framba. A portable multithreading framework. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Kientzle:1995:SBB

[Kie95] Tim Kientzle. Scaling bitmaps with bresenham. C/C++ Users Journal, 13(10):51–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Kientzle:1996:AIC

[Kie96] Tim Kientzle. Approximate inverse color mapping. C/C++ Users Journal, 14(8):19–??, August 1996. CODEN CCUJEX. ISSN 1075-2838.

Kilian:1996:SCM

[Kil96] Marc Kilian. Secrets of the C++ Masters, by Jeff Alger. C/C++ Users Journal, 14(8):79–??, August 1996. CODEN CCUJEX. ISSN 1075-2838. See [Alg95].

Kirkendall:1992:CEV

[Kir92] Steve Kirkendall. CUG365 — Elvis v1.5. C Users Journal, 10 (7):96–??, July 1992. ISSN 0898-

Kiselyov:1996:HMT

[Kis96] Oleg Kiselyov. Handling multiple TCP connections in C++. C/C++ Users Journal, 14(5):17– ??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Kiselyov:1997:SIE

[Kis97] Oleg Kiselyov. Speaking in iostreams-ese. C/C++ Users Journal, 15(5):47–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Kaduch:1998:OOF

[KJV98] Frantisek Kaduch, Damian Jan, and Purificacion Vidal. Objectoriented finite-state machines. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Kay:1989:CPU

[KK89] Judy Kay and Bob Kummerfeld. C programming in a UNIX environment. International computer science series. Addison-Wesley, Reading, MA, USA, 1989. xii + 340 pp. LCCN QA76.73.C15 K39 1988. See book review [Nel90b].

Klein:1993:WMS

[Kle93] Mike Klein. The Windows messaging system. C Users Journal, 11(5):70–??, May 1993. ISSN 0898-9788.

Klementiev:1997:WSL

[Kle97] Dmitri Klementiev. A Windows shell for legacy MS-DOS applications. C/C++ Users Journal, 15 (6):71–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Knauss:1989:PMS

[Kna89] Kevin E. Knauss. A poor man’s solution to the Traveling Salesman Problem. C Users Journal, 7(8):89–??, August 1989. ISSN 0898-9788.

Knoblaugh:1990:USM

[Kno90] Rick Knoblaugh. Using ‘screen machine’. C Users Journal, 8 (2):81–??, February 1990. ISSN 0898-9788.

Knutsson:1998:OGW

[Knu98] Lukas Knutsson. Optimized GIFs from Windows DIBs. C/C++ Users Journal, 16(12):??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Ko:1994:CSM

[KO94] Patrick KO. CUG403 — small matrix toolbox for C. C Users Journal, 12(4):122–??, April 1994. ISSN 0898-9788.

Kochan:1983:PC

[Koc83] Stephen G. Kochan. Programming in C. Hayden Book Co., Rochelle Park, NJ, USA, 1983. 373 pp. LCCN QA76.73.C15 K63 1983. See book review [Kou88].

Kochan:1988:PC

[Koc88] Stephen G. Kochan. Programming in C. Hayden Books C library. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, revised edition, 1988. ISBN 0-672-48420-X (paperback). xi + 445 pp. LCCN QA76.73.C15 K63 1988.

US$24.95. See book review [Kou88].

Kochan:1994:PC

[Koc94] Stephen G. Kochan. Programming in C. Hayden Book Co., Rochelle Park, NJ, USA, 1994. ISBN 0-8104-6261-3 (paperback). 373 pp. LCCN QA76.73.C15 K63 1983. See book review [Kou88].

Kodis:1992:UVA

[Kod92] John Kodis. Using variablelength argument lists in C. C Users Journal, 10(6):31–??, June 1992. ISSN 0898-9788.

Kostrzewa:1999:SSI

[Kos99] Kevin Kostrzewa. STL-style iterators for MFC. C/C++ Users Journal, 17(9):30–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Kotz:1991:CG

[Kot91] David Kotz. CUG334 GNU-PLOT. C Users Journal, 9(4): 110–??, April 1991. ISSN 0898-9788.

Koutros:1988:TPC

[Kou88] Stephen A. Koutros. Programming in C, by Stephen Kochan. C Users Journal, 6(3):15–??, March 1988. ISSN 0898-9788. See [Koc83, Koc88, Koc94].

Kougiouris:1997:YAC

[Kou97] Panos Kougiouris. Yet another command-line parser. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Kernighan:1978:CPL

[KR78] Brian W. Kernighan and Dennis M. Ritchie. The C programming language. Prentice-Hall software series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1978. ISBN 0-13-110163-3. x + 228 pp. LCCN QA76.73.C15 .K47.

Kernighan:1988:CPL

[KR88] Brian W. Kernighan and Dennis M. Ritchie. The C Programming Language. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1988. ISBN 0-13-110362-8 (paperback), 0-13-110370-9 (hardcover), 0-13-115817-1, 0-13-308621-6 (e-book), 0-13-308624-0. xii + 272 pp. LCCN QA76.73.C15 K47 1988. URL

http://wiki.tuhs.org/doku.
php?id=publications:c_prog_

language. This book is 200% a must for any C programmers …. The answers to the exercises can be found in [TG89].

Krantz:1989:PRC

[Kra89] Don Krantz. Programming realtime controllers in C. C Users Journal, 7(11):119–??, November 1989. ISSN 0898-9788.

Krehbiel:1996:PVB

[Kre96] Chris Krehbiel. Programming with VESA BIOS extensions. C/C++ Users Journal, 14(8):37– ??, August 1996. CODEN CCU-JEX. ISSN 1075-2838.

Kristoffersen:1997:PD

[Kri97] Jan Kristoffersen. Portable I/O drivers. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Krupansky:1992:LWP

[Kru92] Jack Krupansky. Liana: A Windows programming language based on C and C++. C Users Journal, 10(7):27–??, July 1992. ISSN 0898-9788.

Kruzdall:1995:TNV

[Kru95] James A. Kruzdall. Transferring numeric values between computers. C/C++ Users Journal, 13 (1):33–??, January 1995. CO-DEN CCUJEX. ISSN 1075-2838. Konstantinidis:1998:ASP

[KS98] Thanos D. Konstantinidis and M. G. Strintzis. Advanced serial port communication under Win32. C/C++ Users Journal, 16(9):??, September 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Kuo:1997:CIA

[Kuo97] Benjamin Kuo. Create Internetaware apps with the Internet SDK. C/C++ Users Journal, 15 (8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Kuzdrall:1990:ES

[Kuz90] James A. Kuzdrall. Executable strings. C Users Journal, 8(5): 23–??, May 1990. ISSN 0898-9788.

Kwok:1989:CGM

[Kwo89a] Conrad Kwok. CUG286 — GRAD for MSC. C Users Journal, 7(7):119–??, July 1989. ISSN 0898-9788.

Kwok:1989:CGT

[Kwo89b] Conrad Kwok. CUG287 — GRAD for Turbo C. C Users Journal, 7(7):119–??, July 1989. ISSN 0898-9788.

Labrosse:1992:MRK

[Lab92]  Jean J. Labrosse.  μC/OS:

the real-time kernel. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-031352-1. xviii + 266 pp. LCCN QA76.76.O63 L23 1992. See book review [Gin93a].

LaBrocca:1993:DTA

[LaB93a] P. J. LaBrocca. Dynamic twodimensional arrays. C Users Journal, 11(11):72–??, November 1993. ISSN 0898-9788.

LaBrocca:1993:MNC

[LaB93b] P. J. LaBrocca. Mixed numbers in C. C Users Journal, 11(4):69– ??, April 1993. ISSN 0898-9788.

LaBrocca:1994:SAM

[LaB94] P. J. LaBrocca. A self-extracting archive for MS-DOS. C Users Journal, 12(2):75–??, February 1994. ISSN 0898-9788.

Labrosse:1998:FPA

[Lab98] Jean J. Labrosse. Fixed-point arithmetic for embedded systems. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

LaCombe:1994:SAF

[LaC94] Andy LaCombe. Spectral analysis and FFT analysis sources. C Users Journal, 12(5):118–??, May 1994. ISSN 0898-9788.

Ladd:1989:AMM

[Lad89a] Scott Robert Ladd. Accessing the MS-DOS master environment. C Users Journal, 7(7):59–??, July 1989. ISSN 0898-9788.

Ladd:1989:TSC

[Lad89b] Scott Robert Ladd. A tool for simplifying complex logic. C Users Journal, 7(2):69–??, February 1989. ISSN 0898-9788.

Ladd:1991:PAM

[Lad91] Scott Robert Ladd. Portability across MS-DOS C compilers. C Users Journal, 9(1):53–??, January 1991. ISSN 0898-9788.

Ladd:1992:CCA

[Lad92] Scott Ladd. C++ components and algorithms: a comprehensive reference for designing and implementing algorithms in C++. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1992. ISBN 1-55851-227-6. xv + 779 pp. LCCN QA76.73.C153 .L33 1992 Accompanying disk shelved in Reserve. See book review [Hal93b]. System requirements for computer disk: Borland C++ v.3 or Microsoft C/ C++ v.7; PC.

Ladd:1994:CCA

[Lad94]  Scott Ladd.  C++ compo-

nents and algorithms. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 1994. ISBN 1-55851-408-2. xx + 476 pp. LCCN QA76.73.C153 L3 1994 Accompanying disk shelved in Reserves. US$39.95. System requirements for computer disk: IBM-compatible PC; PC/MS-DOS; Borland C++ compiler.

Lambert:1989:MME

[Lam89] John K. Lambert. Modifying the MS-DOS environment: A bibliography. C Users Journal, 7(10): 58–??, October 1989. ISSN 0898-9788.

Lam:1997:CCF

[Lam97] Richard B. Lam. A C++ CGI framework. C/C++ Users Journal, 15(1):39–??, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Lang:1989:ITR

[Lan89] Phyllis K. Lang. Improving the timing resolution of an MS-DOS PC. C Users Journal, 7(7):85–??, July 1989. ISSN 0898-9788.

Lane:1990:DOT

[Lan90] Alex Lane. Debugging objects with Turbo debugger. C Users Journal, 8(10):53–??, October 1990. ISSN 0898-9788.

Lane:1991:TOD

[Lan91a] Alex Lane. Object-Oriented Design with Applications, by Grady Booch. C Users Journal, 9(2): 125–??, February 1991. ISSN 0898-9788. See [Boo91].

Lang:1991:CSC

[Lan91b] Don Lang. CUG163 — small C for MS-DOS. C Users Journal, 9(5):135–??, May 1991. ISSN 0898-9788.

Lang:1991:ESC

[Lan91c] Don Lang. Enhancing small C V2.0. C Users Journal, 9(7):61– ??, July 1991. ISSN 0898-9788.

Lane:1993:CJ

[Lan93] Thomas G. Lane. CUG381 — JPEG. C Users Journal, 11(9): 120–??, September 1993. ISSN 0898-9788.

Lansinger:1996:ECA

[Lan96] Ed Lansinger. An embedded C++ application. C/C++ Users Journal, 14(2):19–??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Langseth:1999:UCW

[Lan99a] Jim Langseth. Using C++ with Word97 COM interfaces. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Lansinger:1999:LCL

[Lan99b] Edward J. Lansinger. A lightweight C++ library for embedded LCDs. C/C++ Users Journal, 17(2):??, February 1999.

CODEN CCUJEX. ISSN 1075-2838.

Larsson:1990:CP

[Lar90] Bjorn Larsson. CUG332 PCcurses. C Users Journal, 8(12): 126–??, December 1990. ISSN 0898-9788.

Larbi:1997:BRM

[Lar97] Michael Larbi. Book review:

Multithreading Applications in Win32. C/C++ Users Journal, 15(7):65–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Langsam:1996:DSU

[LAT96] Yedidyah Langsam, Moshe Augenstein, and Aaron M. Tenenbaum. Data structures using C and C++. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-036997-7. xvi + 672 pp. LCCN QA76.73.C15 T46 1996 Bar. Revised edition of [TLA90].

Lawrence:1992:CMU

[Law92a] Daniel Lawrence. CUG366 — MicroEMACS update. C Users Journal, 10(11):74–??, November 1992. ISSN 0898-9788.

Lawrence:1992:CMV

[Law92b] Daniel Lawrence. CUG374 — MicroSpell v2.0. C Users Journal, 10(12):121–??, December 1992. ISSN 0898-9788.

Lawrence:1992:MDD

[Law92c] Danny Lawrence. A MIDI device driver for XENIX. C Users Journal, 10(12):61–??, December 1992. ISSN 0898-9788.

Lawless:1995:IFC

[Law95] James K. Lawless. Interfacing to a FOSSIL communications driver. C/C++ Users Journal, 13(6):59–??, June 1995. CODEN CCUJEX. ISSN 1075-2838.

Leatherbarrow:1996:NSR

[Lea96] Robin J. Leatherbarrow. Nested for statements revisited. C/C++ Users Journal, 14(11):57–??, November 1996. CODEN CCU-JEX. ISSN 1075-2838.

Letendre:1991:CCL

[Let91] William Letendre. CUG339 — ctrl-C library. C Users Journal, 9(4):110–??, April 1991. ISSN 0898-9788.

Ladd:1989:CTS

[LFK89] Scott Robert Ladd, Brian Faquhar, and Kevin E. Knauss. CUG288 — traveling salesman, SD, and master environment. C Users Journal, 7(7):119–??, July 1989. ISSN 0898-9788.

Libes:1989:FTT

[Lib89a]  Don Libes.  Faster than

malloc(). C Users Journal, 7(5): 8–??, May 1989. ISSN 0898-9788.

Libes:1989:IOC

[Lib89b] Don Libes. The international obfuscated C code contest. C Users Journal, 7(7):93–??, July 1989. ISSN 0898-9788.

Libes:1989:UEV

[Lib89c] Don Libes. Updating environment variables. C Users Journal, 7(10):39–??, October 1989. ISSN 0898-9788.

Libes:1990:FSC

[Lib90a] Don Libes. The far side of C. C Users Journal, 8(5):125–??, May 1990. ISSN 0898-9788.

Libes:1990:IST

[Lib90b] Don Libes. Implementing software timers. C Users Journal, 8(11):108–??, November 1990. ISSN 0898-9788.

Libes:1990:LSBa

[Lib90c] Don Libes. Life with static buffers. C Users Journal, 8(1): 28–??, January 1990. ISSN 0898-9788.

Libes:1990:LSBb

[Lib90d] Don Libes. Life with static buffers, part 2. C Users Journal, 8(3):103–??, March 1990. ISSN 0898-9788.

Libes:1990:TCL

[Lib90e] Don Libes. Tool command language. C Users Journal, 8(9):57– ??, September 1990. ISSN 0898-9788.

Libes:1990:UTY

[Lib90f] Don Libes. Using yacc or lex twice in one program. C Users Journal, 8(7):55–??, July 1990. ISSN 0898-9788.

Libes:1991:DPY

[Lib91a] Don Libes. Don’t put this on your resumé. C Users Journal, 9 (5):89–??, May 1991. ISSN 0898-

Libes:1991:E

[Lib91b] Don Libes. Expect. C Users Journal, 9(1):103–??, January 1991. ISSN 0898-9788.

Libes:1991:ITT

[Lib91c] Don Libes. Implementing a trap command. C Users Journal, 9(3): 123–??, March 1991. ISSN 0898-9788.

Libes:1992:OCO

[Lib92] Don Libes. Obfuscated C and other mysteries. Wiley professional computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1992. xii + 413 pp. LCCN QA76.73.C15 L5 1993.

Lienau:1992:CG

[Lie92]  Sara Lienau.  CUG370 —

GATool. C Users Journal, 10 (11):74–??, November 1992. ISSN 0898-9788.

Langer:1997:IUS

[LK97] Angelika Langer and Klaus Kreft. Internationalization using Standard C++. C/C++ Users Journal, 15(9):57–??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Lo:1990:CFA

[LL90] Dennis Lo and David Lo. CUG323 fireworks and adventure. C Users Journal, 8(8):111– ??, August 1990. ISSN 0898-9788.

Laurel:1990:AHI

[LM90] Brenda Laurel and S. Joy Mountford. The Art of human-computer interface design. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51797-3. xvi + 523 + 16 pp. LCCN QA76.9.H85 A78 1990. See book review [Rom91b]. S. Joy Mountford, Manager of the Human Interface Group, Apple Computer Inc., conceived of and technically supported the development of this book.

Little:1994:CS

[LM94a] M. C. Little and D. L. McCue.

C++ SIM. C Users Journal, 12 (3):119–??, March 1994. ISSN 0898-9788.

Little:1994:CCD

[LM94b] M. C. Little and D. L. McCue.

CUG394 — C++SIM discrete simulations. C Users Journal, 12 (2):111–??, February 1994. ISSN 0898-9788.

Long:1991:COP

[Lon91a] Rodney Long. CUG341 — orbit propagation. C Users Journal, 9(5):135–??, May 1991. ISSN 0898-9788.

Long:1991:OP

[Lon91b] Rodney Long. Orbit propagation. C Users Journal, 9(5):97– ??, May 1991. ISSN 0898-9788.

Lopez:1993:TCI

[Lop93]  Russel Lopez.  C++: An

Introduction for Experienced C Programmers, by Rex Jaeschke. C Users Journal, 11(11):117–??, November 1993. ISSN 0898-9788. See [Jae93].

Lopez:1994:PVN

[Lop94] Manuel Lopez. PC/VI: Notes on its origins and current support. C Users Journal, 12(1):117–??, January 1994. ISSN 0898-9788.

Lopez:1997:PLU

[Lop97] Lionel Lopez. Persistent lists using ISAM files. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Lorde:1999:GIS

[Lor99] David Lorde. A generic iterator for strings. C/C++ Users Journal, 17(4):??, April 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Louis:1992:CR

[Lou92] B. Gregory Louis. Cyclic redundancy. C Users Journal, 10(10): 55–??, October 1992. ISSN 0898-9788.

Lowerre:1995:MPC

[Low95] David T. Lowerre. Map projections in C++. C/C++ Users Journal, 13(6):45–??, June 1995. CODEN CCUJEX. ISSN 1075-2838.

Lawrence:1992:CMW

[LP92] Daniel Lawrence and Pierre Perret. CUG373 — MicroEMACS for Windows. C Users Journal, 10(12):121–??, December 1992. ISSN 0898-9788.

Laird:1998:MPL

[LS98] Cameron Laird and Kathryn Soraiz. Motif programming with the LessTif freeware library. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Mallett:1988:BLR

[MA88] Mark Mallett and Jonathan Arnold. The BOX library: A reasonably portable WIMP interface. C Users Journal, 6(11):55– ??, November 1988. ISSN 0898-9788.

Maguire:1993:WSC

[Mag93] Steve Maguire. Writing solid code: Microsoft’s techniques for developing bug-free C programs. Microsoft Press, Bellevue, WA, USA, 1993. ISBN 1-55615-551-4. xxix + 256 pp. LCCN QA76.73.C15 M335 1993 Sci-Eng. US$24.95, CDN$32.95. See book review [Hal93a].

Mahoney:1999:CCC

[Mah99] Colin Mahoney. Calculating CRC checksums in C++. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Maley:1990:WCL

[Mal90] Scott Maley. The world of command line options. C Users Journal, 8(3):25–??, March 1990. ISSN 0898-9788.

Manzai:1988:CMV

[Man88] James H. Manzai. CUG197 — MicroEmacs v3.9 exec and doc. C Users Journal, 6(9):92– ??, September 1988. ISSN 0898-

Mansky:1989:WCI

[Man89] Art Mansky. Writing the C interface for a real-time executive. C Users Journal, 7(7):69–??, July 1989. ISSN 0898-9788.

Mansky:1990:UCP

[Man90] Art Mansky. Using the C preprocessor for device control. C Users Journal, 8(12):33–??, December 1990. ISSN 0898-9788.

Man:1991:MLC

[Man91] Richard F. Man. A multithreading library in C for subsumption architecture. C Users Journal, 9 (11):42–??, November 1991. ISSN 0898-9788.

Manning:1996:FS

[Man96a] Evan Manning. Floating-point summation. C/C++ Users Journal, 14(9):51–??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Manning:1996:UPC

[Man96b] Evan Manning. Uncertainty propagation in C+. C/C++ Users Journal, 14(3):29, 32–38, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Manley:1998:GPT

[Man98] Kevin T. Manley. Generalpurpose threads with I/O completion ports. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Manley:1999:IPT

[Man99] Kevin Manley. Improving performance with thread-private heaps.

C/C++ Users Journal, 17(9):50– ??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Margison:1988:CCS

[Mar88a] Steven E. Margison. CUG271 and CUG272 — Steven’s library. C Users Journal, 6(9):91– ??, September 1988. ISSN 0898-9788.

Markley:1988:CMM

[Mar88b] Andrew Markley. Calling the Microsoft mouse interface from C. C Users Journal, 6(9):88– ??, September 1988. ISSN 0898-9788.

Marking:1990:CGI

[Mar90a] Michael P. Marking. CUG317 — group 3 image processing. C Users Journal, 8(7):111–??, July 1990. ISSN 0898-9788.

Marking:1990:DGI

[Mar90b] Michael P. Marking. Decoding group 3 images. C Users Journal, 8(6):45–??, June 1990. ISSN 0898-9788.

Martensen:1990:PE

[Mar90c] Brett Martensen. Prototyping experiences. C Users Journal, 8 (2):91–??, February 1990. ISSN 0898-9788.

Mariam:1991:PSS

[Mar91] P. Lyle Mariam. Puzzle style:

Some criticisms. C Users Journal, 9(4):118–??, April 1991. ISSN 0898-9788.

Mare:1992:YAC

[Mar92] Adolfo Di Mare. Yet another C++ money class. C Users Journal, 10(4):58–??, April 1992. ISSN 0898-9788.

Martin:1995:DOC

[Mar95] Robert C. Martin. Designing object-oriented C++ applications using the Booch method. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-203837-4. xxxi + 528 pp. LCCN QA76.64 .M384 1995. See book review [Row95].

Marini:1996:CHR

[Mar96] Marc Marini. A class hierarchy for random number generation. C/C++ Users Journal, 14(10): 51–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Martinka:1997:TIS

[Mar97]  Michael Martinka.  A tem-

plate implementation of skip lists. C/C++ Users Journal, 15 (1):49–??, January 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Marcok:1998:DAE

[Mar98] Zlatko Marcok. Dynamic arrays with expression templates. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Mashlan:1995:CPC

[Mas95] Robert Mashlan. Checked pointers for C++. C/C++ Users Journal, 13(10):37–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Mashlan:1997:CMT

[Mas97] Robert Mashlan. A C++ matrix template class. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Maternowski:1988:ASM

[Mat88] Paul Maternowski. Atari ST’s MS-DOS disk-reading ability. C Users Journal, 6(3):12–??, March 1988. ISSN 0898-9788.

Mataya:1989:SDK

[Mat89] Darryl Mataya. SHERLOCK: A different kind of debugger. C Users Journal, 7(1):25–??, January 1989. ISSN 0898-9788.

Matsche:1991:PM

[Mat91] John J. Matsche. Point-and-shoot menus. C Users Journal, 9 (4):57–??, April 1991. ISSN 0898-9788.

McIntosh:1988:CLC

[MC88] Harold V. McIntosh and Geraldo Cisneros. CUG275 — linear cellular automata II. C Users Journal, 6(9):91–??, September 1988. ISSN 0898-9788.

McCarthy:1989:RPT

[McC89] Patrick McCarthy. Recommended practices for TSR’s using blaise computing’s Turbo C tools. C Users Journal, 7(8):9–??, August 1989. ISSN 0898-9788.

McCullough:1992:CU

[McC92] William McCullough. CUG360 — uspell. C Users Journal, 10(3): 118–??, March 1992. ISSN 0898-9788.

McConnell:1993:CCP

[McC93] Steve McConnell. Code complete:

a practical handbook of software construction. Microsoft Press, Bellevue, WA, USA, 1993. ISBN 1-55615-484-4. xviii + 857 pp. LCCN QA76.76.D47 M39 1993. US$35.00, CDN$47.95. See book review [Ush94].

McCauley:1994:FLT

[McC94] Jack J. McCauley. A fuzzy logic torque servo. C Users Journal, 12(3):47–??, March 1994. ISSN 0898-9788.

McClung:1995:UAA

[McC95] Michael McClung. Using associative arrays. C/C++ Users Journal, 13(2):51–??, February 1995. CODEN CCUJEX. ISSN 1075-2838.

McDaniel:1988:ACC

[McD88] Michael McDaniel. Availability of C compiler for Concurrent CP/M. C Users Journal, 6(9):12– ??, September 1988. ISSN 0898-9788.

McKeeman:1998:LJC

[McK98] Bill McKeeman. A lexer for Java in C++. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

McLaughlin:1991:AST

[McL91] Robert McLaughlin. Automated software testing. C Users Journal, 9(2):103–??, February 1991. ISSN 0898-9788.

McLaughlin:1994:TEC

[McL94] Phil McLaughlin. Effective C++, by Scott Meyers. C Users Journal, 12(6):105–??, June 1994. ISSN 0898-9788. See [Mey92].

McMahon:1991:GCS

[McM91] William McMahon. A generic command-line switch. C Users Journal, 9(10):95–??, October 1991. ISSN 0898-9788.

Measday:1991:ECL

[Mea91] Charles Measday. Enhanced command line processing. C Users Journal, 9(6):75–??, June 1991. ISSN 0898-9788.

Measday:1992:CP

[Mea92] Charles Measday. C packages. C Users Journal, 10(6):70–??, June 1992. ISSN 0898-9788.

Meadows:1993:AUT

[Mea93] Roger Meadows. Automated unit testing. C Users Journal, 11(8): 53–??, August 1993. ISSN 0898-9788.

Merrill:1996:CDE

[Mer96] Ted Merrill. A C declarator extractor. C/C++ Users Journal, 14(5):47–??, May 1996. CODEN CCUJEX. ISSN 1075-2838.

Messer:1996:CCC

[Mes96] Greg Messer. A C++ chronograph class. C/C++ Users Journal, 14(9):35–??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Metzger:1992:PTT

[Met92] Robert C. Metzger. Pointer target tracking. C Users Journal, 10 (6):85–??, June 1992. ISSN 0898-

Meyers:1992:ECS

[Mey92] Scott (Scott Douglas) Meyers. Effective C++: 50 specific ways to improve your programs and designs. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56364-9. xvi + 206 pp. LCCN QA76.73.C153 M48 1992. See book review [McL94].

Meyers:1995:MUC

[Mey95] Scott Meyers. Mastering userdefined conversion functions. C/C++ Users Journal, 13(8):57– ??, August 1995. CODEN CCU-JEX. ISSN 1075-2838.

Meyers:1996:CCC

[Mey96] Scott Meyers. Combining C++ and C in the same program. C/C++ Users Journal, 14(10): 35–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Meyers:1998:COC

[Mey98] Scott Meyers. Counting objects in C++. C/C++ Users Journal, 16(4):??, April 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Michaels:1996:SPS

[Mic96]  Laura Michaels.  In search

of a portable screen library. C/C++ Users Journal, 14(9):57– ??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Michaels:1997:CCJ

[Mic97] Laura Michaels. C/C++ and Java library equivalents. C/C++ Users Journal, 15(6):53–55, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Miller:1988:CUL

[Mil88] David A. Miller. CUG268 and 269 — unicorn library v4.0. C Users Journal, 6(8):63–??, August 1988. ISSN 0898-9788.

Miller:1989:CUL

[Mil89] David A. Miller. CUG281 — unicorn library v5.1. C Users Journal, 7(2):103–??, February 1989. ISSN 0898-9788.

Miller:1991:CII

[Mil91] Blake Miller. CUG342 — I8255 interface library. C Users Journal, 9(5):135–??, May 1991. ISSN 0898-9788.

Milam:1994:EDL

[Mil94] Stan Milam. An extended date library for C. C/C++ Users Journal, 12(10):67–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Miller:1997:BRR

[Mil97] William M. Miller. Book review:

Ruminations in C++. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Mirho:1993:DEA

[Mir93] Charles Mirho. Designing an extensible API in C. C Users Journal, 11(1):21–??, January 1993. ISSN 0898-9788.

Mitchell:1999:ECA

[Mit99] Stan Mitchell. Easier console apps in Win32. C/C++ Users Journal, 17(9):63–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

MacKenzie:1992:CFT

[MO92] David J. MacKenzie and Thorsten Ohl. CUG367 — file and text utilities for MS-DOS. C Users Journal, 10(11):74–??, November 1992. ISSN 0898-9788.

Monagle:1997:TRL

[Mon97] Mike Monagle. Tracking down resource leaks with HandleTracker. C/C++ Users Journal, 15(6):19– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Mooallem:1991:SM

[Moo91] Saul Mooallem. 3-D solid modeling. C Users Journal, 9(12):35– ??, December 1991. ISSN 0898-9788.

Moolenaar:1994:CV

[Moo94] Bram Moolenaar. CUG411 — VIM. C/C++ Users Journal, 12 (7):83–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Moreland:1992:CMS

[Mor92] Carl Moreland. CUG372 — Mouse++, String++, and Z++ classes. C Users Journal, 10(12): 122–??, December 1992. ISSN 0898-9788.

Morley:1993:CV

[Mor93a] Mark Morley. CUG383 — VGL. C Users Journal, 11(10):111–??, October 1993. ISSN 0898-9788.

Morris:1993:SCL

[Mor93b] Jim Morris. The SPLASH class library. C Users Journal, 11(10):

49–??, October 1993. ISSN 0898-9788.

Moreno:1998:EDGa

[Mor98a] Carlos Moreno. Efficient 2-D geometric operations, part 1. C/C++ Users Journal, 16(11): 25–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Moreno:1998:EDGb

[Mor98b] Carlos Moreno. Efficient 2-D geometric operations, part 2. C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Moreno:1999:CFD

[Mor99] Carlos Moreno. Constructing FIR digital filters with valarray. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Meyers:1990:STC

[MP90] Larry Meyers and Kenneth Pugh.

Stuffing two characters into an integer. C Users Journal, 8(7):88– ??, July 1990. ISSN 0898-9788.

Miller:1987:CPL

[MQ87] Lawrence H. Miller and Alexander E. Quilici. C programming language: a self-teaching guide. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1987. ISBN 0-471-82560-3. ???? pp. LCCN QA76.73.C15 M55 1987 Bar.

Mann:1991:RUS

[MS91] Daniel Mann and Brett Stewart. Register usage strategies.

C Users Journal, 9(11):66–??, November 1991. ISSN 0898-9788.

Mueller:1994:PAC

[Mue94a] Harald M. Mueller. Powerful assertions for C++. C/C++ Users Journal, 12(10):21–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Mueller:1994:CCE

[Mue94b] Juergen Mueller. CUG391 — C/C++ exploration tools v2.12. C Users Journal, 12(1):87–??, January 1994. ISSN 0898-9788.

Mullin:1989:OOP

[Mul89] Mark Mullin. Object oriented program design: with examples in C++. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-51722-1. xvi + 303 pp. LCCN QA76.73.C15 M855 1989. See book review [Bur90b].

Muller:1993:CNT

[Mul93] Hans G. W. Muller. Changing the name of C Users Journal. C Users Journal, 11(7):135–??, July 1993. ISSN 0898-9788.

Murphy:1993:HIF

[Mur93] Thomas Murphy. Hiding ISAM function libraries with OOP. C Users Journal, 11(1):81–??, January 1993. ISSN 0898-9788.

Mustakos:1990:DCS

[Mus90] Richard Mustakos. Developing a C simulation library. C Users Journal, 8(10):119–??, October 1990. ISSN 0898-9788.

Nadelson:1998:RTE

[Nad98] Mark Nadelson. Real-time error processing on a Unix network. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Naro:1989:CSI

[Nar89] Rick Naro. C to silicon: The in’s and out’s of embedded system design. C Users Journal, 7(5):77– ??, May 1989. ISSN 0898-9788.

Nelson:1989:TCC

[Nel89a] Phyllis Nelson. C Chest and Other C Treasures from Dr. Dobb’s Journal, by Allen Holub. C Users Journal, 7(10):126–??, October 1989. ISSN 0898-9788.

Nelson:1989:TCP

[Nel89b] Phyllis Nelson. C, Problem Solving and Programming, by Kenneth A. Barclay. C Users Journal, 7(11):61–??, November 1989. ISSN 0898-9788. See [Bar89].

Nelson:1990:WYO

[Nel90a] Mark Nelson. Writing your own quicksort. C Users Journal, 8(8): 63–??, August 1990. ISSN 0898-9788.

Nelson:1990:TCP

[Nel90b] Phyllis Nelson. C Programming In A UNIX Environment, by Judy Kay and Bob Kummerfeld. C Users Journal, 8(8):119– ??, August 1990. ISSN 0898-9788. See [KK89].

Nelson:1992:BIH

[Nel92a] Mark Nelson. Bimodal interrupt handling under 80286 DOS extenders. C Users Journal, 10 (12):20–??, December 1992. ISSN 0898-9788.

Nelson:1992:DDS

[Nel92b] Thomas W. Nelson. The device driver as state machine. C Users Journal, 10(3):41–??, March 1992. ISSN 0898-9788.

Nelson:1994:EII

[Nel94] Thomas Nelson. Encapsulating the DOS IOCTL interface. C/C++ Users Journal, 12(11): 27–??, November 1994. CODEN CCUJEX. ISSN 1075-2838.

Nelson:1995:TPP

[Nel95] Thomas W. Nelson. Plug and Play Programming by William Wong. C/C++ Users Journal, 13 (9):79–??, September 1995. CO-DEN CCUJEX. ISSN 1075-2838. See [Won93].

Nelson:1996:MCD

[Nel96] Tom Nelson. Memory caching for disk-based objects. C/C++ Users Journal, 14(10):59–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Nelson:1997:PMS

[Nel97] Tom Nelson. Polyphase merge sorting. C/C++ Users Journal, 15(7):51–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Nelson:1998:DBC

[Nel98] Tom Nelson. Disk-based container objects. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Newcomer:1989:ALP

[New89] Joseph M. Newcomer. Avoiding and locating pointer bugs. C Users Journal, 7(11):67–??, November 1989. ISSN 0898-9788.

Newman:1991:PPS

[New91] Dave Newman. A portable PC screen I/O library. C Users Journal, 9(8):31–??, August 1991. ISSN 0898-9788.

Ngai:1997:TRC

[Nga97]  Kenneth Ngai.  A template

for reference counting. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Nickerson:1996:HJ

[Nic96] Douglas A. Nickerson. Hooked on Java. C/C++ Users Journal, 14 (11):79–80, November 1996. CO-DEN CCUJEX. ISSN 1075-2838. Extensive review of [vSS96].

Niemann:1994:T

[Nie94]  Thomas Niemann.  Today.

C/C++ Users Journal, 12(9):53– ??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Noll:1994:CIO

[Nol94] Landon Noll. CUG397 — international obfuscated C code contest 1984 — 1993. C Users Journal, 12(3):117–??, March 1994. ISSN 0898-9788.

Norton:1992:WDD

[Nor92] Daniel A. Norton. Windows device drivers. C Users Journal, 10 (12):43–??, December 1992. ISSN 0898-9788.

Nadelson:1999:RGU

[NW99] Mark Nadelson and Colleen Woodward. Report generation using an integrated Web browser. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Obenchain:1991:BMC

[Obe91] Robert L. Obenchain. Bug in Microsoft C versions 6.00 and 6.00A. C Users Journal, 9(4): 120–??, April 1991. ISSN 0898-9788.

Ockrassa:1989:SSC

[Ock89] David Ockrassa. Sending special characters to the printer under CP/M on a Commodore 128. C Users Journal, 7(5):25–??, May 1989. ISSN 0898-9788.

ODonnell:1990:CT

[O’D90] Gus O’Donnell. CUG302 — 3-D transforms. C Users Journal, 8(3):130–??, March 1990. ISSN 0898-9788.

ODell:1991:BGB

[O’D91] Robert T. O’Dell. Brief graphics bibliography. C Users Journal, 9(9):137–??, September 1991. ISSN 0898-9788.

Ogg:1990:CPR

[Ogg90a] Harold C. Ogg. The C programmer’s reference: A bibliography of periodicals. C Users Journal, 8(1):133–??, January 1990. ISSN 0898-9788.

Ogg:1990:TCP

[Ogg90b] Harold C. Ogg. C Programmer’s Toolkit, by Jack Purdum. C Users Journal, 8(8):125–??, August 1990. ISSN 0898-9788. See [Pur89b].

Ogg:1990:TUC

[Ogg90c] Harold C. Ogg. Understanding C, by Carl Townsend. C Users Journal, 8(3):79–??, March 1990. ISSN 0898-9788. See [Tow88].

Olsen:1994:SAE

[OH94] Odd A. S. Olsen and Petter H. Heyerdahl. Symbolic access to embedded controllers. C Users Journal, 12(3):21–??, March 1994. ISSN 0898-9788.

Ohl:1992:CGM

[Ohl92]  Thorsten Ohl.  CUG368 —

GNUlib for MS-DOS. C Users Journal, 10(11):74–??, November 1992. ISSN 0898-9788.

Olhoeft:1990:ASV

[Olh90] Gary R. Olhoeft. Addressing super VGA modes from protected mode C. C Users Journal, 8(10): 99–??, October 1990. ISSN 0898-9788.

Olsen:1993:TMU

[Ols93]  Thomas W. Olsen.  Three-

dimensional modeling under Windows 3.1. C Users Journal, 11(3):53–??, March 1993. ISSN 0898-9788.

ONeil:1975:TC

[O’N75] W. M. (William Matthew) O’Neil. Time and the calendars. Sydney University Press, Sydney, NSW, Australia, 1975. ISBN 0-424-00003-2. x + 138 pp. LCCN CE6 .O531. See book review [Pla93g].

ONeil:1978:TC

[O’N78] W. M. (William Matthew) O’Neil. Time and the calendars. Sydney University Press, Sydney, NSW, Australia, 1978. x + 138 pp. LCCN CE6 .O53 1978. See book review [Pla93g].

Osborn:1990:CSE

[Osb90] Gary Osborn. CUG331 SE editor. C Users Journal, 8(12):126– ??, December 1990. ISSN 0898-9788.

Oshel:1989:CF

[Osh89] David C. Oshel. CUG283 — FAFNIR. C Users Journal, 7 (2):103–??, February 1989. ISSN 0898-9788.

Otto:1994:TFP

[Ott94] Erick Otto. Two fast patternmatching algorithms. C Users Journal, 12(2):39–??, February 1994. ISSN 0898-9788.

Oualline:1992:CES

[Oua92] Steve Oualline. C elements of style: the programmers style manual for Elegant C and C++ programs. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1992. ISBN 1-55851-291-8. xi + 265 pp. LCCN QA76.73.C153 O83 1992. US$21.95. See book review [Phi94a].

Owen:1995:DDP

[Owe95] Gregor Owen. Doing “32-bit”

DMA on a PC. C/C++ Users Journal, 13(12):27–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1991:STI

[PA91] Kenneth Pugh and Jamshid Afshar. The sizes of int * and void * pointers. C Users Journal, 9(11):113–??, November 1991. ISSN 0898-9788.

Plauger:1993:NIA

[PA93a] P. J. Plauger and R. F. Anthracite. Notes on initializing arrays and generic swap functions. C Users Journal, 11(8):136–??, August 1993. ISSN 0898-9788.

Pugh:1993:MCS

[PA93b] Kenneth Pugh and Andrew W.

Ackard. Modem command string compendium needed. C Users Journal, 11(5):93–??, May 1993. ISSN 0898-9788.

Pajari:1988:MUW

[Paj88] George E. Pajari. Of mice and UNIX — writing a UNIX device driver for the Microsoft bus mouse. C Users Journal, 6(9):54– ??, September 1988. ISSN 0898-

Panas:1991:CDZ

[Pan91] Michael G. Panas. CUG348 — 8048 Disassembler/Z80 cross assembler. C Users Journal, 9(8): 120–??, August 1991. ISSN 0898-9788.

Panattoni:1993:PF

[Pan93] Dale A. Panattoni. A prompting function. C Users Journal, 11 (4):45–??, April 1993. ISSN 0898-9788.

Panzer:1999:ACI

[Pan99] John Panzer. Automatic code instrumentation. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Paradis:1988:TCP

[Par88] Art Paradis. C Primer Plus. C Users Journal, 6(11):14–??, November 1988. ISSN 0898-9788.

Parker:1995:EMC

[Par95] Mark R. Parker. Encapsulating math coprocessor BCD arithmetic. C/C++ Users Journal, 13 (6):27–??, June 1995. CODEN CCUJEX. ISSN 1075-2838.

Parent:1998:ALM

[Par98] Michael Parent. ActiveX licensing with MD5 encryption. C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Pass:1988:UST

[Pas88] E. M. Pass. UNIX system 5 terminal raw mode setting. C Users Journal, 6(2):69–??, February 1988. ISSN 0898-9788.

Pass:1989:ESC

[Pas89] E. M. Pass. Embedding SQL commands in your C source. C Users Journal, 7(5):105–??, May 1989. ISSN 0898-9788.

Patton:1988:CMC

[Pat88a] Robert L. Patton. CUG266 — microPLOX — chart description language for epson-compatible printers. C Users Journal, 6(8): 63–??, August 1988. ISSN 0898-9788.

Patton:1988:PLM

[Pat88b] Robert L. Patton. Plotting language makes dot matrix graphics more accessible. C Users Journal, 6(8):69–??, August 1988. ISSN 0898-9788.

Anonymous:1992:TCE

[Pat92a] Stephen Patten. C Express:

250+ Ready-To-Run Assembly-Language Routines for Turbo C, Microsoft. C Users Journal, 10 (4):95–??, April 1992. ISSN 0898-9788.

Patton:1992:LSS

[Pat92b] Robert L. Patton. On the lack of a standard screen-clearing function. C Users Journal, 10(6):58– ??, June 1992. ISSN 0898-9788.

Patapis:1995:EGR

[Pat95] George Patapis. Design Patterns, Elements of Reusable Object-Oriented Software by Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. C/C++ Users Journal, 13(10): 78–??, October 1995. CODEN CCUJEX. ISSN 1075-2838. See [Gam95].

Patapi:1996:OCL

[Pat96] George Patapi. The OSE C++ libraries. C/C++ Users Journal, 14(4):19–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Paternostro:1999:JSC

[Pat99] Pat Paternostro. A Java skeleton code generator. C/C++ Users Journal, 17(7):34, 36–38, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Pavlovich:1989:TCP

[Pav89] Donald Pavlovich. C Primer Plus, by Mitchell Waite, Stephen Prata, and Donald Martin. C Users Journal, 7(7):39–??, July 1989. ISSN 0898-9788. See [WPM84].

Pavlov:1990:TCV

[Pav90] Alexander Vladimirovich Pavlov. Turbo C v2.0 bug. C Users Journal, 8(4):139–??, April 1990. ISSN 0898-9788.

Pugh:1991:DBE

[PB91] Kenneth Pugh and Garrett J. Boni. Defining ‘{’ As BEGIN and ‘}’ As END With #define. C Users Journal, 9(3):103–??, March 1991. ISSN 0898-9788.

Plauger:1992:PDS

[PB92] P. J. Plauger and John Beach. Passing data to a spawned process. C Users Journal, 10(1):142– ??, January 1992. ISSN 0898-9788.

Plauger:1994:UTV

[PB94] P. J. Plauger and David Brown.

Using void *. C/C++ Users Journal, 12(7):103–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Passe:1991:CCC

[PBBD91] Steve Passe, Matthew Brandt, Brian Brown, and Stuart Dole. CUG338 — 68000 C compiler and assembler. C Users Journal, 9(4): 110–??, April 1991. ISSN 0898-9788.

Pugh:1993:CDG

[PBS93] Kenneth Pugh, Walter Beck, and Fred C. Smith. Check digit generation algorithms. C Users Journal, 11(3):88–??, March 1993. ISSN 0898-9788.

Plauger:1992:SNS

[PBW92] P. J. Plauger, Elvin E. Birth, and Robert L. Ward. Some notes on software copyrights. C Users Journal, 10(8):140–??, August 1992. ISSN 0898-9788.

Pugh:1991:TFV

[PC91] Kenneth Pugh and Hugo Calleens. floats versus doubles. C Users Journal, 9(6):117–??, June 1991. ISSN 0898-9788.

Pugh:1992:RFR

[PC92] Kenneth Pugh and Tom Crosman. Reading file records into structures. C Users Journal, 10 (4):105–??, April 1992. ISSN 0898-9788.

Pugh:1993:SLL

[PC93] Kenneth Pugh and David Chapman. Sorting linked lists. C Users Journal, 11(1):103–??, January 1993. ISSN 0898-9788.

Passe:1988:CCA

[PD88] Steve Passe and Stuart Dole. CUG261 — 68000 cross assembler under MS-DOS. C Users Journal, 6(3):65–??, March 1988. ISSN 0898-9788.

Pugh:1993:MD

[PD93] Kenneth Pugh and David A. Dennerline. Macros and debugging. C Users Journal, 11(2): 115–??, February 1993. ISSN 0898-9788.

Pugh:1992:PLP

[PdW92] Kenneth Pugh and Gerrit M. de Wit. Printing to laser printers with C. C Users Journal, 10 (6):57–??, June 1992. ISSN 0898-9788.

Pearson:1990:CMC

[Pea90] Gregory Pearson. CUG314 — MNP C library. C Users Journal, 8(6):127–??, June 1990. ISSN 0898-9788.

Peckham:1989:CCC

[Pec89] Martin E. Peckham. Colley CUG cross assemblers compile on Apple IIgs. C Users Journal, 7(7): 122–??, July 1989. ISSN 0898-9788.

Penoyer:1991:DPC

[Pen91] Robert H. Penoyer. Defining pi in C. C Users Journal, 9(7):136–??, July 1991. ISSN 0898-9788.

Penoyer:1993:AF

[Pen93] Robert H. Penoyer. The alphabeta filter. C Users Journal, 11 (7):73–??, July 1993. ISSN 0898-9788.

Perkins:1989:CPG

[Per89] W. Michael Perkins. C Programmer’s Guide to NetBIOS, by W. David Schwaderer. C Users Journal, 7(4):29–??, April 1989. ISSN 0898-9788. See [Sch88c].

Percy:1992:SA

[Per92a] Michael Percy. Simulated annealing. C Users Journal, 10(7):71– ??, July 1992. ISSN 0898-9788.

Perry:1992:CEa

[Per92b] Greg M. Perry. C by example.

Programming series. Que Corporation, Indianapolis, IN, USA, 1992. ISBN 0-88022-813-X. xix + 754 pp. LCCN QA76.73.C15 P45 1992. US$21.95, CDN$28.95. See book review [Tho92].

Perry:1992:CEb

[Per92c] Greg M. Perry. C++ by example. Programming series. Que Corporation, Indianapolis, IN, USA, 1992. ISBN 1-56529-038-0. xxi + 790 pp. LCCN QA 76.73 C153 P443 1992. US$24.95, CDN$31.95.

Persi:1999:VCG

[Per99] Fred Persi. A vector class for geometric computations. C/C++ Users Journal, 17(8):62–??, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Pescio:1997:BCU

[Pes97a]  Carlo Pescio.  Binary con-

stants using template metaprogramming. C/C++ Users Journal, 15(2):51–??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Pescio:1997:STA

[Pes97b] Carlo Pescio. Stack trace assertions using COFF. C/C++ Users Journal, 15(6):41–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Pescio:1998:STE

[Pes98] Carlo Pescio. Stack trace exceptions in Win32. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Peters:1988:UPM

[Pet88] James F. Peters. UNIX programming: methods and tools. Harcourt Brace Jovanovich, San Diego, CA, USA, 1988. ISBN 0-15-593021-4 (paperback). xv + 447 pp. LCCN QA76.76.O63 P52 1988. See book review [Fla90]. System requirements for computer disk: IBM PC.

Petz:1998:CTI

[Pet98] John Petz. A C++ telephony interface. C/C++ Users Journal, 16(10):55–??, October 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Peterson:1999:MSM

[Pet99] Mark C. Peterson. Manipulating sparse matrices. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1992:IA

[PF92a] Kenneth Pugh and Bruno Fassino. Initializing arrays. C Users Journal, 10(6):55–??, June 1992. ISSN 0898-9788.

Pugh:1992:PMM

[PF92b] Kenneth Pugh and Tom Frank.

PC memory model dependencies with %u in printf statements. C Users Journal, 10(3):62–??, March 1992. ISSN 0898-9788.

Plauger:1993:RBD

[PF93a] P. J. Plauger and David Fong.

Reading binary data exceeding 7F using char. C Users Journal, 11(4):125–??, April 1993. ISSN 0898-9788.

Pugh:1993:I

[PF93b] Kenneth Pugh and Sean Furlong.

Identifiers. C Users Journal, 11 (5):91–??, May 1993. ISSN 0898-9788.

Plauger:1994:FFC

[PF94] P. J. Plauger and R. W. J.

Ford. File-handling functions in C: Some questions and answers. C Users Journal, 12(1):119–??, January 1994. ISSN 0898-9788.

Plauger:1991:MEW

[PG91] P. J. Plauger and John K. Gotwals. Multiple exits from within a C code block. C Users Journal, 9(10):137–??, October 1991. ISSN 0898-9788.

Pugh:1992:SHG

[PG92] Kenneth Pugh and James A. Gant. Saving Hercules graphics screens to disk. C Users Journal, 10(3):63–??, March 1992. ISSN 0898-9788.

Pugh:1993:MOT

[PG93] Kenneth Pugh and Stefan Ganev. Multi-lingual output techniques. C Users Journal, 11(4):101–??, April 1993. ISSN 0898-9788.

Pugh:1990:HTD

[PH90] Kenneth Pugh and James Howell. Handling of #define’s by compilers. C Users Journal, 8(7):93–??, July 1990. ISSN 0898-9788.

Pugh:1991:DAM

[PH91a] Kenneth Pugh and Lyle O. Haga. Data alignment in memory. C Users Journal, 9(7):107–??, July 1991. ISSN 0898-9788.

Pugh:1991:OFS

[PH91b] Kenneth Pugh and Lyle O. Haga. Order of functions in source code. C Users Journal, 9(7):106–??, July 1991. ISSN 0898-9788.

Plauger:1993:PUT

[PH93] P. J. Plauger and Sheldon Hoffman. The proper use of write. C Users Journal, 11(7):138–??, July 1993. ISSN 0898-9788.

Perelman-Hall:1996:OSC

[PH96] David Perelman-Hall. Objectoriented SDI-12 communications. C/C++ Users Journal, 14(1):33– ??, January 1996. CODEN CCU-JEX. ISSN 1075-2838.

Phillips:1990:BMU

[Phi90a] Dwayne Phillips. Belief maintenance using the Dempster-Shafer theory of evidence. C Users Journal, 8(3):67–??, March 1990. ISSN 0898-9788.

Phillips:1990:DTB

[Phi90b] Dwayne Phillips. Documentation tool builds action diagrams. C Users Journal, 8(6):107–??, June 1990. ISSN 0898-9788.

Phillips:1990:TNR

[Phi90c]  Dwayne Phillips.  Numerical

Recipes in C — The Art of Scientific Computing, by Press, et al. C Users Journal, 8(12):103– ??, December 1990. ISSN 0898-9788. See [Pre88].

Phillips:1991:CCI

[Phi91a] Dwayne Phillips. CUG343 — C image processing system. C Users Journal, 9(5):135–??, May 1991. ISSN 0898-9788.

Phillips:1991:IPPa

[Phi91b] Dwayne Phillips. Image processing — part 1: Reading the Tag Image File Format. C Users Journal, 9(3):92–??, March 1991. ISSN 0898-9788.

Phillips:1991:IPPb

[Phi91c] Dwayne Phillips. Image processing — part 2: Displaying images and printing numbers. C Users Journal, 9(5):117–??, May 1991. ISSN 0898-9788.

Phillips:1991:IPPc

[Phi91d] Dwayne Phillips. Image processing — part 3: Displaying and printing images using halftoning. C Users Journal, 9(6):89–??, June 1991. ISSN 0898-9788.

Phillips:1991:IPPd

[Phi91e] Dwayne Phillips. Image processing — part 4: Histograms and histogram equalization. C Users Journal, 9(8):59–??, August 1991. ISSN 0898-9788.

Phillips:1991:IPPe

[Phi91f] Dwayne Phillips. Image processing — part 5: Writing images to files and basic edge detection. C Users Journal, 9(11):75– ??, November 1991. ISSN 0898-9788.

Phillips:1992:CCP

[Phi92a] Dwayne Phillips. A $19.95 C compiler? — the power C compiler from MIX software. C Users Journal, 10(9):53–??, September 1992. ISSN 0898-9788.

Phillips:1992:DCU

[Phi92b] Dwayne Phillips. Data compression using Huffman coding. C Users Journal, 10(2):55–??, February 1992. ISSN 0898-9788.

Phillips:1992:TNN

[Phi92c] Dwayne Phillips. Neural Networks in C++, by Adam Blum. C Users Journal, 10(10):101–??, October 1992. ISSN 0898-9788. See [Blu92].

Phillips:1992:FNN

[Phi92d] Dwayne Phillips. The foundation of neural networks: The adaline and madaline. C Users Journal, 10(9):69–??, September 1992. ISSN 0898-9788.

Phillips:1992:IPPa

[Phi92e] Dwayne Phillips. Image processing — part 6: Advanced edge detection. C Users Journal, 10(1): 47–??, January 1992. ISSN 0898-9788.

Phillips:1992:IPPb

[Phi92f] Dwayne Phillips. Image processing — part 7: Spatial frequency filtering. C Users Journal, 10 (10):71–??, October 1992. ISSN 0898-9788.

Phillips:1992:IPPc

[Phi92g] Dwayne Phillips. Image processing — part 8: Image operations. C Users Journal, 10(11):89–??, November 1992. ISSN 0898-9788.

Phillips:1993:CIP

[Phi93a] Dwayne Phillips. C image processing — part 11: Manipulating shapes. C Users Journal, 11(8): 59–??, August 1993. ISSN 0898-9788.

Phillips:1993:TCPb

[Phi93b] Dwayne Phillips. C Programming Guidelines, 2nd edition, by Thomas Plum. C Users Journal, 11(1):119–??, January 1993. ISSN 0898-9788.

Phillips:1993:TCPa

[Phi93c] Dwayne Phillips. C++ Programming Guidelines, by Thomas Plum and Dan Saks. C Users Journal, 11(1):119–??, January 1993. ISSN 0898-9788.

Phillips:1993:IPPb

[Phi93d] Dwayne Phillips. Image processing — part 10: Segmentation using edges and gray shades. C Users Journal, 11(6):67–??, June 1993. ISSN 0898-9788.

Phillips:1993:IPPa

[Phi93e] Dwayne Phillips. Image processing — part 9: Histogram-based image segmentation. C Users Journal, 11(2):63–??, February 1993. ISSN 0898-9788.

Phillips:1993:RCS

[Phi93f] Dwayne Phillips. A revision control system for MS-DOS. C Users Journal, 11(7):33–??, July 1993. ISSN 0898-9788.

Phillips:1994:TCE

[Phi94a] Dwayne Phillips. C Elements of Style: The Programmer’s Style Manual for Elegant C and C++ Programs, by Steve Oualline. C Users Journal, 12(3):115–??, March 1994. ISSN 0898-9788. See [Oua92].

Phillips:1994:IPC

[Phi94b] Dwayne Phillips. Image processing in C, part 12: Boolean and overlay operations. C/C++ Users Journal, 12(11):53–??, November 1994. CODEN CCU-JEX. ISSN 1075-2838.

Phillips:1994:ASM

[Phi94c] Thomas Phillips. Approximate string matching. C Users Journal, 12(4):77–??, April 1994. ISSN 0898-9788.

Phillips:1995:IPCa

[Phi95a] Dwayne Phillips. Image processing in C, part 13: Geometric operations. C/C++ Users Journal, 13(8):23–??, August 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Phillips:1995:IPCb

[Phi95b] Dwayne Phillips. Image processing in C, part 14: Warping and morphing. C/C++ Users Journal, 13(10):55–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1995:IPCc

[Phi95c] Dwayne Phillips. Image processing in C, part 15: Basic texture operations. C/C++ Users Journal, 13(11):55–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1996:CCD

[Phi96a] Andrew Phillips. C++ and C debugging, testing and reliability. C/C++ Users Journal, 14 (10):19–??, October 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Phillips:1996:BNN

[Phi96b] Dwayne Phillips. The backpropagation neural network. C/C++ Users Journal, 14(1):49–??, January 1996. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1996:DSE

[Phi96c] Dwayne Phillips. A Discipline for Software Engineering, by Watts S. Humphrey. C/C++ Users Journal, 14(5):61–??, May 1996. CODEN CCUJEX. ISSN 1075-2838. See [Hum95].

Phillips:1996:IPC

[Phi96d] Dwayne Phillips. Image processing in C, part 16, random dot stereograms. C/C++ Users Journal, 14(4):53–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1998:DMP

[Phi98a]  Dwayne Phillips.  Decision-

making with production systems. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Phillips:1998:IHC

[Phi98b] Dwayne Phillips. Information hiding in C via modular programming. C/C++ Users Journal, 16 (1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1998:SHI

[Phi98c] Dwayne Phillips. Steganography: Hiding information in plain sight. C/C++ Users Journal, 16(11): 49–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1999:CSR

[Phi99a] Andrew Phillips. A container for a set of ranges. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1999:BRP

[Phi99b] Dwayne Phillips. Book review: The Practice of Programming. C/C++ Users Journal, 17(8):69– ??, August 1999. CODEN CCU-JEX. ISSN 1075-2838.

Phillips:1999:SSA

[Phi99c] Dwayne Phillips. Searching for solutions with AI. C/C++ Users Journal, 17(9):18–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Phillips:1999:VDI

[Phi99d] Dwayne Phillips. Visualizing depth images. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Pickover:1990:CPC

[Pic90] Clifford A. Pickover. Computers, pattern, chaos, and beauty: graphics from an unseen world. St. Martin’s Press, New York, NY, USA, 1990. ISBN 0-312-04123-3. xvi + 394 + 10 pp. LCCN T385 .P5 1990. US$29.95. See book review [Hob91b].

Pierce:1988:ACC

[Pie88] Victor Pierce. Availability of C compiler for Concurrent DOS. C Users Journal, 6(9):12–??, September 1988. ISSN 0898-9788.

Pinson:1991:CDS

[Pin91] James Pinson. CUG337 — designing screen interfaces in C. C Users Journal, 9(4):110–??, April 1991. ISSN 0898-9788.

Pinkerton:1993:WMC

[Pin93]  Kevin Pinkerton.  A Win-

dows and MS-DOS C++ output stream for a secondary monitor. C Users Journal, 11(7):25– ??, July 1993. ISSN 0898-9788.

Pipkins:1994:NSW

[Pip94] Jeffrey D. Pipkins. Notes on some weaknesses of Standard C. C/C++ Users Journal, 12(8): 102–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Pipkins:1995:EB

[Pip95] Jeff D. Pipkins. Embedding on a budget. C/C++ Users Journal, 13(2):25–??, February 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Pugh:1992:FVL

[PJ92] Kenneth Pugh and Brett Johnson. Functions with variable length argument lists to process strings. C Users Journal, 10(12): 109–??, December 1992. ISSN 0898-9788.

Pugh:1992:WFD

[PK92] Kenneth Pugh and Harold Kilts. Writing to floppy disks without a DOS. C Users Journal, 10(12): 109–??, December 1992. ISSN 0898-9788.

Plauger:1988:DFS

[Pla88a] P. J. Plauger. Declaring functions in Standard C. C Users Journal, 6(1):15–??, January 1988. ISSN 0898-9788.

Plauger:1988:HSC

[Pla88b] P. J. Plauger. How Standard C groups subexpressions. C Users Journal, 6(9):39–??, September 1988. ISSN 0898-9788.

Plauger:1988:RCE

[Pla88c] P. J. Plauger. The rules controlling expression types have changed somewhat under Standard C. C Users Journal, 6(11): 17–??, November 1988. ISSN 0898-9788.

Plauger:1988:SCP

[Pla88d] P. J. Plauger. Standard C promotes types according to value preserving rules. C Users Journal, 6(8):9–??, August 1988. ISSN 0898-9788.

Plauger:1988:SCD

[Pla88e] P. J. Plauger. Standard C’s ‘doeverything’ expressions form four distinct classes. C Users Journal, 6(6):27–??, June 1988. ISSN 0898-9788.

Plauger:1988:TMS

[Pla88f] P. J. Plauger. Touching memory: Standard C makes the act more precise. C Users Journal, 6 (5):25–??, May 1988. ISSN 0898-9788.

Plauger:1988:TPC

[Pla88g] P. J. Plauger. Types play central role in new Standard C. C Users Journal, 6(3):17–??, March 1988. ISSN 0898-9788.

Plauger:1988:WN

[Pla88h] P. J. Plauger. What’s in a name? C Users Journal, 6(2):21– ??, February 1988. ISSN 0898-9788.

Plauger:1989:CT

[Pla89a] P. J. Plauger. Conditional translation. C Users Journal, 7(7):31– ??, July 1989. ISSN 0898-9788.

Plauger:1989:DEM

[Pla89b] P. J. Plauger. Defining and expanding macros. C Users Journal, 7(5):13–??, May 1989. ISSN 0898-9788.

Plauger:1989:ECI

[Pla89c] P. J. Plauger. Evolution of the C I/O model. C Users Journal, 7(8):17–??, August 1989. ISSN 0898-9788.

Plauger:1989:FSC

[Pla89d] P. J. Plauger. Flow Standard C supports multi-byte character sets. C Users Journal, 7(2):13– ??, February 1989. ISSN 0898-9788.

Plauger:1989:FO

[Pla89e] P. J. Plauger. Formatted output. C Users Journal, 7(11):37– ??, November 1989. ISSN 0898-9788.

Plauger:1989:PT

[Pla89f] P. J. Plauger. Phases of translation. C Users Journal, 7(4):21– ??, April 1989. ISSN 0898-9788.

Plauger:1989:SCA

[Pla89g] P. J. Plauger. Standard C addresses challenge of the international marketplace. C Users Journal, 7(1):9–??, January 1989. ISSN 0898-9788.

Plauger:1989:SIP

[Pla89h] P. J. Plauger. Streams and I/O primitives in Standard C. C Users Journal, 7(10):9–??, October 1989. ISSN 0898-9788.

Plauger:1990:CCF

[Pla90a] P. J. Plauger. Character classification functions. C Users Journal, 8(10):9–??, October 1990. ISSN 0898-9788.

Plauger:1990:DNC

[Pla90b] P. J. Plauger. Depicting nondisplayable characters. C Users Journal, 8(11):130–??, November 1990. ISSN 0898-9788.

Plauger:1990:TCC

[Pla90c] P. J. Plauger. Concurrent C Programming Language, The, by Narain Gehani and William D. Roome. C Users Journal, 8(5): 96–??, May 1990. ISSN 0898-9788. See [GR89].

Plauger:1990:TPPb

[Pla90d] P. J. Plauger. Portable Programs for Parallel Processors. C Users Journal, 8(5):96–??, May 1990. ISSN 0898-9788.

Plauger:1990:TPPa

[Pla90e] P. J. Plauger. Programming Parallel Processors, by Robert G. Babb II. C Users Journal, 8(5): 95–??, May 1990. ISSN 0898-9788. See [Bab87, Bab88].

Plauger:1990:TE

[Pla90f] P. J. Plauger. <errno.h>. C Users Journal, 8(12):9–??, December 1990. ISSN 0898-9788.

Plauger:1990:FI

[Pla90g] P. J. Plauger. Formatted input.

C Users Journal, 8(1):9–??, January 1990. ISSN 0898-9788.

Plauger:1990:BE

[Pla90h] P. J. Plauger. I begin as editor. C Users Journal, 8(11):6–??, November 1990. ISSN 0898-9788.

Plauger:1990:ITC

[Pla90i] P. J. Plauger. Implementing <ctype.h>. C Users Journal, 8 (11):9–??, November 1990. ISSN 0898-9788.

Plauger:1990:IN

[Pla90j] P. J. Plauger. Interpreting the nasties. C Users Journal, 8(7): 9–??, July 1990. ISSN 0898-9788.

Plauger:1990:LGR

[Pla90k] P. J. Plauger. Library ground rules. C Users Journal, 8(8):9–??, August 1990. ISSN 0898-9788.

Plauger:1990:MI

[Pla90l] P. J. Plauger. A matter of interpretation. C Users Journal, 8(6): 9–??, June 1990. ISSN 0898-9788.

Plauger:1990:NSP

[Pla90m] P. J. Plauger. Note on standardizing parallel programming in C. C Users Journal, 8(5):95– ??, May 1990. ISSN 0898-9788.

Plauger:1990:QCPa

[Pla90n] P. J. Plauger. Quiet changes, part I. C Users Journal, 8(2):9–??, February 1990. ISSN 0898-9788.

Plauger:1990:QCPb

[Pla90o] P. J. Plauger. Quiet changes, part II. C Users Journal, 8(3):9–??, March 1990. ISSN 0898-9788.

Plauger:1990:WGAa

[Pla90p] P. J. Plauger. Wha gang agley.

C Users Journal, 8(4):9–??, April 1990. ISSN 0898-9788.

Plauger:1990:WGAb

[Pla90q] P. J. Plauger. Wha gang agley — part II. C Users Journal, 8(5):9– ??, May 1990. ISSN 0898-9788.

Plauger:1990:GR

[Pla90r] P. J. Plauger. With gun and reel. C Users Journal, 8(9):9– ??, September 1990. ISSN 0898-9788.

Plauger:1991:ACI

[Pla91a] P. J. Plauger. ANSI C and ISO C:

‘standard C’? C Users Journal, 9(1):6–??, January 1991. ISSN 0898-9788.

Plauger:1991:BYO

[Pla91b] P. J. Plauger. Build your own locales. C Users Journal, 9(5):8–??, May 1991. ISSN 0898-9788.

Plauger:1991:CSS

[Pla91c] P. J. Plauger. A character set standard: The fight. C Users Journal, 9(11):6–??, November 1991. ISSN 0898-9788.

Plauger:1991:TCUa

[Pla91d] P. J. Plauger. C Users Journal conversion to C++ Users Journal not imminent. C Users Journal, 9(5):6–??, May 1991. ISSN 0898-9788.

Plauger:1991:TCUb

[Pla91e] P. J. Plauger. C Users Journal too ‘academic’? C Users Journal, 9(9):6–??, September 1991. ISSN 0898-9788.

Plauger:1991:FMV

[Pla91f] P. J. Plauger. Formatting monetary values. C Users Journal, 9 (6):9–??, June 1991. ISSN 0898-9788.

Plauger:1991:HTF

[Pla91g] P. J. Plauger. The header <float.h>. C Users Journal, 9 (1):9–??, January 1991. ISSN 0898-9788.

Plauger:1991:HTLa

[Pla91h] P. J. Plauger. The header <limits.h>. C Users Journal, 9(2):7–??, February 1991. ISSN 0898-9788.

Plauger:1991:HTLb

[Pla91i] P. J. Plauger. The header <locale.h>. C Users Journal, 9 (3):7–??, March 1991. ISSN 0898-9788.

Plauger:1991:HTSa

[Pla91j] P. J. Plauger. The header <setjmp.h>. C Users Journal, 9(10):9–??, October 1991. ISSN 0898-9788.

Plauger:1991:HTSb

[Pla91k] P. J. Plauger. The header <stdarg.h>. C Users Journal, 9 (11):9–??, November 1991. ISSN 0898-9788.

Plauger:1991:HTSc

[Pla91l] P. J. Plauger. The header <stddef.h>. C Users Journal, 9 (12):9–??, December 1991. ISSN 0898-9788.

Plauger:1991:ITL

[Pla91m] P. J. Plauger. Implementing <locale.h>. C Users Journal, 9 (4):7–??, April 1991. ISSN 0898-9788.

Plauger:1991:MF

[Pla91n] P. J. Plauger. Math functions.

C Users Journal, 9(7):8–??, July 1991. ISSN 0898-9788.

Plauger:1991:MP

[Pla91o] P. J. Plauger. Math primitives. C Users Journal, 9(8):9–??, August 1991. ISSN 0898-9788.

Plauger:1991:MS

[Pla91p] P. J. Plauger. A math sampler. C Users Journal, 9(9):8– ??, September 1991. ISSN 0898-9788.

Plauger:1992:B

[Pla92a] P. J. Plauger. Bugs. C Users Journal, 10(9):8–??, September 1992. ISSN 0898-9788.

Plauger:1992:CE

[Pla92b] P. J. Plauger. C and the environment. C Users Journal, 10(8):8– ??, August 1992. ISSN 0898-9788.

Plauger:1992:CS

[Pla92c] P. J. Plauger. Comparing strings.

C Users Journal, 10(11):8–??, November 1992. ISSN 0898-9788.

Plauger:1992:TJC

[Pla92d] P. J. Plauger. Journal of C Language Translation ceases publication. C Users Journal, 10(6):6– ??, June 1992. ISSN 0898-9788.

Plauger:1992:HTSa

[Pla92e] P. J. Plauger. The header <signal.h>. C Users Journal, 10 (3):8–??, March 1992. ISSN 0898-9788.

Plauger:1992:HTSb

[Pla92f] P. J. Plauger. The header <stdlib.h>. C Users Journal, 10 (4):8–??, April 1992. ISSN 0898-9788.

Plauger:1992:HTSc

[Pla92g] P. J. Plauger. The header <string.h>. C Users Journal, 10(10):8–??, October 1992. ISSN 0898-9788.

Plauger:1992:ITS

[Pla92h] P. J. Plauger. Implementing <stdio.h>. C Users Journal, 10 (1):8–??, January 1992. ISSN 0898-9788.

Plauger:1992:MF

[Pla92i] P. J. Plauger. Multibyte functions. C Users Journal, 10(7):8– ??, July 1992. ISSN 0898-9788.

Plauger:1992:NCS

[Pla92j] P. J. Plauger. Notes on C++ standardization efforts. C Users Journal, 10(8):6–??, August 1992. ISSN 0898-9788.

Plauger:1992:PTS

[Pla92k] P. J. Plauger. Primitives for <stdio.h>. C Users Journal, 10 (2):8–??, February 1992. ISSN 0898-9788.

Plauger:1992:SS

[Pla92l] P. J. Plauger. Searching strings.

C Users Journal, 10(12):8–??, December 1992. ISSN 0898-9788.

Plauger:1992:SMV

[Pla92m] P. J. Plauger. Source of machinereadable version of the ANSI C standard. C Users Journal, 10(11):144–??, November 1992. ISSN 0898-9788.

Plauger:1992:SAF

[Pla92n] P. J. Plauger. Storage allocation functions. C Users Journal, 10 (6):8–??, June 1992. ISSN 0898-9788.

Plauger:1992:TNC

[Pla92o] P. J. Plauger. Text to numeric conversions. C Users Journal, 10 (5):8–??, May 1992. ISSN 0898-9788.

Plauger:1993:MCS

[Pla93a] Dave Plauger. Making C++ save for threads. C Users Journal, 11 (2):58–??, February 1993. ISSN 0898-9788.

Plauger:1993:AIO

[Pla93b] P. J. Plauger. Applying the increment operator to rvalues. C Users Journal, 11(3):121–??, March 1993. ISSN 0898-9788.

Plauger:1993:CLG

[Pla93c] P. J. Plauger. C++ library ground rules. C Users Journal, 11 (11):10–??, November 1993. ISSN 0898-9788.

Plauger:1993:CLC

[Pla93d] P. J. Plauger. The C library in C++. C Users Journal, 11 (12):10–??, December 1993. ISSN 0898-9788.

Plauger:1993:CIF

[Pla93e] P. J. Plauger. CBEMA increases fees for standards volunteers. C Users Journal, 11(1):8–??, January 1993. ISSN 0898-9788.

Plauger:1993:DSC

[Pla93f] P. J. Plauger. Developing the Standard C++ library. C Users Journal, 11(10):10–??, October 1993. ISSN 0898-9788.

Plauger:1993:TTC

[Pla93g] P. J. Plauger. Time and the Calendars, by William Matthew O’Neill. C Users Journal, 11(1): 20–??, January 1993. ISSN 0898-9788. See [O’N75, O’N78].

Plauger:1993:FCE

[Pla93h] P. J. Plauger. Floating-point C extensions. C Users Journal, 11 (9):10–??, September 1993. ISSN 0898-9788.

Plauger:1993:FCC

[Pla93i] P. J. Plauger. Formal changes to C. C Users Journal, 11(4):10–??, April 1993. ISSN 0898-9788.

Plauger:1993:HTT

[Pla93j] P. J. Plauger. The header <time.h>. C Users Journal, 11 (1):10–??, January 1993. ISSN 0898-9788.

Plauger:1993:KNT

[Pla93k] P. J. Plauger. Keeping the name of C Users Journal. C Users Journal, 11(2):8–??, February 1993. ISSN 0898-9788.

Plauger:1993:LCSb

[Pla93l] P. J. Plauger. Large character set functions. C Users Journal, 11 (6):10–??, June 1993. ISSN 0898-9788.

Plauger:1993:LCSa

[Pla93m] P. J. Plauger. Large character set support. C Users Journal, 11 (5):10–??, May 1993. ISSN 0898-9788.

Plauger:1993:NCE

[Pla93n] P. J. Plauger. Numerical C extensions group. C Users Journal, 11(8):10–??, August 1993. ISSN 0898-9788.

Plauger:1993:TCF

[Pla93o] P. J. Plauger. Time conversion functions. C Users Journal, 11 (2):10–??, February 1993. ISSN 0898-9788.

Plauger:1993:TFF

[Pla93p] P. J. Plauger. Time formatting functions. C Users Journal, 11 (3):10–??, March 1993. ISSN 0898-9788.

Plauger:1993:WCS

[Pla93q] P. J. Plauger. Wide character streams. C Users Journal, 11 (7):10–??, July 1993. ISSN 0898-

Plauger:1994:CLS

[Pla94a] P. J. Plauger. C++ language support library. C Users Journal, 12(3):10–??, March 1994. ISSN 0898-9788.

Plauger:1994:TCM

[Pla94b] P. J. Plauger. C++ Memory Management, by Len Dorfman and Marc J. Neuberger. C Users Journal, 12(2):95–??, February 1994. ISSN 0898-9788. See [DN93].

Plauger:1994:TCU

[Pla94c] P. J. Plauger. C Users Journal to become C/C++ Users Journal. C Users Journal, 12(6):8–??, June 1994. ISSN 0898-9788.

Plauger:1994:TSI

[Pla94d] P. J. Plauger. Software Internationalization and Localization: an Introduction, by E. Uren, R. Howard, and T. Perinotti. C Users Journal, 12(5):105–??, May 1994. ISSN 0898-9788. See [UHP93].

Plauger:1994:TAA

[Pla94e] P. J. Plauger. The Annotated ANSI C Standard, annotated by Herbert Schildt. C Users Journal, 12(1):34–??, January 1994. ISSN 0898-9788. See [Sch90b].

Plauger:1994:E

[Pla94f]  P. J. Plauger.  Extractors.

C/C++ Users Journal, 12(8):10– ??, August 1994. CODEN CCU-JEX. ISSN 1075-2838.

Plauger:1994:HTE

[Pla94g] P. J. Plauger. The header <exception>. C Users Journal, 12(2):10–??, February 1994. ISSN 0898-9788.

Plauger:1994:HTIc

[Pla94h] P. J. Plauger. The header iomanip. C/C++ Users Journal, 12(12):10–??, December 1994. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1994:HTIa

[Pla94i] P. J. Plauger. The header <ios>.

C Users Journal, 12(5):10–??, May 1994. ISSN 0898-9788.

Plauger:1994:HTIb

[Pla94j] P. J. Plauger. The header <istream>. C/C++ Users Journal, 12(7):10–??, July 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1994:HTO

[Pla94k] P. J. Plauger. The header <ostream>. C/C++ Users Journal, 12(9):10–??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1994:HTS

[Pla94l] P. J. Plauger. The header <streambuf>. C Users Journal, 12(6):10–??, June 1994. ISSN 0898-9788.

Plauger:1994:ITO

[Pla94m] P. J. Plauger. Implementing ostream. C/C++ Users Journal, 12(11):10–??, November 1994. CODEN CCUJEX. ISSN 1075-

Plauger:1994:IDC

[Pla94n] P. J. Plauger. Informal draft of C++ standard imminent. C Users Journal, 12(1):8–??, January 1994. ISSN 0898-9788.

Plauger:1994:I

[Pla94o] P. J. Plauger. Inserters. C/C++ Users Journal, 12(10):10–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1994:II

[Pla94p] P. J. Plauger. Introduction to iostreams. C Users Journal, 12 (4):10–??, April 1994. ISSN 0898-9788.

Plauger:1994:RCS

[Pla94q] P. J. Plauger. Revision of the C standard now underway. C Users Journal, 12(2):8–??, February 1994. ISSN 0898-9788.

Plauger:1994:TC

[Pla94r] P. J. Plauger. Technical corrigendum 1. C Users Journal, 12(1): 10–??, January 1994. ISSN 0898-9788.

Plauger:1994:TST

[Pla94s] P. J. Plauger. Thoughts on software testing. C/C++ Users Journal, 12(8):8–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1994:WTC

[Pla94t] P. J. Plauger. Welcome to C/C++ Users Journal. C/C++ Users Journal, 12(7):8–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:TBC

[Pla95a] P. J. Plauger. The Best C/C++ Tips Ever by Anthony Porter. C/C++ Users Journal, 13(1):67– ??, January 1995. CODEN CCU-JEX. ISSN 1075-2838. See [Por93].

Plauger:1995:HTB

[Pla95b] P. J. Plauger. The header bitset. C/C++ Users Journal, 13(9):8–??, September 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1995:HTC

[Pla95c] P. J. Plauger. The header complex. C/C++ Users Journal, 13(10):8–??, October 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1995:HTF

[Pla95d] P. J. Plauger. The header fstream. C/C++ Users Journal, 13(4):8–??, April 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:HTI

[Pla95e] P. J. Plauger. The header iostream. C/C++ Users Journal, 13(6):8–??, June 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1995:HTSb

[Pla95f] P. J. Plauger. The header sstream. C/C++ Users Journal, 13(3):10–??, March 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1995:HTSc

[Pla95g] P. J. Plauger. The header string. C/C++ Users Journal, 13(7):8–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:HTSa

[Pla95h] P. J. Plauger. The header strstream. C/C++ Users Journal, 13(1):10–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:ITC

[Pla95i] P. J. Plauger. Implementing complex. C/C++ Users Journal, 13(11):8–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:ITF

[Pla95j] P. J. Plauger. Implementing fstream. C/C++ Users Journal, 13(5):8–??, May 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:ITSb

[Pla95k] P. J. Plauger. Implementing string. C/C++ Users Journal, 13(8):8–??, August 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1995:ITSa

[Pla95l] P. J. Plauger. Implementing strstream. C/C++ Users Journal, 13(2):10–??, February 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1995:STL

[Pla95m] P. J. Plauger. The standard template library. C/C++ Users Journal, 13(12):10–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:CCEb

[Pla96a] P. J. Plauger. C/C++ Editor’s forum: Transliterating C and C++ code to Java. C/C++ Users Journal, 14(9):6, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:CCEa

[Pla96b] P. J. Plauger. C/C++ Editor’s forum: Uproar over Java. C/C++ Users Journal, 14(5):6, May 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:HTIa

[Pla96c] P. J. Plauger. The header iterator, part 1. C/C++ Users Journal, 14(4):8–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:HTIb

[Pla96d] P. J. Plauger. The header iterator, part 2. C/C++ Users Journal, 14(5):8–??, May 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:HTU

[Pla96e] P. J. Plauger. The header utility. C/C++ Users Journal, 14(3):8–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:I

[Pla96f] P. J. Plauger. Iterators. C/C++ Users Journal, 14(2):8–??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SCAb

[Pla96g] P. J. Plauger. Standard C/C++:

Algorithms. C/C++ Users Journal, 14(8):8–??, August 1996. CODEN CCUJEX. ISSN 1075-

Plauger:1996:SCAa

[Pla96h] P. J. Plauger. Standard C/C++: Allocators. C/C++ Users Journal, 14(6):8–??, June 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SCH

[Pla96i] P. J. Plauger. Standard C/C++: The header memory. C/C++ Users Journal, 14(7):8–??, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SCI

[Pla96j] P. J. Plauger. Standard C/C++: Introduction to <algorithm>. C/C++ Users Journal, 14(9):8– ??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SCL

[Pla96k] P. J. Plauger. Standard C/ C++: Last of the algorithms. C/C++ Users Journal, 14(11): 8–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SCO

[Pla96l] P. J. Plauger. Standard C/C++: Ordering algorithms. C/C++ Users Journal, 14(10):8–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SHO

[Pla96m] P. J. Plauger. STL headers — an overview. C/C++ Users Journal, 14(1):8–??, January 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1997:EFa

[Pla97a] P. J. Plauger. Editor’s forum. C/C++ Users Journal, 15(1):6, January 1997. CODEN CCU-JEX. ISSN 1075-2838.

Plauger:1997:EFb

[Pla97b] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(2):6, February 1997. CODEN CCU-JEX. ISSN 1075-2838.

Plauger:1997:EFc

[Pla97c] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(3):6– ??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Plauger:1997:EFd

[Pla97d] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:EFe

[Pla97e] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(5):6, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:EFf

[Pla97f] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(7):6, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:EFg

[Pla97g] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:EFh

[Pla97h] P. J. Plauger. Editor’s forum.

C/C++ Users Journal, 15(9):6, September 1997. CODEN CCU-JEX. ISSN 1075-2838.

Plauger:1997:EFJ

[Pla97i] P. J. Plauger. Editor’s forum: A Java to C translator. C/C++ Users Journal, 15(6): 6, June 1997. CODEN CCU-JEX. ISSN 1075-2838. URL

http://www.dinkumware.com/;
 http://www.tiac.net/users/
pjp.

Plauger:1997:EC

[Pla97j] P. J. Plauger. Embedded C++. C/C++ Users Journal, 15(2):35– ??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCd

[Pla97k] P. J. Plauger. Standard C/C++: Associative containers. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCf

[Pla97l] P. J. Plauger. Standard C/C++ container adapters. C/C++ Users Journal, 15(6):8–??, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCe

[Pla97m] P. J. Plauger. Standard C/C++: Implementing associative containers. C/C++ Users Journal, 15(5):8–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCg

[Pla97n] P. J. Plauger. Standard C/C++: Library update. C/C++ Users Journal, 15(7):8, 10, 12, 14, 16, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCc

[Pla97o] P. J. Plauger. Standard C/C++:

The header <deque>. C/C++ Users Journal, 15(3):8–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCi

[Pla97p]  P. J. Plauger.  Standard

C/C++: The header <limits>. C/C++ Users Journal, 15(9):8– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCb

[Pla97q] P. J. Plauger. Standard C/C++:

The header <list>. C/C++ Users Journal, 15(2):8–??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCh

[Pla97r] P. J. Plauger. Standard C/C++:

The header <valarray>. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1997:SCCa

[Pla97s] P. J. Plauger. Standard C/C++:

The header <vector>. C/C++ Users Journal, 15(1):8–??, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCj

[Pla98a] P. J. Plauger. Standard C/C++:

Exception safety in STL. C/C++ Users Journal, 16(10):10–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCk

[Pla98b] P. J. Plauger. Standard C/C++: Hash tables. C/C++ Users Journal, 16(11):10–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCi

[Pla98c] P. J. Plauger. Standard C/C++: Last of the facets. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCh

[Pla98d] P. J. Plauger. Standard C/C++: The facet collate. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCe

[Pla98e] P. J. Plauger. Standard C/C++: The facet money get. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCc

[Pla98f] P. J. Plauger. Standard C/C++: The facet moneypunct. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCd

[Pla98g] P. J. Plauger. Standard C/C++: The facet money put. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCb

[Pla98h] P. J. Plauger. Standard C/C++:

The facet num get. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCg

[Pla98i] P. J. Plauger. Standard C/C++:

The facet time get. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCf

[Pla98j] P. J. Plauger. Standard C/C++:

The facet time put. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCa

[Pla98k]  P. J. Plauger.  Standard

C/C++: The facets num put and numpunct. C/C++ Users Journal, 16(1):??, January 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1998:SCCl

[Pla98l] P. J. Plauger. Standard C/C++:

Thread safety. C/C++ Users Journal, 16(12):??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCf

[Pla99a] P. J. Plauger. Standard C/C++:

A better deque. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-

Plauger:1999:SCCh

[Pla99b] P. J. Plauger. Standard C/C++: A better list. C/C++ Users Journal, 17(8):10–??, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCg

[Pla99c] P. J. Plauger. Standard C/C++: A better red-black tree. C/C++ Users Journal, 17(7):10–??, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCj

[Pla99d] P. J. Plauger. Standard C/C++: A better sort. C/C++ Users Journal, 17(10):??, October 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCb

[Pla99e] P. J. Plauger. Standard C/C++: Embedded C++ update. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCk

[Pla99f] P. J. Plauger. Standard C/C++: Frequently answered questions. C/C++ Users Journal, 17(11): ??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCl

[Pla99g] P. J. Plauger. Standard C/C++: Frequently answered questions: STL. C/C++ Users Journal, 17 (12):??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCe

[Pla99h] P. J. Plauger. Standard C/C++:

Multibyte files. C/C++ Users Journal, 17(5):??, May 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCa

[Pla99i] P. J. Plauger. Standard C/C++:

Revising standard C. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCc

[Pla99j] P. J. Plauger. Standard C/C++:

Simple iostreams. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCd

[Pla99k] P. J. Plauger. Standard C/C++:

Unicode files. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCi

[Pla99l] P. J. Plauger. Standard C/C++:

Why 2K? C/C++ Users Journal, 17(9):10–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1992:UVM

[PLBL92] P. J. Plauger, Marty Leisner, Tim Berens, and Andy Levinson. UNIX versus MS-DOS — some letters. C Users Journal, 10(6):130–??, June 1992. ISSN 0898-9788. See additional comments, November 1992, p. 140; April 1993, p. 126.

Plum:1984:CPG

[Plu84] Thomas Plum. C programming guidelines. Plum Hall, 1 Spruce Ave, Cardiff, NJ 08232, USA, 1984. ISBN 0-911537-03-1 (paperback), 0-13-109992-2. 145 pp. LCCN QA76.73.C15 P57 1984 Sci-Eng. US$25.00.

Plum:1988:ACR

[Plu88a] Thomas Plum. ANSI committee refines support for international users. C Users Journal, 6(1):27– ??, January 1988. ISSN 0898-9788.

Plum:1988:ADN

[Plu88b] Thomas Plum. ANSI drops ‘no alias’ from draft standard. C Users Journal, 6(6):34–??, June 1988. ISSN 0898-9788.

Plum:1988:SSB

[Plu88c] Thomas Plum. Six simple benchmark programs give results especially useful to C programmers. C Users Journal, 6(5):44–??, May 1988. ISSN 0898-9788.

Plum:1988:XAR

[Plu88d] Thomas Plum. X3J11 approves revised draft for second public review. C Users Journal, 6(2):51– ??, February 1988. ISSN 0898-9788.

Plum:1989:CPG

[Plu89a]  Thomas Plum.  C program-

ming guidelines. Plum Hall, 1 Spruce Ave, Cardiff, NJ 08232, USA, second edition, 1989. ISBN 0-911537-07-4. 211 pp. LCCN QA76.73.C15 P57 1989. US$30.00.

Plum:1989:FCR

[Plu89b] Thomas Plum. First-order correctness review and test. C Users Journal, 7(1):75–??, January 1989. ISSN 0898-9788.

Plum:1991:RPC

[Plu91] Thomas Plum. Resources for portable C applications. C Users Journal, 9(1):39–??, January 1991. ISSN 0898-9788.

Pugh:1989:PUD

[PM89] Kenneth Pugh and Raymond J.

Martini. Problems using dynamic link libraries under OS/2. C Users Journal, 7(7):29–??, July 1989. ISSN 0898-9788.

Pugh:1991:TCI

[PM91a] Kenneth Pugh and Larry Meyer.

curses, input, and screen refresh. C Users Journal, 9(11): 114–??, November 1991. ISSN 0898-9788.

Pugh:1991:PSR

[PM91b] Kenneth Pugh and A. E. Molony.

Printer setup routine. C Users Journal, 9(5):65–??, May 1991. ISSN 0898-9788.

Pugh:1991:KIC

[PN91] Kenneth Pugh and Peter Nunn.

‘ON KEY’ interrupts in C under MS-DOS. C Users Journal, 9 (10):120–??, October 1991. ISSN 0898-9788.

Plauger:1993:NOB

[PN93] P. J. Plauger and Anthony Naggs. Notes on optimizing the bubble sort. C Users Journal, 11(3):121–??, March 1993. ISSN 0898-9788.

Poindexter:1989:CPW

[Poi89a] Tom Poindexter. CROBOTS players write their own strategy. C Users Journal, 7(1):105– ??, January 1989. ISSN 0898-9788.

Poindexter:1989:CC

[Poi89b] Tom Poindexter. CUG279 — CROBOTS. C Users Journal, 7 (1):101–??, January 1989. ISSN 0898-9788.

Pokorny:1992:SFS

[Pok92] Philip D. Pokorny. Specifying the field of a structure to sizeof(). C Users Journal, 10(1):140–??, January 1992. ISSN 0898-9788.

Pollatschek:1991:CSS

[Pol91] M. A. Pollatschek. CUG349 — simulation subroutine library. C Users Journal, 9(8):120–??, August 1991. ISSN 0898-9788.

Pomerantz:1998:TEH

[Pom98] Dave Pomerantz. Testing error handlers by simulating errors. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Poole:1993:SEP

[Poo93] Stephen M. Poole. Suggestions to enhance the portability of ANSI C. C Users Journal, 11(2):136– ??, February 1993. ISSN 0898-9788.

Popenfoose:1992:WSS

[Pop92] Toby Popenfoose. Wildcard subdirectory searches. C Users Journal, 10(2):90–??, February 1992. ISSN 0898-9788.

Porter:1993:BCT

[Por93]  Anthony Porter.  The best

C/C++ tips ever. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1993. ISBN 0-07-881820-6. various pp. LCCN QA76.73.C15 P68 1993. US$29.95. See book review [Pla95a].

Pottorff:1988:PSG

[Pot88] Brian Pottorff. Programming standard GEM displays in C. C Users Journal, 6(11):75–??, November 1988. ISSN 0898-9788.

Petrovic:1990:TPD

[PP90] Mark Petrovic and Kenneth Pugh. printf() and debugging. C Users Journal, 8(9):111– ??, September 1990. ISSN 0898-9788.

Pugh:1992:UFS

[PP92] Kenneth Pugh and G. Prodasla.

Using the [ format specifier with sscanf(). C Users Journal, 10(9):114–??, September 1992. ISSN 0898-9788.

Pugh:1991:IPR

[PR91] Kenneth Pugh and Tim Riley. Intercepting process return codes. C Users Journal, 9(4):88– ??, April 1991. ISSN 0898-9788.

Press:1988:NRC

[Pre88] William H. Press. Numerical recipes in C: the art of scientific computing. Cambridge University Press, New York, NY, USA, 1988. ISBN 0-521-35465-X, 0-521-35466-8 (diskette). xxii + 735 pp. LCCN QA76.73.C15 N865 1988. See book review [Phi90c].

Press:1992:NRC

[Pre92] William H. Press. Numerical recipes in C: the art of scientific computing. Cambridge University Press, New York, NY, USA, second edition, 1992. ISBN 0-521-43108-5, 0-521-43720-2 (book in C). xxvi + 994 pp. LCCN QA297 .N866 1994b.

Prince:1988:TNS

[Pri88]  Timothy Prince.  Numerical

Software Tools in C, by James Kempf. C Users Journal, 6(5): 61–??, May 1988. ISSN 0898-9788.

Prince:1989:EMC

[Pri89a] Timothy Prince. Efficient matrix coding in C. C Users Journal, 7 (5):59–??, May 1989. ISSN 0898-9788.

Prince:1989:FTF

[Pri89b] Timothy Prince. Float trig functions for the C preprocessor. C Users Journal, 7(8):103–??, August 1989. ISSN 0898-9788.

Prince:1990:CSA

[Pri90a] Timothy Prince. Coding for superscalar architectures. C Users Journal, 8(12):67–??, December 1990. ISSN 0898-9788.

Prince:1990:GST

[Pri90b] Timothy Prince. Generating source for <float.h>. C Users Journal, 8(6):119–??, June 1990. ISSN 0898-9788.

Prince:1990:LTM

[Pri90c] Timothy Prince. Language translation and multi-language development. C Users Journal, 8(4): 137–??, April 1990. ISSN 0898-9788.

Prince:1990:QSM

[Pri90d] Timothy Prince. The quick sort for micros and pipeline processors. C Users Journal, 8(5):97– ??, May 1990. ISSN 0898-9788.

Prince:1991:DSC

[Pri91a] Timothy Prince. Double to (and from) string conversion. C Users Journal, 9(7):75–??, July 1991. ISSN 0898-9788.

Prince:1991:PLD

[Pri91b] Timothy Prince. Pennies in long double. C Users Journal, 9(1): 77–??, January 1991. ISSN 0898-9788.

Prince:1992:CFC

[Pri92a] Timothy Prince. Curve fitting by Chebyshef and other methods. C Users Journal, 10(2):99– ??, February 1992. ISSN 0898-9788.

Prince:1992:MCM

[Pri92b] Timothy Prince. Multiple copy math functions. C Users Journal, 10(4):21–??, April 1992. ISSN 0898-9788.

Prince:1992:TMF

[Pri92c] Timothy Prince. Tuning up math functions. C Users Journal, 10 (12):83–??, December 1992. ISSN 0898-9788.

Prince:1993:PTC

[Pri93] Timothy Prince. Performance tuning a complex FFT. C Users Journal, 11(5):61–??, May 1993. ISSN 0898-9788.

Prince:1994:TFM

[Pri94] Timothy Prince. float-precision math library. C Users Journal, 12 (6):45–??, June 1994. ISSN 0898-9788.

Prince:1996:QMFa

[Pri96a]  Tim Prince.  Quad-precision

math functions, part 1. C/C++ Users Journal, 14(2):39–??, February 1996. CODEN CCU-JEX. ISSN 1075-2838.

Prince:1996:QMFb

[Pri96b]  Tim Prince.  Quad-precision

math functions, part 2. C/C++ Users Journal, 14(3):45–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Provenzano:1989:HTM

[Pro89] Tom Provenzano. A hash table manager in C++. C Users Journal, 7(11):83–??, November 1989. ISSN 0898-9788.

Prodehl:1990:TPE

[Pro90] Tom Prodehl. Programmer’s Essential OS/2 Handbook, by David Cortesi. C Users Journal, 8(9): 135–??, September 1990. ISSN 0898-9788. See [Cor88].

Probert:1992:NNA

[Pro92] Geoffrey Probert. Nearest neighbor algorithm for color matching. C Users Journal, 10(2):30– ??, February 1992. ISSN 0898-9788.

Pruett:1991:MCS

[Pru91] Mark Pruett. Mixing C with SQL. C Users Journal, 9(6):25– ??, June 1991. ISSN 0898-9788.

Pugh:1991:COV

[PS91a] Kenneth Pugh and Arthur Shipman. Code organization and variable placement: The nature of C. C Users Journal, 9(2):121–??, February 1991. ISSN 0898-9788.

Pugh:1991:CCS

[PS91b] Kenneth Pugh and R. Smithers.

Counting the characters and spaces in a string. C Users Journal, 9(12):121–??, December 1991. ISSN 0898-9788.

Anonymous:1992:TTC

[PS92a] Geoffrey Probert and Richard B.

Shepard. Turbo C: Memory Resident Utilities, Screen I/O, and Programming, by Al Stevens. C Users Journal, 10(3):65–??, March 1992. ISSN 0898-9788.

Pugh:1992:MIE

[PS92b] Kenneth Pugh and Jack Steiner.

Modifying IBM extended characters. C Users Journal, 10(5):109– ??, May 1992. ISSN 0898-9788.

Plauger:1993:UTS

[PT93] P. J. Plauger and Stephen Thomas. Using sizeof in preprocessor statements. C Users Journal, 11(11):130–??, November 1993. ISSN 0898-9788.

Pugh:1988:ASL

[Pug88a] Kenneth Pugh. ANSI standards list for UNIX. C Users Journal, 6(1):22–??, January 1988. ISSN 0898-9788.

Pugh:1988:BTS

[Pug88b] Kenneth Pugh. Bug in setvideomode()? C Users Journal, 6(3):12–??, March 1988. ISSN 0898-9788.

Pugh:1988:CCS

[Pug88c] Kenneth Pugh. C compiler source code availability. C Users Journal, 6(8):21–??, August 1988. ISSN 0898-9788.

Pugh:1988:CCC

[Pug88d] Kenneth Pugh. C on Commodore computers. C Users Journal, 6 (6):68–??, June 1988. ISSN 0898-9788.

Pugh:1988:CVD

[Pug88e] Kenneth Pugh. ‘const’ vs. ‘#define’. C Users Journal, 6 (9):10–??, September 1988. ISSN 0898-9788.

Pugh:1988:CPD

[Pug88f] Kenneth Pugh. Conversion packages for dBASE to C translation. C Users Journal, 6(5):15– ??, May 1988. ISSN 0898-9788.

Pugh:1988:CSU

[Pug88g] Kenneth Pugh. Converting a string to uppercase using toupper(). C Users Journal, 6(3):11–??, March 1988. ISSN 0898-9788.

Pugh:1988:CDD

[Pug88h] Kenneth Pugh. Converting dBase data files for C applications. C Users Journal, 6(3):9–??, March 1988. ISSN 0898-9788.

Pugh:1988:TFE

[Pug88i] Kenneth Pugh. format() in Ecosoft C. C Users Journal, 6 (2):11–??, February 1988. ISSN 0898-9788.

Pugh:1988:FDO

[Pug88j] Kenneth Pugh. Formatting disks for other DOS’s under MS-DOS. C Users Journal, 6(3):10–??, March 1988. ISSN 0898-9788.

Pugh:1988:FRV

[Pug88k] Kenneth Pugh. ‘function return values mismatches’ error. C Users Journal, 6(9):10–??, September 1988. ISSN 0898-9788.

Pugh:1988:GLL

[Pug88l] Kenneth Pugh. GRAD library linking. C Users Journal, 6(1): 21–??, January 1988. ISSN 0898-9788.

Pugh:1988:HPR

[Pug88m] Kenneth Pugh. How to perform record-oriented file I/O. C Users Journal, 6(2):10–??, February 1988. ISSN 0898-9788.

Pugh:1988:IL

[Pug88n] Kenneth Pugh. Indirection levels. C Users Journal, 6(9):10– ??, September 1988. ISSN 0898-

Pugh:1988:ISP

[Pug88o] Kenneth Pugh. Information sources for programming TSR’s. C Users Journal, 6(11):14–??, November 1988. ISSN 0898-9788.

Pugh:1988:IHT

[Pug88p] Kenneth Pugh. INIT.C in HOC6 from The UNIX Programming Environment. C Users Journal, 6 (11):10–??, November 1988. ISSN 0898-9788.

Pugh:1988:MAS

[Pug88q] Kenneth Pugh. Memory allocation for strings and numbers in C. C Users Journal, 6(11):12–??, November 1988. ISSN 0898-9788.

Pugh:1988:MR

[Pug88r] Kenneth Pugh. Menu routines. C Users Journal, 6(11):9– ??, November 1988. ISSN 0898-9788.

Pugh:1988:MEI

[Pug88s] Kenneth Pugh. MS-DOS error interception by application programs. C Users Journal, 6(3):9– ??, March 1988. ISSN 0898-9788.

Pugh:1988:MMA

[Pug88t] Kenneth Pugh. MS-DOS memory allocation and free space. C Users Journal, 6(11):10–??, November 1988. ISSN 0898-9788.

Pugh:1988:RDS

[Pug88u] Kenneth Pugh. Reading disk sectors into memory in C. C Users Journal, 6(11):9–??, November 1988. ISSN 0898-9788.

Pugh:1988:RCG

[Pug88v] Kenneth Pugh. ROMable code generation: Some notes. C Users Journal, 6(9):9–??, September 1988. ISSN 0898-9788.

Pugh:1988:SCP

[Pug88w] Kenneth Pugh. Sending Control/PrtSc to a printer from C. C Users Journal, 6(11):10–??, November 1988. ISSN 0898-9788.

Pugh:1988:SEV

[Pug88x] Kenneth Pugh. Setting environment variables in MS-DOS. C Users Journal, 6(3):12–??, March 1988. ISSN 0898-9788.

Pugh:1988:TST

[Pug88y] Kenneth Pugh. sizeof() in #if expressions prohibited in ANSI Draft C Standard. C Users Journal, 6(5):15–??, May 1988. ISSN 0898-9788.

Pugh:1988:SCC

[Pug88z] Kenneth Pugh. Small C compiling problems. C Users Journal, 6(1):21–??, January 1988. ISSN 0898-9788.

Pugh:1988:SEI

[Pug88-27] Kenneth Pugh. Some examples illustrate use of ‘extern’ keyword. C Users Journal, 6(5):13– ??, May 1988. ISSN 0898-9788.

Pugh:1988:TSP

[Pug88-28] Kenneth Pugh. stdarg package. C Users Journal, 6(6):69–??, June 1988. ISSN 0898-9788.

Pugh:1988:TCT

[Pug88-29] Kenneth Pugh. Test compilers for treatment of two-byte character constants. C Users Journal, 6 (5):14–??, May 1988. ISSN 0898-9788.

Pugh:1988:UAO

[Pug88-30] Kenneth Pugh. UNIX alternatives to overlays. C Users Journal, 6(2):9–??, February 1988. ISSN 0898-9788.

Pugh:1988:WLP

[Pug88-31] Kenneth Pugh. Window library packages for C. C Users Journal, 6(5):15–??, May 1988. ISSN 0898-9788.

Pugh:1989:ASA

[Pug89a] Kenneth Pugh. Accessing ‘struct’ from assembly. C Users Journal, 7(7):28–??, July 1989. ISSN 0898-9788.

Pugh:1989:TAT

[Pug89b] Kenneth Pugh. atof() and strtod() compared. C Users Journal, 7(7):26–??, July 1989. ISSN 0898-9788.

Pugh:1989:CCA

[Pug89c] Kenneth Pugh. C compilers for the Apple II. C Users Journal, 7 (10):58–??, October 1989. ISSN 0898-9788.

Pugh:1989:CPA

[Pug89d] Kenneth Pugh. Changing pointers-to-functions in an array of structures. C Users Journal, 7 (4):9–??, April 1989. ISSN 0898-9788.

Pugh:1989:COU

[Pug89e] Kenneth Pugh. Character output without using printf(). C Users Journal, 7(7):23–??, July 1989. ISSN 0898-9788.

Pugh:1989:DPK

[Pug89f] Kenneth Pugh. Deactivating the PrintScreen key. C Users Journal, 7(11):48–??, November 1989. ISSN 0898-9788.

Pugh:1989:DBP

[Pug89g] Kenneth Pugh. Differences between pointers and arrays. C Users Journal, 7(1):20–??, January 1989. ISSN 0898-9788.

Pugh:1989:DPF

[Pug89h] Kenneth Pugh. Double pointers to functions. C Users Journal, 7 (4):10–??, April 1989. ISSN 0898-9788.

Pugh:1989:DXL

[Pug89i] Kenneth Pugh. Drawing with XOR lines using Turbo C 2.0. C Users Journal, 7(11):49–??, November 1989. ISSN 0898-9788.

Pugh:1989:IUC

[Pug89j] Kenneth Pugh. Initializing unions in C. C Users Journal, 7(2):38–??, February 1989. ISSN 0898-9788.

Pugh:1989:MMC

[Pug89k] Kenneth Pugh. ‘memcpy’ and ‘memmove’ compared. C Users Journal, 7(7):27–??, July 1989. ISSN 0898-9788.

Pugh:1989:MMM

[Pug89l] Kenneth Pugh. Mixed memory model pointers. C Users Journal, 7(11):45–??, November 1989. ISSN 0898-9788.

Pugh:1989:MFO

[Pug89m] Kenneth Pugh. MS-DOS file opening for both reading and writing. C Users Journal, 7(10): 56–??, October 1989. ISSN 0898-9788.

Pugh:1989:OSB

[Pug89n] Kenneth Pugh. Optimizing a screen blanker. C Users Journal, 7(4):12–??, April 1989. ISSN 0898-9788.

Pugh:1989:TPR

[Pug89o] Kenneth Pugh. pointer-to-a-function

return types. C Users Journal, 7 (5):22–??, May 1989. ISSN 0898-9788.

Pugh:1989:PWI

[Pug89p] Kenneth Pugh. Preprocessing: What it is. C Users Journal, 7 (10):54–??, October 1989. ISSN 0898-9788.

Pugh:1989:SSQ

[Pug89q] Kenneth Pugh. Some semantic questions. C Users Journal, 7 (7):27–??, July 1989. ISSN 0898-9788.

Pugh:1989:TAL

[Pug89r] Kenneth Pugh. Translating assembly language to C. C Users Journal, 7(10):55–??, October 1989. ISSN 0898-9788.

Pugh:1989:TSC

[Pug89s] Kenneth Pugh. Transporting small C to the Apple. C Users Journal, 7(1):19–??, January 1989. ISSN 0898-9788.

Pugh:1989:UEF

[Pug89t] Kenneth Pugh. Using ‘ecvt,’ ‘fcvt,’ and ‘gcvt’. C Users Journal, 7(7):24–??, July 1989. ISSN 0898-9788.

Pugh:1989:WVI

[Pug89u] Kenneth Pugh. What is a virtual interface? C Users Journal, 7(8): 10–??, August 1989. ISSN 0898-9788.

Pugh:1990:ASE

[Pug90a] Kenneth Pugh. Access to structure elements elucidated. C Users Journal, 8(6):72–??, June 1990. ISSN 0898-9788.

Pugh:1990:AOM

[Pug90b] Kenneth Pugh. Accidental opening of MS-DOS files twice. C Users Journal, 8(9):109–??, September 1990. ISSN 0898-9788.

Pugh:1990:AVP

[Pug90c] Kenneth Pugh. Array vs. pointer names. C Users Journal, 8(8):75– ??, August 1990. ISSN 0898-9788.

Pugh:1990:ASM

[Pug90d] Kenneth Pugh. Arrays of structures and memory limitations with Turbo C 2.0. C Users Journal, 8(11):105–??, November 1990. ISSN 0898-9788.

Pugh:1990:AOP

[Pug90e] Kenneth Pugh. Assigning one pointer value to another of different data type. C Users Journal, 8(11):101–??, November 1990. ISSN 0898-9788.

Pugh:1990:BFS

[Pug90f] Kenneth Pugh. Back to fundamentals: Simple text output to printers. C Users Journal, 8(6): 71–??, June 1990. ISSN 0898-9788.

Pugh:1990:CLN

[Pug90g] Kenneth Pugh. Casts and lvalues: A note. C Users Journal, 8(11):102–??, November 1990. ISSN 0898-9788.

Pugh:1990:CMA

[Pug90h] Kenneth Pugh. Contiguous memory access for structures. C Users Journal, 8(11):104–??, November 1990. ISSN 0898-9788.

Pugh:1990:CBF

[Pug90i] Kenneth Pugh. Converting BA-SIC floating point files to C. C Users Journal, 8(5):69–??, May 1990. ISSN 0898-9788.

Pugh:1990:TDV

[Pug90j] Kenneth Pugh. #defines vs. typedefs. C Users Journal, 8 (1):113–??, January 1990. ISSN 0898-9788.

Pugh:1990:DMC

[Pug90k] Kenneth Pugh. Defining manifest constants. C Users Journal, 8(3): 34–??, March 1990. ISSN 0898-9788.

Pugh:1990:DCS

[Pug90l]  Kenneth Pugh.  Displaying

control-characters on the screen. C Users Journal, 8(10):83–??, October 1990. ISSN 0898-9788.

Pugh:1990:EAS

[Pug90m] Kenneth Pugh. Efficient access to structure elements. C Users Journal, 8(2):70–??, February 1990. ISSN 0898-9788.

Pugh:1990:EFR

[Pug90n] Kenneth Pugh. Emulating Fortran’s ‘repeat’ format construct. C Users Journal, 8(3):36–??, March 1990. ISSN 0898-9788.

Pugh:1990:GNO

[Pug90o] Kenneth Pugh. Great name/ obscure code contest announcement. C Users Journal, 8(4):49– ??, April 1990. ISSN 0898-9788.

Pugh:1990:HPK

[Pug90p] Kenneth Pugh. Halt printing by a keypress. C Users Journal, 8 (12):89–??, December 1990. ISSN 0898-9788.

Pugh:1990:IUT

[Pug90q] Kenneth Pugh. Implementation

and use of va arg in <stdarg.h>.

C Users Journal, 8(5):69–??, May 1990. ISSN 0898-9788.

Pugh:1990:LLF

[Pug90r] Kenneth Pugh. Linking library functions at runtime under MS-DOS. C Users Journal, 8(7):92– ??, July 1990. ISSN 0898-9788.

Pugh:1990:LUF

[Pug90s] Kenneth Pugh. Locking UNIX files. C Users Journal, 8(9):109– ??, September 1990. ISSN 0898-9788.

Pugh:1990:TMP

[Pug90t] Kenneth Pugh. malloc() problems under UNIX disappear on PC. C Users Journal, 8(2):69–??, February 1990. ISSN 0898-9788.

Pugh:1990:MPA

[Pug90u] Kenneth Pugh. More on passing arrays. C Users Journal, 8(4):49– ??, April 1990. ISSN 0898-9788.

Pugh:1990:NNC

[Pug90v] Kenneth Pugh. ‘new’ naming in C. C Users Journal, 8(1):115–??, January 1990. ISSN 0898-9788.

Pugh:1990:NAP

[Pug90w] Kenneth Pugh. Notes on arrays of pointers. C Users Journal, 8 (11):99–??, November 1990. ISSN 0898-9788.

Pugh:1990:NP

[Pug90x] Kenneth Pugh. Notes on preprocessing. C Users Journal, 8 (7):92–??, July 1990. ISSN 0898-9788.

Pugh:1990:TOM

[Pug90y] Kenneth Pugh. offsetof() macro explained. C Users Journal, 8(2):72–??, February 1990. ISSN 0898-9788.

Pugh:1990:IEV

[Pug90z] Kenneth Pugh. On initializing external variables in header files. C Users Journal, 8(10):83– ??, October 1990. ISSN 0898-9788.

Pugh:1990:PBF

[Pug90-27] Kenneth Pugh. Porting binary files between a PC and a Macintosh. C Users Journal, 8(2):74– ??, February 1990. ISSN 0898-9788.

Pugh:1990:PCS

[Pug90-28] Kenneth Pugh. Porting C source from MS-DOS to CP/M. C Users Journal, 8(3):34–??, March 1990. ISSN 0898-9788.

Pugh:1990:PUD

[Pug90-29] Kenneth Pugh. Possibility of user-defined data types. C Users Journal, 8(11):101–??, November 1990. ISSN 0898-9788.

Pugh:1990:PIS

[Pug90-30] Kenneth Pugh. Processing interrupts and stack overflow in C. C Users Journal, 8(2):77–??, February 1990. ISSN 0898-9788.

Pugh:1990:TQP

[Pug90-31] Kenneth Pugh. qsort() problem in Turbo C++ v1.04. C Users Journal, 8(12):92–??, December 1990. ISSN 0898-9788.

Pugh:1990:RPO

[Pug90-32] Kenneth Pugh. Re-directing printer output under MS-DOS. C Users Journal, 8(7):87–??, July 1990. ISSN 0898-9788.

Pugh:1990:RMF

[Pug90-33] Kenneth Pugh. Reading MS-DOS files into a UNIX system. C Users Journal, 8(6):75–??, June 1990. ISSN 0898-9788.

Pugh:1990:RS

[Pug90-34] Kenneth Pugh. Replies on stringizing. C Users Journal, 8 (7):93–??, July 1990. ISSN 0898-9788.

Pugh:1990:SCP

[Pug90-35] Kenneth Pugh. Simple C printing demo. C Users Journal, 8(6):71– ??, June 1990. ISSN 0898-9788.

Pugh:1990:SBX

[Pug90-36] Kenneth Pugh. Some books on X Window System. C Users Journal, 8(6):78–??, June 1990. ISSN 0898-9788.

Pugh:1990:SNC

[Pug90-37] Kenneth Pugh. Some notes on C style. C Users Journal, 8(4):53– ??, April 1990. ISSN 0898-9788.

Pugh:1990:SNF

[Pug90-38] Kenneth Pugh. Some notes on function pointers. C Users Journal, 8(7):87–??, July 1990. ISSN 0898-9788.

Pugh:1990:SPI

[Pug90-39] Kenneth Pugh. Some precedence issues. C Users Journal, 8(4):50– ??, April 1990. ISSN 0898-9788.

Pugh:1990:TUT

[Pug90-40] Kenneth Pugh. Tips on using printf(). C Users Journal, 8 (7):87–??, July 1990. ISSN 0898-9788.

Pugh:1990:TFK

[Pug90-41] Kenneth Pugh. Trapping function keys. C Users Journal, 8(3): 33–??, March 1990. ISSN 0898-9788.

Pugh:1990:TCS

[Pug90-42] Kenneth Pugh. Type ‘char’ — signed or unsigned? C Users Journal, 8(1):113–??, January 1990. ISSN 0898-9788.

Pugh:1991:CSC

[Pug91a] Kenneth Pugh. Coding style in C. C Users Journal, 9(11):116– ??, November 1991. ISSN 0898-9788.

Pugh:1991:D

[Pug91b] Kenneth Pugh. DLLs for DOS.

C Users Journal, 9(9):99–??, September 1991. ISSN 0898-9788.

Pugh:1991:MMT

[Pug91c] Kenneth Pugh. Maximum and minimum of three integers. C Users Journal, 9(12):120–??, December 1991. ISSN 0898-9788.

Pugh:1991:MTC

[Pug91d] Kenneth Pugh. More on const. C Users Journal, 9(1):95–??, January 1991. ISSN 0898-9788.

Pugh:1991:MPP

[Pug91e] Kenneth Pugh. More pointer problems. C Users Journal, 9(4): 85–??, April 1991. ISSN 0898-9788.

Pugh:1991:NCS

[Pug91f] Kenneth Pugh. Naming conflicts with standard library function names. C Users Journal, 9(6):119–??, June 1991. ISSN 0898-9788.

Pugh:1991:RFK

[Pug91g] Kenneth Pugh. Reading function keys of IBM PC in C programs. C Users Journal, 9(5):65–??, May 1991. ISSN 0898-9788.

Pugh:1991:RCT

[Pug91h] Kenneth Pugh. Recursive calls to main(). C Users Journal, 9 (10):120–??, October 1991. ISSN 0898-9788.

Pugh:1991:TTA

[Pug91i] Kenneth Pugh. typedef: Advantages and disadvantages. C Users Journal, 9(12):119–??, December 1991. ISSN 0898-9788.

Pugh:1991:UTT

[Pug91j] Kenneth Pugh. Using TLIB, the Turbo C Librarian. C Users Journal, 9(10):117–??, October 1991. ISSN 0898-9788.

Pugh:1991:YL

[Pug91k] Kenneth Pugh. YACC and Lex. C Users Journal, 9(2):117–??, February 1991. ISSN 0898-9788.

Pugh:1992:NO

[Pug92a] Kenneth Pugh. =- as Non-ANSI operator. C Users Journal, 10(5): 110–??, May 1992. ISSN 0898-9788.

Pugh:1992:CFW

[Pug92b] Kenneth Pugh. Calling functions from within a function. C Users Journal, 10(4):105–??, April 1992. ISSN 0898-9788.

Pugh:1992:CEH

[Pug92c] Kenneth Pugh. Critical error handling. C Users Journal, 10(5): 111–??, May 1992. ISSN 0898-9788.

Pugh:1992:DSA

[Pug92d] Kenneth Pugh. Declaring static and automatic pointer variables. C Users Journal, 10(2):108–??, February 1992. ISSN 0898-9788.

Pugh:1992:EDL

[Pug92e] Kenneth Pugh. Exceeding DOS’s 640K limit. C Users Journal, 10(9):115–??, September 1992. ISSN 0898-9788.

Pugh:1992:GCD

[Pug92f] Kenneth Pugh. Generating check digits for error detection. C Users Journal, 10(11):83–??, November 1992. ISSN 0898-9788.

Pugh:1992:GAS

[Pug92g] Kenneth Pugh. Globally accessing strings. C Users Journal, 10 (8):132–??, August 1992. ISSN 0898-9788.

Pugh:1992:LLF

[Pug92h] Kenneth Pugh. Linked list functions to work with multiple data types. C Users Journal, 10(10): 63–??, October 1992. ISSN 0898-9788.

Pugh:1992:OEP

[Pug92i] Kenneth Pugh. Order of evaluation and post-increment evaluation time. C Users Journal, 10 (8):131–??, August 1992. ISSN 0898-9788.

Pugh:1992:PVB

[Pug92j] Kenneth Pugh. Passing values between programs. C Users Journal, 10(8):130–??, August 1992. ISSN 0898-9788.

Pugh:1992:PFD

[Pug92k] Kenneth Pugh. Pointers to functions and double pointers. C Users Journal, 10(9):111–??, September 1992. ISSN 0898-9788.

Pugh:1992:PAI

[Pug92l] Kenneth Pugh. Problems in array initialization. C Users Journal, 10(8):131–??, August 1992. ISSN 0898-9788.

Pugh:1992:RCL

[Pug92m] Kenneth Pugh. Reading in command line arguments. C Users Journal, 10(8):129–??, August 1992. ISSN 0898-9788.

Pugh:1992:RTT

[Pug92n] Kenneth Pugh. Return types for main(). C Users Journal, 10(8): 129–??, August 1992. ISSN 0898-9788.

Pugh:1992:SFS

[Pug92o] Kenneth Pugh. Setting file sizes under UNIX. C Users Journal, 10 (2):109–??, February 1992. ISSN 0898-9788.

Pugh:1992:SGK

[Pug92p] Kenneth Pugh. Some guidelines for keeping maintenance simpler. C Users Journal, 10(7):88–??, July 1992. ISSN 0898-9788.

Pugh:1992:SAT

[Pug92q] Kenneth Pugh. Speeding up access time when using linked lists. C Users Journal, 10(8):127–??, August 1992. ISSN 0898-9788.

Pugh:1992:TTS

[Pug92r] Kenneth Pugh. typedefs and strong type-checking. C Users Journal, 10(4):113–??, April 1992. ISSN 0898-9788.

Pugh:1992:UCG

[Pug92s] Kenneth Pugh. Using code generators for creating C code. C Users Journal, 10(3):61–??, March 1992. ISSN 0898-9788.

Pugh:1992:UGV

[Pug92t] Kenneth Pugh. Using gotos versus multiple returns. C Users Journal, 10(2):109–??, February 1992. ISSN 0898-9788.

Pugh:1992:UCO

[Pug92u] Kenneth Pugh. Using the conditional operator ?:. C Users Journal, 10(2):107–??, February 1992. ISSN 0898-9788.

Pugh:1992:VSN

[Pug92v] Kenneth Pugh. Volume serial numbers under MS-DOS. C Users Journal, 10(2):109–??, February 1992. ISSN 0898-9788.

Pugh:1993:CO

[Pug93a] Kenneth Pugh. C++ operators.

C Users Journal, 11(11):106–??, November 1993. ISSN 0898-9788.

Pugh:1993:CP

[Pug93b] Kenneth Pugh. Character pointers. C Users Journal, 11(12):94– ??, December 1993. ISSN 0898-9788.

Pugh:1993:CCT

[Pug93c] Kenneth Pugh. Compiling C++ templates. C Users Journal, 11 (12):91–??, December 1993. ISSN 0898-9788.

Pugh:1993:DC

[Pug93d] Kenneth Pugh. Documentation of classes. C Users Journal, 11 (8):103–??, August 1993. ISSN 0898-9788.

Pugh:1993:DPT

[Pug93e] Kenneth Pugh. Double pointers: Their uses. C Users Journal, 11 (8):104–??, August 1993. ISSN 0898-9788.

Pugh:1993:DPS

[Pug93f] Kenneth Pugh. Dynamic program suspension: Some notes on DESQview. C Users Journal, 11(6):111–??, June 1993. ISSN 0898-9788.

Pugh:1993:EDA

[Pug93g] Kenneth Pugh. External declarations in ANSI C. C Users Journal, 11(2):117–??, February 1993. ISSN 0898-9788.

Pugh:1993:FPC

[Pug93h] Kenneth Pugh. Floating point constants. C Users Journal, 11 (10):130–??, October 1993. ISSN 0898-9788.

Pugh:1993:FRV

[Pug93i] Kenneth Pugh. Function return values. C Users Journal, 11(3): 87–??, March 1993. ISSN 0898-9788.

Pugh:1993:HSW

[Pug93j] Kenneth Pugh. Heap or stack — which should you use? C Users Journal, 11(10):123–??, October 1993. ISSN 0898-9788.

Pugh:1993:IL

[Pug93k] Kenneth Pugh. Internationalization and localization. C Users Journal, 11(7):111–??, July 1993. ISSN 0898-9788.

Pugh:1993:IFS

[Pug93l] Kenneth Pugh. Internationalization functions specified under X/open portability guidelines. C Users Journal, 11(1): 103–??, January 1993. ISSN 0898-9788.

Pugh:1993:LRV

[Pug93m] Kenneth Pugh. Lack of return values in constructors and destructors. C Users Journal, 11 (3):87–??, March 1993. ISSN 0898-9788.

Pugh:1993:LC

[Pug93n] Kenneth Pugh. Lint for C++?

C Users Journal, 11(9):123–??, September 1993. ISSN 0898-9788.

Pugh:1993:MNF

[Pug93o] Kenneth Pugh. More notes on function pointers. C Users Journal, 11(7):113–??, July 1993. ISSN 0898-9788.

Pugh:1993:NAO

[Pug93p] Kenneth Pugh. A note on the auto-increment operator. C Users Journal, 11(11):107–??, November 1993. ISSN 0898-9788.

Pugh:1993:PMA

[Pug93q] Kenneth Pugh. Pointers and multi-dimensional arrays. C Users Journal, 11(11):103–??, November 1993. ISSN 0898-9788.

Pugh:1993:PPO

[Pug93r] Kenneth Pugh. Postfix and prefix operators. C Users Journal, 11(8):101–??, August 1993. ISSN 0898-9788.

Pugh:1993:SNF

[Pug93s] Kenneth Pugh. Some notes on functions using derived classes as arguments or return values. C Users Journal, 11(11):106–??, November 1993. ISSN 0898-9788.

Pugh:1993:SNI

[Pug93t] Kenneth Pugh. Some notes on identifiers. C Users Journal, 11 (2):115–??, February 1993. ISSN 0898-9788.

Pugh:1993:SF

[Pug93u] Kenneth Pugh. Structures and files. C Users Journal, 11(9):124– ??, September 1993. ISSN 0898-9788.

Pugh:1993:VCS

[Pug93v] Kenneth Pugh. Version control: Some notes. C Users Journal, 11(9):126–??, September 1993. ISSN 0898-9788.

Pugh:1993:WFT

[Pug93w] Kenneth Pugh. Wrapper functions for malloc and free. C Users Journal, 11(3):91–??, March 1993. ISSN 0898-9788.

Pugh:1994:COC

[Pug94a] Kenneth Pugh. C++ overloaded constructors and operators. C/C++ Users Journal, 12 (12):93–??, December 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Pugh:1994:CDS

[Pug94b] Kenneth Pugh. Class design:

Some notes on principles. C Users Journal, 12(1):75–??, January 1994. ISSN 0898-9788.

Pugh:1994:CSS

[Pug94c] Kenneth Pugh. A class for string storage. C Users Journal, 12 (2):84–??, February 1994. ISSN 0898-9788.

Pugh:1994:CTD

[Pug94d] Kenneth Pugh. Converting doubles to strings. C/C++ Users Journal, 12(7):61–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1994:FRV

[Pug94e] Kenneth Pugh. Function return value conventions. C/C++ Users Journal, 12(11):103–??, November 1994. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1994:HCF

[Pug94f] Kenneth Pugh. Handling constructor failures. C/C++ Users Journal, 12(10):93–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1994:LC

[Pug94g] Kenneth Pugh. Lint for C++. C Users Journal, 12(1):75–??, January 1994. ISSN 0898-9788.

Pugh:1994:MFC

[Pug94h] Kenneth Pugh. Moving from FORTRAN to C. C Users Journal, 12(6):75–??, June 1994. ISSN 0898-9788.

Pugh:1994:NFP

[Pug94i] Kenneth Pugh. Notes on function prototypes. C Users Journal, 12 (5):59–??, May 1994. ISSN 0898-9788.

Pugh:1994:PTN

[Pug94j] Kenneth Pugh. Parentheses with new operator. C/C++ Users Journal, 12(9):55–??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1994:PCC

[Pug94k] Kenneth Pugh. The placement of constants in code: Some notes. C Users Journal, 12(2):83–??, February 1994. ISSN 0898-9788.

Pugh:1994:PT

[Pug94l] Kenneth Pugh. Pointer types. C Users Journal, 12(1):77–??, January 1994. ISSN 0898-9788.

Pugh:1994:PA

[Pug94m] Kenneth Pugh. Pointers and arrays. C Users Journal, 12(4):113– ??, April 1994. ISSN 0898-9788.

Pugh:1994:RTC

[Pug94n] Kenneth Pugh. Run-time type checking in C++. C Users Journal, 12(3):65–??, March 1994. ISSN 0898-9788.

Pugh:1994:SA

[Pug94o] Kenneth Pugh. Structure alignment. C Users Journal, 12(5):61– ??, May 1994. ISSN 0898-9788.

Pugh:1994:UCL

[Pug94p] Kenneth Pugh. Using C libraries in C++. C Users Journal, 12 (5):57–??, May 1994. ISSN 0898-9788.

Pugh:1994:WUPb

[Pug94q] Kenneth Pugh. When to use pointers. C/C++ Users Journal, 12(7):59–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1994:WUPa

[Pug94r] Kenneth Pugh. When to use pointers: Calling by reference. C Users Journal, 12(6):78–??, June 1994. ISSN 0898-9788.

Pugh:1995:MPR

[Pug95a] Kenneth Pugh. Are marching pointers really faster? C/C++ Users Journal, 13(1):91–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Pugh:1995:M

[Pug95b] Kenneth Pugh. Moving on.

C/C++ Users Journal, 13(3):89– ??, March 1995. CODEN CCU-JEX. ISSN 1075-2838.

Pugh:1995:QSF

[Pug95c] Kenneth Pugh. qsort and static functions. C/C++ Users Journal, 13(2):85–??, February 1995. CODEN CCUJEX. ISSN 1075-

Purdum:1989:PP

[Pur89a] Jack Purdum. Problems with pointers. C Users Journal, 7(8): 59–??, August 1989. ISSN 0898-9788.

Purdum:1989:CPT

[Pur89b] Jack J. (Jack Jay) Purdum. C programmer’s toolkit. Programming series. Que Corporation, Indianapolis, IN, USA, 1989. ISBN 0-88022-457-6. xvi + 350 pp. LCCN QA 76.73 C15 P868 1989. US$39.95, CDN$59.95. See book review [Ogg90b].

Purdum:1991:STP

[Pur91] Jack Purdum. Some thoughts on portability. C Users Journal, 9 (1):45–??, January 1991. ISSN 0898-9788.

Purdum:1992:CPT

[Pur92] Jack J. (Jack Jay) Purdum. C programmer’s toolkit. Programming series. Que Corporation, Indianapolis, IN, USA, second edition, 1992. ISBN 0-88022-788-5. xxix + 601 pp. LCCN QA 76.73 C15 P868 1992. US$39.95, CDN$49.95.

Pugh:1991:HE

[PW91] Kenneth Pugh and Glen Watson. Hiding and encapsulation. C Users Journal, 9(11):115–??, November 1991. ISSN 0898-9788.

Pugh:1990:SDT

[PY90] Kenneth Pugh and Ken Yerves. Storing data in .EXE file segments. C Users Journal, 8(12):

95–??, December 1990. ISSN 0898-9788.

Quackenbush:1988:AVS

[Qua88] Edward M. Quackenbush. American vendors’ service to Japan lacks. C Users Journal, 6(6):90– ??, June 1988. ISSN 0898-9788.

Rabung:1989:TCT

[Rab89] John Rabung. C Trainer, The, by Alan Feuer. C Users Journal, 7(5):103–??, May 1989. ISSN 0898-9788. See [Feu86].

Ragen:1995:CDP

[Rag95] Alex Ragen. C database programming with ODBC. C/C++ Users Journal, 13(11):73–??, November 1995. CODEN CCU-JEX. ISSN 1075-2838.

Rahimi:1997:MAC

[Rah97] Ali Rahimi. A multidimensional array class. C/C++ Users Journal, 15(7):39–??, July 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Ramey:1988:CRT

[Ram88] Robert Ramey. CUG262 — ramey tools. C Users Journal, 6 (5):75–??, May 1988. ISSN 0898-9788.

Ramey:1992:PS

[Ram92] Robert Ramey. The postman’s sort. C Users Journal, 10(8):59– ??, August 1992. ISSN 0898-9788.

Rathe:1990:UOM

[Rat90] Richard Rathe. Using OOPs in Mac designs. C Users Journal, 8(8):103–??, August 1990. ISSN 0898-9788.

Rathe:1991:XCM

[Rat91] Richard Rathe. ‘X’ commands for Mac HyperCard. C Users Journal, 9(7):25–??, July 1991. ISSN 0898-9788.

Raynor:1992:TAM

[Ray92] Gary J. Raynor. assert macro. C Users Journal, 10(5):128–??, May 1992. ISSN 0898-9788.

Ream:1990:CRV

[Rea90a] Edward K. Ream. CUG318 RED v7.0. C Users Journal, 8(7):111– ??, July 1990. ISSN 0898-9788.

Ream:1990:CCV

[Rea90b] Edward K. Ream. CUG319 CPP v5.3. C Users Journal, 8(7):111– ??, July 1990. ISSN 0898-9788.

Ream:1991:CSMb

[Rea91a] Edward K. Ream. CUG355 — Sherlock for MS-DOS. C Users Journal, 9(12):100–??, December 1991. ISSN 0898-9788.

Ream:1991:CSMa

[Rea91b] Edward K. Ream. CUG356 — Sherlock for Macintosh. C Users Journal, 9(12):100–??, December 1991. ISSN 0898-9788.

Ream:1992:DS

[Rea92] Edward K. Ream. Debugging with sherlock. C Users Journal, 10(6):121–??, June 1992. ISSN 0898-9788.

Reck:1998:TSR

[Rec98] Bill Reck. Thread synchronization with reference-counting handles. C/C++ Users Journal, 16 (2):??, February 1998. CODEN CCUJEX. ISSN 1075-2838.

Remy:1997:PSH

[Rem97] Martin Remy. Portable signal handling under UNIX. C/C++ Users Journal, 15(3):23–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Retzlaff:1996:CCS

[Ret96] Don Retzlaff. C/C++ source to HTML converter: Use those ubiquitous Web browsers to advantage and show C or C++ code in a more readable format. C/C++ Users Journal, 14(10): 57–58, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Revesz:1993:CPS

[Rév93] Thomas Révész. Clipping polygons with Sutherland-Hodgman’s algorithm. C Users Journal, 11 (8):23, 26, 29, 30, 33, 34, August 1993. ISSN 0898-9788.

Ribar:1989:SSD

[Rib89] John Ribar. A survey of system differences affecting ports between UNIX and VAX environments. C Users Journal, 7(1): 97–??, January 1989. ISSN 0898-9788.

Rice:1991:TNM

[Ric91] Jerry Rice. Numerical Methods for Scientists and Engineers still in print. C Users Journal, 9 (10):137–??, October 1991. ISSN 0898-9788.

Richards:1999:ALT

[Ric99] Etienne Richards. Adding level-2 thread safety to existing objects.

C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Ridge:1991:TDS

[Rid91] John Ridge. Data Structures Using C, by A. Tenenbaum, Y. Langsam, and M. J. Augenstein. C Users Journal, 9(10):100–??, October 1991. ISSN 0898-9788. See [TLA90].

Riemersma:1998:BDT

[Rie98]  Thiadmer Riemersma.  A

balanced dithering technique. C/C++ Users Journal, 16(12): ??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Ringle:1999:SCT

[Rin99]  Jonathan Ringle.  Singleton

creation the thread-safe way. C/C++ Users Journal, 17(10): ??, October 1999. CODEN CCU-JEX. ISSN 1075-2838.

Reichard:1991:XWP

[RJ91] Kevin Reichard and Eric F. Johnson. X Window programming — part 3: More Xlib programming. C Users Journal, 9(7):93–??, July 1991. ISSN 0898-9788.

Resendes:1998:ICD

[RL98] Robert Resendes and Marc Laukien. Introduction to CORBA distributed objects. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Romer:1992:CCA

[RM92] Andrew E. Romer and Paul Mc-Kee. CUG363 — 68020 cross assembler. C Users Journal, 10(5): 121–??, May 1992. ISSN 0898-9788.

Ranade:1993:ECP

[RN93] Jay Ranade and Alan Nash.

The elements of C programming style. McGraw-Hill, Inc., New York, NY, USA, 1993. ISBN 0-07-051278-7. xx + 340 pp. LCCN QA76.73.C15 R36 1993 Bar. US$29.95. See book review [All93f].

Robison:1994:PFF

[Rob94] Steve Robison. PCX file format documentation source noted. C Users Journal, 12(2):128–??, February 1994. ISSN 0898-9788.

Roberts:1995:LLC

[Rob95] Rick Roberts. Linux — the low cost UNIX. C/C++ Users Journal, 13(1):43–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Rodley:1995:TDO

[Rod95a] John Rodley. Distributed Operating Systems by Andrew Tanenbaum. C/C++ Users Journal, 13 (9):77–??, September 1995. CO-DEN CCUJEX. ISSN 1075-2838. See [Tan95].

Rodley:1995:TEC

[Rod95b] John Rodley. Efficient C/C++ Programming by Steve Heller. C/C++ Users Journal, 13(12): 69–??, December 1995. CODEN CCUJEX. ISSN 1075-2838. See [Hel95].

Rodley:1996:TLD

[Rod96] John Rodley. Len Dorfman’s C++ By Example by Len Dorfman. C/C++ Users Journal, 14 (2):65–??, February 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Roe:1999:WCD

[Roe99] Eric Roe. A wrapper class for dynamically linked plug-ins. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Roggenkamp:1989:MTR

[Rog89] Steve Roggenkamp. Monitor task reveals execution patterns. C Users Journal, 7(2):77–??, February 1989. ISSN 0898-9788.

Rogers:1995:LD

[Rog95] John Rogers. Locking without deadlocks. C/C++ Users Journal, 13(11):35–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Rogers:1996:OEE

[Rog96] Joey Rogers. An object-oriented expression evaluator. C/C++ Users Journal, 14(4):43–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Rollison:1992:BCL

[Rol92] Leon Rollison. Borland compiler license agreements. C Users Journal, 10(1):138–??, January 1992. ISSN 0898-9788.

Rombouts:1991:TAS

[Rom91a] Tom Rombouts. Algorithmics: The Spirit of Computing, by David Harel. C Users Journal, 9 (2):131–??, February 1991. ISSN 0898-9788. See [HR92].

Rombouts:1991:TAH

[Rom91b] Tom Rombouts. Art of Human-Computer Interface Design, The, edited by Brenda Laurel. C Users Journal, 9(4):111–??, April 1991. ISSN 0898-9788. See [LM90].

Ross:1992:SDT

[Ros92a] Ed Ross. A simple datacompression technique. C Users Journal, 10(10):113–??, October 1992. ISSN 0898-9788.

Roswell:1992:LCS

[Ros92b] Ronald L. Roswell, Jr. Learning C by studying libraries. C Users Journal, 10(5):115–??, May 1992. ISSN 0898-9788.

Rosen:1994:SSP

[Ros94a] Jeff Rosen. A simple soundex program. C/C++ Users Journal, 12(9):49–??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Ross:1994:RDC

[Ros94b] John W. Ross. Record-oriented data compression. C Users Journal, 12(4):83–??, April 1994. ISSN 0898-9788.

Ross:1995:FSS

[Ros95] John W. Ross. Fast string searching. C/C++ Users Journal, 13 (7):63–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Ross:1997:SDP

[Ros97] John W. Ross. Sockets: Down and dirty programming for the Web. C/C++ Users Journal, 15 (5):39–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Rothkin:1991:PUD

[Rot91] Steven Rothkin. PC UART device driver. C Users Journal, 9 (12):62–??, December 1991. ISSN 0898-9788.

Rowe:1995:RCM

[Row95] Robin Rowe. Designing Object-Oriented C++ Applications Using the Booch Method by Robert C. Martin. C/C++ Users Journal, 13(7):77–??, July 1995. CO-DEN CCUJEX. ISSN 1075-2838. See [Mar95].

Rees:1991:UMS

[RP91] Joel Rees and Kenneth Pugh. Use of microprocessor string and loop instructions by C compilers. C Users Journal, 9(7):??, July 1991. ISSN 0898-9788.

Reid:1992:SFD

[RP92] Christopher E. Reid and Thomas B. Passin. A simple formula for digital filtering. C Users Journal, 10 (5):46–??, May 1992. ISSN 0898-9788.

Rogers:1990:CDT

[RSSS90] William M. Rogers, Dean Lance Smith, Eric R. Smith, and Michael Schwartz. CUG322 doctor’s tools. C Users Journal, 8(8): 111–??, August 1990. ISSN 0898-9788.

Ruark:1997:TBN

[Rua97]  John Ruark.  A template-

based network implementation. C/C++ Users Journal, 15(9):33– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Rumore:1998:BST

[Rum98] Mike Rumore. A binary search tree with iterators. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Rumore:1999:LMD

[Rum99] Mike Rumore. LUP matrix decomposition using templates. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Rupley:1988:UCM

[Rup88] John Rupley. Usefulness of CUG membership. C Users Journal, 6 (2):69–??, February 1988. ISSN 0898-9788.

Reis:1991:PC

[RY91] Anthony Dos Reis and Li Yun.

Pointer-pointers in C. C Users Journal, 9(3):83–??, March 1991. ISSN 0898-9788.

Sainsbury:1992:IT

[Sai92] J. Sainsbury. Incomplete types.

C Users Journal, 10(12):129–??, December 1992. ISSN 0898-9788.

Saks:1989:PCS

[Sak89] Daniel Saks. Portability in communications software. C Users Journal, 7(11):101–??, November 1989. ISSN 0898-9788.

Saks:1990:WSH

[Sak90a] Daniel Saks. Writing standard headers: The string functions. C Users Journal, 8(1):83–??, January 1990. ISSN 0898-9788.

Saks:1990:WYO

[Sak90b] Daniel Saks. Writing your own standard headers: <stdlib.h>,

                                        <stddef.h>, <stdarg.h>, and

<limits.h>. C Users Journal, 8(3):95–??, March 1990. ISSN 0898-9788.

Saks:1991:FNO

[Sak91a] Daniel Saks. Function name overloading. C Users Journal, 9(11): 103–??, November 1991. ISSN 0898-9788.

Saks:1991:PMP

[Sak91b] Daniel Saks. Paving the migration path. C Users Journal, 9(1): 87–??, January 1991. ISSN 0898-9788.

Saks:1991:RT

[Sak91c] Daniel Saks. Reference types. C Users Journal, 9(9):91–??, September 1991. ISSN 0898-9788.

Saks:1991:RMC

[Sak91d] Daniel Saks. Rewriting modules as classes. C Users Journal, 9(7): 115–??, July 1991. ISSN 0898-9788.

Saks:1991:WYF

[Sak91e] Daniel Saks. Writing your first class. C Users Journal, 9(3):115– ??, March 1991. ISSN 0898-9788.

Saks:1991:YFC

[Sak91f] Daniel Saks. Your first class. C Users Journal, 9(5):107–??, May 1991. ISSN 0898-9788.

Saks:1992:DA

[Sak92a] Daniel Saks. Dynamic arrays.

C Users Journal, 10(11):65–??, November 1992. ISSN 0898-9788.

Saks:1992:IVA

[Sak92b] Daniel Saks. Initialization vs. assignment. C Users Journal, 10 (9):93–??, September 1992. ISSN 0898-9788.

Saks:1992:OO

[Sak92c] Daniel Saks. Operator overloading. C Users Journal, 10(1):91– ??, January 1992. ISSN 0898-9788.

Saks:1992:OOPa

[Sak92d] Daniel Saks. Operator overloading — part 2. C Users Journal, 10(3):67–??, March 1992. ISSN 0898-9788.

Saks:1992:OOPb

[Sak92e] Daniel Saks. Operator overloading — part 3. C Users Journal, 10 (5):85–??, May 1992. ISSN 0898-9788.

Saks:1992:OOPc

[Sak92f] Daniel Saks. Operator overloading — part 4. C Users Journal, 10 (7):51–??, July 1992. ISSN 0898-9788.

Saks:1993:FTO

[Sak93a] Daniel Saks. The function operator[]. C Users Journal, 11(1):93–??, January 1993. ISSN 0898-9788.

Saks:1993:IPa

[Sak93b] Daniel Saks. Inheritance, part 1. C Users Journal, 11(3):103– ??, March 1993. ISSN 0898-9788.

Saks:1993:IPb

[Sak93c] Daniel Saks. Inheritance, part 2. C Users Journal, 11(5):81–??, May 1993. ISSN 0898-9788.

Saks:1993:LN

[Sak93d] Daniel Saks. Looking up names. C Users Journal, 11(8):117–??, August 1993. ISSN 0898-9788.

Saks:1993:NC

[Sak93e] Daniel Saks. Nested classes. C Users Journal, 11(7):117–??, July 1993. ISSN 0898-9788.

Saks:1993:RLE

[Sak93f] Daniel Saks. Recent language extensions to C++. C Users Journal, 11(6):117–??, June 1993. ISSN 0898-9788.

Saks:1993:RR

[Sak93g] Daniel Saks. Rewriting and reconsidering. C Users Journal, 11 (9):99–??, September 1993. ISSN 0898-9788.

Saks:1993:TIPa

[Sak93h] Daniel Saks. Temporary inconvenience, part 1. C Users Journal, 11(10):103–??, October 1993. ISSN 0898-9788.

Saks:1993:TIPb

[Sak93i] Daniel Saks. Temporary inconvenience, part 2. C Users Journal, 11(11):109–??, November 1993. ISSN 0898-9788.

Saks:1993:VF

[Sak93j] Daniel Saks. Virtual functions.

C Users Journal, 11(12):103–??, December 1993. ISSN 0898-9788.

Saks:1994:CFPa

[Sak94a] Daniel Saks. Compilation firewalls, part 1. C Users Journal, 12 (4):99–??, April 1994. ISSN 0898-9788.

Saks:1994:CFPb

[Sak94b] Daniel Saks. Compilation firewalls, part 2. C Users Journal, 12 (5):67–??, May 1994. ISSN 0898-9788.

Saks:1994:DGCa

[Sak94c] Daniel Saks. Designing generic container classes, part 1. C Users Journal, 12(6):61–??, June 1994. ISSN 0898-9788.

Saks:1994:DGCb

[Sak94d] Daniel Saks. Designing generic container classes, part 2. C/C++ Users Journal, 12(8):77–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Saks:1994:DGCc

[Sak94e] Daniel Saks. Designing generic container classes, part 3 iterators. C/C++ Users Journal, 12(9):87– ??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Saks:1994:DGCd

[Sak94f] Daniel Saks. Designing generic container classes, part 4. C/C++ Users Journal, 12(10):81–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Saks:1994:DGCe

[Sak94g] Daniel Saks. Designing generic container classes, part 5. C/C++ Users Journal, 12(11):91–??, November 1994. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1994:DGCf

[Sak94h] Daniel Saks. Designing generic container classes, part 6: Templates. C/C++ Users Journal, 12 (12):81–??, December 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1994:HVF

[Sak94i] Daniel Saks. How virtual functions work. C Users Journal, 12 (1):61–??, January 1994. ISSN 0898-9788.

Saks:1994:OO

[Sak94j] Daniel Saks. Overloading and overriding. C Users Journal, 12 (2):87–??, February 1994. ISSN 0898-9788.

Saks:1994:RTV

[Sak94k] Daniel Saks. The return types of virtual functions. C Users Journal, 12(3):91–??, March 1994. ISSN 0898-9788.

Saks:1995:SGN

[Sak95a] Dan Saks. A sensible grammar notation. C/C++ Users Journal, 13(11):85–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:SS

[Sak95b] Dan Saks. Style and syntax.

C/C++ Users Journal, 13(10): 81–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:UCD

[Sak95c] Dan Saks. Understanding C++ declarations. C/C++ Users Journal, 13(12):87–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:CAC

[Sak95d] Daniel Saks. C++ at CD registration. C/C++ Users Journal, 13(1):83–??, January 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1995:CSR

[Sak95e] Daniel Saks. Changes in the scope rules. C/C++ Users Journal, 13(6):79–??, June 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1995:EMM

[Sak95f] Daniel Saks. Even more minor enhancements. C/C++ Users Journal, 13(5):89–??, May 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:MEC

[Sak95g] Daniel Saks. Minor enhancements to C++ as of CD registration. C/C++ Users Journal, 13 (2):101–??, February 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1995:MME

[Sak95h] Daniel Saks. More minor enhancements as of CD registration. C/C++ Users Journal, 13 (3):79–??, March 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:MCM

[Sak95i] Daniel Saks. Mutable class members. C/C++ Users Journal, 13 (4):79–??, April 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:OACb

[Sak95j] Daniel Saks. Other assorted changes, part 1. C/C++ Users Journal, 13(7):81–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:OACa

[Sak95k] Daniel Saks. Other assorted changes, part 2. C/C++ Users Journal, 13(8):81–??, August 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1995:OACc

[Sak95l] Daniel Saks. Other assorted changes, part 3. C/C++ Users Journal, 13(9):81–??, September 1995. CODEN CCUJEX. ISSN 1075-2838.

Saks:1996:CTPa

[Sak96a] Dan Saks. C++ theory and practice: Abstract declarators, part 1. C/C++ Users Journal, 14(6): 69–??, June 1996. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1996:CTPb

[Sak96b] Dan Saks. C++ theory and practice: Abstract declarators, part 2. C/C++ Users Journal, 14(7): 61–??, July 1996. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1996:CTPc

[Sak96c] Dan Saks. C++ theory and practice: Abstract declarators part 3. C/C++ Users Journal, 14(9): 61–??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Saks:1996:CTPe

[Sak96d] Dan Saks. C++ theory and practice: const as a promise. C/C++ Users Journal, 14(11): 81–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Saks:1996:CTPd

[Sak96e] Dan Saks. C++ theory and practice: Declarators, finale. C/C++ Users Journal, 14(10):79–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Saks:1996:PCDa

[Sak96f] Dan Saks. Parsing C++ declarations, part 1. C/C++ Users Journal, 14(2):77–??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Saks:1996:PCDb

[Sak96g] Dan Saks. Parsing C++ declarations, part 2. C/C++ Users Journal, 14(3):63–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Saks:1996:PGP

[Sak96h] Dan Saks. Perspectives on grammars and parsers. C/C++ Users Journal, 14(5):63–??, May 1996. CODEN CCUJEX. ISSN 1075-

Saks:1996:RPE

[Sak96i] Dan Saks. Recovering from parsing errors. C/C++ Users Journal, 14(4):63–??, April 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1996:UCD

[Sak96j] Dan Saks. Understanding C++ declarators. C/C++ Users Journal, 14(1):59–??, January 1996. CODEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPb

[Sak97a] Dan Saks. C++ theory and practice: C++ at its second CD ballot. C/C++ Users Journal, 15 (2):67–??, February 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPc

[Sak97b] Dan Saks. C++ theory and practice: Class-specific new and delete. C/C++ Users Journal, 15(3):73–??, March 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPh

[Sak97c] Dan Saks. C++ theory and practice: Maybe it wasn’t such a good idea after all. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPa

[Sak97d] Dan Saks. C++ theory and practice: new and delete. C/C++ Users Journal, 15(1):59–??, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPe

[Sak97e] Dan Saks. C++ theory and practice: Placement delete. C/C++ Users Journal, 15(5):79–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPd

[Sak97f] Dan Saks. C++ theory and practice: Placement new. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Saks:1997:CTPf

[Sak97g] Dan Saks. C++ theory and practice: The syntax of newand delete-expressions, part 1. C/C++ Users Journal, 15(6):87– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1997:CTPg

[Sak97h] Dan Saks. C++ theory and practice: The syntax of newand delete-expressions, part 2. C/C++ Users Journal, 15(7):73– ??, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1997:CTPi

[Sak97i] Dan Saks. C++ theory and practice: Work-arounds for a mistake. C/C++ Users Journal, 15(9):71– ??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Saks:1998:CTPa

[Sak98a] Dan Saks. C++ theory and practice: An introduction to namespaces. C/C++ Users Journal, 16 (1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Saks:1998:CTPb

[Sak98b] Dan Saks. C++ theory and practice: Basing style on design principles. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Saks:1998:CTPf

[Sak98c] Dan Saks. C++ theory and practice: Classes vs. namespaces. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Saks:1998:CTPc

[Sak98d] Dan Saks. C++ theory and practice: Partitioning with namespaces, part 1. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Saks:1998:CTPd

[Sak98e] Dan Saks. C++ theory and practice: Partitioning with namespaces, part 2. C/C++ Users Journal, 16(5):??, May 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1998:CTPe

[Sak98f] Dan Saks. C++ theory and practice: Partitioning with namespaces, part 3. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Saks:1999:CTPf

[Sak99a] Dan Saks. C++ theory and practice: Isolating design decisions, part 1. C/C++ Users Journal, 17(8):81–??, August 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1999:CTPg

[Sak99b] Dan Saks. C++ theory and practice: Isolating design decisions, part 2. C/C++ Users Journal, 17 (9):73–??, September 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1999:CTPa

[Sak99c] Dan Saks. C++ theory and practice: Partitioning with classes. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1999:CTPh

[Sak99d] Dan Saks. C++ theory and practice: Standard C++ as a high-level language? C/C++ Users Journal, 17(11):??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Saks:1999:CTPd

[Sak99e] Dan Saks. C++ theory and practice: Thinking deeper. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Saks:1999:CTPc

[Sak99f] Dan Saks. C++ theory and practice: Thinking deeply. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Saks:1999:CTPe

[Sak99g] Dan Saks. C++ theory and practice: Thinking even deeper. C/C++ Users Journal, 17(7):69– ??, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Saks:1999:CTPb

[Sak99h] Dan Saks. C++ theory and practice: Trimming excess fat. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Sant:1989:WDA

[San89] Joseph Sant. Writing data acquisition software. C Users Journal, 7(11):55–??, November 1989. ISSN 0898-9788.

Sanders:1993:CLR

[San93] Sid Sanders. Controlling the last row of screen on line feed from printf: The problem. C Users Journal, 11(3):124–??, March 1993. ISSN 0898-9788.

Savard:1996:SMF

[Sav96] Marco Savard. Shared memories as files. C/C++ Users Journal, 14(4):23–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Sawtell:1994:CTT

[Saw94] Christopher Sawtell. CUG395C — typing tutor. C Users Journal, 12(2):112–??, February 1994. ISSN 0898-9788.

Schechter:1989:CQG

[SB89] Dan Schechter and David Bryant. CUG282 — quip and graphics. C Users Journal, 7(2):103–??, February 1989. ISSN 0898-9788.

Schildt:1985:CME

[Sch85] Herbert Schildt. C made easy. Osborne/McGraw-Hill, Berkeley, CA, USA, 1985. ISBN 0-07-881178-3 (paperback). x + 292 pp. LCCN QA76.73.C15 S35 1985. See book review [Haw88].

Schechter:1988:TCG

[Sch88a] Dan Schechter. Computer Graphics for the IBM PC, by Leendert Ammeraal. C Users Journal, 6 (9):49–??, September 1988. ISSN 0898-9788. See [Amm87a].

Schechter:1988:HAH

[Sch88b] Dan Schechter. How to access Hercules graphics plane. C Users Journal, 6(8):81–??, August 1988. ISSN 0898-9788.

Schwaderer:1988:CPG

[Sch88c] W. David Schwaderer. C programmer’s guide to NetBIOS. Howard W. Sams, Indianapolis, IN 46268, USA, 1988. ISBN 0-672-22638-3. xvi + 314 pp. LCCN TK5105.7 .S378 1988. US$24.95. See book review [Per89].

Schaefges:1989:OGE

[Sch89a] Tom Schaefges. An objectoriented graphics editor. C Users Journal, 7(2):51–??, February 1989. ISSN 0898-9788.

Schrader:1989:AFT

[Sch89b] Denis Schrader. Automated FORTRAN-C translation saves programmer effort. C Users Journal, 7(8):55–??, August 1989. ISSN 0898-9788.

Schechter:1990:CHM

[Sch90a] Dan Schechter. CUG305 HGA Mandelbrot explorer and card games. C Users Journal, 8(3):

131–??, March 1990. ISSN 0898-9788.

Schildt:1990:AAC

[Sch90b] Herbert Schildt. The Annotated ANSI C Standard: American National Standard for Programming Languages: C. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881952-0. xvi + 219 pp. LCCN QA76.73.C15S356 1990. US$39.95. See book review [Pla94e].

Schimandle:1990:MBC

[Sch90c] James Schimandle. Microsoft BASIC’s and C’s floating point formats. C Users Journal, 8(7): 139–??, July 1990. ISSN 0898-9788.

Schimandle:1990:MCV

[Sch90d] James Schimandle. Microsoft C v5.10 memory allocation routine problems. C Users Journal, 8 (1):136–??, January 1990. ISSN 0898-9788.

Schmitt:1990:CEM

[Sch90e] David A. Schmitt. C extensions for multi-threading. C Users Journal, 8(8):33–??, August 1990. ISSN 0898-9788.

Schumann:1990:CSD

[Sch90f] Jan Schumann. CUG326 — SoftC database library. C Users Journal, 8(9):130–??, September 1990. ISSN 0898-9788.

Schell:1991:SSC

[Sch91a] Joe Schell. Some small C++ classes. C Users Journal, 9(4):

73–??, April 1991. ISSN 0898-9788.

Schermerhorn:1991:WC

[Sch91b] L. J. G. Schermerhorn. Word counting. C Users Journal, 9(6): 105–??, June 1991. ISSN 0898-9788.

Schmitt:1991:RUT

[Sch91c] David A. Schmitt. Reviving the UNIX sbrk function. C Users Journal, 9(2):97–??, February 1991. ISSN 0898-9788.

Schore:1991:OMC

[Sch91d] Michael Schore. Octree method of color matching. C Users Journal, 9(8):43–??, August 1991. ISSN 0898-9788.

Schwan:1991:CTT

[Sch91e] Eduard Schwan. CUG345 — TLC/TLP. C Users Journal, 9(7):126–??, July 1991. ISSN 0898-9788.

Schelin:1992:GSF

[Sch92a] Carl Schelin. Generating screen fonts. C Users Journal, 10(11): 84–??, November 1992. ISSN 0898-9788.

Schumann:1992:SDL

[Sch92b] Mark W. Schumann. SoftC database library. C Users Journal, 10(9):117–??, September 1992. ISSN 0898-9788.

Schwaderer:1992:CPG

[Sch92c] W. David Schwaderer. C programmer’s guide to NetBIOS, IPX, and SPX. Howard W. Sams, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30050-8. xvi + 505 pp. LCCN QA76.73.C15S38 1992. US$49.95, CDN$62.95. System requirements for computer disk: IBM-compatible PC; DOS; Microsoft C/C++ 7 or Borland C++ 3.1; NetBIOS.

Schumann:1993:UCC

[Sch93a] Mark W. Schumann. Using C/C++ with Clipper. C Users Journal, 11(12):19–??, December 1993. ISSN 0898-9788.

Schwartz:1993:EMO

[Sch93b] David W. Schwartz. An efficient method for optimizing binary trees. C Users Journal, 11 (2):45–??, February 1993. ISSN 0898-9788.

Schmidt:1995:CF

[Sch95a] Bobby Schmidt. Completing the foundation. C/C++ Users Journal, 13(12):73–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1995:LTW

[Sch95b] Bobby Schmidt. Living in two worlds. C/C++ Users Journal, 13(11):99–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:BTS

[Sch96a] Bobby Schmidt. Balancing type safety with sociability. C/C++ Users Journal, 14(2): 67–??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:BDT

[Sch96b] Bobby Schmidt. Building a data type in C and C++. C/C++ Users Journal, 14(1):69–??, January 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:CSC

[Sch96c] Bobby Schmidt. Controlling silent conversions. C/C++ Users Journal, 14(4):79–??, April 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:CBI

[Sch96d] Bobby Schmidt. Creating a Boolean inserter. C/C++ Users Journal, 14(3):79–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:LCCc

[Sch96e] Bobby Schmidt. The learning C/C++ curve: C → C++ mutations, part 3. C/C++ Users Journal, 14(8):87–??, August 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:LCCa

[Sch96f] Bobby Schmidt. The learning C/C++urve: C → C++ mutations, part 1. C/C++ Users Journal, 14(6):75–??, June 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:LCCb

[Sch96g] Bobby Schmidt. The learning C/C++urve: C → C++ mutations, part 2. C/C++ Users Journal, 14(7):79–??, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:LCCd

[Sch96h] Bobby Schmidt. The learning C/ C++urve: C → C++ mutations, part 4. C/C++ Users Journal, 14 (9):69–??, September 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:LCW

[Sch96i] Bobby Schmidt. The learning C/ C++urve: The why files, part two. C/C++ Users Journal, 14 (10):69–??, October 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:WF

[Sch96j] Bobby Schmidt. The why files. C/C++ Users Journal, 14(5):71– ??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Schroeder:1996:VSD

[Sch96k] Will Schroeder. Visualizing scientific data. C/C++ Users Journal, 14(6):33–??, June 1996. CODEN CCUJEX. ISSN 1075-2838.

Schulist:1996:SMC

[Sch96l] Stephen Schulist. A simple matrix class. C/C++ Users Journal, 14(6):27–??, June 1996. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCc

[Sch97a] Bobby Schmidt. The learning C/C++urve: And now for a word from our sponsors. C/C++ Users Journal, 15(3):81– ??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:1997:LCCa

[Sch97b] Bobby Schmidt. The learning C/C++urve: Driving you to abstraction. C/C++ Users Journal, 15(1):65–??, January 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCb

[Sch97c] Bobby Schmidt. The learning C/C++urve: Further adventures in abstraction. C/C++ Users Journal, 15(2):71–??, February 1997. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCd

[Sch97d] Bobby Schmidt. The learning C/C++urve: Into the swamp. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCe

[Sch97e] Bobby Schmidt. The learning C/C++urve: Let me say that about this. C/C++ Users Journal, 15(5):73–??, May 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCh

[Sch97f] Bobby Schmidt. The learning C/C++urve: Me and my arrow. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCf

[Sch97g] Bobby Schmidt. The learning C/C++urve: Sargasso Sea. C/C++ Users Journal, 15(6):77– ??, June 1997. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:1997:LCCi

[Sch97h] Bobby Schmidt. The learning C/C++urve: The pointer variations. C/C++ Users Journal, 15 (9):79–??, September 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:1997:LCCg

[Sch97i] Bobby Schmidt. The learning C/C++urve: Getting to the point(er). C/C++ Users Journal, 15(7):67–??, July 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCa

[Sch98a] Bobby Schmidt. The learning C/C++urve: All this and C++ too! C/C++ Users Journal, 16 (1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCg

[Sch98b] Bobby Schmidt. The learning C/C++urve: C9X VLAs. C/C++ Users Journal, 16(7):??, July 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCi

[Sch98c] Bobby Schmidt. The learning C/C++urve: C/C++urve ball. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:1998:LCCf

[Sch98d] Bobby Schmidt. The learning C/C++urve: Dis-array. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCb

[Sch98e] Bobby Schmidt. The learning C/C++urve: Morte d’autopointer. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:1998:LCCj

[Sch98f] Bobby Schmidt. The learning C/C++urve: Penumbra. C/C++ Users Journal, 16(10): 88–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCc

[Sch98g] Bobby Schmidt. The learning C/C++urve: Reflective C++. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCk

[Sch98h] Bobby Schmidt. The learning C/C++urve: Second contact. C/C++ Users Journal, 16(11): 85–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCe

[Sch98i] Bobby Schmidt. The learning C/C++urve: Spring cleaning. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCd

[Sch98j] Bobby Schmidt. The learning C/C++urve: Very small array. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:LCCh

[Sch98k] Bobby Schmidt. The learning C/C++urve: Yo quiero VLAs. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1998:UEP

[Sch98l] Bobby Schmidt. Uncaught exceptions: Phoenix rising. C/C++ Users Journal, 16(12):??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UEB

[Sch99a] Bobby Schmidt. Uncaught exceptions: Building sand castles. C/C++ Users Journal, 17(7):89– ??, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:1999:UEE

[Sch99b] Bobby Schmidt. Uncaught exceptions: Eroica. C/C++ Users Journal, 17(9):89–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UEG

[Sch99c] Bobby Schmidt. Uncaught exceptions: Great expectations. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UEJ

[Sch99d] Bobby Schmidt. Uncaught exceptions: Just say no. C/C++ Users Journal, 17(8):89–??, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UEO

[Sch99e] Bobby Schmidt. Uncaught exceptions: One small step. C/C++ Users Journal, 17(10):??, October 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UEPc

[Sch99f] Bobby Schmidt. Uncaught exceptions: Paper or plastic. C/C++ Users Journal, 17(11):??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UEPa

[Sch99g] Bobby Schmidt. Uncaught exceptions: Party like it’s 1999. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:1999:UEPb

[Sch99h] Bobby Schmidt. Uncaught exceptions: Phantom MenaC++. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UESa

[Sch99i] Bobby Schmidt. Uncaught exceptions: September song. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1999:UESb

[Sch99j] Bobby Schmidt. Uncaught exceptions: Singletons, five-liners, and multiple inclusion. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Schwartz:1999:DIB

[Sch99k] Robert Allan Schwartz. Default initialization of built-in types. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Scott:1988:CMM

[Sco88]  Michael Scott.  CUG247 —

M.I.R.A.C.L. — A multi-preci-sion arithmetic library. C Users Journal, 6(5):76–??, May 1988. ISSN 0898-9788.

Scott:1992:CI

[Sco92] Vince Scott. C inside the 8250. C Users Journal, 10(9):39–??, September 1992. ISSN 0898-9788.

Scolnick:1993:RES

[Sco93] Martin Scolnick. Random event simulation for C programmers. C Users Journal, 11(10):69–??, October 1993. ISSN 0898-9788.

Simar:1989:PYD

[SD89] Ray Simar, Jr. and Alan L. Davis. Program your DSP in C. C Users Journal, 7(5):69–??, May 1989. ISSN 0898-9788.

Seekamp:1996:PMG

[SDWB96] Chris Seekamp, Gary Domrow, Tony Wrobel, and Don Bulka. A portable multiprocess gateway. C/C++ Users Journal, 14(4):35– 41, April 1996. CODEN CCU-JEX. ISSN 1075-2838.

Searfoss:1991:EJC

[Sea91]  Glenn Searfoss.  Encoding

Japanese characters. C Users Journal, 9(7):47–??, July 1991. ISSN 0898-9788.

Sear:1998:TWN

[Sea98] Neil Sear. Template wrappers for network messages. C/C++ Users Journal, 16(11):75–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Sekiguchi:1989:CPS

[Sek89]  M. Sekiguchi.  CUG284 —

portable 8080 system. C Users Journal, 7(4):108–??, April 1989. ISSN 0898-9788.

Servies:1990:SWT

[Ser90] Tony Servies. Spiffier Windows for Turbo C. C Users Journal, 8(1):90–??, January 1990. ISSN 0898-9788.

Shammas:1992:AC

[Sha92] Namir Clement Shammas. Advanced C++. Howard W. Sams, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30158-X. xxiii + 784 pp. LCCN QA76.73.C153 S42 1992. US$39.95, CDN$49.95. See book review [Dos95].

Sheppard:1990:EYF

[She90]  Gene Sheppard.  Evaluating

your floating point library. C Users Journal, 8(8):121–??, August 1990. ISSN 0898-9788.

Shemitz:1992:MSM

[She92] Jon Shemitz. Multitasking state machines. C Users Journal, 10 (11):23–??, November 1992. ISSN 0898-9788.

Shipman:1990:DTM

[Shi90] Arthur Shipman. Debugging with two monitors. C Users Journal, 8 (12):25–??, December 1990. ISSN 0898-9788.

Shipman:1991:AMT

[Shi91] Arthur Shipman. Allocating multiple char arrays. C Users Journal, 9(5):54–??, May 1991. ISSN 0898-9788.

Sickler:1994:TIV

[Sic94] Clyde Sickler. Inside Visual C++, by David J. Kruglinksi. C/C++ Users Journal, 12(7):93– ??, July 1994. CODEN CCU-JEX. ISSN 1075-2838.

Silver:1988:CVC

[Sil88] Lee Silver. C:LINES/C:TREE (ver. 1.80) code documentation and analysis utilities. C Users Journal, 6(1):28–??, January 1988. ISSN 0898-9788.

Singleton:1993:TCT

[Sin93a]  David Singleton.  cout and

cerr for Windows. C Users Journal, 11(9):49–??, September 1993. ISSN 0898-9788.

Sink:1993:USD

[Sin93b] Eric W. Sink. Using the SEN-TINEL debugging environment. C Users Journal, 11(10):119–??, October 1993. ISSN 0898-9788.

Singleton:1994:UWM

[Sin94] David Singleton. Using Windows memory management services. C/C++ Users Journal, 12 (9):19–??, September 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Siyan:1989:PGI

[Siy89] Karanjit S. Siyan. Portability guideline issues for C. C Users Journal, 7(1):65–??, January 1989. ISSN 0898-9788.

Skelly:1991:CCO

[Ske91] Christopher Skelly. Creating C++-Like “objects” in C. C Users Journal, 9(12):109–??, December 1991. ISSN 0898-9788.

Skelly:1993:PPCa

[Ske93a]  Christopher Skelly.  Pointer

power in C and C++, part 1. C Users Journal, 11(2):93–??, February 1993. ISSN 0898-9788.

Skelly:1993:PPCb

[Ske93b] Christopher Skelly. Pointer power in C and C++, part 2. C Users Journal, 11(3):62–??, March 1993. ISSN 0898-9788.

Skelly:1994:PPM

[Ske94] Christopher Skelly. Powerful pointers to member functions. C/C++ Users Journal, 12(10): 51–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Smaha:1988:PVF

[Sma88] Steve Smaha. PC/VI faithfully replicates capabilities of UNIX original. C Users Journal, 6(5): 64–??, May 1988. ISSN 0898-9788.

Small:1995:SCT

[Sma95] John W. Small. Simulating C++ templates in C and C++. C/C++ Users Journal, 13(2):67– ??, February 1995. CODEN CCUJEX. ISSN 1075-2838.

Smeldley:1989:CCV

[Sme89] Michael Smeldley. CUG278 — CXL v4.0. C Users Journal, 7 (1):101–??, January 1989. ISSN 0898-9788.

Smith:1988:CHP

[Smi88a] Fred C. Smith. CUG236 — highly portable utilities. C Users Journal, 6(8):63–??, August 1988. ISSN 0898-9788.

Smith:1988:ULE

[Smi88b] Stanford Smith. User likes essential graphics library. C Users Journal, 6(8):35–??, August 1988. ISSN 0898-9788.

Smith:1989:TCM

[Smi89] Michael J. Smith. Turbo C: Memory Resident Utilities, Screen I/O, and Programming, by Al Stevens. C Users Journal, 7(5): 57–??, May 1989. ISSN 0898-9788.

Smith:1992:VPR

[Smi92a] Lowell Smith. A versatile plotting routine. C Users Journal, 10(11): 121–??, November 1992. ISSN 0898-9788.

Smith:1992:DMW

[Smi92b] William Smith. Debugging with macro wrappers. C Users Journal, 10(10):31–??, October 1992. ISSN 0898-9788.

Smith:1992:TCM

[Smi92c] William Smith. C Mathematical Function Handbook, The, by Louis Baker. C Users Journal, 10(9):123–??, September 1992. ISSN 0898-9788. See [Bak92].

Smith:1992:ECS

[Smi92d] William Smith. Extending C with stored SQL procedures. C Users Journal, 10(7):39–??, July 1992. ISSN 0898-9788.

Smith:1992:NCC

[Smi92e] William Smith. Number crunching in C. C Users Journal, 10 (5):37–??, May 1992. ISSN 0898-9788.

Smith:1992:PCL

[Smi92f] William Smith. Porting command line user interfaces to GUIs. C Users Journal, 10(4):32– ??, April 1992. ISSN 0898-9788.

Smith:1992:RTR

[Smi92g] William Smith. Ray tracing for rendering 2-D cross-sections of 3-D objects. C Users Journal, 10 (8):45–??, August 1992. ISSN 0898-9788.

Smith:1992:UPP

[Smi92h] William Smith. Using the PC’s parallel port for bidirectional communications. C Users Journal, 10(9):55–??, September 1992. ISSN 0898-9788.

Smith:1993:CFE

[Smi93a] Lowell Smith. Curve fitting with extrapolation. C Users Journal, 11(6):25–??, June 1993. ISSN 0898-9788.

Smith:1993:PIH

[Smi93b] Lowell Smith. Printing images to HP laser printers. C Users Journal, 11(9):137–??, September 1993. ISSN 0898-9788.

Smith:1993:C

[Smi93c] William Smith. C-clearly.

C Users Journal, 11(3):69–??, March 1993. ISSN 0898-9788.

Smith:1993:ESF

[Smi93d] William Smith. An essential string function library. C Users Journal, 11(1):72–??, January 1993. ISSN 0898-9788.

Smith:1993:LFF

[Smi93e] William Smith. A library of financial functions. C Users Journal, 11(2):51–??, February 1993. ISSN 0898-9788.

Smith:1993:VFP

[Smi93f] William Smith. Varying floatingpoint precision. C Users Journal, 11(7):87–??, July 1993. ISSN 0898-9788.

Smith:1993:WDF

[Smi93g] William Smith. Windows, DLLs, and floating point types. C Users Journal, 11(9):21–??, September 1993. ISSN 0898-9788.

Smith:1994:SFT

[Smi94a] William Smith. A short floatingpoint type in C++. C Users Journal, 12(1):23–??, January 1994. ISSN 0898-9788.

Smith:1994:CPC

[Smi94b] Thad Smith III. Coding with printable characters. C Users Journal, 12(2):51–??, February 1994. ISSN 0898-9788.

Sparks:1990:TPG

[Spa90] Robert J. Sparks. Programmer’s Guide To Turbo C 2.0, by Chao C. Chien. C Users Journal, 8 (7):83–??, July 1990. ISSN 0898-9788. See [Chi89].

Spees:1989:AEM

[Spe89a] William Spees. Accessing extended memory. C Users Journal, 7(11):125–??, November 1989. ISSN 0898-9788.

Spees:1989:ADF

[Spe89b] William Spees. AutoFlow-C draws flowcharts directly from C source files. C Users Journal, 7 (2):73–??, February 1989. ISSN 0898-9788.

Spratt:1990:SPI

[Spr90] Lindsey Spratt. A small Prolog interpreter. C Users Journal, 8 (3):133–??, March 1990. ISSN 0898-9788.

Spuler:1992:CC

[Spu92] David Spuler. Comprehensive C.

Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-156514-1. xvii + 397 pp. LCCN QA76.73.C15 S66 1992. US$7.95. See book review [Tho93].

Sicherer-Roetman:1999:IDS

[SR99a] Anneke Sicherer-Roetman. An interactive drawing surface component for C++ Builder. C/C++ Users Journal, 17(10):??, October 1999. CODEN CCUJEX. ISSN 1075-2838.

Sicherer-Roetman:1999:PRT

[SR99b] Anneke Sicherer-Roetman. Portable run-time multilanguage support with lingua. C/C++ Users Journal, 17(3):??, March 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Saks:1989:PAA

[SS89] Daniel Saks and Nancy Saks. Pointer arithmetic at memory segment boundaries. C Users Journal, 7(10):27–??, October 1989. ISSN 0898-9788.

StClair:1993:PFA

[St.93] John St. Clair. A proposal for financing American software development. C Users Journal, 11(6): 144–??, June 1993. ISSN 0898-9788.

Stanley:1989:GVN

[Sta89] John Stanley. Global variable naming in ‘portable’ code. C Users Journal, 7(7):124–??, July 1989. ISSN 0898-9788.

Stevens:1987:TCM

[Ste87] Al Stevens. Turbo C: memoryresident utilities, screen I/O, and programming techniques. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1987. ISBN 0-943518-35-0 (paperback). x + 315 pp. LCCN QA76.73.C15 S734 1987 Bar. US$24.95.

Stevens:1989:CT

[Ste89] Roger T. Stevens. Clipping techniques. C Users Journal, 7(8):49– ??, August 1989. ISSN 0898-9788.

Stevens:1990:RPH

[Ste90] Roger T. Stevens. Resident print handlers using Turbo C. C Users Journal, 8(7):49–??, July 1990. ISSN 0898-9788.

Stevens:1992:LAU

[Ste92a] John W. M. Stevens. Lexical analysis using search tries. C Users Journal, 10(4):67–??, April 1992. ISSN 0898-9788.

Stevens:1992:VMP

[Ste92b] Roger T. Stevens. A versatile menu program for Turbo C. C Users Journal, 10(4):41–??, April 1992. ISSN 0898-9788.

Stern:1993:CA

[Ste93a]  Marc Stern.  CUG390 —

ACTLIB. C Users Journal, 11(12):112–??, December 1993. ISSN 0898-9788.

Stevens:1993:GQP

[Ste93b] Roger T. Stevens. Get quick, professional plots with GNUPLOT. C Users Journal, 11(6):99–??, June 1993. ISSN 0898-9788.

Stevens:1994:CNN

[Ste94] Gregory Stevens. CUG396 — NNUTILS neural network. C Users Journal, 12(2):113–??, February 1994. ISSN 0898-9788.

Stevens:1997:IWS

[Ste97] Keith Stevens. An ISAPI Webserver extension SMTP gateway. C/C++ Users Journal, 15(5):23– ??, May 1997. CODEN CCU-JEX. ISSN 1075-2838.

Stitt:1990:PRB

[Sti90] Martin Stitt. A packetized ring buffer. C Users Journal, 8(5):33– ??, May 1990. ISSN 0898-9788.

Stimpson:1992:NIU

[Sti92a] Michael Stimpson. Numerical integration using adaptive quadrature. C Users Journal, 10(5):31– ??, May 1992. ISSN 0898-9788.

Stitt:1992:DCT

[Sti92b] Martin Stitt. Debugging: creative techniques and tools for software repair. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1992. ISBN 0-471-55829-X. xxiv + 408 pp. LCCN QA76.9.D3 S748 1992. See book review [War92b].

Stout:1995:OFN

[Sto95] Bob Stout. “olympic” filtering for noisy data. C/C++ Users Journal, 13(3):21–??, March 1995. CODEN CCUJEX. ISSN 1075-2838.

Stroustrup:1986:CPL

[Str86] Bjarne Stroustrup. The C++ programming language. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-12078-X (paperback). viii + 327 pp. LCCN QA76.73.C153 S77 1987. US$21.95. See book review [Cof88].

Stroustrup:1987:CPL

[Str87] Bjarne Stroustrup. The C++ programming language. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, reprinted with corrections edition, 1987. ISBN 0-201-12078-X (paperback). viii + 328 pp. LCCN QA 76.73 C153 S77 1987.

Strayer:1988:UCG

[Str88a] Donald M. Strayer. Using C for graphics applications on the Apple II. C Users Journal, 6(11):45– ??, November 1988. ISSN 0898-9788.

Stringer:1988:CCT

[Str88b] Roger J. Stringer. CUG263 — C wndw toolkit. C Users Journal, 6(5):75–??, May 1988. ISSN 0898-9788.

Stringer:1991:OP

[Str91a]  Roger J. Stringer.  Object-

oriented programming. C Users Journal, 9(12):101–??, December 1991. ISSN 0898-9788.

Stroustrup:1991:CPL

[Str91b] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-53992-6. xi + 669 pp. LCCN QA76.73.C15 S79 1991.

Stroustrup:1995:CPL

[Str95] Bjarne Stroustrup. The C++ programming language. Addison-Wesley, Reading, MA, USA, second, reprinted with corrections august, 1995 edition, 1995. ISBN 0-201-53992-6. xi + 699 pp. LCCN QA76.73.C15 S79 1995.

Stroustrup:1999:LSC

[Str99] Bjarne Stroustrup. Learning Standard C++ as a new language. C/C++ Users Journal, 17 (5):??, May 1999. CODEN CCU-JEX. ISSN 1075-2838.

Suereth:1993:NLE

[Sue93a] Russell Suereth. Natural language expansions for tense and number. C Users Journal, 11(6): 51–??, June 1993. ISSN 0898-9788.

Suereth:1993:NLP

[Sue93b] Russell Suereth. A natural language processor. C Users Journal, 11(4):57–??, April 1993. ISSN 0898-9788.

Suereth:1994:ECP

[Sue94] Russell Suereth. Expanding a conversation processor for time. C Users Journal, 12(4):19–??, April 1994. ISSN 0898-9788.

Sutter:1999:OAM

[Sut99a] Herb Sutter. Optimizations that aren’t (in a multithreaded world). C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Sutter:1999:UAE

[Sut99b] Herb Sutter. Using auto ptr effectively. C/C++ Users Journal, 17(10):??, October 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Swartz:1988:CBF

[Swa88] Ray Swartz. CUG155 — B-TREE and fast Fourier transform. C Users Journal, 6(9):92– ??, September 1988. ISSN 0898-9788.

Swaminathan:1994:CS

[Swa94]  Gnanasekaran  Swaminathan.

CUG400 — Socket++. C Users Journal, 12(4):121–??, April 1994. ISSN 0898-9788.

Swart:1995:TSE

[Swa95a] Bob Swart. Scientific and Engineering C++ by John Barton and Lee Nackman. C/C++ Users Journal, 13(6):57–??, June 1995. CODEN CCUJEX. ISSN 1075-2838. See [BN94].

Swart:1995:TUI

[Swa95b] Bob Swart. Uninterrupted Interrupts, by Ralf Brown. C/C++ Users Journal, 13(12):71–??, December 1995. CODEN CCUJEX. ISSN 1075-2838. See [Bro94].

Swart:1995:PCJ

[Swa95c] Bob Swart. Windows Programming Power with Custom Controls by Paul Cilwa and Jeff Duntemann. C/C++ Users Journal, 13(4):61–??, April 1995. CODEN CCUJEX. ISSN 1075-2838. See [CD94].

Swart:1995:EPS

[Swa95d] Bob Swart. EMS professional shareware libraries — utilities for C/C++. C/C++ Users Journal, 13(2):81–??, February 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Swartzendruber:1997:DTB

[Swa97] Michael Swartzendruber. Dynamic table binding with MFC CRecordset classes. C/C++ Users Journal, 15(4):??, April 1997. CODEN CCUJEX. ISSN 1075-2838.

Shah:1998:USL

[SX98] Amal Shah and Hong Xiao. Using shared libraries across platforms. C/C++ Users Journal, 16 (5):??, May 1998. CODEN CCU-JEX. ISSN 1075-2838.

Syck:1991:RRA

[Syc91] Gary Syck. Removing recursion from algorithms. C Users Journal, 9(2):25–??, February 1991. ISSN 0898-9788.

Tackett:1994:UIS

[Tac94] Jack Tackett. Using the international section of WIN.INI. C/C++ Users Journal, 12(12): 25–??, December 1994. CODEN CCUJEX. ISSN 1075-2838.

Tal:1989:SB

[Tal89] David Tal. Screen blanker. C Users Journal, 7(4):12–??, April 1989. ISSN 0898-9788.

Tal:1993:CCL

[Tal93] John Tal. CUG387 — C/C++ lost algorithms. C Users Journal, 11(11):120–??, November 1993. ISSN 0898-9788.

Tamashiro:1997:TPT

[Tam97] David Tamashiro. Three printing techniques for Windows 95 console applications. C/C++ Users Journal, 15(1):45–??, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Tanenbaum:1995:DOS

[Tan95] Andrew S. Tanenbaum. Distributed operating systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-219908-4. xvii + 614 pp. LCCN QA76.76.O63 T357 1995. See book review [Rod95a].

Taylor:1992:OPN

[Tay92] Nigel Taylor. Objective-C programming on NeXT. C Users Journal, 10(12):35–??, December 1992. ISSN 0898-9788.

Taylor:1993:UGP

[Tay93] Ian Lance Taylor. The UUCP g protocol. C Users Journal, 11(1): 63–??, January 1993. ISSN 0898-9788.

Taylor:1995:ICG

[Tay95] Bradford Taylor. An iostream class for the GPIB. C/C++ Users Journal, 13(12):53–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Taylor:1997:IT

[Tay97] David G. Taylor. Interrupt thunking. C/C++ Users Journal, 15(3):41–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Tackett:1994:CSS

[TC94] Walter Alden Tackett and Aviram Carmi. CUG401 — SGPC (simple genetic programming in C). C Users Journal, 12(4):121– ??, April 1994. ISSN 0898-9788.

Thomas:1995:IFC

[TC95] Jim Thomas and Jerome T. Coonen. An introduction to floatingpoint C extensions. C/C++ Users Journal, 13(1):49–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Tearle:1988:CGa

[Tea88a] Mel Tearle. CUG199 — GED. C Users Journal, 6(1):53–??, January 1988. ISSN 0898-9788.

Tearle:1988:CGb

[Tea88b] Mel Tearle. CUG199 — GED. C Users Journal, 6(2):57–??, February 1988. ISSN 0898-9788.

Teeter:1990:TFC

[Tee90] Eric Teeter. Translating Fortran to C. C Users Journal, 8(10): 91–??, October 1990. ISSN 0898-9788.

Tennberg:1998:CAD

[Ten98] Patrick Tennberg. Creating active data types via multithreading. C/C++ Users Journal, 16 (1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Terry:1991:OBT

[Ter91] Bruce M. Terry, Jr. Optimizing binary trees. C Users Journal, 9 (6):65–??, June 1991. ISSN 0898-9788.

Tondo:1989:CAB

[TG89] Clovis L. Tondo and Scott E. Gimpel. The C Answer Book —Solutions to the Exercises in The C Programming Language. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1989. ISBN 0-13-109653-2. 208 pp. LCCN QA76.73 C15 T66 1989. This book contains answers to the problems in [KR88].

Tharp:1988:FOP

[Tha88] Alan L. Tharp. File organization and processing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1988. ISBN 0-471-60521-2. xv + 398 pp. LCCN QA76.9.F5 T48 1988. See book review [Bur92a].

Thaler:1998:CVL

[Tha98] Les Thaler. Creating a virtual listbox with MAPI. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Thewalt:1994:CI

[The94] Chris Thewalt. CUG395A — input-edit. C Users Journal, 12 (2):112–??, February 1994. ISSN 0898-9788.

Thomas:1992:TCE

[Tho92] Kirk Thomas. C By Example, by Greg Perry. C Users Journal, 10(12):127–??, December 1992. ISSN 0898-9788. See [Per92b].

Thomas:1993:TCC

[Tho93] Kirk Thomas. Comprehensive C, by David Spuler. C Users Journal, 11(6):105–??, June 1993. ISSN 0898-9788. See [Spu92].

Tikhonov:1998:SCC

[Tik98] Alexandre Tikhonov. An SQL client class. C/C++ Users Journal, 16(10):21–??, October 1998. CODEN CCUJEX. ISSN 1075-

Troy:1989:CPL

[TK89] Douglas A. Troy and James D. Kiper. The C programming language: including ANSI C, portability, and software engineering. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-39803-X. xviii + 546 pp. LCCN QA76.73.C15 T76 1989 Bar.

Tenenbaum:1990:DSU

[TLA90] Aaron M. Tenenbaum, Yedidyah Langsam, and Moshe Augenstein. Data structures using C. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-199746-7. x + 662 pp. LCCN QA76.73.C15 T46 1990 Bar. See book review [Rid91] and later edition [LAT96].

Tomasik:1991:DBY

[Tom91a] Jerzy Tomasik. Don’t blow your stack. C Users Journal, 9(10): 73–??, October 1991. ISSN 0898-9788.

Tomasik:1991:RDF

[Tom91b] Jerzy Tomasik. Removing duplicate files across disk drives. C Users Journal, 9(2):63–??, February 1991. ISSN 0898-9788.

Tomassini:1991:PS

[Tom91c] Marco Tomassini. Programming with sockets. C Users Journal, 9 (9):39–??, September 1991. ISSN 0898-9788.

Townsend:1988:UC

[Tow88] Carl Townsend. Understanding C. Sams understanding series.

Howard W. Sams, Indianapolis, IN 46268, USA, 1988. ISBN 0-672-27278-4 (paperback). xi + 364 pp. LCCN QA76.73.C15 T68 1988. US$17.95. See book review [Ogg90c].

Taylor:1992:RVL

[TP92] Conrad Taylor and Kenneth Pugh. Reading very large integers into data structures. C Users Journal, 10(5):113–??, May 1992. ISSN 0898-9788.

Thomas:1994:EUK

[TR94] Philip K. Thomas and Shmuel Rotenstreich. Enhancing the UNIX Korn shell using predictor techniques. C Users Journal, 12(3):83–??, March 1994. ISSN 0898-9788.

Treuenfels:1994:EFV

[Tre94] Anton Treuenfels. An efficient flood visit algorithm. C/C++ Users Journal, 12(8):39–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Tribble:1993:SES

[Tri93] David R. Tribble. Some extensions to Standard C proposed. C Users Journal, 11(5):120–??, May 1993. ISSN 0898-9788.

Tsang:1990:NCS

[Tsa90] Kim Tsang. A note on comment style. C Users Journal, 8(12):91– ??, December 1990. ISSN 0898-9788.

Tsai:1994:DOC

[Tsa94] Louis Tsai. Designing an OOP compiler. C Users Journal, 12 (5):37–??, May 1994. ISSN 0898-9788.

Tucker:1996:DDW

[Tuc96] Andrew Tucker. A DBWin32 debugger for Windows. C/C++ Users Journal, 14(10):25–??, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Tucker:1998:RJC

[Tuc98] Andrew Tucker. Reading Java class files in C++. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Turner:1992:CC

[Tur92] Myron Turner. CUG364 — C-ACROSS. C Users Journal, 10 (6):124–??, June 1992. ISSN 0898-9788.

Uren:1993:SIL

[UHP93] Emmanuel Uren, Robert Howard, and Tiziana Perinotti. Software internationalization and localization: an introduction. Van Nostrand Reinhold Co., New York, NY, USA, 1993. ISBN 0-442-01498-8. xx + 300 pp. LCCN QA76.754.U84 1993. See book review [Pla94d].

Usher:1994:TCC

[Ush94] Tommy Usher. Code Complete, by Steve McConnell. C Users Journal, 12(4):123–??, April 1994. ISSN 0898-9788. See [McC93].

VanCamp:1992:QCM

[Van92] Kenneth E. Van Camp. A quadcode class for mapping. C Users Journal, 10(8):17–??, August 1992. ISSN 0898-9788.

VanCamp:1993:UWI

[Van93] Kenneth E. Van Camp. Using wrappers to improve portability of commercial libraries. C Users Journal, 11(1):35–??, January 1993. ISSN 0898-9788.

VanCamp:1995:DIU

[Van95] Kenneth E. Van Camp. Dynamic inheritance using filter classes. C/C++ Users Journal, 13(6):69– ??, June 1995. CODEN CCU-JEX. ISSN 1075-2838.

VanCamp:1996:TBQ

[Van96] Kenneth Van Camp. A templatebased quicksort. C/C++ Users Journal, 14(3):19–??, March 1996. CODEN CCUJEX. ISSN 1075-2838.

Vasilyev:1996:ICG

[Vas96] Kostya Vasilyev. Implementing a cross-platform graphics engine. C/C++ Users Journal, 14(1):19– ??, January 1996. CODEN CCU-JEX. ISSN 1075-2838.

Vibert:1989:HSR

[VC89] Jean-François Vibert and Beatrice Cointot. High speed, realtime data acquisition in C. C Users Journal, 7(1):39–??, January 1989. ISSN 0898-9788.

vandenBerg:1991:PCM

[vdB91] Frank van den Berg. Portable C for the 8051 microcontroller. C Users Journal, 9(3):59–??, March 1991. ISSN 0898-9788.

vanderPoel:1991:VDE

[vdP91] Robert van der Poel. Variable declarations and extern. C Users Journal, 9(8):127–??, August 1991. ISSN 0898-9788.

vanderPoel:1993:OCO

[vdP93] Robert van der Poel. Obfuscated C and Other Mysteries, by Don Libes. C Users Journal, 11 (10):131–??, October 1993. ISSN 0898-9788. See [Lib92].

Volkman:1993:CTD

[VE93] Victor R. Volkman and John English. Class DOSThread: A base class for multithreaded DOS programs. C Users Journal, 11(12): 113–??, December 1993. ISSN 0898-9788.

Vernon:1989:DCC

[Ver89] Vaughn Vernon. Design and coding considerations for UNIX/MS-DOS application portability. C Users Journal, 7(1):59–??, January 1989. ISSN 0898-9788.

Versaw:1990:P

[Ver90] Larry Versaw. PC-METRIC. C Users Journal, 8(1):119–??, January 1990. ISSN 0898-9788.

Vesely:1995:DE

[Ves95] Alessandro Vesely. Debugging with exceptions. C/C++ Users Journal, 13(10):19–??, October 1995. CODEN CCUJEX. ISSN 1075-2838.

Vannier:1989:CCM

[VG89a] Michael W. Vannier and Daniel Geist. CUG293 and CUG294 — 3D medical imaging: Source and 3D medical imaging: Images. C Users Journal, 7(10):123–??, October 1989. ISSN 0898-9788.

Vannier:1989:SR

[VG89b] Michael W. Vannier and Daniel Geist. Shading 3-D reconstructions. C Users Journal, 7(8):71– ??, August 1989. ISSN 0898-9788.

vonHolten:1992:PNP

[vH92] Dieter von Holten. Passing NULL pointers as function arguments. C Users Journal, 10(9):132–??, September 1992. ISSN 0898-9788.

Vidal:1992:EH

[Vid92] Carlos Vidal. Exception handling. C Users Journal, 10(9):19– ??, September 1992. ISSN 0898-9788.

Vigorita:1989:MHC

[Vig89] Howard Vigorita. MetaWare high C 386. C Users Journal, 7(2):87– ??, February 1989. ISSN 0898-9788.

Villani:1991:BES

[Vil91] Pasquale J. Villani. Building embedded systems with Turbo C. C Users Journal, 9(3):67–??, March 1991. ISSN 0898-9788.

Vila:1996:LTC

[Vil96]  Leonardo Zayas Vila.  3-D

lookup table color matching. C/C++ Users Journal, 14(11): 51–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Villani:1988:CNO

[VM88] Pasquale J. Villani and David J. MacKenzie. CUG264 — NRO and other tools. C Users Journal, 6(5):75–??, May 1988. ISSN 0898-9788.

Volkman:1988:DOF

[Vol88a] Victor R. Volkman. Debugger offers flexible macro facility. C Users Journal, 6(9):51– ??, September 1988. ISSN 0898-9788.

Volkman:1988:NPG

[Vol88b] Victor R. Volkman. New parser generator produces fast parsers. C Users Journal, 6(6):38–??, June 1988. ISSN 0898-9788.

Volkman:1989:BGB

[Vol89a]  Victor R. Volkman.  Bison:

A GNU breed of YACC. C Users Journal, 7(8):117–??, August 1989. ISSN 0898-9788.

Volkman:1989:TSC

[Vol89b] Victor R. Volkman. Small C Compiler, A, by James E. Hendrix. C Users Journal, 7(2):9–??, February 1989. ISSN 0898-9788. See [Hen88].

Volkman:1989:MLY

[Vol89c] Victor R. Volkman. MKS Lex and Yacc — A user’s view. C Users Journal, 7(7):77–??, July 1989. ISSN 0898-9788.

Volkman:1989:PDL

[Vol89d] Victor R. Volkman. Public domain Lex is fast and Flex-ible. C Users Journal, 7(10):129–??, October 1989. ISSN 0898-9788.

Volkman:1990:BTC

[Vol90a]  Victor R. Volkman.  Blaise

Turbo C tools V2.0. C Users Journal, 8(12):101–??, December 1990. ISSN 0898-9788.

Volkman:1990:CCF

[Vol90b] Victor R. Volkman. Calling C functions from SQL. C Users Journal, 8(6):83–??, June 1990. ISSN 0898-9788.

Volkman:1990:DML

[Vol90c] Victor R. Volkman. DIVVY multitasking library. C Users Journal, 8(12):100–??, December 1990. ISSN 0898-9788.

Volkman:1990:GGL

[Vol90d] Victor R. Volkman. GraphX graphics library. C Users Journal, 8(8):127–??, August 1990. ISSN 0898-9788.

Volkman:1990:HGL

[Vol90e] Victor R. Volkman. The HALO graphics library. C Users Journal, 8(3):115–??, March 1990. ISSN 0898-9788.

Volkman:1990:MGC

[Vol90f] Victor R. Volkman. Menuet:

A GUI construction set for MS-DOS applications. C Users Journal, 8(12):101–??, December 1990. ISSN 0898-9788.

Volkman:1990:MU

[Vol90g] Victor R. Volkman. MKMK utility. C Users Journal, 8(12):100– ??, December 1990. ISSN 0898-

Volkman:1990:MCM

[Vol90h]  Victor R. Volkman.  MSI’s

CodeRunneR makes your TSRs run. C Users Journal, 8(11):91– ??, November 1990. ISSN 0898-9788.

Volkman:1990:MDA

[Vol90i] Victor R. Volkman. Multitasking with the DESQview API C library. C Users Journal, 8(7):99– ??, July 1990. ISSN 0898-9788.

Volkman:1990:P

[Vol90j] Victor R. Volkman. PC-Tags. C Users Journal, 8(12):102–??, December 1990. ISSN 0898-9788.

Volkman:1990:RDP

[Vol90k] Victor R. Volkman. R and D publications needs on-line services. C Users Journal, 8(12):27–??, December 1990. ISSN 0898-9788.

Volkman:1990:SCC

[Vol90l] Victor R. Volkman. A survey of CUG C compilers. C Users Journal, 8(4):131–??, April 1990. ISSN 0898-9788.

Volkman:1991:CCT

[Vol91a] Victor R. Volkman. C communications toolkit, v1.00B, from magna carta software. C Users Journal, 9(4):103–??, April 1991. ISSN 0898-9788.

Volkman:1991:CMC

[Vol91b] Victor R. Volkman. Conquer multitasking with Conquerrent C 2.01. C Users Journal, 9(10):64– ??, October 1991. ISSN 0898-9788.

Volkman:1993:CCP

[Vol93a] Victor R. Volkman. Convert C programs into multithreaded applications. C Users Journal, 11 (4):87–??, April 1993. ISSN 0898-9788.

Volkman:1993:CJS

[Vol93b] Victor R. Volkman. CUG #381:

JPEG software. C Users Journal, 11(10):113–??, October 1993. ISSN 0898-9788.

Volkman:1993:CLE

[Vol93c] Victor R. Volkman. CUG library editors. C Users Journal, 11(5): 109–??, May 1993. ISSN 0898-9788.

Volkman:1993:JJG

[Vol93d] Victor R. Volkman. JMODEM, JPEG, and GZIP. C Users Journal, 11(9):119–??, September 1993. ISSN 0898-9788.

Volkman:1993:VG

[Vol93e] Victor R. Volkman. VGL and ghostscript. C Users Journal, 11 (10):111–??, October 1993. ISSN 0898-9788.

Volkman:1994:ACS

[Vol94a] Victor R. Volkman. AISEARCH — C++ search class library. C/C++ Users Journal, 12(11): 111–??, November 1994. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1994:CLU

[Vol94b] Victor R. Volkman. CUG417:

LIBFTP for UNIX, CUG418: Rasmol molecular graphics. C/C++ Users Journal, 12(10):

119–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1994:CR

[Vol94c] Victor R. Volkman. CUG421: RFVDEMO. C/C++ Users Journal, 12(11):107–??, November 1994. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1994:MPM

[Vol94d] Victor R. Volkman. MINED: A powerful multi-platform text editor. C Users Journal, 12(5):99– ??, May 1994. ISSN 0898-9788.

Volkman:1994:W

[Vol94e] Victor R. Volkman. WTWG. C/C++ Users Journal, 12(7):85– ??, July 1994. CODEN CCU-JEX. ISSN 1075-2838.

Volkman:1995:C

[Vol95a]  Victor R. Volkman.  Cross-

assembler. C/C++ Users Journal, 13(3):91–??, March 1995. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1995:CMC

[Vol95b] Victor R. Volkman. CUG427: Multijoy, CUG428: Pictor video lib, CUG429: KBFake. C/C++ Users Journal, 13(6):95–??, June 1995. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1995:CM

[Vol95c] Victor R. Volkman. CUG430: m68kdis. C/C++ Users Journal, 13(7):103–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1995:CRI

[Vol95d] Victor R. Volkman. CUG431:

DOS remote installable file system. C/C++ Users Journal, 13 (9):95–??, September 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Volkman:1995:CXG

[Vol95e] Victor R. Volkman. CUG434:

XYZ++ 3-D graphics. C/C++ Users Journal, 13(10):101–??, October 1995. CODEN CCU-JEX. ISSN 1075-2838.

Volkman:1995:RB

[Vol95f] Victor R. Volkman. RDCF, and BSPLINE. C/C++ Users Journal, 13(1):95–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1995:WUC

[Vol95g] Victor R. Volkman. Windows, update to CUG391: C exploration tools. C/C++ Users Journal, 13(12):94–??, December 1995. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1996:AEW

[Vol96a] Victor R. Volkman. Algorithms are everywhere on the Web. C/C++ Users Journal, 14(3):87– ??, March 1996. CODEN CCU-JEX. ISSN 1075-2838.

Volkman:1996:CUG

[Vol96b] Victor R. Volkman. A C Users’ Group update. C/C++ Users Journal, 14(6):49–??, June 1996. CODEN CCUJEX. ISSN 1075-

Volkman:1996:CSX

[Vol96c] Victor R. Volkman. C/C++ sources: X and Windows code from the Web. C/C++ Users Journal, 14(7):75–??, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1996:CSSb

[Vol96d] Victor R. Volkman. C/C++ sources: In search of C++ libraries. C/C++ Users Journal, 14(11):90–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1996:CSSa

[Vol96e] Victor R. Volkman. C/C++ sources: STL help on the Web. C/C++ Users Journal, 14(9):89– ??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1996:MG

[Vol96f] Victor R. Volkman. Micro-C and GNU. C/C++ Users Journal, 14(1):88–??, January 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Volkman:1996:VCS

[Vol96g]  Victor R. Volkman.  Visit

the client/server coffeehouse. C/C++ Users Journal, 14(5):75– ??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Volkman:1997:CCSa

[Vol97a] Victor R. Volkman. C/C++ sources: Cross-platform resources on the Web. C/C++ Users Journal, 15(1):71–??, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1997:CCSg

[Vol97b] Victor R. Volkman. C/C++ sources: Desperately seeking sources. C/C++ Users Journal, 15(7):89–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1997:CCSe

[Vol97c] Victor R. Volkman. C/C++ sources: Internet on the Internet. C/C++ Users Journal, 15(5):87– ??, May 1997. CODEN CCU-JEX. ISSN 1075-2838.

Volkman:1997:CCSc

[Vol97d] Victor R. Volkman. C/C++ sources: UNIX and database resources on the Internet. C/C++ Users Journal, 15(3):89–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1997:CSA

[Vol97e] Victor R. Volkman. CUG semiannual update. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1998:CCS

[Vol98a] Victor R. Volkman. C/C++ sources: Best of reader requests: Part III. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1998:CCU

[Vol98b] Victor R. Volkman. A C/C++ users group (CUG) library update. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1999:CCU

[Vol99] Victor Volkman. C/C++ Users Group (CUG) CD-ROM, April 1999 Edition. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Vossen:1990:RSC

[Vos90] Joseph K. Vossen. Reference and source code source for TIFF routines. C Users Journal, 8(3):38– ??, March 1990. ISSN 0898-9788.

vanHoff:1996:HJ

[vSS96] Arthur van Hoff, Sami Shaio, and Orca Starbuck. Hooked on Java: creating hot Web sites with Java applets. The Java Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48837-X (paperback), 0-201-85274-8 (computer laser optical disk). xviii + 181 pp. LCCN QA76.64.V36 1996. US$29.95. URL http:

//www.aw.com/devpress/java/

index.html. Includes CD-ROM. See [Nic96].

Wagner:1990:CC

[Wag90] Thomas Wagner. CUG330 — CTask. C Users Journal, 8(11): 126–??, November 1990. ISSN 0898-9788.

Wall:1988:TSE

[Wal88] Mary Wall. Software Engineering in C, by Peter A. Darnell and Philip E. Margolis. C Users Journal, 6(5):42–??, May 1988. ISSN 0898-9788. See [DM88].

Wall:1989:TGP

[Wal89] Mary Wall. Graphics Programming in C: A comprehensive resource for every C programmer. C Users Journal, 7(8):109–??, August 1989. ISSN 0898-9788.

Walker:1991:PNG

[Wal91] John Q. Walker II. Positioning nodes for general trees. C Users Journal, 9(2):47–62, February 1991. ISSN 0898-9788.

Wall:1992:TIC

[Wal92] Mary Wall. Illustrated C, by Leor Zolman. C Users Journal, 10(6): 119–??, June 1992. ISSN 0898-9788.

Walker:1999:EEC

[Wal99] Arthur P. Walker. enum++ — an enum class code generator. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Wampler:1996:VPG

[Wam96] Bruce E. Wampler. V: A portable GUI framework. C/C++ Users Journal, 14(8):21–??, August 1996. CODEN CCUJEX. ISSN 1075-2838.

Ward:1986:PCM

[War86] Terry A. Ward. Programming C on the Macintosh. Scott, Foresman and Company, Glenview, IL, USA, 1986. ISBN 0-673-18274-6 (paperback). xv + 411 pp. LCCN QA76.8.M3 W37 1986 Sci-Eng. US$21.95. See book review [And88b].

Ward:1988:BCR

[War88a] Robert L. Ward. Building a circular Or ring buffer with robust data encapsulation. C Users Journal, 6(5):49–??, May 1988. ISSN 0898-9788.

Ward:1988:CLDa

[War88b] Robert L. Ward. CUG library disk bugs and modifications. C Users Journal, 6(1):65–??, January 1988. ISSN 0898-9788.

Ward:1988:CLDb

[War88c] Robert L. Ward. CUG library disk formats. C Users Journal, 6(1):67–??, January 1988. ISSN 0898-9788.

Ward:1988:FNP

[War88d] Robert L. Ward. Finding ‘null pointer errors’ with MSC. C Users Journal, 6(2):27–??, February 1988. ISSN 0898-9788.

Ward:1988:FP

[War88e] Robert L. Ward. Function prototypes. C Users Journal, 6(11):95– ??, November 1988. ISSN 0898-9788.

Ward:1988:H

[War88f] Robert L. Ward. Handles. C Users Journal, 6(11):95–??, November 1988. ISSN 0898-9788.

Ward:1988:KLP

[War88g] Robert L. Ward. Kyocera laser printer as plotter. C Users Journal, 6(2):29–??, February 1988. ISSN 0898-9788.

Ward:1988:NUA

[War88h] Robert L. Ward. Notes on the use of angle brackets in Xerox’s Ventura Publisher. C Users Journal, 6(1):21–??, January 1988. ISSN 0898-9788.

Ward:1988:SMS

[War88i] Robert L. Ward. A simple menu system for MS-DOS and Unix. C Users Journal, 6(3):49–??, March 1988. ISSN 0898-9788.

Ward:1988:VTW

[War88j] Robert L. Ward. Variations on a theme, or are we having fun yet? C Users Journal, 6(1):9–??, January 1988. ISSN 0898-9788.

Ward:1989:BFS

[War89a] Robert L. Ward. Building a finite state machine. C Users Journal, 7(2):94–??, February 1989. ISSN 0898-9788.

Ward:1989:CAC

[War89b] Robert L. Ward. CUJ announces change to monthly schedule. C Users Journal, 7(11):6–??, November 1989. ISSN 0898-9788.

Ward:1989:DRC

[War89c] Robert L. Ward. Distribution rights for CUG library disks. C Users Journal, 7(10):141–??, October 1989. ISSN 0898-9788.

Ward:1989:FSM

[War89d] Robert L. Ward. A finite state machine generator. C Users Journal, 7(4):101–??, April 1989. ISSN 0898-9788.

Ward:1989:NN

[War89e] Robert L. Ward. NeXT’s niche. C Users Journal, 7(1):6–??, January 1989. ISSN 0898-9788.

Ward:1989:NR

[War89f] Robert L. Ward. No religion. C Users Journal, 7(4):6–??, April 1989. ISSN 0898-9788.

Ward:1989:RCU

[War89g] Robert L. Ward. Restrictions on commercial use of code from CUG library disks. C Users Journal, 7(4):112–??, April 1989. ISSN 0898-9788.

Ward:1989:SRC

[War89h] Robert L. Ward. Some references on compiler design. C Users Journal, 7(10):136–??, October 1989. ISSN 0898-9788.

Ward:1989:TCS

[War89i] Robert L. Ward. Training in C: Some sources. C Users Journal, 7(1):112–??, January 1989. ISSN 0898-9788.

Ward:1989:TBC

[War89j] Robert L. Ward. Two books on C compilers. C Users Journal, 7 (10):136–??, October 1989. ISSN 0898-9788.

Ward:1989:WGC

[War89k] Robert L. Ward. Writing good code. C Users Journal, 7(7):6–??, July 1989. ISSN 0898-9788.

Ward:1990:XXX

[War90a] Jonathan Ward. XMODEM, XMODEM — 1K, and XMO-DEM/CRC. C Users Journal, 8 (9):91–??, September 1990. ISSN 0898-9788.

Ward:1990:BCP

[War90b] Robert L. Ward. Bad C pun contest results. C Users Journal, 8 (6):6–??, June 1990. ISSN 0898-9788.

Ward:1990:DAS

[War90c] Robert L. Ward. Documentation:

Advice to software publishers. C Users Journal, 8(2):6–??, February 1990. ISSN 0898-9788.

Ward:1990:PPE

[War90d] Robert L. Ward. P. J. Plauger to edit C Users Journal. C Users Journal, 8(9):6–??, September 1990. ISSN 0898-9788.

Ward:1990:PSR

[War90e] Robert L. Ward. Practical schedulers for real-time applications. C Users Journal, 8(4):95–??, April 1990. ISSN 0898-9788.

Ward:1990:PAA

[War90f] Robert L. Ward. Programming after the advent: Standard C. C Users Journal, 8(7):6–??, July 1990. ISSN 0898-9788.

Ward:1990:SUB

[War90g] Robert L. Ward. Some UNIX book recommendations. C Users Journal, 8(7):138–??, July 1990. ISSN 0898-9788.

Ward:1991:TCM

[War91a] Jonathan Ward. Complete Modem Reference, The, by Gilbert Held. C Users Journal, 9(11): 121–??, November 1991. ISSN 0898-9788. See [Hel91].

Ward:1991:DIW

[War91b] Robert L. Ward. Debugging instrumentation wrappers for heap functions. C Users Journal, 9 (10):41–??, October 1991. ISSN 0898-9788.

Ware:1991:MCI

[War91c] John Ware. Mixing C with informix. C Users Journal, 9(6): 47–??, June 1991. ISSN 0898-9788.

Ward:1992:CTL

[War92a] Robert L. Ward. C as a teaching language: The case against. C Users Journal, 10(2):123–??, February 1992. ISSN 0898-9788.

Ward:1992:TDC

[War92b] Robert L. Ward. Debugging: Creative Techniques and Tools for Software Repair, by Martin Stitt. C Users Journal, 10(10):93–??, October 1992. ISSN 0898-9788. See [Sti92b].

Ward:1992:RMD

[War92c] Robert L. Ward. Reading Macintosh 3.5in disks on a PC. C Users Journal, 10(9):130–??, September 1992. ISSN 0898-9788.

Ward:1998:ASE

[War98] Warren Ward. Applying stream encryption. C/C++ Users Journal, 16(9):??, September 1998. CODEN CCUJEX. ISSN 1075-2838.

Watson:1988:PJV

[Wat88]  Paul D. Watson.  Problems

with Japanese vendors serving the U.S. C Users Journal, 6(9): 99–??, September 1988. ISSN 0898-9788.

Watson:1993:DCP

[Wat93] Robert Watson. DMA controller programming in C. C Users Journal, 11(11):35–??, November 1993. ISSN 0898-9788.

Ward:1988:EDT

[WB88] Robert L. Ward and Steven H.

Bernbaum. Editorial direction of The C Users Journal. C Users Journal, 6(5):84–??, May 1988. ISSN 0898-9788.

Ward:1989:OP

[WC89] Jonathan Ward and Gary Culp.

An Othello player. C Users Journal, 7(4):89–??, April 1989. ISSN 0898-9788.

Ward:1990:CCU

[WC90] Robert L. Ward and Phil E.

Cogar. Criticism of CUJ user reports, and some guidelines. C Users Journal, 8(1):138–??, January 1990. ISSN 0898-9788.

Weber:1993:DTR

[Web93] David Weber. Detecting text regions using cellular automata. C Users Journal, 11(8):35–??, August 1993. ISSN 0898-9788.

Weber:1994:MS

[Web94] David Weber. Minimalist scaling.

C/C++ Users Journal, 12(8):27– ??, August 1994. CODEN CCU-JEX. ISSN 1075-2838.

Weber:1996:TCC

[Web96] David Weber. Two C++ tutorials, Practical C++ Programming, by Steve Oualline, and C++ The Core Language by Gregory Satir and Doug Brown. C/C++ Users Journal, 14(3):53– ??, March 1996. CODEN CCU-JEX. ISSN 1075-2838.

Weeks:1992:GTT

[Wee92] Kevin D. Weeks. Glass-box testing — techniques for preventing software bugs. C Users Journal, 10(10):47–??, October 1992. ISSN 0898-9788.

Weinstein:1988:LPA

[Wei88a] Sydney S. Weinstein. Let the Programmer’s Assistant make control your complex, multi-file projects. C Users Journal, 6(3): 57–??, March 1988. ISSN 0898-9788.

Weinstein:1988:TMA

[Wei88b] Sydney S. Weinstein. A typical ‘make’ application. C Users Journal, 6(5):9–??, May 1988. ISSN 0898-9788.

Weinstein:1988:UST

[Wei88c] Sydney S. Weinstein. UNIX signals and terminal control parameters are more natural fix for BBS problem. C Users Journal, 6 (2):62–??, February 1988. ISSN 0898-9788.

Weinstein:1988:UMS

[Wei88d] Sydney S. Weinstein. Using ‘make’ with ‘sccs’ and with object libraries. C Users Journal, 6 (6):53–??, June 1988. ISSN 0898-9788.

Weinstein:1989:SCL

[Wei89] Sydney S. Weinstein. Source code librarians. C Users Journal, 7 (5):95–??, May 1989. ISSN 0898-9788.

Weinstein:1990:AXW

[Wei90a] Sydney S. Weinstein. Availability of X Window System for MS-DOS. C Users Journal, 8(5):71– ??, May 1990. ISSN 0898-9788.

Weinstein:1990:CBN

[Wei90b] Sydney S. Weinstein. Controversy brews on the nets. C Users Journal, 8(10):131–??, October 1990. ISSN 0898-9788.

Weinstein:1990:GT

[Wei90c] Sydney S. Weinstein. Games and tongues. C Users Journal, 8(2): 131–??, February 1990. ISSN 0898-9788.

Weinstein:1990:HGN

[Wei90d] Sydney S. Weinstein. How to get net software. C Users Journal, 8 (1):124–??, January 1990. ISSN 0898-9788.

Weinstein:1990:IWS

[Wei90e] Sydney S. Weinstein. It was summer vacation. C Users Journal, 8 (12):97–??, December 1990. ISSN 0898-9788.

Weinstein:1990:CB

[Wei90f] Sydney S. Weinstein. comp.

sources.unix is back!  C

Users Journal, 8(8):133–??, August 1990. ISSN 0898-9788.

Weinstein:1990:PGP

[Wei90g] Sydney S. Weinstein. A Perl of great price. C Users Journal, 8(4):125–??, April 1990. ISSN 0898-9788.

Weinstein:1990:PGT

[Wei90h] Sydney S. Weinstein. Portable graphics through image format conversions. C Users Journal, 8(6):129–??, June 1990. ISSN 0898-9788.

Weinstein:1991:AAY

[Wei91a] Sydney S. Weinstein. Archie at your service. C Users Journal, 9(8):113–??, August 1991. ISSN 0898-9788.

Weinstein:1991:NMN

[Wei91b] Sydney S. Weinstein. New moderator needed. C Users Journal, 9(3):129–??, March 1991. ISSN 0898-9788.

Weinstein:1991:TC

[Wei91c] Sydney S. Weinstein. comp.

sources.reviewed. C Users

Journal, 9(6):127–??, June 1991. ISSN 0898-9788.

Weinstein:1991:R

[Wei91d] Sydney S. Weinstein. The “reviews” are in. C Users Journal, 9 (10):111–??, October 1991. ISSN 0898-9788.

Weinstein:1991:SIN

[Wei91e] Sydney S. Weinstein. Special issue: Network news. C Users Journal, 9(2):109–??, February 1991. ISSN 0898-9788.

Weinstein:1991:WWN

[Wei91f] Sydney S. Weinstein. We welcome a new moderator. C Users Journal, 9(4):97–??, April 1991. ISSN 0898-9788.

Weinstein:1992:BFN

[Wei92a] Sydney S. Weinstein. Backlog forces new rules. C Users Journal, 10(10):96–??, October 1992. ISSN 0898-9788.

Weinstein:1992:FIH

[Wei92b] Sydney S. Weinstein. Finally — it’s here! C Users Journal, 10(12):113–??, December 1992. ISSN 0898-9788.

Weinstein:1992:EM

[Wei92c] Sydney S. Weinstein. Is everything miscellaneous? C Users Journal, 10(3):77–??, March 1992. ISSN 0898-9788.

Weinstein:1992:IB

[Wei92d] Sydney S. Weinstein. It’s back! C Users Journal, 10(4):99–??, April 1992. ISSN 0898-9788.

Weinstein:1992:NW

[Wei92e] Sydney S. Weinstein. Normal is wonderful. C Users Journal, 10 (8):101–??, August 1992. ISSN 0898-9788.

Weinstein:1992:WAS

[Wei92f] Sydney S. Weinstein. Where have all the sources gone. C Users Journal, 10(1):85–??, January 1992. ISSN 0898-9788.

Weinstein:1992:WGS

[Wei92g] Sydney S. Weinstein. Where to get the sources. C Users Journal, 10(2):115–??, February 1992. ISSN 0898-9788.

Weinstein:1992:YAI

[Wei92h] Sydney S. Weinstein. You asked for it — X. C Users Journal, 10 (6):63–??, June 1992. ISSN 0898-9788.

Weisfeld:1992:SPU

[Wei92i] Matt Weisfeld. Synchronizing processes using critical sections. C Users Journal, 10(9):61–??, September 1992. ISSN 0898-9788.

Weinstein:1993:ADC

[Wei93a] Sydney S. Weinstein. Automatic distribution control. C Users Journal, 11(12):97–??, December 1993. ISSN 0898-9788.

Weinstein:1993:TWU

[Wei93b] Sydney S. Weinstein. Writing a UNIX Device Driver, 2nd edition, by Janet I. Egan and Thomas J. Teixeira. C Users Journal, 11(5):79–??, May 1993. ISSN 0898-9788. See [ET92].

Weinstein:1993:FFa

[Wei93c] Sydney S. Weinstein. Flip/flop. C Users Journal, 11(6):107–??, June 1993. ISSN 0898-9788.

Weinstein:1993:FFb

[Wei93d] Sydney S. Weinstein. Flop/flip? C Users Journal, 11(8):123–??, August 1993. ISSN 0898-9788.

Weinstein:1993:IB

[Wei93e] Sydney S. Weinstein. It’s back? C Users Journal, 11(4):97–??, April 1993. ISSN 0898-9788.

Weinstein:1993:MII

[Wei93f] Sydney S. Weinstein. Magic images via imagemagick. C Users Journal, 11(10):93–??, October 1993. ISSN 0898-9788.

Weinstein:1993:NSI

[Wei93g] Sydney S. Weinstein. On the networks: Special issue: USENET network news update. C Users Journal, 11(1):1–??, January 1993. ISSN 0898-9788.

Weinstein:1993:WHA

[Wei93h] Sydney S. Weinstein. What happened — again? C Users Journal, 11(2):109–??, February 1993. ISSN 0898-9788.

Weisfeld:1993:PUI

[Wei93i] Matt Weisfeld. A portable user interface using curses. C Users Journal, 11(4):19–??, April 1993. ISSN 0898-9788.

Weisfeld:1993:SLE

[Wei93j] Matt Weisfeld. Solving linear equations using C. C Users Journal, 11(2):99–??, February 1993. ISSN 0898-9788.

Weinstein:1994:MA

[Wei94a] Sydney Weinstein. Mail assistants. C/C++ Users Journal, 12 (12):61–??, December 1994. CO-DEN CCUJEX. ISSN 1075-2838.

Weinstein:1994:CD

[Wei94b] Sydney S. Weinstein. Concurrent development. C Users Journal, 12(2):97–??, February 1994. ISSN 0898-9788.

Weinstein:1994:ECP

[Wei94c] Sydney S. Weinstein. Expensive CD player. C Users Journal, 12(4):109–??, April 1994. ISSN 0898-9788.

Weinstein:1994:SIU

[Wei94d] Sydney S. Weinstein. Special issue: USENET network news update. C Users Journal, 12(1):81– ??, January 1994. ISSN 0898-9788.

Weinstein:1994:UUB

[Wei94e] Sydney S. Weinstein. Using UNIX for a BBS. C/C++ Users Journal, 12(9):101–??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Weinstein:1994:WDE

[Wei94f] Sydney S. Weinstein. Where did everybody go? C Users Journal, 12(6):101–??, June 1994. ISSN 0898-9788.

Weisfeld:1994:ALS

[Wei94g] Matt Weisfeld. An alternative to large switch statements. C Users Journal, 12(4):67–??, April 1994. ISSN 0898-9788.

Weitzel:1994:SHI

[Wei94h] Martin Weitzel. Safe handling of input strings. C/C++ Users Journal, 12(8):97–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Weinstein:1995:LFW

[Wei95a] Sydney Weinstein. So long, fare well …. C/C++ Users Journal, 13(10):105–??, October 1995.

CODEN CCUJEX. ISSN 1075-2838.

Weinstein:1995:WGS

[Wei95b] Sydney Weinstein. Where to get the source. C/C++ Users Journal, 13(2):111–??, February 1995. CODEN CCUJEX. ISSN 1075-2838.

Weinstein:1995:RYW

[Wei95c] Sydney S. Weinstein. Relate your way through the storm. C/C++ Users Journal, 13(4):91–??, April 1995. CODEN CCUJEX. ISSN 1075-2838.

Weinstein:1995:SS

[Wei95d] Sydney S. Weinstein. S’kill is S’nice. C/C++ Users Journal, 13(7):99–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Weisfeld:1995:SCU

[Wei95e] Matt Weisfeld. A synchronization class using critical-sections. C/C++ Users Journal, 13(11): 45–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Weisfeld:1996:BCL

[Wei96] Matt Weisfeld. Building command lines for child processes. C/C++ Users Journal, 14(2):47– ??, February 1996. CODEN CCUJEX. ISSN 1075-2838.

Weller:1990:FMA

[Wel90] Steve Weller. Fast memory allocation scheme. C Users Journal, 8(4):103–??, April 1990. ISSN 0898-9788.

Wells:1991:EOU

[Wel91] Bruce Wells. English as the official U.S. language. C Users Journal, 9(8):124–??, August 1991. ISSN 0898-9788.

Welsh:1993:DES

[Wel93] Tom Welsh. Debugging embedded systems. C Users Journal, 11 (10):19–??, October 1993. ISSN 0898-9788.

Welch:1994:CSB

[Wel94a] David Welch. CUG413A — sound blaster tools. C/C++ Users Journal, 12(8):87–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Welstead:1994:SLD

[Wel94b] Steve Welstead. Scrolling list dialog for scientific programming. C Users Journal, 12(4):51–??, April 1994. ISSN 0898-9788.

Welstead:1995:DOL

[Wel95]  Steve Welstead.  Data ob-

ject list dialog for Windows. C/C++ Users Journal, 13(9):23– ??, September 1995. CODEN CCUJEX. ISSN 1075-2838.

Welstead:1996:CCM

[Wel96] Steve Welstead. C++ classes for MDI Windows management. C/C++ Users Journal, 14(11): 41–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Welstead:1999:JOL

[Wel99] Steve Welstead. A Java object list dialog. C/C++ Users Journal, 17(1):??, January 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Wendel:1997:PTC

[Wen97] J. David Wendel. A polled timer class. C/C++ Users Journal, 15 (2):25–??, February 1997. CO-DEN CCUJEX. ISSN 1075-2838.

Weiskamp:1990:CCP

[WF90] Keith Weiskamp and Bryan Flamig. The complete C++ primer. Academic Press Inc., New York, USA, 1990. ISBN 0-12-742687-6. xvii + 524 pp. LCCN QA76.73.C153 W45 1990. See book review [Hob91a]. Includes features of the new version 2.0.

Weiskamp:1992:CCP

[WF92] Keith Weiskamp and Bryan Flamig. The complete C++ primer. Academic Press Inc., New York, USA, second edition, 1992. ISBN 0-12-742686-8. xvii + 540 pp. LCCN QA76.73.C153 W45 1992. Includes features of the new version 2.1 and templates.

Ward:1988:CNC

[WH88] Robert L. Ward and Kenji Hino.

CUG265 — new ‘cpio’ installation kit supports system independent file transfer. C Users Journal, 6(6):8–??, June 1988. ISSN 0898-9788.

Ward:1989:PMX

[WH89] Robert L. Ward and Kenji Hino.

Porting MicroPlox to XENIX. C Users Journal, 7(5):118–??, May 1989. ISSN 0898-9788.

Wheeler:1997:UVX

[Whe97] Mark Wheeler. Using VC++ 4.x with WinCGI. C/C++ Users Journal, 15(5):33–??, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Whitten:1988:DDI

[Whi88] Robert Whitten. ‘disk doctor’ illustrates direct control of PC floppy drive. C Users Journal, 6 (9):17–??, September 1988. ISSN 0898-9788.

White:1989:CMH

[Whi89a] Eric White. Controlling the malloc heap. C Users Journal, 7(2):45–??, February 1989. ISSN 0898-9788.

White:1989:PIS

[Whi89b] Eric White. A portably implemented stack checking function. C Users Journal, 7(4):69– ??, April 1989. ISSN 0898-9788.

White:1989:SCU

[Whi89c] Eric White. Serial communications for UNIX. C Users Journal, 7(7):41–??, July 1989. ISSN 0898-9788.

White:1990:OPP

[Whi90a] Eric White. Object-oriented programming as a programming style. C Users Journal, 8(2):43– ??, February 1990. ISSN 0898-9788.

Whitten:1990:ETD

[Whi90b] Robert Whitten. External tools for debugging C. C Users Journal, 8(4):89–??, April 1990. ISSN 0898-9788.

White:1991:COC

[Whi91] Ronald G. White. Copy-on-write objects for C++. C Users Journal, 9(8):71–??, August 1991. ISSN 0898-9788.

Whitley:1992:CG

[Whi92] Darrell Whitley. CUG369 — genitor. C Users Journal, 10(11):74– ??, November 1992. ISSN 0898-9788.

Whitney:1998:SLC

[Whi98] Bill Whitney. Skip lists in C++.

C/C++ Users Journal, 16(11): 67–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Willers:1990:CAC

[Wil90] H. G. Willers. CUG316 AS8 cross assembler. C Users Journal, 8(6): 127–??, June 1990. ISSN 0898-9788.

Wilhelms:1991:SSF

[Wil91]  Gerhard Wilhelms.  Source

of stroked fonts for Borland’s graphic interface. C Users Journal, 9(5):136–??, May 1991. ISSN 0898-9788.

Williams:1992:CSC

[Wil92a] Al Williams. A console stream class for Borland C++. C Users Journal, 10(1):125–??, January 1992. ISSN 0898-9788.

Williams:1992:MCI

[Wil92b] Al Williams. Manipulators for console I/O. C Users Journal, 10 (11):45–??, November 1992. ISSN 0898-9788.

Wilde:1993:TYO

[Wil93a] Norman Wilde. Testing your objects. C Users Journal, 11(5):25– ??, May 1993. ISSN 0898-9788.

Wilhelms:1993:STP

[Wil93b] Gerhard Wilhelms. Source of point editor. C Users Journal, 11(1):129–??, January 1993. ISSN 0898-9788.

Wilt:1993:TC

[Wil93c] Nicholas Wilt. Templates in C++. C Users Journal, 11(5):33– ??, May 1993. ISSN 0898-9788.

Wilde:1994:DUI

[Wil94] Norman Wilde. Dealing with uncertain inputs: Exploratory software engineering. C/C++ Users Journal, 12(7):25–??, July 1994. CODEN CCUJEX. ISSN 1075-2838.

Williams:1995:MMM

[Wil95a] K. B. Williams. Magic matrix multiply. C/C++ Users Journal, 13(6):19–??, June 1995. CODEN CCUJEX. ISSN 1075-2838.

Williams:1995:TSF

[Wil95b] K. B. Williams. Testing sort functions. C/C++ Users Journal, 13 (7):33–??, July 1995. CODEN CCUJEX. ISSN 1075-2838.

Williams:1996:SSE

[Wil96a] K. B. Williams. A simple simultaneous equation solver. C/C++ Users Journal, 14(1):45–??, January 1996. CODEN CCUJEX. ISSN 1075-2838.

Williams:1996:TMF

[Wil96b] K. B. Williams. Testing math functions: When requirements are tight, we must carefully examine all potential sources of error. Make sure your math library isn’t the weak link in the chain. C/C++ Users Journal, 14(12): 49–54, 58–65, December 1996. CODEN CCUJEX. ISSN 1075-2838. Describes a package that extends the Cody-Waite-Plauger work on the ELEFUNT package for the testing of the elementary functions, including the inverse hyperbolic functions, cube root, and Bessel functions of the first and second kinds. The C++ package implements 192-bit extended precision versions of all of the functions, so that accurate results are available for comparison with the normal double-precision results.

Williams:1996:VMA

[Wil96c] K. B. Williams. Versatile matrix addition and subtractions. C/C++ Users Journal, 14(5):51– ??, May 1996. CODEN CCU-JEX. ISSN 1075-2838.

Williams:1998:DMC

[Wil98] Jim Williams. Detecting memory corruption with dog tags. C/C++ Users Journal, 16(5):??, May 1998. CODEN CCUJEX. ISSN 1075-2838.

Wiorek:1988:ACC

[Wio88a] Glen Wiorek. Availability of compilers for commodores. C Users Journal, 6(6):72–??, June 1988. ISSN 0898-9788.

Wiorek:1988:RLA

[Wio88b] Glen Wiorek. A real ‘lint’ for Amiga users. C Users Journal, 6 (11):25–??, November 1988. ISSN 0898-9788.

Wiorek:1988:RFC

[Wio88c] Glen Wiorek. Reviewer finds C workshop much improved. C Users Journal, 6(2):55–??, February 1988. ISSN 0898-9788.

Wiorek:1988:SSN

[Wio88d] Glen Wiorek. Some sources on natural language parsing. C Users Journal, 6(6):71–??, June 1988. ISSN 0898-9788.

Wiorek:1989:TCP

[Wio89] Glen Wiorek. C Programmer’s Guide to Serial Communications, The, by Joe Campbell. C Users Journal, 7(1):15–??, January 1989. ISSN 0898-9788. See book review [Cam87].

Wisehart:1997:SFS

[Wis97] Daniel Wisehart. Software for fail-safe applications. C/C++ Users Journal, 15(2):41–??, February 1997. CODEN CCU-JEX. ISSN 1075-2838.

Wisehart:1999:DES

[Wis99] Daniel J. Wisehart. Debugging embedded systems. C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Withers:1988:CME

[Wit88] Robert Withers. CUG259 — MS-DOS execution profiler. C Users Journal, 6(3):66–??, March 1988. ISSN 0898-9788.

Withers:1990:CG

[Wit90a] Robert Withers. CUG333 gAWK. C Users Journal, 8(12): 126–??, December 1990. ISSN 0898-9788.

Withers:1990:DGV

[Wit90b] Robert Withers. Declaring global variables. C Users Journal, 8 (12):93–??, December 1990. ISSN 0898-9788.

Withers:1990:OAP

[Wit90c] Robert Withers. OS/2 anonymous pipes. C Users Journal, 8 (7):31–??, July 1990. ISSN 0898-9788.

Withers:1990:OIC

[Wit90d] Robert Withers. OS/2 interprocess communication features. C Users Journal, 8(11):35–??, November 1990. ISSN 0898-9788.

Wittenberg:1995:LPC

[Wit95] Lee Wittenberg. Literate programming in C and C++ using CWEB. C/C++ Users Journal, 13(5):55, 57–63, May 1995. CO-DEN CCUJEX. ISSN 1075-2838.

Wagner-Krankel:1992:SPC

[WK92] Alan Wagner-Krankel. Smart pointers in C++. C Users Journal, 10(8):77–??, August 1992. ISSN 0898-9788.

Wingo:1994:PMF

[WL94] Scot Wingo and Louis Lu. Porting Microsoft’s foundation class library to UNIX. C Users Journal, 12(1):55–??, January 1994. ISSN 0898-9788.

Ward:1989:LMS

[WO89] Robert L. Ward and David C. Oshel. Library, macros and screen editor produce general purpose forms engine. C Users Journal, 7 (2):105–??, February 1989. ISSN 0898-9788.

Wolff:1994:CME

[Wol94] Thomas Wolff. CUG399 — MINED editor. C Users Journal, 12(3):118–??, March 1994. ISSN 0898-9788.

Wong:1993:PPP

[Won93] William Wong. Plug and play programming: an object-oriented construction kit. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1993. ISBN 1-55851-302-7. xi + 484 pp. LCCN QA76.64 .W663 1993. US$39.95, CDN$49.95. See book review [Nel95].

Woodruff:1997:SCU

[Woo97] Eric Woodruff. Serial communication under Win32. C/C++ Users Journal, 15(7):43–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

Withers:1988:CCI

[WP88] Robert Withers and Jeffrey D. Pipkins. CUG259 — console I/O and withers tools. C Users Journal, 6(3):65–??, March 1988. ISSN 0898-9788.

Waite:1984:CPP

[WPM84] Mitchell Waite, Stephen Prata, and Donald Martin. C primer plus: User friendly guide to the C programming language. Howard W. Sams, Indianapolis, IN 46268, USA, 1984. ISBN 0-672-22090-3 (paperback). 531 pp. LCCN QA76.73.C15 W35 1984. US$19.95. See book review [Pav89].

Waite:1987:CPP

[WPM87] Mitchell Waite, Stephen Prata, and Donald Martin. C primer plus. Howard W. Sams, Indianapolis, IN 46268, USA, revised edition, 1987. ISBN 0-672-22582-4 (paperback). xvii + 558 pp. LCCN QA76.73.C15 W35 1987. US$23.95.

Winroth:1993:EHC

[WR93] Harald Winroth and Matti Rendahl. Exception handling in C. C Users Journal, 11(10):33–??, October 1993. ISSN 0898-9788.

Wright:1990:CFL

[Wri90a] Maynard A. Wright. Complex function library. C Users Journal, 8(9):45–??, September 1990. ISSN 0898-9788.

Wright:1990:PXC

[Wri90b] Maynard A. Wright. Problems with X.25’s CRC-16 code. C Users Journal, 8(8):138–??, August 1990. ISSN 0898-9788.

Wheaton-Smith:1990:CC

[WS90] Simon Wheaton-Smith. C++ criticized. C Users Journal, 8 (2):137–??, February 1990. ISSN 0898-9788.

Ward:1988:EW

[WW88] Robert L. Ward and Donna Stucky Ward. An excited welcome! C Users Journal, 6(1):7–??, January 1988. ISSN 0898-9788.

Xiao:1997:UTA

[Xia97] Hong Xiao. Using templates across UNIX platforms. C/C++ Users Journal, 15(3):33–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Young:1991:CEE

[YEA91] Scott Young and Marwan El-AUGI. CUG336 — EGAPAL/ EDIPAL. C Users Journal, 9(3): 133–??, March 1991. ISSN 0898-9788.

Yehle:1990:CM

[Yeh90] James Yehle. CUG312 — makemaker. C Users Journal, 8(5): 123–??, May 1990. ISSN 0898-9788.

Yokoyama:1988:CYM

[Yok88] Michael M. Yokoyama. CUG254 — yacc for Microsoft C. C Users Journal, 6(1):53–??, January 1988. ISSN 0898-9788.

Yokoyama:1989:LPU

[Yok89] Michael M. Yokoyama. Legal problems from using CUG library disk code. C Users Journal, 7(4):

112–??, April 1989. ISSN 0898-9788.

York:1996:HTP

[Yor96] Kyle A. York. High-speed transfers on a PC parallel port. C/C++ Users Journal, 14(11): 69–??, November 1996. CODEN CCUJEX. ISSN 1075-2838.

Young:1996:TSB

[You96] Warren Young. Two STL books.

C/C++ Users Journal, 14(9):29– ??, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Yuen:1994:HTW

[Yue94] Andy Yuen. Handling timeconsuming Windows tasks. C Users Journal, 12(1):35–??, January 1994. ISSN 0898-9788.

Zeidler:1991:DFC

[Zei91] Steve Zeidler. Doing fractions in C++. C Users Journal, 9(11):33– ??, November 1991. ISSN 0898-9788.

Zeke:1995:DCG

[Zek95] Laszlo Zeke. Designing a crossplatform GUI. C/C++ Users Journal, 13(4):41–??, April 1995. CODEN CCUJEX. ISSN 1075-2838.

Zenier:1991:CFC

[Zen91]  Mark Zenier.  CUG335 —

Frankenstein cross assemblers. C Users Journal, 9(3):133–??, March 1991. ISSN 0898-9788.

Zerman:1989:DFS

[Zer89] Leonard Zerman. Determining free space under MS-DOS. C Users Journal, 7(7):23–??, July 1989. ISSN 0898-9788.

Zia:1990:CVG

[Zia90] Ismail Zia. CUG325 — VGA graphics library. C Users Journal, 8(9):130–??, September 1990. ISSN 0898-9788.

Zimmerman:1991:TST

[Zim91] Dan Zimmerman. Software Testing Techniques. C Users Journal, 9(6):107–??, June 1991. ISSN 0898-9788.

Zmola:1996:DWP

[Zmo96] Carl Zmola. Dragto–A Windows programmer’s utility. C/C++ Users Journal, 14(11):23–??, November 1996. CODEN CCU-JEX. ISSN 1075-2838.

Zolman:1990:AFIb

[Zol90a] Leor Zolman. An applied file I/O tutorial: Text-based disk routines. C Users Journal, 8(6):93– ??, June 1990. ISSN 0898-9788.

Zolman:1990:AFIa

[Zol90b] Leor Zolman. An applied file I/O tutorial: The mini-database system. C Users Journal, 8(5):105– ??, May 1990. ISSN 0898-9788.

Zolman:1990:AFIc

[Zol90c] Leor Zolman. An applied file I/O tutorial: Using binary file I/O. C Users Journal, 8(7):123–??, July 1990. ISSN 0898-9788.

Zolman:1990:PCF

[Zol90d] Leor Zolman. A practical C file I/O tutorial: A mini-database program. C Users Journal, 8(4):

69–??, April 1990. ISSN 0898-9788.

Zolman:1990:TMD

[Zol90e] Leor Zolman. Tools for MS-DOS directory navigation. C Users Journal, 8(2):113–??, February 1990. ISSN 0898-9788.

Zolman:1991:LSM

[Zol91a] Leor Zolman. A login shell for MS-DOS. C Users Journal, 9 (2):127–??, February 1991. ISSN 0898-9788.

Zolman:1991:PJM

[Zol91b] Leor Zolman. A portable journal manager. C Users Journal, 9(6): 109–??, June 1991. ISSN 0898-9788.

Zolman:1991:PCLa

[Zol91c] Leor Zolman. Processing code listings for publication. C Users Journal, 9(8):105–??, August 1991. ISSN 0898-9788.

Zolman:1991:PCLb

[Zol91d] Leor Zolman. Processing code listings for publication — part 2. C Users Journal, 9(10):125–??, October 1991. ISSN 0898-9788.

Zolman:1992:PMCa

[Zol92a] Leor Zolman. A portable menu compiler. C Users Journal, 10 (1):102–??, January 1992. ISSN 0898-9788.

Zolman:1992:PMCb

[Zol92b] Leor Zolman. A portable menu compiler — part 2. C Users Journal, 10(3):87–??, March 1992. ISSN 0898-9788.

Zolman:1992:PMCc

[Zol92c] Leor Zolman. A portable menu compiler — part 3. C Users Journal, 10(5):95–??, May 1992. ISSN 0898-9788.

Zolman:1992:PMCd

[Zol92d] Leor Zolman. A portable menu compiler — part 4. C Users Journal, 10(6):101–??, June 1992. ISSN 0898-9788.

Zolman:1992:PMCe

[Zol92e] Leor Zolman. A portable menu compiler — part 5: CMENU. C Users Journal, 10(8):116–??, August 1992. ISSN 0898-9788.

Zwakenberg:1991:IAC

[Zwa91] Hans G. Zwakenberg. Inexact alphanumeric comparisons. C Users Journal, 9(5):127–??, May 1991. ISSN 0898-9788.